A Bibliography of Publications in ACM SIGPLAN Notices, 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 3.58

Title word cross-reference

(R) [505]. 1 [2376]. 3 [862, 2835]. 7 × 24 × 365 [7]. ^{++} [1416]. ^{5} [2060]. _{light} [2795]. A ∗ BN [445]. α [2321, 217]. D^{2}R [812]. J^{+} = J [1006]. k [986]. k > 1 [1383]. L [1885]. λ [1887, 1414, 1415, 377, 1806]. N [2343]. ∥ [786]. π [2775, 2643, 965, 2777]. R_{Lin} [999].

-body [2343]. -calculus [1887, 2775, 2653, 2643, 1414, 1415, 377, 2777, 2665]. -flow [1806]. -limiting [986]. -metric [2321]. -nodes [2694]. -valued [2835].

.. [1359]. ../SIGPLAN/Hotlist [1359].

//www.cs.rutgers.edu/pldi99= [2252].

/streams [1242].

0 [800]. ’00 [2900, 2302].

1 [1505]. 10514-1 [1510, 1511]. 15th [715]. 1972 [715]. 1990 [276, 195]. 1994 [2874, 2872, 2875, 2873]. 1996 [2887, 1497, 1498]. 1997 [1719]. 1999 [2143, 2252].

2 [468, 226, 1916, 244, 208, 90, 666, 456, 1510, 1883, 1511, 1401, 817, 500, 1605]. 2000 [1928, 2827]. 21st [2870, 2303]. 22nd [2877]. 23rd [2885]. 24th [2888]. 25th [2892]. 26th [2901]. 27th [2302]. 2nd [2865, 1145, 381].

3 [526, 1205, 138]. 32-bit [220]. 3D [863, 844]. 3rd [1874].

4.0 [1279]. 4th [1719, 1119].

5 [1997]. 51 [2448]. 5th [1593].

6 [2040]. 6000 [272]. 64-Bit [599].

80 [560, 146]. 8X [49].

’90 [2852, 2853, 2858, 232, 502, 2440, 1662]. 90/95/HPF [2440]. ’91 [2855, 2856, 2858]. ’92 [2862, 2861, 569, 1323]. ’93 [2867, 1068]. ’94 [2874, 2870, 2875]. ’95 [2877, 2878, 2879, 12, 2883, 2882]. ’96 [2885, 2886, 2887, 1370, 1599, 1600]. ’97 [2888, 2889, 1481, 1475, 1652, 1935, 1875, 1936]. ’98 [2892, 2893, 1947, 1946, 1961, 2053, 2434, 2435, 2037]. ’99 [2895, 2898, 2899, 2901, 2142, 2252, 2052, 2039, 2305]. 9X [837, 1268].

= [1386, 1680, 2767]. =J [1007].

A*B [446]. ABC [931, 223]. Abotec [809]. absorbing [2414]. Abstract [643, 731, 651, 357, 1618, 260, 642, 2384, 2606, 636, 650, 2359, 128, 2360, 1254, 2520, 2525, 1792, 2668, 2583, 2782, 224, 2417, 1759, 1460, 364, 505, 2370, 1022, 2361, 2192, 362, 570, 2657, 1807, 2620, 2566, 2825, 2799, 2330, 2548, 2607, 2419, 2422, 91, 2427, 617]. Abstraction [1562, 274, 1724, 2604, 2605, 907, 517, 2357, 837, 2842, 86, 1723]. abstraction-oriented [837]. Abstractions [516, 1069, 644, 1766, 2191]. Abstracts [195]. AC [1245]. Accelerating [2132]. accept [80]. acceptance [2491]. accepted [880, 881, 420]. Access [993, 257, 982, 428, 2328, 1252, 1323, 559, 266, 1982, 1245, 291, 55, 909, 910, 1114, 1115, 409, 2498]. Access/Execute [257]. Accesses [982]. accommodation [730]. According [2024]. account [2687, 2831]. Accounting [2068]. Accuracy [327, 542, 902, 1108]. Accurate [1894, 1201, 1895, 974, 2758]. Accurately [1431, 109, 111]. Ace [1736]. Achieve [1727, 952, 771, 772]. Achievements [1848]. Acknowledgements [700]. ACM [2143, 2448, 1874, 2876, 2889, 2892, 2893, 2895, 1378, 1379, 2142, 2252, 2304, 2302, 2880, 2881, 1164, 2434, 2435, 2858, 1361, 1362, 1810, 2883]. ACM/SIGAda [2434, 2435]. ACOOL [1144]. acoustic [2495]. Acquiring [801]. Acquisition [1562, 1143]. Across [2467, 1177, 2647, 102]. Active [2117, 1055, 399, 2167, 866, 826, 1260, 1263]. Activities [2208, 2431, 1387, 1944, 1390, 1391, 1361, 1371, 2248]. Actor [2022]. Actors [707, 168, 173, 1505]. ActorSpace [766]. Ada [439, 440, 2445, 238, 684, 1347, 429, 946, 947, 837, 462, 682, 1268]. Ada-based [2445]. Ada95 [1943]. Adaptable [25, 196]. Adaptation [1617]. Adapter [2116]. adapters [2664]. Adapting [1529, 710]. A}daptive [2304, 1681, 2473, 25, 2075, 2076, 1040, 2131, 2349]. Adaptors [1060]. Addendum [2879]. Adding [1826, 2084, 1017, 1178]. Additions [1511]. Address [2127, 1021, 1755, 1868, 1516, 1752, 547, 550, 971]. Address-Based [1516]. addresses [783]. addressing [1117, 1782, 1783]. ADL [1011]. ADT [2762, 1017]. Advance [1612, 2252]. advantages [1107]. Adventure [2364]. Adverbial [154]. aesthetics [1657]. affine [2773]. Affordable [1793, 2365]. After [1673, 1082]. against [493]. Age [2411]. Age-based [2411]. Agenda [819]. Agent [965, 1922, 966, 2405, 1153]. Agent-oriented [1922]. Agents [2473, 1842, 166, 2834]. Aggregate [2836]. aggregates [73, 75]. Aggregation [1737]. Aggressive [1686]. Agora [1125]. Agreement [2261, 2262]. Ahead [1524, 1525]. AI [418, 1346, 1492]. Aid [436]. Aircraft [1901]. Akikazu [1169]. Alan [815]. ALBA [707]. Albuquerque [2860, 2867]. Algebra [2483, 2501].

Algebraic [2571, 2623, 1293, 2478, 2551, 930]. ALGOL [685, 674, 2612]. Algorithm [1701, 718, 1252, 2153, 1677, 1977, 1402, 2244, 2643, 1885, 2515, 515, 2528, 541, 2673, 2780, 2711, 2694, 282]. Algorithmic [2059, 2060, 2601, 311, 278]. Algorithms [2117, 2326, 1757, 423, 2469, 251, 1238, 1231, 1825, 1254, 2690, 385, 55, 2666, 338, 2718, 2341, 1088, 355, 331, 2347, 786, 1113, 1261]. Alias [2072, 2797, 1987, 2695, 730, 986, 2821, 1194, 1195]. aliases [2627]. Aliasing [825, 2027, 2526, 412, 2553, 515, 732]. Alida [705]. Alignment [646, 1690, 1732, 1109, 775, 2609, 2408, 1257, 1691]. alive [2506]. All-Software [1752]. allegro [339]. Alloc [459]. allocated [2526]. Allocation [256, 1676, 2067, 1704, 991, 1990, 1727, 631, 2638, 2695, 747, 1207, 104, 299, 746, 758, 759, 2743, 2246, 1703, 103, 2610, 755, 520, 519, 102, 1096, 1483]. Allocation-intensive [631]. ALLOY [118]. almost [2726]. AlphaAXP [1596]. Alphonse [517]. Also [1658]. Alternative [2513, 468, 775, 1384, 200]. Amadeus [839]. ambients [2833]. Amdahl [1748]. Amendment [1953]. among [737]. amortized [1455]. Analogy [448, 2518]. Analyses [2031, 1762, 286, 287, 2759, 2628]. Analysis [2894, 1457, 1995, 2269, 2450, 1983, 2142, 1572, 2276, 57, 1427, 129, 1526, 603, 271, 1680, 1618, 2382, 1432, 1570, 1987, 2029, 987, 1697, 5, 1349, 2130, 1685, 1746, 2163, 1852, 1986, 1700, 2156, 442, 2316, 1621, 421, 15, 2437, 1522, 451, 2032, 64, 1623, 2331, 1446, 1426, 1158, 2277, 2196, 1629, 967, 1895, 636, 1041, 2026, 2113, 1447, 983, 1042, 998, 1678, 1699, 1633, 992, 2278, 2027, 2549, 1212, 2695, 951, 1292, 2739, 1768, 1758, 2786, 2798, 2387, 2814, 729, 2525, 1333, 115, 2386, 2587, 876, 2645, 2740, 1251, 357, 2417, 2745, 2813, 117, 2797, 986, 2785, 2640]. analysis [733, 147, 988, 2514, 1698, 2390, 2724, 2806, 877, 2750, 2666, 2700, 405, 1774, 1773, 516, 997, 1894, 849, 850, 366, 2722, 2821, 734, 735, 1334, 1336, 758, 759, 732, 2424, 1335, 2394, 2528, 2675, 780, 781, 280, 2608, 2436, 54, 1622, 2348, 2429, 2783, 1326, 1327, 2720, 368, 504, 2848, 2836, 2693, 903, 904, 1194, 1195, 455, 474, 1802, 2725, 2835, 354, 768, 2799, 2430, 2527, 2548, 1454, 1767, 2758, 369, 2726, 2353, 503, 2618, 2398, 1192, 1193, 174, 851]. analysis-level [877]. analytical [338]. analyze [513]. Analyzer [441]. analyzers [831]. Analyzing [1533, 1835, 2669, 1824]. Anchored [1029]. And-and [322]. And-Parallelism [720, 938, 939, 721]. And/Or [720, 721, 938, 939]. AND/OR-parallelism [939, 938]. ANDF [1163]. Andorra [322]. angry [1933, 1934]. animal [623]. animated [862, 863]. Animating [424]. animation [1785]. Annealing [1416]. AnnoDomini [2827]. Annotated [1645, 1043, 144, 1815, 211, 1721, 1420]. Annotating [1282]. Annotation [1905, 1906]. Annotation-Directed [1906, 1905]. Annotations [1009, 594, 1912, 2121, 2728, 1911]. Announce [2459]. Announce-Listen [2459]. Announced [1664]. Announcement [1928, 1961, 2053, 698]. Announcements [1363, 1387]. Announcing [814, 2259, 2260]. Annual [2860, 2863, 2875, 1488, 2302, 276, 815, 1810, 2882, 2853, 2856, 2861, 2858]. Anomalies [428, 909]. Anomaly [266, 55, 1152]. anomolies [910]. Anonymous [711]. ANSI [1506, 445, 446, 386, 2716, 489]. ANSI-X3J16 [489]. ANSI-X3J16/ISO-WG21 [489]. Anthropology [1569, 1568]. Antonio [2901, 2858]. Antony [606]. Anymore [1668, 1667]. AP1000 [1105]. aperiodic [1325]. Apertos [600]. APL [320, 2024]. APL99 [2303]. Apology [2207]. Applets [2451, 2464, 2826]. Appliances [2340]. Applicable [1891, 1890]. Application [2320, 1047, 5, 551, 1751, 425, 2443, 621, 2437, 6, 2195, 859, 622, 1219, 2387, 785, 2736, 907, 379, 138, 2666, 332, 830, 2436, 2681, 2496, 2836, 1342, 1343, 2353, 2545, 2404, 96]. Application-controlled [551]. Application-Specific [425, 830]. Applications [2875, 2878, 2899, 2290, 2095, 1653, 1621, 639, 1733, 885, 705, 1062, 1445, 1742, 1741, 1170, 1228, 2007, 858, 1344, 2882, 713, 725, 24, 641, 1076, 1098, 2354, 712, 726, 916, 2094, 1073, 330, 2246, 638, 1622, 1248, 2599, 2357, 1454, 1063, 2355, 1133, 2345]. Applicative [2701]. Applied [2032]. Applying [10]. Appreciated [2184]. Approach [2198, 1036, 2202, 801, 1014, 1322, 1236, 1400, 1661, 1028, 1061, 1722, 1940, 621, 1347, 2201, 653, 1530, 859, 1483, 434, 711, 1707, 716, 2121, 1351, 1268, 2549, 2199, 363, 1770, 1769, 494, 1660, 48, 1467, 2736, 239, 2414, 1053, 2660, 848, 1235, 618, 2551, 755, 2308, 2309, 792, 1803, 2419, 1706, 2028, 289, 69, 275, 622, 1595]. approaches [2513, 395]. Appropriate [1339]. approximate [515]. approximating [2529]. approximations [2423]. April [2251, 2857, 2869]. APT [1198]. Aquarius [814]. Arbitrary [2130, 279]. Architectural [2873, 2884, 1013, 436, 272, 1516, 22, 1105]. Architecture [1313, 255, 273, 264, 2031, 1354, 2034, 550, 639, 1578, 1579, 2329, 1559, 246, 867, 70, 2218, 2118, 1353, 417, 638, 1314, 868, 971, 2219, 2470]. Architecture-Independent [639, 638]. Architecture-Oriented [1579, 1578]. Architectures [604, 20, 659, 2284, 1742, 1741, 74, 2343, 305, 306, 2607]. area [544, 134, 2352]. Arguments [459, 1269]. Ariadne [907]. Arithmetic [1662, 208, 50]. Arity [2162]. Array [2303, 993, 651, 1747, 1746, 2805, 639, 1981, 2608, 1982, 650, 2000, 114, 2609, 1251, 733, 126, 127, 1253, 1220, 2675, 2348, 1257, 130, 1034]. Array-data [2608]. Arrays [447, 262, 350, 232, 2557, 852, 538, 493]. Art [450]. Artificial [2148, 2149]. ASAP [2448, 2154]. ascent [84]. ASCII [669]. Aspect [1841, 2410, 2409]. aspect-oriented [2410, 2409]. Aspects [1510, 2601]. ASPLOS [2873, 2859, 2884, 1543, 2138]. ASPLOS-V [2859]. ASPLOS-VI [2873, 1543]. ASPLOS-VII [2884]. ASPLOS-VIII [2138]. Assembler [48]. assembly [2846, 2803]. Assertion [2154, 2448]. Assertions [1658, 1590, 958, 949]. Assess [1614]. assessments [2321]. Assets [1567]. Assignment [889, 1986, 1155, 2288, 2281, 2593, 1217, 2570, 1213, 2611, 456, 2629]. assignments [2612]. Assisted [438, 1092, 1093]. Association [2886, 2210, 2211]. Associations [1069]. Associativity [2130]. Asynchronous [659, 654, 1991, 1992, 2777, 2531]. Atlanta [2895, 2896, 2897, 2252, 2251]. Atlas [2409]. ATOM [983]. Atomic [2023, 2366, 901, 2779]. Attribute [2318, 1004, 1032, 1794, 147, 2595, 490, 121, 234, 390, 119, 120]. attribution [2594]. attributions [2680]. audio [2495]. Austin [2878]. Author [884, 1542, 1865, 2108, 2443]. Authoring [2201]. Authors [1938, 2312, 1656, 1954, 2058, 2150, 2220, 2263]. Auto [1750]. Auto-Blocking [1750]. Automata [1919, 2536, 342, 534, 1040, 2540, 467]. Automata-driven [2536]. Automated [147, 234, 2664, 1249, 1250]. Automatic [128, 1341, 1407, 1408, 2609, 1680, 1566, 1676, 2067, 2550, 2291, 656, 1743, 1696, 2784, 1181, 1565, 1257, 2347, 1786, 2355, 290, 1060, 2628, 1467, 1182, 2348, 809, 1133]. Automatically [432, 2009, 1740, 1739, 100]. automatically-generated [100]. Automating [2205]. Autonomous [423, 592, 1505]. auxiliary [2737]. avoid [380]. Avoidance [2163]. Avoiding [1102, 522, 1200]. Awards [1387, 1635, 1664, 2050]. Aware [2128]. Away [1305, 1304, 2047]. Axiom [1150].

Axiomatization [1781].

B [890]. Back [1861, 1451, 809]. back-ends [1451]. back-of-the-envelope [809]. Backtracking [1659, 999, 234]. backward [1122]. Bad [1654, 1972, 1057, 1655]. Bagof [138]. Bagof/3 [138]. Bags [1177]. Balanced [760, 1540, 1209, 976]. Balancing [422, 794, 60]. BaLinda [137]. Bandwidth [1523, 250, 2118]. Bank [1719]. Banks [1536]. Barnes [440]. barotropic [66]. Barrier [2822, 572, 270, 2230, 771, 772, 1256]. barriers [1464]. Bart [919]. Base [2157]. Based [2864, 2871, 2900, 1583, 1720, 1157, 1987, 583, 624, 1697, 2467, 2163, 707, 1236, 985, 1516, 1860, 1828, 2103, 1084, 1535, 2170, 1295, 266, 1599, 2077, 1051, 2120, 1015, 1054, 1483, 2333, 1842, 1574, 2375, 973, 1351, 1642, 1832, 2902, 2449, 1212, 2450, 1582, 1758, 72, 2786, 191, 1202, 1203, 1117, 1776, 1701, 2740, 1617, 907, 2445, 2393, 749, 750, 147, 2594, 614, 1698, 1831, 2504, 374, 2556, 1774, 1773, 332, 2412, 1235, 734, 735, 2063, 835, 2409, 2428, 1859, 1346, 2102, 1534, 333, 965, 618, 2493, 1340, 2128, 368, 2848, 2651, 1977, 1342, 1343, 2548, 1206, 1280, 1281]. based [2479, 2411, 569, 443, 2334, 1505, 2787, 966, 2484, 1641, 611, 1732, 937, 98, 2482, 1002]. Basic [1554, 1168, 1030, 2341, 2390, 362, 1816]. Basis [584, 1607, 898, 899, 203]. basket [409]. Bass [1171]. Batched [1075]. Battlefield [1321]. Bayanihan [2477]. BDD [1745, 1744]. Be [2339, 575, 387]. Beach [2885]. Beetle [1639, 1640]. Before [1146, 1673, 1147, 1082]. Before/After [1082]. beginning [679]. Behavior [1835, 18, 1822, 2131, 2111, 27, 864, 865, 2330, 284]. Behavioral [2776, 178]. behold [167]. Belts [1397, 1394, 1392]. Benchmark [2339, 205]. benchmarking [1081]. Benchmarks [271, 27]. Benefits [2078, 1619, 226, 1260]. best [829]. beta [2733, 2819]. beta-reduction [2733]. Better [1212, 1589, 1588, 2767, 1133]. Between [1737, 1155, 2212, 2213, 2525, 2762, 1339, 123, 198, 2472, 2566, 2233, 2239]. Beyond [1013, 1404, 1863, 1864, 1620, 169, 1627, 508, 1403, 986]. bi [2640]. bi-directional [2640]. bibliographies [150]. Bibliography [1645, 1043, 144, 1815, 1044, 211, 51, 1721, 1420, 935, 383, 212, 917, 2649, 476]. Bidirectional [2316, 1287, 2315]. bifurcation [365]. Bigot [2224]. BigTypes [1799, 1798]. bimodal [2400]. Binary [483, 2098, 2173, 1319, 2123, 1699, 1320, 2239]. Binding [366, 1895, 350, 1894, 2191, 286, 287, 2618]. Binding-Time [1895, 1894, 286, 287, 2618]. bindings [848, 1449, 2190, 2647]. Biological [1374, 1375]. Biology [1853, 416]. Birkhäuser [480]. birth [671]. Bisimilarity [2755]. bisimulation [1468]. Bit [599, 1941, 1782, 1783, 220, 2233, 971]. bit-addressing [1782, 1783]. bits [2239]. BLAS3 [1750]. Blending [2373]. Block [719, 1898, 1674, 1524, 711, 1897, 531, 1525, 1107]. Blocked [251]. Blocking [1750, 1710, 540, 1709]. Blocks [896, 2652, 362]. Blueprints [1846, 1847]. BlurFit [2436, 2437]. BNF [2152]. Bobrow [450]. body [2343]. Bollmann [622, 621]. Bollmann-Zuse [622, 621]. Book [662, 1602, 1637, 1167, 466, 1168, 381, 439, 1166, 1226, 1881, 607, 440, 274, 481, 606, 525, 465, 451, 480, 885, 1227, 1169, 705, 1231, 1229, 1230, 157, 1170, 1228, 450, 796, 1171]. books [699]. Boolean [142]. BOOM [2508]. boost [243]. boosting [556]. Borland [607]. Borrow [2073]. Boston [2900, 2902, 480]. Both [603, 1149, 322]. bottleneck [547]. Bottom [1366, 1186]. Bottom-up [1366]. Boudillet [481]. Bound [2000, 126, 127]. Boundaries [2467, 102, 963]. boundary [1224]. Bounded [813, 1908, 2368, 2582, 2602, 2662, 2723, 1907, 314, 2703]. Bounding [2279, 473]. bounds [1776, 2704]. boxcharts [155]. boxing [2667]. Bradfield [480]. Branch [763, 1687, 1526, 247, 975, 2284, 1524, 2125, 1989, 1690, 1109, 1202, 1203, 543, 542, 1201, 1525, 1108, 1691]. branches [524, 1200]. branching [243]. Breadth [1745, 1031, 1744]. Breadth-First [1745, 1031, 1744]. break [507]. Breaking [669]. breakpoint [2651]. breakpoints [106, 552, 727]. Brew [1318]. Bridging [2800, 2472]. Brief [1146, 1234, 696, 684, 690, 688, 692, 689, 697, 691, 143, 694, 686, 695, 1147, 687, 693, 2247]. Brinch [1953, 1881, 2046]. bringing [1493]. Broek [1166, 1226]. browsing [395, 348]. Bruce [1170, 1228]. Buffer [2327, 1104]. Bug [951]. Bugfind [43]. Bugs [1423]. Build [2253, 2254, 80]. Builder [797]. Building [224, 877, 1053, 398, 2688, 2104, 2105, 6, 2093, 2477, 983, 360, 33, 2161, 2064, 2409, 501, 1199]. Builds [2383]. BURG [475]. BURS [523]. Business [18, 1057, 29, 1344, 843]. bylaws [1362, 1361]. Bytecode [2452, 2096, 2396, 2832, 2808]. Bytecodes [2171, 2465, 1164].

C [1509, 802, 1167, 607, 1966, 274, 481, 525, 1499, 1500, 667, 796, 2735, 1506, 1130, 464, 468, 1572, 445, 446, 1009, 1588, 1589, 163, 139, 1412, 1413, 2682, 655, 632, 798, 1725, 1660, 1303, 2736, 933, 1024, 2461, 1601, 1571, 934, 2580, 1174, 1932, 472, 386, 1349, 2724, 1479, 1917, 524, 1905, 1132, 717, 2490, 1653, 1963, 1661, 835, 1414, 1415, 1671, 1137, 2191, 77, 639, 1814, 242, 1906, 1705, 229, 1590, 626, 627, 686, 2042, 492, 1350, 1080, 1495, 1496, 2716, 1677, 1348, 403, 678, 950, 485, 2512, 2047, 687, 681, 1, 92, 2007, 387, 443, 1560, 1042]. C [1662, 581, 1133, 616, 1344, 1877, 1597, 1192, 1193, 1493, 1494, 1357, 1405, 1358, 1672, 2064, 937, 1996, 964, 631, 1268, 489]. C* [321]. C-to-Forth [2512]. C/C [1509, 934, 1662]. CA [2873, 73, 75, 2853]. Cache [2122, 265, 1520, 2270, 2271, 2290, 1689, 251, 266, 252, 1540, 2131, 1519, 984, 2289, 2232, 547, 1221, 746, 551, 1101, 1340, 792, 1113, 548, 1107]. Cache-Based [266]. cache-coherent [1107]. Cache-Conscious [2122, 2270, 2271, 2232]. Caches [2130, 574, 573, 1102, 540, 285, 546]. Caching [263, 1457, 1244, 1225, 2416, 549]. CACL [587]. Calculating [2425, 2267]. Calculation [1023, 629, 1775, 971]. calculational [2820]. calculator [809]. Calculi [1910, 2555, 2655, 2567, 1909, 2589, 2677]. Calculus [2202, 2081, 1891, 2189, 2838, 2709, 2801, 1887, 2775, 2565, 2847, 2754, 2653, 2643, 1890, 2755, 1451, 2556, 2824, 2395, 1414, 1415, 2753, 2515, 1458, 1471, 377, 2681, 2611, 2777, 2776, 2665]. Calendar [1409, 1478, 1503, 1713, 2011, 1811, 1914, 1927, 2038, 2051, 2141, 2209, 2250, 2301, 1398]. Calendrical [1238]. California [2862, 2877, 2892, 2868, 1612]. Call [1378, 1379, 1481, 1410, 1475, 1502, 1501, 1652, 2142, 2143, 1947, 1915, 1946, 2012, 2052, 2039, 2303, 2304, 2302, 2040, 1075, 10, 1370, 2683, 1836, 2189, 1703, 1704, 1273, 2765, 2611, 2709, 2801, 1887, 2682, 2630, 1272, 2830, 1450, 2665]. Call-by-Need [2189, 2709, 2801, 2830]. call-by-value [1887, 1450, 2665]. Call-cost [1703, 1704]. Calling [1444, 805, 2372, 2714, 806, 2346]. Calls [1873, 1572, 1571, 996, 1152, 229]. Cambridge [2884, 439, 440, 606, 525, 465, 451, 480, 2869]. Caml [2361]. Campenhout [157]. Can [2360, 1867, 1310, 1850, 575, 387]. Canada [2893, 2855]. canonical [2073]. Capabilities [1234, 2847]. capability [1117]. capability-based [1117]. Capturing [2131]. Car [1840, 1843, 1839]. card [2229]. Care [36]. cares [1657]. Carolina [2863]. carrying [2766]. Case [10, 1720, 1432, 1038, 270, 1849, 1513, 597, 1476, 1324, 168, 2409, 333, 1326, 1327, 1512, 945, 576, 26]. Cases [1573, 2088, 2087]. Casting [1493, 2278, 802]. Castle [619]. CAT [488]. catamorphisms [2747]. Catching [1423]. categorical [79]. categories [2691, 1230]. categorized [2649]. Catenable [1764, 1765]. caused [728]. Cayenne [2181]. CBN [2422]. CC [1541]. CC-NUMA [1541]. CCG [283]. CCP [1902, 1903]. Cecil [2102, 2103]. Cedar [331]. Cellang [535]. Cells [1374, 1375]. cellular [342, 534]. Center [10, 32]. Centered [16, 396]. Centric [1710, 1709]. Century [2303]. Certain [2369]. Certifying [2008]. CEV [2455]. CFG [2206]. CGI [2204]. Chair [918, 920, 921, 923, 924, 925, 1613, 1634, 1371, 1487, 2431, 1361, 1362, 1663, 1809, 1872, 1647]. Challenge [1825]. Challenges [2503, 2489, 1619, 1624, 2340, 2351, 2427]. CHAM [2754]. Change [1573, 1144]. Changes [2023, 1361, 1511, 2414, 1362]. Changing [1839, 1840, 2508]. channels [67]. ChaosMON [425]. Characteristics [272, 1101]. Characterizing [549, 902]. Chare [65]. Charles [480, 1229]. Charleston [2863]. CHARM [835]. chart [462]. Chasm [1569, 1568]. cheap [2810]. Checkable [1015, 1766]. checkers [1271]. Checking [2384, 1507, 1611, 2000, 400, 2781, 828, 2631, 2605, 1462, 2771, 126, 127, 2730, 2641, 2799]. checkpoint [63]. checks [1220]. chemical [2520, 623]. Cheney [1269]. CHI’91 [620]. Chip [2115, 1513, 1512]. CHOCOLATE [2677]. choosing [1127]. Church [2817]. Church-style [2817]. Cichelli [40, 41]. Cilk [1262, 1997]. Cilk-5 [1997]. Circle [2019, 2018]. Circular [302]. Circularity [1004]. CISC [553, 273]. City [2861, 1667, 1668]. CL [339]. CL/TB [339]. CLANG [156]. Clarification [825]. Clarifications [1511]. Clarity [1165]. Class [2157, 1566, 1300, 2274, 1286, 601, 639, 2069, 1084, 1822, 2292, 2023, 1299, 1054, 1560, 1855, 740, 741, 414, 179, 2403, 2813, 2559, 877, 1285, 395, 105, 2794, 602, 330, 1823, 618, 1472, 2540, 1464, 2319]. Class-Free [1054]. Classes [2217, 560, 2809, 2378, 1240, 1921, 1818, 1942, 1594, 933, 2097, 2361, 1817, 2394, 2641, 753, 2478, 2664, 616, 1344, 1876, 1239]. Classification [1369, 585]. Classroom [1589, 1372, 1388, 2861, 1373, 1389]. clause [53]. Clauses [1290, 2536]. clean [2309, 1038, 2308]. CLIB [797]. Client [1529, 1018, 1073]. Client-Side [1018]. client/server [1073]. clock [1096]. CLOS [180]. Closing [1812, 2009, 558, 1813]. closure [1758, 2746, 2783, 2685]. Closures [2216, 2217, 2191]. CLP [1902, 356, 310, 505, 1903, 999]. CLU [688, 675]. Cluster [2473, 2128, 1337, 215]. Cluster-based [2128]. Clustered [1533, 2352]. Clustering [2089, 906]. Clusters [2054, 2055]. CML [1468, 1762, 309]. Co [2304]. coagulating [283]. Coalescing [982, 2741]. Coarse [323]. Coarse-grain [323]. COBOL [833, 834]. Coco [93, 244]. Coco-2 [93, 244]. Code [1550, 1424, 1958, 257, 1702, 1688, 2031, 432, 2282, 2326, 722, 1554, 1436, 1711, 2291, 2294, 1750, 979, 2342, 975, 1434, 2288, 430, 2004, 1684, 2281, 2286, 1506, 2003, 739, 1246, 553, 1341, 494, 483, 2408, 1218, 2797, 208, 2393, 1111, 2735, 2845, 1959, 220, 99, 497, 744, 745, 234, 514, 1797, 1213, 90, 836, 608, 283, 522, 1200, 2766, 101, 305, 306, 1472, 2691, 2028, 100, 1998, 976]. Codes [2267]. coding [1128]. Cogen [1466]. Coherence [1532, 1735, 265, 1443, 1749, 72, 792, 1113]. coherency [548]. Coherent [405, 1751, 1107]. coinductive [1459]. Collaboration [1574, 2070, 2071]. Collaboration-Based [1574]. Collaborative [2455, 2474]. collect [394]. Collected [984, 297, 298]. Collecting [1843, 2579, 2783]. Collection [2002, 714, 2279, 1993, 1683, 560, 2331, 1044, 2280, 1756, 463, 957, 1224, 294, 748, 2232, 2534, 518, 2650, 2656, 1991, 295, 296, 338, 1788, 2245, 173, 1992, 751, 752, 2237, 393, 2235, 2411].

Collections [1828]. Collective [18, 2352]. Collector [1429, 2332, 2178, 1055, 2616, 749, 750, 2580, 2266, 773, 2228, 2240, 2236, 1092, 1093, 2241]. collide [179]. Colorado [2899]. Coloring [1538, 1689, 574, 1483, 2283, 573, 1537, 299, 103]. Column [1372, 1373]. Combination [562]. Combinations [2668]. Combinator [2183]. Combinators [2375]. Combinatory [2678]. Combined [2027, 938, 939]. Combining [2229, 2534, 2206, 1084, 630]. Comfortable [1717, 1716]. COMFY [1880, 1716, 1717]. Comment [1504, 818, 886, 2223, 718, 2443, 1128]. Comments [621, 1483, 1943, 622, 1609]. commercial [1101, 1900]. Committee [2139, 1916, 1387, 1913, 1926, 2300, 1944, 1390, 1361, 1371, 1544, 1474, 2208, 1960, 1390, 1391]. Common [2487, 1932, 1607, 1931, 2800, 339, 797, 2065, 469]. commonality [956]. Communicating [2614, 329, 166]. COmmunication [2677, 739, 408, 1532, 1427, 1738, 592, 160, 2564, 1516, 1752, 1485, 1515, 2484, 1732, 1996, 1341, 783, 2352, 2351, 2657, 2615, 1257, 1263]. Communications [1743]. Commutativity [1426]. Compact [132, 2281, 1076, 462]. compacting [2228]. Companion [440]. Comparative [572, 81]. Compare [803, 205]. Compared [2024]. Comparing [273, 2777, 2235]. Comparison [1284, 1385, 634, 1175, 213, 2447, 55, 208, 47]. Compatibility [2098, 1319, 1320]. Compatible [2085, 50]. compendium [2890]. Compilation [114, 968, 1435, 1899, 970, 1886, 439, 1165, 1287, 1748, 2172, 1440, 1695, 1219, 1241, 1791, 1790, 2659, 2589, 2376, 2792, 1452]. Compile [643, 385, 2587, 1531, 912, 913, 1629, 2350, 973, 782, 1259, 840]. Compile-Time [643, 1531, 973, 385, 2587, 912, 913, 1629, 782, 840]. Compile-Time/Run-Time [1531]. Compile/run [2350]. Compile/run-time [2350]. Compiler [1550, 1482, 1880, 1537, 1538, 1508, 1110, 2244, 2110, 260, 2298, 518, 624, 1679, 1610, 2285, 1606, 657, 1028, 1733, 1534, 1535, 390, 1340, 2317, 2008, 1982, 1421, 679, 799, 150, 44, 2280, 1438, 1707, 1256, 1837, 1506, 2003, 206, 1120, 292, 293, 1509, 1552, 93, 208, 934, 147, 386, 524, 527, 321, 1105, 122, 2598, 2228, 2341, 1209, 955, 90, 362, 541, 65, 156, 84, 2512, 1206, 1113, 1706, 219, 500, 444, 1638, 1337, 1637]. Compiler-Based [1535, 1534]. Compiler-Controlled [2110, 260]. Compiler-Directed [1538, 1537]. Compilers [2896, 1915, 2881, 1874, 647, 1426, 43, 502, 1339, 2742, 2243, 1232, 559, 1123]. Compiling [2171, 1747, 1848, 642, 2700, 744, 745, 371, 1304, 1305, 377, 810, 636, 503, 130, 754]. Complete [1976, 1004, 1991, 1992, 2544]. completeness [2623]. Complex [891, 1242, 18, 1069, 609, 978, 185]. Complexity [2640, 1462, 2160, 1834, 2733, 2785, 952, 1774, 1773, 2555, 2588, 1384, 2321, 2634, 349, 1833, 1271]. Compliance [1548]. Component [2479, 2483, 1342, 1343, 2413]. Component-based [2479, 1342, 1343]. Componential [1697, 1698]. Components [1586, 1587, 1163, 34, 801, 1850, 2075, 1574, 1254, 378, 2076, 218, 163]. Composable [2595]. Composing [2680, 2752, 1695, 2688]. Composite [2091, 399, 2032, 1851, 876]. Composition [1082, 1306, 1297, 1277, 2573, 2697, 1243, 1298, 1276]. Compositional [2645, 1780, 2398, 2831]. compositions [178]. Compound [2100, 1594]. compoundness [357]. comprehension [453]. comprehensions [114]. comprehensive [338, 2588, 2832]. Compressed [1066, 1067]. Compressing [2292]. Compression [263, 1526, 2282, 1711, 2291, 536]. Computable [2519, 1922, 2540]. Computation [1616, 2290, 2176, 2466, 793, 718, 1154, 1258, 2518, 940, 1244, 2627, 764, 2745, 2692, 2764, 517, 2737, 314, 2751, 2774, 2421, 1198, 2649, 786, 2545, 2422, 352, 1263]. Computational [1628, 1721, 929, 2634]. Computations [2450, 1734, 433, 1748, 660, 1732, 737, 901, 2416, 2765]. Compute [1541, 1087]. Computer [1374, 698, 2177, 1868, 1231, 2607, 1614, 577, 2483, 195, 1375, 553, 927, 928, 2496, 233, 2501]. Computers [1381, 973, 1915]. COMPUTING [1376, 2886, 2891, 1379, 2251, 445, 446, 2451, 1968, 1969, 2432, 1719, 1681, 2462, 2467, 427, 981, 1252, 2473, 2299, 2477, 593, 1377, 1175, 2486, 2359, 114, 2433, 1718, 1970, 2772, 1302, 2343, 1153, 2607]. concatenation [2556, 2590]. Concept [2026, 600, 134]. Concepts [1940, 577, 1125, 2390, 383]. Conceptual [1037, 1385, 210, 216]. Concerning [889]. Concerns [2046]. Concert [1010]. Conclude [2261, 2262]. Concrete [1284, 1074]. Concreteness [1279]. Concurrency [1009, 594, 389, 431, 2586, 68, 56, 780, 781, 2572]. Concurrent [73, 75, 820, 1315, 1661, 21, 1048, 1814, 1966, 1893, 1039, 575, 1305, 2032, 1025, 1445, 2532, 593, 146, 316, 1660, 2616, 2841, 76, 452, 1324, 773, 2245, 499, 168, 835, 1049, 63, 1892, 1304, 836, 171, 1127, 2751, 2657, 2715, 309, 233, 2574, 2712, 1131, 2540, 791, 500, 2658, 167, 670, 689, 2748, 47, 667]. concurrent/distributed [1127]. Conditional [1687, 2663, 543, 2812, 1200]. Conference [2852, 2855, 2860, 2862, 2865, 2863, 2867, 2869, 2874, 2870, 2875, 2873, 2877, 2878, 2884, 2885, 2886, 2887, 2888, 2889, 2892, 2893, 2895, 2898, 2899, 1264, 1363, 1409, 1376, 1478, 1503, 1378, 1713, 1928, 2139, 2143, 2252, 2303, 2251, 1812, 1811, 1945, 1914, 1927, 2038, 2051, 2141, 2209, 2250, 2301, 1651, 2208, 2432, 1370, 1398, 1497, 1477, 2435, 2305, 1377, 2883, 2882, 195, 1718, 1970, 1936, 2434, 1719, 1544, 1968, 1969, 1498, 1935]. Conferences [2300, 2054, 1476, 2055, 1371, 1647]. Configurable [959]. configurations [409]. configuring [378]. Confined [2392]. Conflict [561, 1980, 1102, 1104]. confluence [1797]. confuse [2192]. Confusion [1643]. conjugate [787, 788, 331]. connected [1254]. connection [2762, 292, 293]. conquer [2347, 385]. CONS [459, 1269]. Conscious [2122, 2270, 2271, 2232]. Consequences [262]. Conservative [1644, 748, 2534, 200, 2235]. Considerations [1720, 1127]. considered [1918, 447, 491, 150]. Consistency [267, 1316, 1822, 1514, 546, 1749, 1317, 1823, 182]. Consistent [435, 433]. constant [1196, 1197, 736, 319]. Constants [1172, 1151, 958, 1173, 949]. Constraining [1290]. Constraint [1984, 1929, 584, 1862, 2103, 1068, 2849, 1985, 2731, 2841, 168, 610, 2635, 1930, 2532, 2574, 2712, 2711, 358, 335, 413, 2658]. Constraint-Based [2103]. ConstraintLisp [610]. Constraints [1722, 1575, 1896, 2788, 1135, 1422, 170, 225, 510, 189, 116, 310, 1336, 1800, 1461, 511]. construct [140]. Constructing [2630]. Construction [1482, 1444, 34, 1638, 1836, 1028, 612, 1745, 1060, 2446, 147, 2550, 527, 2511, 1471, 1744, 1637]. Constructive [583, 1136, 110, 2633, 91]. Constructor [1372, 1373]. Constructors [717, 485]. Constructs [1735, 659, 213, 1330, 2810, 2203, 189, 83]. consumer [950]. Consummating [1725]. Contain [1921]. Containment [821]. Contention [1757, 269]. Context [1682, 987, 988, 1896, 252, 1194, 1195, 1895, 2837, 1894, 1797, 2830, 1040, 279, 1192, 1193, 476]. context-dependent [1040]. Context-Free [1896, 279]. Context-insensitive [1194, 1195]. Context-Sensitive [987, 988, 1192, 1193]. context-sensitivity [1797]. contexts [1471, 1118]. Continuation [1155, 967, 2687, 2617].

Continuation-Passing [967, 2687, 2617]. Continuations [1386, 1430, 962, 68, 2182, 224, 2559, 2578, 754, 105, 2626, 2560, 2561]. Continuous [2850, 538]. Contraction [1981]. Contracts [1084, 1567, 2414, 178]. Contraint [2102]. Contraint-based [2102]. Contrasting [1101]. Contravariance [1367]. Contributes [567]. Control [1717, 125, 1448, 1430, 2459, 1700, 981, 634, 2284, 2333, 1716, 2584, 1215, 132, 1669, 1762, 2517, 510, 105, 155, 228, 118, 2844, 2783, 2774, 2495, 2491, 1198, 2560, 455, 1114, 1115, 369, 409, 742, 1086, 2334, 2498, 1006, 511]. Control- [125]. control-flow [369]. Controlled [2110, 260, 2193, 551]. controllers [1328, 1329]. ControlShell [1342, 1343]. conventional [198, 2243]. conventions [2714]. Conversations [1025]. Conversion [1819, 1351, 88, 2827, 2625, 2746, 2685, 1820, 761, 762]. Conversions [1019]. convert [141]. Converting [2461]. convolution [292, 293]. cooC [667]. COOL [1144, 1999, 794, 187, 1039, 1482]. Cooper [718]. Cooperation [1650]. Cooperative [557]. Coordinating [2563]. Coordination [268, 2562]. Copies [1747]. Coprocessing [157]. coprocessors [2496]. Copy [2336, 2335, 2073]. copying [962, 2650, 1788, 2412, 773, 2240, 2235]. CORBA [1974, 1845, 1558, 2486]. Core [423, 2166, 2838, 1241, 1468, 2395]. Corner [1409, 1376, 1713, 1812, 1811, 1945, 1914, 1927, 2038, 2051, 2141, 2209, 2250, 2301, 1651, 2432, 1370, 1398, 2305]. coroutines [198]. Corporate [32, 2248, 2249]. Corpus [1202, 1203]. Corpus-based [1202, 1203]. corral [829]. correct [133, 2658]. Correction [2207, 2448, 1966]. Correctness [2801, 2651, 2798, 346, 2538, 2757, 362, 2708, 2618]. correlation [349, 542, 1108]. Correspondence [1399, 1411, 2312, 1656, 1954, 2058, 2150, 2220, 2263, 1412, 2265, 1155, 2443, 2314, 2397, 2151, 2226, 1380, 1925, 2325].

Corrigendum [2723]. Cost

[1571, 2328, 1707, 2118, 1788, 989, 990, 2743, 1458, 1703, 1704, 1706, 2499, 1931, 1932]. cost-effective [2118]. Costs [422, 1864, 1109, 1863, 2750]. Could [2019, 1918, 1937, 2018]. Count [1029]. Counterpoint [1396, 1397, 1392, 1393]. Counters [1682, 2121]. Counting [977, 2234, 368]. counts [2233]. coupled [545]. Coupling [148]. course [943, 2510, 461, 2501]. Covariance [1367]. Covariant [1126, 1187]. coverage [2652]. CPR [2284]. CPS [1610, 2625]. Cracking [2464]. Crashes [1520]. Creating [1313, 1636, 1314, 1548]. Creative [1227]. Crew [1901]. Critical [2207, 428, 2158, 2535, 946, 947, 275]. critique [878, 879]. CRL [712, 713]. CRL/Pascal [712, 713]. Crole [1230]. Crop [2083, 2082]. Cross [2005, 1075, 48, 2172, 713, 712, 1770, 1769]. Cross-Module [2005, 2172, 1770, 1769]. Crossing [1569, 1568]. CST [1348]. Cult [13, 1296]. Culture [13, 1296]. Curcio [2448]. Current [624, 472]. Curricula [2432, 2433]. Curricular [1404, 1372, 1388]. Curriculum [1968, 1969, 1719, 1718, 1970]. Curry [2774]. Currying [1442]. Customizable [905, 1736, 499]. Customization [1084]. customize [2244]. Customized [983]. Customizing [1316, 1317]. Cut [2383]. Cycle [1984, 1070, 1005, 1985, 1857, 1858, 1324]. cycle-stealing [1324]. cyclic [2724].

D [2448, 1166, 1226, 862]. D2R [811]. DAG [2724]. DAGs [2845]. Dagstuhl [619]. Dallas [2858]. Dangling [2214, 2215]. Daniel [450]. DARP [2450]. Data [1995, 2450, 1983, 1258, 1161, 640, 1029, 643, 125, 2122, 2458, 794, 1708, 1526, 2290, 1349, 1026, 1242, 642, 784, 1746, 656, 1322, 2115, 2368, 985, 1252, 2316, 2474, 1441, 1709, 1710, 2079, 1274, 1535, 646, 9, 1143, 2032, 1238, 327, 1546, 1853, 1540, 647, 634, 1446, 1980, 2120, 967, 1536, 2151, 2799, 645, 636, 250, 1447, 1728, 2007, 1541, 992, 740, 741, 1210, 1211, 782, 334, 776, 1241, 536, 1245, 1110, 291, 2609, 783, 2232, 328, 1467, 1215, 1221, 953, 954, 2526, 2640, 134, 733, 2692, 2550, 1122, 1416, 1460, 2750, 516, 76, 1775, 773, 1259]. data [320, 74, 1534, 349, 2528, 2675, 280, 2608, 912, 913, 2802, 2348, 1545, 1455, 2792, 1257, 1198, 778, 779, 506, 777, 474, 1131, 765, 1255, 1086, 552, 727, 2239, 581, 282, 1107, 274]. Data- [125]. Data-Centric [1710, 1709]. Data-Flow [1161, 1983, 1210, 1211, 2675, 2348]. Data-Memory [1536]. Data-Parallel [1995, 1252, 1546, 645, 776, 291, 2609, 783, 953, 954, 2750, 1545, 778, 779]. data-parallelism [777]. Data-structures [2368]. data-type [740, 741]. Database [1080, 2336, 2137, 1079, 1062, 859, 1714, 1382, 27, 416, 1307, 202, 415, 2335, 169, 183, 182, 278]. Databases [1077, 28]. Dataflow [1685, 812, 1158, 1251, 1027, 2693, 811, 2759, 162, 503, 1808, 1806]. datatypes [2747, 151, 1045]. David [1229]. Davie [606]. Dawes [439]. Days [1003]. DBMS [415, 1830]. DBMSs [1829]. DCG [1111]. DDD [1365]. Dead [1688, 979, 2007]. deadlines [2241]. Deadlock [429]. Debaere [157]. Debate [1585]. Debates [1397, 1394, 1392]. Debuggable [431]. Debugger [1188, 822, 710, 312, 1129, 907, 908, 496]. Debuggers [1365, 212, 917]. Debugging [432, 426, 2490, 497, 1860, 995, 2365, 430, 992, 2286, 424, 2638, 1424, 731, 494, 919, 43, 328, 914, 911, 311, 56, 898, 899, 906, 1859, 897, 915, 1227]. DEC [2222]. Decay [1683]. December [2448]. Decentralized [2134]. Decidability [1047]. Decidable [2662, 828, 2723, 2839, 2703]. Deciding [2544]. Decision [22, 100]. decisions [334]. declarations [116, 492]. Declarative [1619, 582, 2225, 2226, 1856]. declustering [538]. Decoding [2267]. Decomposition [1274, 1416]. decrease [1213]. Deep [1187]. Default [2712]. defect [581]. defects [2389]. defense [336]. Deferred [1029, 1024]. Defined [652]. defining [827, 616]. definitely [2795]. Definition [2872, 1013, 2271, 1019, 709, 665, 197, 526, 238, 472, 1011]. Definition/Use [1013]. Definitions [2147, 1722, 2146, 2583, 1456, 89]. Deforestation [2383, 1794, 2788]. defunctionalization [1761, 1760]. defusing [802]. delay [2403]. delayed [305, 306]. delayed-load [305, 306]. Delays [434]. Delegation [1555, 1295, 1277, 2197, 1276]. Delegation-Based [1295]. Delinearization [507]. Delivering [1076]. Demand [125, 2692, 1529, 2786]. Demand-Driven [125, 2692, 2786]. Dennis [1602]. Denotational [1486, 1495, 2787, 1496]. dentistry [1937]. Denver [2899]. deoptimization [497]. dependable [2415]. Dependence [125, 1448, 1349, 985, 734, 735, 991, 2551, 2120, 636, 2639, 132, 1122, 281, 2630, 507, 2675, 280, 1198, 2674]. Dependence-based [734, 735]. dependences [319, 506]. Dependencies [1300, 1457, 2551]. dependency [2838]. Dependent [2181, 2000, 2843, 720, 112, 113, 1040, 721]. depth [2718]. depth-first [2718]. Derivation [2205]. derivations [2567]. Deriving [2666, 1456]. descent [2063, 84]. Describing [1374, 1375]. Description [1561, 1628, 725, 726, 1422, 2243, 203]. Desert [2033]. Design [2852, 2855, 2862, 2867, 2874, 2886, 2889, 2893, 2895, 2078, 1313, 59, 255, 2252, 2022, 2360, 2175, 2276, 34, 1071, 2274, 1852, 1368, 1845, 1885, 121, 21, 2338, 106, 601, 1323, 1575, 2, 1310, 2367, 440, 1849, 2351, 15, 259, 22, 451, 541, 2008, 1851, 2077, 1051, 653, 2196, 859, 1170, 1228, 29, 2883, 1548, 1642, 565, 1292, 2066, 1328, 1329, 163, 1072, 407, 48, 876, 2408, 202, 2203, 2595, 1465, 1339, 1479, 1480, 164, 471, 189, 602, 1337, 1657, 498, 235, 1314, 44, 727, 1641, 617]. DesignFest [1869]. Designing [1910]. Designs [1824, 1574, 2410, 2389]. destructive [940, 2547, 456, 2725, 387]. destructors [387]. Detecting [428, 2029, 901, 355, 2672, 900, 2389, 2430]. Detection [728, 993, 435, 1425, 1005, 429, 266, 327, 347, 55, 909, 910, 912, 913]. Determinacy [354]. Determination [628]. Determining [2770, 2526]. Determinism [2193]. Deterministic [2189, 1919, 1445, 2690, 952, 2538, 2544]. Develop [2459]. Developing [832, 2454, 916, 1720, 397, 955, 2042, 620, 705, 2340, 838]. Development [595, 10, 1508, 1564, 576, 890, 16, 1073, 1477, 381, 1607, 639, 1274, 2075, 1062, 1184, 1642, 597, 605, 404, 1778, 1129, 1509, 1563, 1476, 830, 638, 2076, 623, 678, 679, 703, 2497, 581, 2499, 683, 1641]. Devices [1671, 1672]. Dewan [1171]. Diagnostic [1444]. diagrams [1862]. Dialect [2485, 207]. Did [623]. Diego [2868, 2892]. differences [123]. Differential [2637]. Differentiation [2177]. Diffusion [2178]. Digital [1231, 1536, 2495, 1937]. Dimensional [1042]. dimensioned [1876]. Dimensions [460]. Diminishing [1643]. Direct [1571, 2119, 2004, 1102, 2775, 2173]. direct-mapped [1102]. Direct-threaded [2004]. Directed [1538, 1906, 1704, 2160, 2317, 1675, 1889, 1438, 1537, 2744, 1759, 2724, 1905, 2420, 1439, 2564, 1703, 2376, 1888]. directing [495]. direction [235]. directional [2640, 107, 108]. Directions [1233, 1615, 1625, 1626, 1842, 543, 1232]. Directories [265]. disambiguation [1104]. Disciplined [1555, 1350]. Discover [623]. Discovering [2737]. Discovery [1864, 1863]. Discrete [2453, 2460, 2482, 697, 676, 64, 237, 465]. Discrete-Event [2453, 2460, 697, 64]. Discretionary [1712]. Discussion [2041, 851, 857]. Disk [262, 538]. Disks [2117, 1521]. Dispatch [1066, 1289, 2377, 1837, 1067, 2101]. Dispatched [996]. dispatching [2401]. Displacement [1289]. Display [425]. Disposable [1792]. disposal [529]. Disposition [1275]. disruptiveness [2241]. Dissemination [1832, 1831]. Dissemination-Based [1832, 1831]. Distillation [1529]. distinction [2541]. distinguish [773]. Distributed [2450, 1737, 648, 437, 1071, 1245, 2457, 655, 1708, 1738, 2459, 662, 1531, 1533, 586, 637, 820, 433, 1242, 2467, 423, 2095, 1549, 2156, 2468, 21, 2338, 2472, 425, 2105, 1521, 1733, 6, 2178, 871, 2093, 647, 1055, 636, 660, 2479, 434, 1803, 588, 1756, 33, 650, 1052, 1727, 2482, 766, 1219, 739, 59, 85, 72, 1072, 839, 2713, 785, 1215, 134, 325, 1991, 2094, 1073, 793, 901, 74, 906, 1992, 2104, 2416, 1302, 2343, 2234, 1248, 1127, 1331, 2825, 194, 1114, 1115, 1086, 900, 1787, 902, 347, 2415, 2449, 2470]. Distributed-Memory [637, 636, 1756, 793, 1248]. Distribution [1708, 2128, 408, 840, 775]. Distributions [656]. Diverse [89]. Diversity [1295]. divide [2347, 385]. Divide-and-Conquer [385]. Division [2132, 972]. DJ [1447]. DLP [662]. DMA [1324]. DMPPs [1257]. Do [2214, 1085, 1864, 1397, 1394, 1392, 2215, 1863, 1396, 1395, 1393, 318]. DOACROSS [319]. Doctoral [12, 1593, 1119]. Document [578]. Documentation [601]. Documenting [566, 453]. DoD [682]. Does [1854, 2071, 1812, 943, 208, 803, 2767, 2070, 1813, 1090]. dog [2144, 2145]. DOGMA [2470]. Doing [843]. Domain [1075, 580, 5, 357, 579, 1308, 395, 903, 904]. domains [2668, 2527]. Dominator [1154, 2673]. Dominators [2652, 2597]. Don [796]. don’t [482]. Double [247, 1764, 1765]. double-ended [1764, 1765]. Double-Width [247]. DOVE [2449]. down [244, 147, 490, 952]. dozen [1933, 1934]. Dr [1170, 1228]. drag [1452]. drawing [922]. Dreaming [1388, 1389]. Driven [125, 1993, 1761, 1760, 2786, 2692, 2536, 1099, 1100]. Driving [1840, 1839]. DROL [588]. Dropping [1898, 1897]. DrScheme [2014, 2015]. DSM [1749]. DSMs [1532]. DSP [2328, 2329]. DSP-Processors [2328]. DSPs [2281]. Dual [1536, 467]. During [1822, 2423, 1823]. DyC [2297]. Dylan [1551]. Dyn}amic [2304, 124, 1435, 719, 1899, 613, 1747, 2290, 1681, 1436, 1425, 325, 1529, 1104, 2368, 985, 370, 2069, 1828, 1903, 2336, 1818, 2131, 1684, 812, 2818, 660, 2092, 2345, 1837, 1732, 1594, 2695, 1224, 831, 2851, 2546, 388, 1793, 1111, 2735, 1902, 497, 326, 1325, 2335, 1817, 512, 2669, 2816, 542, 1821, 2560, 811, 963, 1453]. Dynamical [158]. Dynamically [996, 1818, 1172, 1173, 2569, 1102, 115, 2683, 2526, 1817]. Dynamically-Dispatched [996]. dynamically-typed [115]. Dynamically-Valued [1172, 1173]. dynamics [1098, 1777]. Dynamo’00 [2304]. Dynascope [495].

eager [1207, 529]. EAPLS [2210, 2211]. Early [1003, 2266, 658, 673, 769, 770]. Ease [661]. Easter [1023, 629]. Eat [2144, 2145]. economics [411]. ECOOP [1961, 1599]. Ed [381, 1231]. Eddy [157]. Edge [2807, 1278]. Edicates [805, 806]. Edison [443]. Edited [439, 1166, 1226, 1171]. editing [346, 1222, 1223, 119, 120]. edition [212, 917]. Editor [2207, 1023, 578, 2045, 209, 2223, 2221, 2222, 767, 2224, 704]. Editorial [1971, 1352]. Education [1607, 2195]. Educational [1665]. Educator [1866]. EEL [1222, 1223]. Effect [252, 732]. Effective [1435, 980, 1681, 1570, 1590, 1739, 1740, 1756, 1707, 2327, 297, 298, 2003, 1205, 2118, 1706]. effectiveness [1768, 71, 1118]. Effects [2164, 1887, 2627, 2571, 2689, 2165]. Efficiency [1994, 646, 81, 2715]. Efficient [1433, 993, 258, 1179, 1899, 853, 854, 2401, 2627, 730, 1692, 1734, 1517, 1746, 1345, 1689, 2642, 1347, 659, 2295, 280, 266, 1248, 1731, 2032, 2281, 587, 653, 556, 1707, 1199, 1838, 552, 1192, 1193, 1837, 2399, 1469, 1262, 748, 2710, 1111, 2735, 2749, 1518, 2240, 2356, 2528, 507, 836, 912, 913, 1730, 523, 1452, 2402, 791, 1706, 1097, 790]. Efficiently [575, 445, 446, 513, 2475, 58]. Effort [23, 1081]. Efforts [1665, 2140]. Eiffel [207, 1837, 1268]. Eight [798, 1031]. eighteenth [2856]. Elected [2431, 1872]. electrically [2502]. elegance [1657]. elementary [2819, 140, 609]. ELFS [1061]. Eliens [662]. Eliminate [1134]. eliminates [1775]. Eliminating [547, 982, 524, 506, 1980, 2000, 2730, 1256]. Elimination [1984, 1687, 1688, 980, 1701, 1978, 979, 1220, 2153, 2315, 1219, 1985, 2428, 1979, 2849]. Ellie [197]. Else [2214, 2215]. Embedded [2896, 2282, 2339, 2368, 2338, 2332, 1267, 2331, 2491, 11, 2281, 2340, 2337, 2333, 2747, 1053, 2490, 452, 1335, 1266, 2334, 2499, 1338, 611]. Embedding [2367, 647, 45, 142]. embeddings [2851]. emerald [1667, 1668]. Empirical [2036, 461, 2134, 2113, 631, 943, 55, 406]. Employing [67]. Encapsulate [1299]. encapsulated [836]. Encapsulation [1484, 1240, 1724, 2362, 926, 1239, 1723]. Encoding [2267, 2186]. End [2479, 2064, 1365, 245]. endangered [728]. ended [1764, 1765]. ends [1451]. Engine [1840, 20, 1839, 2511]. Engineering [2894, 2142, 10, 580, 1621, 2337, 2393, 579, 1308, 1622, 390, 1646]. Engineers [465, 2510]. engines [2496]. Enhance [2283, 943]. Enhanced [2282]. enhancement [388]. Enhancing [453, 793]. enough [1301]. Enriching [1471]. ensemble [233]. Ensuring [570, 591]. Entities [2080]. entry [461]. enumerated [226, 470, 336]. enumeration [1335]. enumerations [135]. envelope [809]. Environment [2450, 2455, 1508, 1659, 970, 1539, 1818, 2033, 1051, 668, 708, 645, 1076, 847, 1509, 201, 325, 396, 2409, 1817, 118, 53, 162, 2497, 1063, 611, 2449]. Environmental [1562]. Environments [437, 1166, 1226, 187, 619, 1472, 119, 120]. eNVy [1094, 1095]. ephemera [698]. Ephemerons [1844]. EPIC [2284]. Epsilon [929]. EQL [2159]. Equality [1386]. Equate [413]. Equational [1546, 1032, 373, 1545, 2670, 777, 359, 1595]. Equations [1898, 1897, 784, 527, 507]. equivalence [2776]. Equivalent [1942, 2720]. Erasure [2187]. Erhard [275, 622]. Erlang [1778, 2384, 1772]. error [476]. Errors [993, 2029, 1425, 729, 2192, 172, 2530]. Escape [2798, 2387, 2386, 504, 2785, 368, 2398]. ESSAY [1234]. Essays [1953, 1882, 1881]. essence [754, 2576]. Essential [217, 869, 870, 1344]. estimated [284]. Estimation [23]. estimator [334]. Estimators [974]. ETAPS [2305]. eureka [1857, 1858]. EuroForth [1935, 1936]. European [2210, 2211, 2054]. EVA [149]. Evaluate [1184, 214]. Evaluating [1260, 2418]. Evaluation [2864, 2871, 2900, 2117, 1901, 1899, 2276, 651, 1908, 2067, 267, 2296, 1056, 1956, 1957, 1886, 2468, 572, 1236, 1891, 1516, 1981, 1893, 1305, 2348, 1514, 1675, 1440, 1695, 1631, 1559, 1749, 2902, 1887, 288, 2648, 2744, 1907, 147, 2550, 1890, 2297, 2420, 2405, 356, 372, 855, 1235, 2598, 121, 371, 2619, 355, 1892, 1304, 2343, 361, 541, 1900, 2599, 1455, 456, 485, 359, 358, 2419, 360, 548, 1632, 529]. Evaluator [1889, 1888, 2170]. evalution [2423]. Event [1275, 2453, 437, 589, 2460, 1845, 333, 434, 2482, 57, 590, 907, 697, 676, 64]. Event-based [333, 907]. Event-Orientation [589, 590]. Events [2459, 589, 590]. Every [2280]. Everyone [2471]. Everywhere [2471]. Evicted [2638]. evidence [493]. Evolution [2207, 1601, 1021, 2158, 1822, 2023, 1567, 31, 2221, 2414, 61, 169, 1823, 677, 680, 2498]. Evolutionary [2075, 2076]. Evolving [1318, 874, 875, 613]. Exact [2176, 280]. examining [2028]. example [532, 133, 138]. examples [1128]. Exception [193, 2047, 1097]. Exceptions [2475, 2273, 2272, 2848]. exclusion [554]. Executable [1685, 665, 1351, 2797, 952, 89, 1222, 1223]. Execute [257]. Executing [238, 1098]. Execution [1179, 2268, 1734, 1549, 2034, 2123, 2134, 1988, 720, 2710, 1324, 313, 2124, 1326, 1327, 841, 842, 2679, 2350, 721]. Executive [1387, 1913, 1926, 2300, 1944, 1390, 1361, 1371, 77, 1474, 2208, 1960]. executives [2487]. Exemplar [1084]. Exemplar-Based [1084]. Exercise [724]. Exhaustive [1447, 60]. Existing [1610, 1548]. exit [948]. expansion [2804, 1797]. Experience [1746, 1323, 658, 1853, 1080, 1281, 787, 788, 424, 964, 2029, 398, 769, 770, 90, 1280, 2404, 581]. Experiences [832, 767, 415, 1278, 1742, 1741, 2064, 179, 2417, 2427]. experiencing [180]. Experimental [585, 1254, 2798, 54]. experimenting [230]. Experiments [2456, 2027]. Expert [9, 229]. Explained [466]. Explicit [2516, 643, 782, 2006, 2625, 1306, 1912, 149, 848, 1214, 2655, 1801, 1911]. explicitly [784, 2342, 2629]. Exploitation [320, 2327, 67, 938, 939]. Exploiting [1682, 2463, 326, 243, 1536, 1988, 765, 720, 58, 721]. exploits [322]. Explorative [723]. Exploratory [471, 1076]. Explorer [2344]. Exploring [1833, 1834, 577, 1125]. Exponential [1004, 2544]. Exponential-Time [1004].

Exponentiation [2364]. express [2729]. expressible [98]. Expressing [1920]. expression [138, 473, 217]. Expressions [2061, 651, 1024, 724, 1976, 348]. expressive [800, 2777]. expressiveness [2690, 2370]. extend [950]. Extended [643, 651, 642, 1020, 2079, 636, 650, 2024, 2695, 1254, 115, 1792, 1759, 2370, 370, 1022, 2657, 2825, 2419, 2422, 895]. Extending [1275, 2646, 2091, 1323, 1748, 2842, 930]. Extensible [1436, 2477, 1124, 908, 410, 1061]. Extension [1191, 2081, 1661, 1146, 246, 1190, 1660, 2793, 2568, 1147, 2761, 489]. Extensional [2699]. Extensions [669, 2458, 2466, 1368, 666, 1883, 431, 1121, 941, 942, 136, 2703]. extent [2560, 1876]. External [1265, 551]. extractor [2404].

F [1726, 2803]. Facilities [585, 345]. factoring [2710]. Factory [2442, 381, 1714]. Fail [799]. Fail-Safe [799]. Failure [152, 145, 46]. Fair [1400, 408]. Fairness [2537]. False [1737, 1259, 506]. family [553]. Far [2017, 1850, 2016]. Fast [2003, 1435, 1572, 554, 2813, 1436, 2285, 106, 2364, 1684, 2135, 2758, 100, 350, 537, 1117, 475, 2492, 2351, 2527, 2499, 38]. Faster [2730, 1449]. fault [855]. Feasibility [2469, 214]. Featherweight [2001, 2395]. Feature [1172, 628, 1173]. Features [272, 1486, 946, 947, 347]. February [1719, 2858]. Federated [2251]. Feedback [1284, 1681, 996]. Feeling [1668, 1667]. Fetch [268]. Fetch-and-Increment [268]. FFTs [2465]. FGCS [1003]. Ficl [2310, 2311]. Fiction [564, 563]. Field [1614]. Fields [1921, 2400]. Fifth [2880, 2497]. FIG [2044]. FIG-Forth [2044]. File [263, 1520, 1322, 1061, 2441, 537, 536, 2403, 78]. Files [2292, 2507]. Filesystem [2467]. Filter [2734]. filtered [470]. Finalization [1844]. Financial [580, 37, 579]. find [1032]. finding [2492].

Fine [1738, 260, 2089, 587, 1740, 1530, 1114, 1115, 2366, 1706, 1707, 960, 961, 300, 301, 2245, 1739, 787, 788]. Fine-Grain [1738, 260, 1740, 1530, 2366, 1114, 1115, 1706, 1707, 300, 301, 2245, 1739, 787, 788]. Fine-Grained [2089, 587, 960, 961]. Finite [2193, 2294, 1333, 2529, 2839, 2370, 2657, 467]. finite-rank [2839, 2370]. Finite-State [2294, 1333]. Fips [723]. Firmware [2442]. First [740, 741, 1920, 1300, 2794, 2511, 1464, 1745, 2890, 2559, 2755, 105, 2563, 2718, 1472, 1744, 1031]. First-class [740, 741, 2794, 1464, 2559, 105, 1472]. first-order [2755, 2563]. Firth [440]. Five [804]. fixed [2582]. Fixpoints [2371]. FL [2874, 2856]. FLASH [1089, 1112]. flashlite [1878]. Flattening [1247]. flaws [472]. Flexibility [1279, 1112]. Flexible [2072, 2382, 1679, 1549, 2089, 1684, 1767, 2739]. Flick [1679]. Flight [436, 2493]. Floating [1431, 109, 2471, 248, 1988, 111, 1662, 2065, 208, 1449]. Floating-Point [1431, 2471, 248, 1988, 111, 1662]. Florida [2885]. Flow [1682, 1161, 1196, 1197, 1026, 1746, 1986, 1700, 718, 1439, 2316, 2032, 1446, 967, 1895, 1447, 992, 1210, 1211, 1983, 2739, 1768, 2814, 2627, 2640, 1759, 733, 2692, 2550, 1762, 784, 510, 1894, 2722, 155, 2528, 2675, 2608, 2348, 2844, 2783, 2720, 2811, 2551, 455, 474, 2799, 2758, 369, 2823, 1006, 1806, 511, 1808]. Flow-directed [1439, 1759]. Flow-Insensitive [1986, 718, 2758]. Flow-sensitive [1196, 1197, 2627]. flowgraph [2673]. Fly [253, 909, 910, 898, 899, 218]. FNC [121]. FNC-2 [121]. Focus [705]. Fold [2185]. folding [1796, 1795]. food [2144, 2145]. FOOL [2040]. Foreign [2173]. Forever [1816]. fork [131]. fork-join [131]. Form [1701, 1986, 1155, 2288, 1977, 2695, 87, 214, 730, 2190, 2805, 528, 2765, 2073]. FORMAC [690, 679]. Formal [2621, 404, 890, 804, 1036, 1940, 1575, 1630, 1787, 1351, 1002, 2714, 204, 97, 238, 2396, 313, 2321, 91]. Formalism [1275, 603, 593]. Formally [1720, 2667]. Formally-Based [1720]. Forman [1319]. Format [1161]. formatting [453, 950]. Former [1809]. FORML [2310, 2311]. forms [2820]. FORmula [679]. formulae [2518, 2517]. formulae-as-type [2517]. Formulas [977]. Formulating [1237]. Formulation [1692]. Fornax [2482]. FORTH [1498, 2858, 2861, 148, 2322, 2500, 2503, 2509, 2510, 2442, 2504, 1497, 1604, 1879, 1715, 1639, 1878, 2017, 2056, 1935, 1950, 2148, 2016, 1951, 2057, 2149, 2044, 2218, 2219, 2257, 2258, 2505, 2513, 2506, 691, 2511, 2493, 2494, 2495, 2491, 2496, 677, 2310, 2311, 2512, 2507, 1640, 2501, 1419]. Forth-based [2504, 2493]. Forths [1417, 1418]. Fortran [1229, 457, 49, 232, 292, 293, 1412, 1413, 775, 652, 502, 466, 1956, 1729, 1253, 1180, 2440, 2441, 2046, 633, 1662]. Fortran-90 [502]. FORTRAN77 [1188]. Forum [2041, 919]. forwarding [2683]. Foundation [20, 2319, 424, 1125, 2522, 2774]. Foundations [944, 1145, 1771, 1048, 1049, 2421, 2585, 2574]. Four [460]. Fourier [2285]. Fourth [2866, 2868, 2861, 1915]. FP [201, 614]. Fractal [2021, 2020]. Fragile [2157]. fragmentation [2231, 1134]. fragmenting [2236]. fragments [2778]. frame [963]. Framework [1313, 1636, 1448, 2457, 1581, 16, 1311, 1310, 451, 1910, 2180, 2077, 1426, 2477, 2026, 1447, 2480, 2286, 1548, 1832, 831, 2676, 2593, 2534, 733, 1580, 1831, 2396, 184, 1312, 313, 317, 2610, 1496, 509, 1314, 2636, 2415, 976]. Framework-Centered [16]. Frameworks [1864, 566, 1166, 1226, 26, 1863, 1301, 1909]. France [2888, 2898]. Francisco [2853, 2862, 2877]. Free [2295, 1896, 1054, 1365, 763, 295, 296, 279, 2590, 476]. freeness [357, 2424]. Frequency [1446]. fresh [126, 127]. friend [1352]. friendly [1336, 443]. Frob [2200]. Front [2479, 2064, 1365, 245]. Front-End [2479, 1365]. Frontier [328]. Fudgets [2166]. full [2604, 2749, 354]. fuller [372]. Fully [2566, 2381, 1042, 2701]. fun [795, 1462]. Function [1572, 42, 1571, 987, 384, 612, 1837, 2682, 988, 736, 229, 963]. Functional [2887, 2898, 1904, 2385, 1378, 1379, 2323, 1964, 1029, 2212, 2157, 201, 1794, 1785, 1763, 2015, 1026, 1451, 894, 657, 2177, 1722, 2191, 969, 2200, 2358, 2201, 2437, 481, 606, 608, 1561, 2751, 2365, 723, 2261, 2308, 2196, 1159, 1507, 1611, 1934, 2043, 1270, 1789, 350, 2213, 66, 2526, 134, 1465, 952, 1027, 2764, 479, 198, 2606, 89, 2613, 222, 1780, 370, 1795, 1797, 2634, 2436, 277, 2802, 365, 2654, 2774, 1807, 2679, 2309, 2566, 1802, 2790, 2719, 1803, 82, 352, 2576, 1808, 2239, 391, 948, 231, 1806, 1595]. Functional-Imperative [723]. Functions [2198, 40, 39, 2206, 41, 2042, 2169, 2314, 2267, 341, 1178, 1175, 2549, 2199, 609, 1792, 2747, 2429, 2313, 1800, 1199]. functor [2691]. functors [2702, 2701]. Fund [1712]. Fundamentals [816]. further [382]. Fusing [1122]. Fusion [1981, 2734]. Futamura [2420]. Future [1233, 1812, 2084, 1573, 1867, 1625, 1626, 2299, 1057, 1614, 1813, 2707, 2489, 1861, 1232, 1302, 1463]. Futures [1075, 2346, 790]. Fuzzy [1251].

G [1166, 1226, 450, 2174, 456]. G-2 [456]. G-machine [2174]. G2L [1796]. GA [2251]. gaining [1797]. Galois [877]. Game [2099, 1038]. Gap [2472]. Garbage [2002, 1923, 714, 1654, 1668, 1717, 2019, 1224, 2279, 1429, 1683, 382, 1967, 1843, 173, 2332, 2579, 2178, 2237, 2331, 1055, 984, 1044, 2280, 1756, 533, 463, 957, 294, 748, 2232, 2534, 518, 2650, 2616, 2656, 749, 750, 1991, 295, 296, 2266, 338, 1788, 773, 2245, 1992, 2240, 2062, 751, 752, 2236, 1092, 1093, 393, 2241, 2411, 297, 298, 1923]. Garbage-Collected [984, 297, 298]. Garbage-Collector-Safety [1429].

Garbage-further [382]. Garfinkel [885]. Garment [1724, 1723]. gating [1199]. Gaussian [891]. GAWK [1492]. GC [1161]. GCODE [222]. gdb [1188]. General [1432, 985, 701, 960, 961, 2736, 2351, 81, 509]. general-purpose [2351]. generalisation [2423]. generalization [448]. Generalized [1448, 311, 2597, 352, 724]. Generate [2281, 208, 2691]. generated [208, 100]. Generating [1246, 1899, 783, 2742, 2598, 94, 220]. Generation [1958, 257, 204, 20, 1436, 2294, 656, 1065, 1434, 1684, 2003, 739, 1341, 749, 750, 1111, 2735, 1959, 234, 1797, 523, 100, 290, 1998, 2628]. generation-based [749, 750]. Generational [714, 1993, 1683, 2232, 2534, 2616, 393]. generative [48]. Generator [1189, 1506, 99, 2063, 283]. Generators [1444, 118]. Generic [1148, 1722, 1178, 2375, 1646, 839, 398, 2687, 2237, 2413]. Genericity [2084, 2085, 1277, 1058, 1276]. Generics [1058]. Genetic [2326]. gentle [478]. geometrical [953, 954]. geometry [2577, 2706, 2572]. Geophysical [2469]. Georgia [2895, 2896, 2897, 2252]. Get [2206, 880, 881, 420, 1105]. gigaflops [292, 293]. GILENA [2066]. Give [888, 887, 1449, 976]. GIVE-N-TAKE [976]. GJ [2395]. GKS [148]. Global [738, 304, 1156, 1616, 1427, 2459, 1218, 435, 433, 1539, 1516, 1752, 1609, 2638, 758, 759, 1777, 900, 290]. Globalization [367]. Globally [992, 2286]. GNU [1188, 524]. Go [1850, 1090]. Goal [1675]. Goal-Directed [1675]. Going [2010]. gold [1076]. Good [1654, 1972, 888, 1057, 1655, 2727, 887]. Gorlen [274]. Got [623]. governments [843]. gradient [787, 788, 331]. grail [448]. Grain [1738, 260, 1740, 1530, 2366, 300, 301, 2245, 323, 1739, 1114, 1115, 1706, 1707, 787, 788]. Grained [2089, 587, 960, 961]. Grammar [2318, 1149, 97, 121]. Grammars [719, 2152, 1004, 204, 613, 196, 1032, 1794, 2595, 490, 800, 234, 1796, 390, 279, 236, 2850, 119, 120]. Granularity [1749, 117]. Graph [1161, 1157, 1001, 1836, 2624, 103, 991, 2283, 299, 2724, 222, 1777, 2693, 2413, 297, 298]. Graph-Based [1157]. Graphic [891]. Graphical [165, 1365, 1784, 202]. Graphics [607, 2488, 862, 863, 844]. Graphs [1984, 1791, 1790, 1447, 998, 2639, 1763, 1985, 2550, 997, 2630, 2551, 2674, 1006]. greater [956]. greedy [1207]. Gregor [450]. Grid [639, 638]. GROOP [862, 863]. groundness [357]. Group [2886, 2470, 2858, 1559, 682]. Groups [2090, 2079, 1605]. Growing [1950, 1951]. GRT [839]. Guaranteeing [2241]. Guarded [1152, 490, 2190]. Guardians [749, 750]. GUI [1789]. Guide [95, 334, 397, 1056]. Guided [1286, 1051, 1285, 101]. gulf [2800]. GUM [1428]. Gupta [481]. GURRR [1156]. Guy [1229].

H [451, 157, 2173]. H/Direct [2173]. half [1933, 1934]. half-dozen [1933, 1934]. Halstead [2024]. Hamming [529]. Handbook [1229]. Handling [742, 853, 854, 193, 855, 172, 2047, 1097, 476]. Hansen [1953, 1881, 2046]. Happen [1654, 1655]. Happened [1868]. Hard [1564, 1563, 1330, 310, 1325, 2599]. Hardened [1554]. Hardware [1682, 438, 273, 2175, 1117, 260, 1751, 2119, 1097, 557, 1092, 1093, 1118]. Hardware-Assisted [438, 1092, 1093]. Hardware-Coherent [1751]. Hardware-Software [2119]. HARE [219]. harmful [1918, 447, 491, 150]. Hash [40, 39, 41, 612]. Hashing [42, 1175, 2557]. Haskell [606, 114, 2175, 2173, 478, 479, 2800, 2718, 2367, 2261, 2262, 2748, 1428, 2375, 724]. Hasn’t [1868]. having [795]. hazards [1208, 2672]. hcc [1506]. head [1821].

Header [2025]. Health [36]. Heap [1452, 2111, 2366, 2724, 1134, 240]. heap-directed [2724]. Heard [2159]. Heaven [2372]. Height [2284]. Heiny [607]. held [1719, 1650]. Hell [2372]. Helland [919]. Help [1867, 1310, 1064, 2070, 2071]. Heresy [589, 590]. heretical [45]. Hermes [528]. Heterogeneous [660]. Heuristic [1421]. heuristics [1479, 1480]. Hewlett [1278]. Hiding [1532, 1084]. Hierarchical [2377, 2234, 299, 331]. hierarchical-memory [331]. Hierarchies [562, 1560, 179, 853, 854, 877, 240]. Hierarchy [2119, 1565, 1855]. High [2891, 2450, 544, 258, 2207, 1249, 1250, 775, 2133, 2463, 1061, 2158, 2792, 1684, 1229, 250, 915, 2092, 682, 1261, 2485, 916, 2735, 2118, 71, 330, 2599, 778, 779, 615, 768, 2355, 1096, 1956, 1729, 2046, 1170, 1228, 633]. High-Bandwidth [250, 2118]. High-level [1249, 1250, 1684, 2092, 2735, 71, 2599, 778, 779, 615]. High-Performance [2891, 2463, 2485, 916]. high-speed [330, 1170, 1228]. Higher [2702, 2677, 2796, 2162, 2541, 1000, 2657, 2670, 119, 120, 2035, 2549, 1212, 2739, 1770, 1769, 731, 2781, 2603, 1771, 2526, 2660, 2555, 2606, 2722, 2821, 2701, 309, 2719, 2788, 2592]. Higher-Order [2162, 2035, 2702, 2796, 2541, 2657, 2670, 119, 120, 2549, 1212, 2739, 1770, 1769, 731, 2781, 2603, 1771, 2526, 2660, 2555, 2606, 2722, 2821, 2701, 2719, 2788, 2592]. Highly [1729, 836, 76, 1131]. Hilton [2252]. hinder [2070, 2071]. Hindrance [1064]. histogramming [1254]. Historical [1948]. History [2865, 2869, 1812, 672, 1813, 670, 673, 674, 675, 702, 676, 681, 1812]. hole [2802]. holes [1134]. HoME [498]. HOPL [2865, 2869, 699, 1812, 1813, 795]. HOPL-II [2865, 2869, 795]. Horn [53]. Horn-clause [53]. Host [1608, 1548]. hot [2256, 1999]. Hotlist [1359]. Hour [1840, 1839]. Howard [2774]. HPC [2456, 61]. HPC/VORX [61]. HPF [1738, 2440]. HPJava [2458]. Hu [818, 886]. Human [577]. Hurts [2471]. hybrid [2504, 1325, 2528, 548]. Hydrology [1636]. hylomorphisms [1456]. Hype [1864, 1863]. hypercube [2506, 321]. hyperfairness [2537]. Hypermedia [1053, 2201].

I/O [1242, 1324, 1061, 950]. IBM [272]. ICFP [2887, 2898, 1947]. ichbiah [440]. Icon [672, 692]. idea [1857]. ideas [210, 216]. identification [2836]. Identifier [2439, 2438]. Identifiers [86]. Identify [1573]. Identifying [123]. Idiom [2439]. Idiomatic [2]. idioms [2438, 2552]. IDL [1010, 1679]. Idle [1988, 2770]. IEC [1510, 1511]. IEEE [803, 1665]. If [761, 762]. If-Conversion [761, 762]. IFAD [1036]. IFIP [1916, 1605]. II [2865, 2869, 1269, 1937, 1812, 1813, 1720, 795, 1099, 1100]. III [1238]. illustrating [122]. ILP [1514, 2134]. im [773]. image [1254, 1231]. Images [1052]. impact [1112]. Impacts [872, 873]. Imperative [2212, 125, 1608, 894, 657, 2613, 1068, 723, 1895, 2183, 82, 1159, 1186, 2763, 2801, 2213, 731, 239, 1779, 170, 516, 1894, 361, 2599, 53, 2823]. Implement [1574, 2232, 2475]. Implementation [2852, 2855, 2862, 2867, 2874, 2886, 2889, 2893, 2895, 1210, 1211, 1928, 1929, 2252, 258, 776, 1827, 578, 1618, 1950, 1997, 1852, 1065, 1323, 1981, 1347, 2119, 1181, 2318, 1731, 2008, 1714, 644, 29, 716, 2665, 1428, 2883, 600, 1642, 661, 59, 2066, 2798, 1469, 66, 201, 1771, 2616, 134, 2595, 1991, 1465, 1951, 823, 2749, 415, 89, 121, 106, 1106, 1992, 2048, 2356, 965, 1182, 1730, 498, 1930, 65, 791, 727, 966, 1641, 481]. Implementations [820, 572, 205, 2534, 736, 2410, 2651]. Implemented [1078, 575, 2490]. Implementing [2406, 2407, 648, 1188, 785, 2132, 1782, 1783, 2203, 1026, 1368, 1558, 1068, 229, 753, 11, 2188, 1707, 2399, 1027, 76, 313, 1131, 1706, 790]. Implications [1752]. Implicative [2518]. implicit [2022, 1335]. Imprecise [2273, 2272]. Improve [1986, 2289, 747, 297, 298]. improved [2408, 2636]. Improvement [2830, 2708]. Improving [1983, 104, 2290, 746, 516, 1755, 975, 774, 1209, 327, 542, 1540, 1541, 1989, 1108, 1212, 1110]. impure [2732]. In-Cooperation [1650]. in-flight [2493]. In-memory [9]. In/Garbage [1654, 1668, 1717, 2019, 1923]. Inclusion [1984, 1838, 1985]. incomplete [80]. increase [1209, 2389]. Increment [268]. Incremental [2639, 225, 517, 2295, 995, 574, 1154, 1447, 1678, 2205, 573, 2228, 1471, 2737, 2528, 2230, 2649, 2673, 2241, 360, 2545]. independence [460]. Independent [1433, 639, 810, 2112, 960, 961, 2735, 1222, 1223, 638]. Index [1542, 1865, 1871, 2108, 2109, 2152, 884, 883]. Indexed [1504, 1402, 546]. indexing [852, 2536]. indirect [2682]. Induced [825, 2627, 2553]. induction [128, 508]. Inductive [2583, 1459, 1807]. Inductively [2385]. Industrial [527, 2299]. industry [1073]. inefficiency [1458]. Infant [714]. Inference [1284, 2383, 1291, 2291, 2812, 1074, 2180, 2569, 2822, 2738, 291, 2837, 1776, 112, 113, 2643, 2666, 2555, 2704, 2794, 2839, 2570, 2601, 2376, 402, 2815, 2817, 2818]. Infinitary [2783]. infinite [2697]. Infix [88]. Inflicted [2081]. informal [215]. Information [1938, 2312, 36, 1656, 1954, 2058, 2150, 2220, 2263, 1017, 381, 718, 425, 2441, 1080, 2340, 2092, 730, 395, 771, 772, 2737, 2844, 2811, 2823, 1332]. informatively [513]. Infrastructure [2451, 1062, 1123, 2298]. Inheritance [1191, 2523, 561, 1059, 1562, 1021, 848, 1022, 821, 2318, 1565, 562, 1058, 716, 177, 1190, 191, 192, 853, 854, 133, 1152, 823, 398, 371, 2522, 2568, 107, 108, 2780, 869, 870, 2402]. Inheritance-Like [1562]. Initialization [2096]. Inline [1676, 2067, 822]. Inline-Debugger [822]. Inlining [1686, 625, 1439, 2004, 1768]. Inner [2217]. input [1149, 80]. Insecure [2265, 2264]. Insensitive [1986, 718, 1194, 1195, 2758]. Insertion [1566]. insights [623]. Instance [1240, 582, 396, 1239]. instance-centered [396]. Instantiation [626, 627]. Institutions [37]. Instruction [256, 1162, 271, 253, 2114, 131, 2113, 157, 2280, 261, 246, 728, 304, 303, 475, 67, 760, 1209, 285, 755]. Instruction-Level [2114, 261, 1209]. Instructions [247, 1442, 1324, 2535]. instrument [2509, 2495]. instrumentation [768, 2345]. instrumenting [483]. Integer [1995, 1988, 1151, 2492, 142]. Integers [891, 972]. Integrated [1531, 1400, 668, 708, 1728]. Integrating [256, 395, 658, 769, 770, 456, 856, 857, 239, 277, 2346, 1107]. Integration [263, 1047, 1619, 1089, 18, 1166, 1226, 2077, 2319, 2446, 332, 53, 1063]. Integrity [591, 2824]. Intelligence [2148, 2149]. Intelligent [1659, 1153, 2416]. Intensional [1386, 2187, 1801, 2690, 2700]. Intensive [9, 631]. inter [1122]. Interaction [255, 1735, 2521, 2170, 577, 2638, 198, 2588, 2706]. interactions [2537, 737, 2563, 77, 2415]. Interactive [1508, 1043, 1310, 1052, 1509, 767, 898, 899, 2344, 2357, 1133]. Interactors} [77]. Interconnection [1013]. Interconvertibility [1896]. Interdependent [1828]. Interest [2886, 2858]. Interface [2872, 276, 2068, 2173, 1019, 1021, 797, 253, 1176, 1015, 1279, 1281, 1016, 725, 1171, 1130, 726, 398, 1105, 545, 339, 845, 846, 1022, 1011, 1280, 2664, 348, 1957]. Interfaces [2127, 1017, 560, 1020, 1012, 1060, 2066, 1339, 620]. interfacing [2509]. Interference [1702]. Interlanguage [2373]. Interleaving [1116]. Intermediate [2876, 1010, 1157, 2378, 1165, 1160, 2188, 1164, 2800, 2691]. International [2873, 2884, 2887, 2898, 1376, 1378, 2143, 2303, 2435, 1499, 2041, 1377, 195, 2434, 619, 1500]. Internet [2116, 1586, 1587, 2094, 2095, 2486].

Interoperability [2456, 1322]. Interpretation [125, 1680, 2380, 2384, 2756, 128, 2525, 357, 2197, 362, 2330, 2548, 157]. interpretations [2583, 2782, 2538, 2799]. Interpreter [1018, 1849, 1053, 484, 2716, 137]. Interpreters [1528, 1225, 614, 2514, 2688, 2717]. Interpreting [1486]. Interprocedural [1219, 2269, 1687, 986, 987, 1685, 736, 718, 732, 1196, 1197, 2627, 2683, 2813, 2692, 988, 2743, 515, 2344, 955, 2693, 2628]. Interrupts [2366]. intersection [2839, 2370, 2811]. Intraprocedural [1690, 1691]. Intrinsic [1523]. Introducing [1610, 1240, 1274, 1239]. Introduction [1639, 895, 606, 465, 685, 696, 486, 684, 690, 688, 692, 689, 478, 697, 691, 694, 701, 686, 223, 695, 2620, 687, 693, 1640, 2247, 704, 38]. Introductory [1379, 943]. Invariant [972]. Invariants [1423, 1862]. Inventor [844]. Inversion [353, 628]. Investigating [2129]. Investigation [2036]. investigations [382]. Invitation [1357, 2037, 1358]. Invited [2026, 2359, 2360, 2417, 2361, 2421, 2427]. invocation [2356, 2481]. invocations [654]. Involving [198]. IR’95 [1164, 2876]. Ira [1319]. IRIS [844]. Irregular [1734, 1729, 1733, 641, 785, 1247, 512, 1248]. Irregularly [660]. Ishikawa [1499, 1500]. Islands [412]. ISMM [2053]. isn’t [1458]. ISO [530, 1510, 1511]. ISO-WG21 [489]. ISO/IEC [1510, 1511]. ISO/VDM [530]. ISO/VDM-SL [530]. Isolating [2689]. Isolation [729, 2126]. isomorphisms [2593]. isomorphs [2730]. Issues [1564, 601, 602, 403, 2340, 183, 1563, 240]. ISWIM [709]. ISWIM-Like [709]. Iterated [2741]. Iteration [1137, 2371, 2426, 2582, 509]. iteration-reordering [509]. Iterative [115, 1653, 164, 581]. iterative-design [164]. iterator [2063]. iterator-based [2063]. iterators [224]. itself [1861]. IV [457]. iWarp [642]. IWPC [1499, 1500].

J

[1167, 466, 439, 440, 481, 606, 815, 1958, 1007]. Jade [323, 2585]. jalapeño [2407, 2406]. James [381]. Jan [157]. January [2853, 2856, 2872, 2870, 2877, 2885, 2888, 2892, 2900, 2901, 2902, 2860, 2863]. Japan [1499, 1500]. Jatich [705]. Java [2891, 1968, 1969, 2795, 2003, 1826, 2002, 2450, 2406, 2407, 1958, 2001, 2082, 2768, 2451, 2145, 2217, 2256, 2171, 2373, 1588, 1589, 2453, 2387, 2388, 2454, 1827, 2455, 2084, 2418, 2456, 2100, 2403, 2457, 2458, 2085, 2432, 2433, 2386, 2460, 1718, 1971, 1970, 2068, 2461, 2098, 2393, 1959, 2462, 2096, 2396, 2463, 1973, 2464, 2465, 1164, 2466, 2264, 2265, 2156, 2468, 2395, 2469, 2097, 2471, 2473, 2475, 2378, 2048, 2049, 2069, 2356, 1972, 2062, 1921, 2476, 2832, 2769, 2292, 2151, 2477, 2478, 2086, 2479, 2808, 2280, 2480, 2404, 2083, 2398, 2484, 2485, 2831, 2486, 2482, 1719]. Java-based [2450]. Java-RMI [2455]. Java/HPC [2456]. javab [2452]. JavaScript [2255, 2256, 1884]. JDBC [1974]. Jersey [789]. JFlow [2844]. JICC [1968]. Jim [450]. JMLC [1481]. John [662, 607, 274, 481, 796]. Johnson [381]. join [2754, 131]. join-calculus [2754]. Joint [2140]. Jose [2873, 1612]. journey [2655]. JPVM [2462]. Jr [1229]. JRes [2068]. JSpace [2048]. JSPICE [2479]. JTED [2460]. Julian [480]. July [715, 1605]. Jump [994, 736]. jumps [522]. June [2862, 2874, 2889, 2893, 2852, 2855, 2867]. just [2003, 2256, 152, 2767, 1]. just-in-time [2003]. Justifying [1971]. JWarp [2453].

Kahn [1470, 2539]. Kala [409].

Kaleidoscope [170, 1068]. Kanazawa [1499, 1500]. Kansas [2861]. KDOPS [1035]. Keep [2308, 2309]. Keith [274]. Kennedy [718]. Kernel [634, 1462, 187, 65, 410]. Kernel-Control [634]. key [2488, 394]. Keynote [1868]. Keywords [2438, 2306, 2307, 2439]. Kiczales [450]. Kinds [1975]. King [706]. Kingdom [1968]. KL1 [1003]. Knoop [722]. Knowledge [1561, 408]. Koelbel [1229]. Konrad [622, 275]. Krell [1170, 1228].

L [885, 1230, 1191, 1190, 1808]. l-flow [1808]. L-systems [1190, 1191, 1190]. label [2653]. label-selective [2653]. Lag [1452]. LAmbda [2677, 1770, 1769, 1897, 1898, 2081, 2189, 2170, 1910, 2709, 2565, 2514, 2653, 2577, 1451, 2555, 2515, 1458, 1471, 2655, 2567, 1909, 2611, 2665, 2191]. Lambda-Calculi [1910, 2567, 1909]. Lambda-Dropping [1898, 1897]. lambda-sigma [1451]. Lambda-splitting [1770, 1769]. lambdas [2239]. LAND* [2190]. landmine [802]. Language [2852, 2855, 2862, 2867, 2874, 2886, 2889, 2893, 2895, 2059, 1433, 1924, 2252, 1010, 2181, 719, 1717, 2084, 1975, 1608, 1720, 2085, 1443, 1330, 662, 1846, 2445, 1581, 637, 2173, 1715, 1950, 2096, 1997, 1726, 1019, 1479, 2467, 2206, 1368, 2159, 1354, 707, 895, 1606, 1236, 657, 709, 1963, 2060, 1137, 1940, 1381, 2367, 2378, 440, 1035, 1346, 1039, 1896, 1181, 1546, 723, 1955, 1628, 653, 162, 1015, 1172, 1889, 1281, 644, 1, 588, 1016, 665, 2883, 682, 1357, 1405, 628, 1922, 713, 725, 2446, 1646, 1789, 1826, 2569, 1210, 1211, 59, 177, 197, 960, 961, 2022, 2066, 206]. language [140, 2714, 1716, 85, 712, 726, 1120, 776, 2698, 828, 2631, 526, 2713, 2221, 487, 66, 1125, 1847, 486, 388, 802, 149, 953, 954, 518, 1580, 535, 2735, 2203, 472, 1465, 1951, 2396, 2846, 672, 415, 452, 2167, 1235, 479, 471, 189, 198, 2606, 2793, 1885, 2800, 2598, 337, 571, 2197, 370, 228, 1106, 2416, 610, 528, 1657, 965, 1182, 570, 2803, 1127, 365, 1545, 2615, 2391, 50, 223, 235, 456, 79, 777, 932, 1331, 309, 2825, 678, 460, 1011, 2055, 1173, 1888, 1280, 2823, 352, 1185, 1505, 1808, 2239, 1133, 217, 966, 611, 210, 231, 98, 1806, 489].

Language [278, 216]. Language-Independent [1433]. Languages [2860, 2865, 2863, 2869, 2872, 2870, 2875, 2873, 2877, 2878, 2884, 2885, 2886, 2888, 2892, 2896, 2899, 669, 1374, 1284, 1379, 1410, 1915, 2303, 2302, 2212, 125, 1812, 2207, 568, 583, 1848, 1670, 2881, 1026, 1836, 1615, 803, 2210, 2211, 1037, 1385, 1000, 1981, 1068, 1295, 2158, 575, 1305, 1874, 620, 1724, 1238, 2365, 2179, 1160, 1074, 1025, 1895, 1948, 1949, 2054, 2188, 1054, 1507, 1756, 1707, 1139, 1614, 1662, 1631, 2043, 1678, 2882, 2186, 2379, 195, 2024, 2853, 2856, 2901, 1375, 1212, 2359, 2739, 2596, 944, 2387, 731, 204, 213, 392, 2683, 1204, 1552, 1779, 340, 1999, 2374, 1669]. languages [1339, 1027, 2771, 295, 296, 2731, 356, 412, 1894, 855, 773, 499, 2722, 2821, 1232, 2661, 697, 228, 175, 83, 1304, 836, 141, 277, 956, 361, 676, 1040, 2599, 2715, 2566, 2623, 2689, 2397, 2725, 615, 2790, 2719, 358, 569, 791, 1706, 58, 1138, 1632, 467, 1723, 476, 2890, 1145, 815, 1611]. larceny [2073]. Large [34, 2469, 1516, 37, 1756, 33, 2354, 1102, 2417, 513, 2242, 330, 874, 875, 343, 838, 963, 581]. Large-Scale [2469, 1516, 34, 838]. Largeness [1834, 1833]. late [2191, 2767]. Latency [1532, 1539, 540, 760]. Lattice [640]. Lattices [877]. Laurence [525]. Lava [2175]. Law [1748]. Layer [1561, 2636]. layered [2840]. layers [1090]. Layout [2270, 1287, 1221, 107, 108, 2402]. laziness [2514, 372]. Lazy [1904, 459, 722, 2176, 514, 969, 2675, 2365, 485, 1207, 2598, 2204, 2619, 277, 2830, 1455, 456, 352, 2239, 529]. LCAS [442]. LCM [1106]. LCR98 [1915]. LCTES [1875]. LCTES’99 [2896]. Leading [1278]. Leading-Edge [1278]. leaks [2430]. Leapfrogging [790]. Learn [2019, 1918, 1937, 2018]. Learned [1558, 2093, 571]. Learning [1939, 1975, 567, 1051, 2409]. lecture [2359, 2360, 2361]. Left [1919, 490].

Left-to-Right [1919]. Legacy [30, 31]. Len [1171]. Length [2125]. Leopard [623]. Less [609, 2547, 1103, 91]. Lessons [1558, 1278, 1853, 2093, 1054, 1321, 1288]. Let [1449]. LET* [2190]. Let-floating [1449]. Letter [1487, 2207, 1023, 42, 2431, 2045, 1361, 1362, 1613, 1634, 1663, 1809, 1371, 1872, 1647, 1175]. Letters [2223, 2221, 2222, 2224]. Level [1424, 719, 258, 2207, 263, 432, 1710, 2114, 2158, 1305, 1910, 261, 1249, 1250, 1129, 2735, 60, 877, 800, 67, 374, 71, 1315, 1709, 1209, 461, 1304, 1909, 2599, 1153, 2792, 1684, 778, 779, 615, 768, 915, 2092]. Lex [1132]. lexical [2647]. lexically [617]. Lezere [1603]. Liberation [430]. Libes [796]. Libraries [2463, 601, 1822, 2454, 395, 602, 1823, 2237]. Library [832, 2453, 632, 1601, 1242, 638, 639, 2183, 464, 2413, 1406, 1405, 2319]. License [2261, 2262]. Life [1858, 1857, 1038]. Life-cycle [1858, 1857]. Lifetime [756, 757, 2111, 747, 2526, 368]. Lifetime-sensitive [756, 757]. Lightweight [599, 1852, 2094, 2095, 2034, 2685]. Like [1562, 709, 2186]. limitation [2548]. limiting [986]. LimitLESS [265]. Limits [261]. Linda [207, 2048]. Line [1689, 1889, 536, 1888]. Linear [1142, 2578, 1700, 981, 1252, 305, 306, 455, 474, 473, 1507, 1611, 1990, 533, 936, 1141, 2596, 941, 2762, 214, 2829, 2784, 2424, 2694, 2726, 1271, 130, 375]. Linear-scan [1990]. Linear-Time [1700, 1252, 305, 306]. Linearity [2753]. Linearization [1551, 1059]. lines [1466]. Linguistic [1736, 654]. Link [971, 1205]. Link-time [971, 1205]. Linkage [341]. Linked [2120, 963]. linking [2778, 2846]. links [1453]. LISA [1181, 1182, 2318]. Lisp [2374, 797, 469, 373, 391, 533, 339, 52, 2732, 680, 693, 2065, 137, 617]. List [1364, 963]. Listen [2459]. Lists [1177, 504]. Literate [49, 1490, 1591, 1355, 1408, 2155, 211]. Live [2331, 367, 898, 899]. Lively [533]. Liveness [2002]. living [2506]. LL [1383]. Load [2276, 758, 759, 1527, 794, 60, 305, 306]. Load-Reuse [2276]. Load/store [758, 759]. Loading [2069]. Loads [1978, 1979, 539]. loans [2073]. Local [2002, 2328, 2815, 2129, 544, 783, 134, 2190, 2622, 2654, 2708]. Locality [1527, 1522, 2128, 1519, 2289, 1541, 2121, 738, 1110, 794, 746, 793, 1260, 297, 298, 282]. Locality-Aware [2128]. lock [2399]. Locking [1079, 2400]. Locks [2001, 789]. log [536]. log-structured [536]. LogC [611]. LogGP [2353]. Logic [1904, 2385, 1929, 1608, 1659, 662, 715, 1432, 1696, 1169, 2227, 720, 59, 936, 1328, 1329, 2637, 2645, 357, 2668, 2710, 117, 239, 2731, 2764, 2529, 116, 2606, 168, 2564, 2633, 2635, 2669, 373, 2679, 2789, 1930, 2670, 53, 2533, 2835, 358, 503, 58, 375, 721, 938, 939]. Logical [2080, 1135, 2573, 171, 2421, 2620]. Logics [1561]. LogP [764]. London [1719]. Long [995, 2246]. Long-Running [995, 2246]. Look [2557, 2136, 1146, 533, 1352, 1460, 126, 127, 1147, 946, 947]. Lookahead [1383]. Looking [1829, 1830]. Lookup [1677, 574, 573]. Loop [635, 1522, 2327, 1270, 1122, 300, 301, 1247, 510, 559, 317, 509, 948, 511]. Loops [2112, 318, 1122, 909, 910, 319, 519, 1216]. LoPC [1757]. Loren [607]. losing [1797]. Lot [2019, 1918, 1937, 2018]. Lotos [1148]. Loveman [1229]. loving [2821]. Low [1315, 1530, 2499]. Low-cost [2499]. Low-level [1315]. Lower [2704]. LR [2596, 1383, 80]. LRPD [1216]. Lucinda [344].

M [466, 439, 440, 157, 373]. M-LISP [373]. M.T.A. [1269]. ma [533, 2557, 2902]. Machine [1717, 1162, 260, 2193, 1950, 1026, 2114, 2288, 1068, 810, 647, 1442, 1033, 960, 961, 1716, 2520, 1422, 2735, 1951, 1027, 2742, 505, 320, 1222, 1223, 2706, 2174, 932, 2831, 617, 1958, 292, 293, 2374, 2069].

Machine-Independent [810, 2735, 1222, 1223]. Machinery [2886]. Machines [1666, 651, 1516, 635, 1277, 636, 1756, 650, 739, 738, 2082, 304, 2738, 502, 1215, 1959, 364, 1248, 2535, 1276, 2350, 787, 788, 2002, 2083, 2484]. Macintosh [2496]. Macro [615, 1797]. Macros [2558, 616, 743]. Made [39]. Madness [1952]. MagPIe [2352]. Mahoney [885]. main [1094, 1095]. Mainframe [705, 1062]. Maintaining [1822, 1823, 877, 2673]. Maintenance [381]. Make [1854, 2547]. makes [350, 702]. Making [2084, 392, 1814, 1966, 22, 2531, 1344]. Malik [2207]. Man [1140]. Manage [9]. Management [1142, 2099, 2006, 14, 2136, 2470, 2317, 1559, 1212, 1141, 2847, 2238, 551, 830, 194, 1786, 183, 1787, 546, 240, 410, 2628]. Manager [580, 1133, 797]. Managing [419, 1283, 381, 1567, 310]. MANEVRO [618]. Manifest [2659]. Manifesting [1279]. Manipulating [640]. MAnipulation [679, 2864, 2871, 2900, 832, 1490, 2902, 1491]. manipulators [950]. Manual [2452, 1720, 1606, 535, 458]. manually [138]. Manufacturing [1313, 2810, 7, 1314]. MAP [2196]. Maple [786]. mapped [1102]. Mappers [647]. Mapping [316, 1689, 2112, 1208, 1255]. Mappings [652, 1747]. Maps [1071, 1072]. March [1952, 2858, 2854, 2861, 1952]. Marion [303]. mark [2244, 2580, 2245, 2235]. mark-&-sweep [2245]. mark-and-sweep [2580]. mark-sweep [2235]. market [2508]. marketing [2500]. marking [2229]. marriage [2165]. Marsokhod [1342, 1343]. Martian [1342, 1343]. Martin [1227]. Mary [1229]. Maspar [646]. MASPLAS [1652, 1370]. MASPLAS’96 [1363]. MASPLAS98 [2013]. Massachusetts [2869, 2884, 2900, 480]. Masses [2217, 2216]. Massively [651, 1381, 502, 64]. Matching [1919, 2226, 1460, 358, 2225].

Mathematical [1955, 1246]. Mathematics [465, 477]. Matrices [2364]. Matrix [1490, 1591, 1750, 1491, 1592]. Matrix-Multiplication [1750]. Maximal [2804]. Maximizing [2773]. May [2886, 2887, 2895, 2896, 2897, 2252, 2251, 730, 986, 623, 2868]. may-alias [730, 986]. MBURG [1366]. mcc [1508, 1509]. MCode [1165]. MDC [201]. Meaning [2061]. measure [2791]. measured [1788]. Measurement [621, 241, 275, 622]. Measurements [631]. Measuring [1824]. Mechanism [257, 2127, 2274, 1562, 1844, 592, 1137, 2089, 1724, 926, 1263, 1723]. Mechanisms [561, 2136, 1752, 1736, 1788]. Mechanization [801]. Media [2132, 158]. Medical [36]. Meeting [603, 1605]. Meetings [1650, 1649]. Meets [1561, 1795]. Megaprogramming [598]. Member [1677]. Members [2007]. memo [1792]. Memoing [2132]. Memoization [2135, 1776, 952]. memoizing [2042]. Memory [1737, 254, 648, 264, 1443, 1708, 1738, 2110, 982, 2238, 2650, 2029, 1531, 1533, 1425, 586, 637, 2006, 267, 2136, 1539, 1751, 2472, 247, 1252, 658, 1316, 2246, 1733, 2119, 1514, 2131, 2331, 647, 2137, 1445, 1536, 1530, 636, 250, 1756, 2129, 650, 2126, 1754, 1219, 1212, 739, 537, 747, 72, 785, 1091, 540, 2587, 1215, 2847, 1104, 746, 551, 1089, 557, 793, 1259, 2231, 320, 760, 74, 769, 770, 1317, 830, 558, 326, 1106, 2240, 9, 2343, 331, 324, 912, 913, 1248, 897, 2357, 1114, 1115, 2430, 2402, 2350, 1787, 1133, 1134, 240, 1094, 1095, 1753]. memory-efficient [2240]. mentor [1476]. Mentoring [1477]. Merging [2283, 285]. Mesh [1743]. Mesh-Partitioning [1743]. Message [1819, 2472, 658, 659, 172, 1820, 115, 1091, 1125, 911, 401, 1089, 769, 770, 1957]. Message-Passing [658, 1091, 769, 770, 1957]. Messages [1237, 2816, 1260, 1263]. messaging [1090]. Meta [1305, 450, 2399, 374, 1304].

Meta-Level [1305, 374, 1304]. meta-lock [2399]. Meta-Object [450]. Metaclass [2074, 1083]. metaclasses [1082]. metacomputation [353]. metalanguage [315]. Metamethods [1673]. MetaML [2422]. Metamodeling [17]. Metaobject [1303, 180]. MetaObjects [1306]. metatheory [2790]. Metcalf [466]. Method [1066, 1150, 40, 41, 1565, 1818, 574, 2196, 2481, 1642, 1067, 573, 1776, 1152, 2155, 1817, 512, 2492, 2356, 829, 2522, 2589, 963, 1641, 937]. methodologies [2575]. Methodology [643, 1349, 1064, 1306, 782, 221, 76, 310, 1131]. Methods [1827, 10, 1046, 576, 15, 582, 1421, 1170, 1228, 577, 892, 400, 1292, 2781, 2447, 1152, 893, 903, 904, 1002]. Metric [624, 2321, 349]. Metrics [804, 1643, 144, 1189, 1176, 621, 1420, 23, 407, 221, 332, 241, 54, 275, 622]. Mexico [2860, 2867]. MI [1063]. Micro [705]. Microarchitectural [2367]. microprocessor [1098, 1096]. Microprocessors [2360, 1129]. Microsoft [2476, 2261, 2262, 2319]. Mid [1870]. Mid-Year [1870]. Midyear [1962]. Migratable [2480]. Migrating [553, 1853]. migration [1244, 1087, 793]. Mike [465, 486]. Miles [1840, 1839]. Miller [919]. MIMD [320, 787, 788]. Mind.Forth [2148, 2149]. minds [1602, 1603]. Minimal [40, 39, 42, 260, 41, 1919, 612, 2779, 1707, 1175, 2395, 94, 1706, 1185]. Minimization [1702]. Minimizing [1029, 1289, 2328, 2773]. Minimum [2743]. minority [795]. minutes [1605]. MIPS [271]. Miranda [138]. Miss [2130]. Misses [1980, 1102]. mixed [2419, 391]. Mixin [191, 1473, 823]. Mixin-based [191]. mixing [170]. Mixins [2168, 2809]. ML [968, 648, 2171, 2586, 2616, 2559, 307, 1620, 2368, 2800, 1020, 1434, 2544, 1799, 1798, 525, 2568, 789, 2589, 2376, 2761, 1437, 1627, 2362, 1206, 1438, 431, 1998, 2186]. ML-like [2186].

ML-style [2589]. MLS [859]. MO [2861]. Mobile [1433, 2178, 2833, 2405, 2678, 2124, 2825, 2834]. Möbius [1952]. Moby [2274]. mod [446]. modal [2745, 1998]. Modal-ML [1998]. MoDE [186]. Model [2117, 1854, 1920, 640, 2091, 2456, 2605, 1683, 1300, 2771, 1001, 1368, 1549, 2384, 1819, 1030, 429, 1546, 2077, 1728, 661, 2714, 1241, 66, 775, 764, 300, 301, 164, 2764, 338, 2502, 94, 2410, 228, 1458, 965, 1545, 2799, 62, 2787, 966, 391, 1820]. modeled [848]. Modeling [596, 2445, 1757, 1835, 56, 1069, 1037, 1828, 2572, 858, 1594, 1249, 1250, 2447, 2390, 335]. Modelling [2584, 1796]. Models [1636, 804, 2291, 267, 894, 2156, 421, 451, 1514, 2134, 2561, 2092, 641, 1246, 2603, 2851, 202, 405, 1862, 277, 2607, 1787, 469]. Moderator [919]. Modern [2359, 1326, 1327]. Modest [889, 1597]. modifiable [204, 236]. Modification [253, 2079, 97, 732]. Modifying [1818, 1817, 608]. Modula [468, 526, 226, 208, 1205, 90, 1510, 1883, 1511, 817, 500]. Modula-2 [468, 226, 208, 90, 1510, 1883, 1511, 817, 500]. Modula-3 [526, 1205]. MODULA2 [452]. Modular [2393, 2168, 1606, 1955, 1294, 1758, 2645, 2846, 2729, 2717]. Modularity [1727, 2595]. modularization [2778]. Module [2005, 2275, 1693, 1694, 926, 2172, 1770, 1769, 213, 449, 102]. Module-Sensitive [1693, 1694]. Modules [1046, 2380, 645, 2381, 2646, 1473, 1999, 2541, 2660, 2659, 2701, 2361, 2568, 2592]. Modulo [1692, 2329, 756, 757]. Modulus [612]. Molecular [1853]. Moment [715]. Monad [2717]. Monadic [1696, 2164, 1781, 2183, 2362, 2801]. monads [2762, 2686, 2840, 2688, 2165]. Monitoring [436, 423, 2034, 1671, 425, 313, 2093, 1098, 55, 2493, 148, 1672]. Monitors [670, 2475, 70, 313]. monotone [370]. Monotonic [1551, 561, 1059]. Montréal [2893]. Morel [1609]. Mortality [714].

Mostly [294, 2412, 2844, 2235]. Mostly-copying [2412, 2235]. mostly-static [2844]. motion [463, 1218, 514, 1217, 2342, 722]. motivation [461]. Movement [643, 782]. Moving [1449, 403, 2236]. Mozart [2259, 2260]. MP [2346]. mpC [1381]. MPI [2352, 2350]. m}pilation [2304]. MUCKing [2056, 2057]. Multi [1066, 1142, 1827, 2457, 1046, 2132, 1710, 1381, 1347, 1910, 2679, 62, 1805, 1912, 1804, 1911, 1364, 400, 1067, 2537, 1141, 2781, 60, 1709, 1101, 103, 1909, 1153, 243, 2823, 2422]. Multi-Level [1710, 1910, 60, 1709, 1909]. Multi-List [1364]. Multi-Media [2132]. Multi-Method [1066, 1067]. Multi-Methods [1827, 1046, 400, 2781]. Multi-model [62]. Multi-Paradigm [1381]. multi-party [2537]. Multi-pass [2679]. multi-PSI [60]. multi-register [103]. Multi-Stage [1912, 1805, 1804, 1911, 2422]. Multi-Tasking [1347]. Multi-threaded [2457, 2823]. Multi-Threading [1142, 1141]. multi-user [1101]. multi-way [243]. Multibox [1008]. Multicomputer [1148, 487, 765]. multicomputers [321, 791]. multigraphs [2630]. Multilevel [861]. multiloop [507]. Multimedia [1748, 399, 2357]. Multiparadigm [1028, 235, 456]. multiparty [2563, 2588, 77, 2415]. multiphase [558]. multipipeline [1422]. Multiple [2031, 1059, 340, 2101, 2318, 903, 904, 645, 1524, 1525, 716, 179, 192, 853, 854, 2401, 133, 379, 398, 1775, 2647, 2633, 185, 107, 108, 869, 870, 2402, 1118]. Multiple-Block [1524, 1525]. Multiple-domain [903, 904]. Multiplication [1591, 2132, 1750, 972, 1592, 945]. Multiplications [1151]. multiprocessing [789]. Multiprocessor [438, 2279, 637, 2115, 1513, 1515, 785, 1087, 2656, 1089, 1112, 557, 326, 331, 1512, 915, 549]. Multiprocessors [648, 1538, 1708, 267, 1751, 1755, 481, 2126, 1258, 1537, 1259, 1116, 1088, 324, 792, 1261, 1107]. multiprogrammed [2350, 1261]. Multitasking [1170, 1228]. Multithreaded [1997, 2277, 1742, 1741, 1262, 2616]. Multithreading [1174, 1707, 1116, 1706]. Munin [72]. museum [698]. Musical [1595, 1032]. must [2695, 2821]. must-alias [2821]. mutable [2762, 773]. Mutation [1680]. mutual [554]. Mysteries [796]. Myth [1585, 1584]. Myths [2316].

N [446, 976]. name [2626, 1144, 2611, 86]. Named [2377]. names [2654]. namespace [2204]. Naming [95]. narrowing [2671]. National [2298]. Native [1958, 2463, 2467, 1959]. Native-Language-Based [2467]. Natural [95, 2066, 2619, 373]. naturally [2174]. Nature [1653]. Navigation [8]. navigational [948]. NDC [1660, 1661]. Near [1690, 1691, 2017, 2016]. Near-optimal [1690, 1691]. Need [1397, 2189, 1383, 1630, 1394, 1662, 1392, 2709, 2801, 1396, 2765, 2830, 1395, 1393]. needed [2671]. Needs [603, 1646]. Negation [1920]. neither [2557]. nemesis [2107, 2106]. NESL [1469]. Nest [1522]. Nested [1731, 1728, 776, 1152, 225, 909, 910, 1730, 92]. nests [1247]. net [300, 1948, 1949, 301]. Nets [2170, 2504, 2521]. Netstation [2116]. Network [2891, 2451, 2127, 2462, 1529, 2467, 1311, 1539, 2128, 1842, 134, 545, 1312]. Network-Based [1842]. Networked [1867]. Networks [649, 1231, 544, 1470, 2167, 330, 2784, 2351, 2539, 2355]. Neumann [2738]. Neural [1231, 2504]. neuron [2504]. Nevada [2889]. Never [2159, 2767]. Newly [1872, 2431]. News [1474, 2208, 1960, 1387, 1913, 1926, 2300, 1944, 1390, 1361, 1371]. Newton [2019, 2018, 1280, 1281]. Next [20, 867, 868]. NeXTStep [885]. Niklaus [1637]. Nineteenth [2860]. Ninflet [2480]. Ninth [2875]. nirvana [2107, 2106]. No [2043, 533, 2557]. Node [2283, 325, 2355]. nodes [1821, 2694]. Noised [470]. Non [640, 537, 941, 2193, 2189, 2424, 780, 781, 1445, 1742, 1741, 540, 952, 2538, 479, 2124, 2240, 2236, 2719, 2711, 2241, 2422, 1094, 1095]. non-blocking [540]. Non-concurrency [780, 781]. non-copying [2240]. Non-Determinism [2193]. Non-Deterministic [2189, 1445, 952, 2538]. non-disruptiveness [2241]. non-fragmenting [2236]. Non-linear [941]. non-moving [2236]. Non-numeric [1742, 1741]. non-strict [479, 2124, 2719, 2711]. non-trivial [2422]. Non-Uniform [640]. Non-volatile [537, 1094, 1095]. Nondeterminancy [911]. nondeterminism [2789]. Nonlinear [942]. Nook [1602, 1637, 1881]. normal [87]. Normalization [2202, 559]. Notation [2320, 2443]. notations [155]. Note [1367, 1507, 1673, 236, 1611]. Notes [1948, 159, 2648]. Notices [1023, 2448, 1352, 2010, 2221, 2222]. notion [192, 2517]. Novel [2269, 2610]. November [2899]. Nowak [705]. NOWs [2349]. NUMA [264, 326, 559, 635, 1541]. Number [664, 663]. numbering [1218]. Numbers [1431, 109, 2176, 111, 529]. numeric [1742, 1741, 1876]. Numerical [2466, 2454, 2478, 66]. NUML [965, 966]. NUSL [89]. NY [705].

O [481, 1242, 1324, 1061, 950]. O.LISP [203]. Oberon [1596, 1941, 135, 666]. Oberon-2 [666]. Obfuscated [796]. Object [2875, 2878, 2899, 1284, 595, 1854, 1636, 1830, 1920, 36, 2070, 2071, 1583, 840, 1589, 414, 2091, 816, 832, 1191, 1150, 2456, 596, 1071, 10, 578, 1484, 603, 1050, 20, 2445, 1570, 2067, 1554, 580, 1848, 1932, 2168, 5, 2096, 1001, 1836, 2090, 454, 1056, 1368, 1835, 2095, 19, 2156, 14, 2468, 584, 855, 189, 30, 1070, 2470, 21, 1061, 1558, 1607, 1048, 1069, 1316, 1385, 1000, 1860, 1346, 2416, 567, 1267, 1266, 1278, 1039, 575, 1305, 2089, 274, 885, 1561, 824, 582, 1287, 2317, 26, 1546, 1599, 841, 842, 2093, 1080, 402, 1818, 2632, 1074, 1824, 1674, 1079]. Object [1051, 1417, 1276, 1277, 593, 859, 1576, 1041, 1054, 218, 1184, 1, 29, 31, 588, 667, 32, 1052, 1856, 1559, 598, 1825, 450, 1420, 35, 2882, 1727, 1419, 27, 597, 565, 2756, 1076, 177, 960, 961, 867, 1829, 1246, 553, 1582, 165, 1121, 944, 2387, 1190, 828, 2631, 1072, 483, 115, 392, 407, 876, 388, 2447, 864, 1204, 1552, 193, 579, 1931, 188, 1479, 1480, 164, 184, 187, 415, 2094, 405, 394, 419, 178, 395, 1073, 2242, 1283, 412, 849, 850, 417, 207, 835, 1862, 1593, 1049, 397, 862, 863, 1317, 384, 181, 1859, 406, 874]. object [875, 242, 610, 1817, 175, 1124, 1302, 1304, 836, 570, 866, 389, 1127, 166, 1545, 1153, 2391, 176, 865, 2715, 172, 2761, 2397, 1418, 194, 1600, 868, 409, 1063, 872, 873, 183, 182, 837, 791, 856, 857, 2389, 833, 413, 444, 410, 469, 964, 2415, 174, 882, 98, 834, 1002, 1145, 1119, 2310, 2311, 851, 1133]. object-based [98, 1002]. Object-Orientation [1607]. Object-Oriented [2875, 2878, 2899, 1284, 595, 36, 2071, 1583, 832, 1191, 596, 1071, 10, 578, 1484, 603, 2445, 1570, 1848, 2168, 5, 1836, 1368, 2095, 19, 14, 584, 1070, 21, 1061, 1048, 1069, 1385, 1000, 1860, 1346, 567, 1267, 1039, 575, 1305, 274, 885, 1561, 824, 582, 2317, 26, 1546, 1080, 1818, 1074, 1824, 1079, 1051, 1417, 1277, 593, 1184, 1, 31, 588, 1052, 1856, 598, 1825, 1420, 35, 2882, 1419, 597, 565, 2070, 454, 189, 1266, 841, 842, 402, 2632, 1276, 1076, 177, 960, 961, 1582, 1121, 944, 2387, 1190, 828, 2631, 1072, 115, 392, 876, 388, 2447]. object-oriented [864, 1204, 1552, 193, 188, 1479, 1480, 164, 184, 187, 415, 2094, 405, 419, 178, 395, 1073, 412, 849, 850, 417, 207, 1862, 1593, 1049, 397, 862, 863, 181, 1859, 406, 874, 875, 610, 1817, 175, 1124, 1304, 836, 570, 389, 1127, 1545, 176, 865, 2715, 2761, 2397, 1418, 872, 873, 182, 837, 791, 856, 2389, 833, 413, 444, 469, 964, 2415, 882, 834, 1002, 1145, 1119, 851, 1133]. object-orientedness [242]. Object-preserving [414]. Object-Relational [1830, 1829]. Objective [2761, 2361]. objectivity [1130]. Objects [2449, 1830, 1557, 1077, 1142, 1555, 2091, 1307, 599, 28, 589, 1676, 1291, 586, 1867, 1085, 2081, 1308, 1549, 592, 1315, 861, 383, 1000, 2361, 1828, 9, 2178, 1055, 630, 1437, 587, 2111, 1577, 591, 711, 2480, 33, 1282, 577, 1135, 2227, 2756, 1829, 1141, 1318, 179, 876, 590, 170, 399, 2755, 1301, 878, 879, 394, 370, 2591, 185, 171, 531, 826, 2400, 1556, 963]. Objects-Storage [1142]. ObjectWindows [1167]. Observable [2604, 2080]. observation [464]. Observations [1935, 1936]. Obstacles [595]. Obtaining [2715]. OCCAM [159]. Occurrences [312, 434]. October [2875, 2873, 2878, 2884, 276, 1612]. OCTOPUS [2447]. ODBMS [1853]. ODE [1051]. ODMG [1854, 2091]. offer [699, 623]. offline [2419]. Offs [2119]. old [1352, 394, 417]. Olden [1244]. OMA [1559]. Omega [506]. OMG [1559]. OMT [596]. On-Demand [1529]. On-Line [1889, 536, 1888]. On-the-Fly [253, 909, 910, 898, 899]. Once [1142, 2828, 1141]. One [1430, 1843, 885, 2233, 1695, 2043, 1149, 2488, 1669, 286, 287, 152]. One-bit [2233]. One-Shot [1430]. Online [1984, 7, 1359, 2419, 1985]. Online-and-offline [2419]. Only [1530, 2359, 2360, 2417, 2361, 2427]. Ontario [2855]. onto [553]. OO [404, 1292, 1372, 1388, 1373, 1389, 1125, 576, 1568, 1569, 1322, 1819, 37, 15, 1288, 829, 17, 1025, 1062, 1820, 605, 23]. OO7 [1081]. OODB [7, 1323]. OODBMS [1081, 6, 860]. OOP [1403, 1404, 418, 568, 190, 1047, 869, 870, 569]. OOPSLA [2875, 2878, 2899, 880, 2879, 881, 1962, 1598, 12, 194, 1600, 420, 569, 2882, 2037]. OOT [1833, 1834]. opaque [2810]. Open [919, 2009, 1852, 766, 2218, 2834, 2219]. OpenMP [2349]. operands [103]. Operating [2873, 2884, 255, 599, 1520, 1315, 1541, 2427, 600, 1098, 2374, 2489, 550, 326, 1786, 1103, 1086, 549]. operating-system [1786]. Operation [7, 1549, 2093, 538, 1253, 1821]. operational [2698, 2676, 2830, 2790]. operationally [2787]. operationally-based [2787]. Operations [1030, 1324, 2352, 778, 779, 278]. operator [2560]. operators [2063]. opportunism [394]. opportunities [2427]. Opposition [1920]. OPS5 [841, 842]. Optimal [651, 1692, 2845, 2567, 897, 995, 314, 1421, 1255, 2129, 475, 2577, 2667, 771, 772, 2515, 1777, 1198, 305, 306, 548, 1690, 1691]. Optimality [1458, 2514]. Optimally [2581]. Optimisation [2163]. Optimistic [329, 1740, 1263, 1739]. O}ptimization [2304, 1995, 1284, 2005, 643, 1427, 2731, 1517, 251, 799, 2284, 359, 1609, 2226, 974, 739, 782, 729, 1249, 1250, 385, 2683, 1205, 2707, 1518, 2552, 2478, 2225, 971]. Optimizations [2296, 1982, 2281, 1732, 1996, 738, 1770, 1769, 1110, 733, 2297, 2426, 1209, 84, 1256]. optimize [1253]. Optimized [1424, 710, 992, 2286, 494, 497, 2028]. Optimizer [1610, 1685]. Optimizers [2161, 214, 501, 290]. Optimizing [1550, 1066, 1067, 573, 115, 1738, 2326, 1679, 996, 1214, 110, 1434, 2004, 2716, 574, 1158, 2047, 1438, 43, 1552, 126, 127, 2742, 1797, 955, 90, 2635, 1900, 368, 219, 69, 1123, 282]. OPTVIEW [2028]. OR-parallel [60]. or-parallelism [322, 939, 938]. oraclizable [2539]. Orca [85, 840, 644]. Orchestrating [737]. Order [2677, 2162, 1696, 1000, 2123, 2137, 2135, 2035, 682, 2549, 1212, 2739, 2702, 1770, 1769, 731, 2781, 2603, 1771, 2526, 2755, 2796, 2541, 2660, 2555, 2606, 2722, 2821, 2563, 2701, 2657, 2670, 2719, 2817, 2788, 119, 120, 2592]. ordered [654]. ordering [69]. Ordinary [1523]. Oregon [2870]. Organization [1264, 273, 1221, 153, 161, 351]. Organizations [2136, 874, 875]. Orientation [589, 1607, 590, 2495]. Oriented [2875, 2878, 2899, 1284, 595, 1636, 36, 2071, 1583, 816, 832, 1191, 1150, 596, 1071, 10, 578, 1484, 603, 42, 2445, 1570, 1848, 2168, 5, 1001, 1836, 1368, 2095, 19, 14, 584, 1070, 21, 1061, 1841, 1048, 1069, 1316, 1385, 1000, 1860, 1346, 567, 1267, 1039, 575, 1305, 274, 885, 1561, 824, 582, 2317, 26, 1546, 1599, 1080, 1297, 612, 1818, 1074, 1824, 1079, 1051, 1417, 1277, 593, 1579, 859, 1184, 1, 31, 588, 1139, 667, 1175, 1052, 1856, 598, 1825, 1420, 35, 2882, 1419, 597, 565, 1076, 177, 960, 961, 1246, 2070, 1582, 165, 1121, 944, 2387]. oriented [1190, 828, 2631, 1072, 115, 392, 407, 876, 2408, 388, 2447, 864, 1204, 1552, 193, 188, 2390, 1479, 1480, 164, 454, 184, 187, 415, 2094, 878, 879, 405, 419, 178, 395, 1073, 1283, 412, 849, 850, 189, 417, 207, 835, 2410, 1862, 2409, 1593, 1049, 397, 862, 863, 1317, 384, 181, 1859, 406, 874, 875, 610, 1817, 1266, 175, 1124, 1304, 836, 570, 389, 1127, 1545, 841, 842, 1298, 176, 402, 865, 2632, 2715, 172, 2761, 2397, 1418, 1276, 1578, 1063, 218, 872, 873, 182, 837, 791, 856, 1138, 2389, 833, 413, 444, 1922, 713, 725, 469, 964, 2415]. oriented [174, 882, 834, 1002, 1145, 1119, 851, 1133, 712, 726]. orientedness [242]. Orlando [2856, 2874]. Orlow [274]. orthodox [2073]. Orthogonal [136, 711, 2412, 2567]. orthogonality [2612]. OS/2 [1401]. OT [1857, 1858]. Other [803, 1282, 796, 2024]. Our [1614, 1812, 1813]. out-of-core [1241]. Out-Of-Order [2135, 2123, 2137]. outer [2747]. outline [528]. output [1149]. Overhead [1532, 1075, 1530, 2327, 2682, 1090, 1113]. Overlapping [2146, 2147, 2124].

Overloading [1566, 1918, 2542]. Overview [2015, 1405, 344, 2014, 1153, 530, 1406]. OVIDE [1143]. own [2144, 2145, 2762]. Ownership [2072, 1113]. Oxford [466, 1605].

P [481, 639, 90]. P-code [90]. PA [2886]. package [188]. Packaging [1183]. Packard [1278]. packet [2167]. PaCT [1376, 1377]. PaCT-95 [1376, 1377]. Padded [1033]. Page [1538, 2136, 1537, 1087, 551, 326]. page-cache [551]. Pages [1884, 2444]. Painting [185]. pair [2424]. pair-sharing [2424]. palimpsest [2502]. Pancode [155, 200]. Panel [1830, 2071, 1077, 880, 1812, 1587, 568, 2208, 1547, 1564, 2099, 1846, 589, 1840, 576, 1866, 2088, 1585, 1569, 1864, 2106, 1858, 861, 1869, 2105, 874, 1302, 1870, 564, 871, 1834, 2299, 860, 859, 870, 1577, 868, 1054, 591, 858, 2083, 598, 605, 867, 1829, 1292, 2070, 2082, 1813, 1586, 1584, 1553, 1307, 1563, 1050, 1847, 590, 1839, 1568, 1863, 2107, 1339, 1308, 1301, 1283, 1857, 881, 2104, 875, 563, 1288, 1833, 838, 869, 1057, 1576, 569, 857, 833, 882]. Panorama [908]. Paper [1629, 2026, 880, 2223, 881, 420, 569]. Papers [1481, 1410, 1475, 1502, 1501, 1873, 2142, 2143, 1947, 1915, 1946, 2012, 2052, 2039, 2303, 2304, 2302, 2040, 1968, 1969, 2860, 2863, 2870, 2877, 2885, 2888, 2892, 1970, 143]. Paradigm [1381, 766, 2405, 406]. Paradigms [1619, 1784, 239, 2343, 53]. PARALLEL [1376, 2854, 2857, 2866, 2868, 2897, 1995, 813, 2819, 1583, 640, 1254, 258, 1953, 1735, 2453, 1666, 2458, 651, 2460, 1243, 2880, 2462, 426, 1757, 2257, 1734, 820, 2463, 642, 1345, 1882, 427, 707, 2469, 1755, 21, 1252, 2473, 425, 1516, 1881, 1381, 639, 2562, 1546, 1499, 1500, 449, 810, 1748, 634, 653, 645, 393, 1377, 660, 1728, 2480, 1756, 1428, 2483, 315, 1745, 1744, 945, 1364, 720, 424, 1732, 1996, 2359, 935, 59, 177, 738, 1246, 2022, 1582, 140, 2354, 1341, 776, 294, 1241, 2697, 916, 57, 385, 291, 2609, 783, 502, 328, 2587, 201, 322, 914]. parallel [764, 911, 953, 954, 2772, 325, 1465, 60, 737, 2749, 784, 2750, 767, 332, 56, 793, 909, 910, 898, 899, 2342, 1214, 323, 1106, 2341, 63, 2416, 512, 908, 912, 913, 118, 897, 64, 314, 1545, 212, 917, 778, 779, 172, 2357, 65, 162, 62, 2607, 2629, 1255, 1505, 1808, 617, 721, 1806, 2258, 1169, 1231, 2484]. parallel-distributed [325]. Parallel-program [315, 328]. Parallelism [640, 260, 2265, 2114, 1731, 1930, 2533, 261, 720, 738, 2690, 322, 67, 2264, 71, 320, 2773, 1209, 1730, 777, 131, 2349, 2531, 765, 58, 938, 939, 1929, 721, 939]. Parallelization [2452, 1408, 1743, 2820, 331, 329, 1407, 2805, 774, 2348, 2552, 1216, 2347]. Parallelized [432, 430, 1740, 1739]. parallelizer [2344]. Parallelizing [978, 1349, 330, 1426, 786, 1247, 1105, 1123, 937]. Parameter [1183, 2565, 112, 113]. Parameter-passing [2565]. parameterization [1826]. Parameterized [2768, 288, 2393, 2729, 2784]. Parameters [719, 664, 663]. Parametric [1290, 2721, 2835, 2086, 2621, 2646, 2620]. Parametricity [2369, 2600, 2791, 2622]. Parametrized [893, 892]. ParaScope [767]. Parasitic [1827]. ParCeL [1505]. ParCeL-1 [1505]. Paris [2888, 2898]. Parity [538, 262]. PARLOG [47]. Parse [2254, 97, 2253]. parser [2063]. Parsers [1008, 613, 952, 234, 81, 80]. ParsesraP [1149]. Parsing [2147, 2214, 2254, 2307, 2439, 933, 1024, 717, 935, 2324, 244, 475, 490, 279]. Part [1715, 1269, 1669]. Partial [2864, 2871, 2900, 2902, 1984, 1887, 1901, 1899, 1688, 980, 1701, 1978, 1908, 1985, 1886, 372, 1891, 979, 1892, 1893, 1305, 2599, 2153, 1889, 358, 1609, 1440, 1695, 2315, 1219, 288, 2648, 2744, 1907, 2423, 1890, 2420, 2538, 356, 2598, 371, 355, 1979, 1304, 361, 1900, 2429, 2560, 2834, 1888, 359, 2419, 360, 352]. Partial-Evaluation [1893, 1892]. partially [2418]. Participation [1378, 1379, 1652, 1370, 1357, 1358]. Partitioning [1743, 2759, 334, 1340, 2711, 1453]. party [2537]. PASCAL [803, 951, 712, 726, 1121, 670, 689, 895, 694, 2494, 53, 44, 92, 683, 713, 725]. Pascal-oriented [712, 726, 713, 725]. PASCAL-XSC [803]. pass [2679, 245]. Passing [1183, 2472, 1155, 658, 659, 967, 963, 1091, 2676, 1125, 2565, 911, 401, 2687, 1089, 769, 770, 2617, 1957]. Past [1812, 2084, 1614, 1813]. PASTE [2142]. Pat [919]. patents [532]. Path [1983, 2814, 214, 1828, 157, 2125, 2807, 1335, 2491, 937]. Path-Based [1828, 937]. Path-sensitive [2814]. Paths [2212, 2213, 2293, 1540]. Pattern [2594, 1849, 1919, 2226, 2446, 1766, 1460, 172, 358, 2225, 1876]. Pattern-based [2594]. Pattern-Matching [1919]. Patterns [2078, 1313, 1404, 1372, 1388, 2147, 2214, 2254, 2307, 2439, 13, 1296, 1864, 1309, 1353, 1354, 584, 566, 21, 1310, 1849, 22, 1982, 1851, 1548, 2324, 2203, 1863, 166, 623, 1314, 1600]. Paulson [525]. PC [705, 1242]. PCCTS [458]. PCLOS [180]. Pedagogical [1600]. Peephole [2226, 2225]. Penalty [247]. People [1654, 1655]. PEPM [2900]. PEPM’93 [2864]. Perfect [40, 39, 42, 41, 612, 1175]. perform [302]. Performance [2891, 2450, 1682, 2354, 273, 2456, 1666, 1017, 914, 2290, 1533, 1750, 2133, 7, 2463, 267, 272, 1956, 1845, 2468, 1065, 572, 1729, 1751, 1061, 425, 251, 248, 262, 1335, 1752, 421, 6, 252, 2046, 1540, 1229, 2137, 984, 1528, 2337, 1092, 1093, 633, 2129, 2126, 2065, 2035, 973, 2121, 1754, 2484, 1989, 2485, 1749, 27, 334, 747, 916, 775, 2650, 188, 2405, 1112, 332, 855, 771, 772, 2228, 774, 333, 1101, 331, 243, 556, 2355, 1103, 549, 548, 1261, 1107, 1753].

Perlis [815]. Perry [274]. persistence [2412, 277]. Persistent [1078, 1056, 1558, 839, 415, 855]. Personal [2486, 670, 233]. Perspective [2207, 1618, 1385, 2158, 1834, 1633, 1833, 366]. Perspectives [1547, 1553, 1623]. Perturbation [421, 768, 333]. Petal [1521]. Peter [2224]. Petersburg [2885, 1376, 1377]. Petri [300, 301]. Petri-net [301, 300]. Phase [644, 2541]. phi [2694]. Phil [705]. Philadelphia [2886, 1388, 1389]. pHluid [1465]. physical [547, 551]. Pi [1891, 1890, 2753, 2776, 966]. Pi-Calculus [1891, 1890, 2753, 2776]. Pickett [439]. Pictures [604, 2195, 922]. PIDL [726, 725]. PIDL/Pascal [726]. Piece [2365]. Piff [465]. Pipeline [2330, 2672]. pipelined [143, 519]. Pipelining [1421, 1208, 302, 2610]. Pipes [2505, 654]. Pitas [1231]. Pizza [2769]. PL [1401]. PL/I [1401]. Placement [643, 2122, 1743, 2473, 262, 782, 2232, 326, 976]. placing [2694, 1199]. Plains [2852]. PLAN [2167]. Planning [1901]. Platform [437, 789]. Play [2075, 2076]. PLDI [2874, 2886, 2889, 2893, 2895, 1475, 2052, 2039, 2883]. pldi99= [2252]. plea [227, 199]. pleas [227]. Plexico [274]. Plug [2075, 2076]. Plug-and-Play [2075, 2076]. Point [1431, 109, 2471, 248, 1988, 111, 1394, 1395, 1662, 2065, 1876, 1392, 1393, 192, 208, 2582]. Point-extent [1876]. Point/Counterpoint [1392, 1393]. pointed [1459]. Pointer [993, 2029, 1986, 985, 1696, 2553, 825, 2277, 1015, 2278, 1732, 2027, 2627, 2806, 516, 732, 2430, 2398, 1192, 1193]. Pointer-Based [985, 1015, 1732]. Pointer-Induced [825, 2553, 2627]. Pointers [1029, 2461, 987, 129, 802, 986, 988, 2724, 447, 493, 1453]. Points [987, 2726, 1359, 988, 2758, 1360]. Points-to [987, 2726, 988, 2758]. Policies [264]. polyadic [2643]. Polycyclic [259]. polyhedra [318]. Polymorphic [1291, 2570, 2179, 1889, 1177, 2549, 2781, 2653, 2812, 2591, 377, 2589, 2615, 2776, 1888, 2828, 38]. polymorphically [365]. Polymorphism [1725, 2187, 1290, 1012, 2626, 2103, 491, 2086, 2621, 2603, 2699, 2625, 2700, 2794, 2102, 2376, 2620, 2397, 2817]. Polynomial [1952, 429, 952]. PolyP [2793]. polytypic [2793]. polyvariant [1758, 2811]. Pool [1034]. Poor [1140, 1128]. POPL [2853, 2856, 2870, 2877, 2885, 2888, 2890, 2892, 2901, 2302, 477]. POPL99 [2012]. Popular [2320, 2468, 2443, 1630]. Portability [1645, 1751, 2513, 90]. Portable [1482, 1412, 437, 2656, 822, 2243, 253, 1428, 1548, 1506, 776, 1413, 835, 908, 789, 219, 790, 2499]. Portia [396]. Porting [1508, 1509, 1596]. Portland [2872, 2870, 2875]. portrayals [905]. Position [1615, 1629, 569]. positioning [101]. Positive [2705]. post [2597]. post-dominators [2597]. poster [1789, 1790, 1793, 1794, 1796, 1795, 1797, 1798, 2192, 1807, 1800, 1801, 1802, 1803, 1804, 1806]. PostScript [1952]. power [1217, 2777, 2548, 929]. PowerEpsilon [930]. powerloop [83]. PowerPC [1129, 1509]. PowerPCs [1506]. pp [2448, 525]. PPOPP [2854, 2857, 2868, 1502, 2880]. PPOPP’99 [2897]. Practical [2324, 2147, 2214, 2254, 2307, 2439, 1432, 1036, 281, 2103, 2404, 727, 1344, 2739, 392, 733, 1861, 2102, 1772, 2844, 2843]. Practice [2854, 2857, 2866, 2868, 2897, 2880, 1309, 628, 2070, 2071, 2769, 1128]. Practitioners [603]. Pragmatic [1014, 2179, 1268]. Prasun [1171]. Pre [2154, 2448, 1950, 1951]. Pre-Forth [1950, 1951]. Pre-Processor [2448, 2154]. Precise [2130, 1074, 2693, 973]. Precision [2002]. preconditioned [787, 788]. preconditions [384]. predicated [2401, 2348]. Predicates [435, 914, 901, 900]. Predictable [1662]. Predicting [1666, 543, 1324, 284]. Prediction [1526, 1286, 975, 1527, 2125, 973, 2354, 763, 1202, 1203, 1285, 542, 1201, 2330, 1108, 66].

Predictive [2353]. Predictors [1524, 747, 1525]. prefecture [1499]. Prefecturem [1500]. Prefetching [249, 1535, 2120, 2403, 540, 1534, 541]. Prefix [2267, 88, 314]. Preliminary [2039, 813, 2029, 581, 210, 216, 943, 2677, 2097, 2681, 1153]. premature [948]. preparing [934]. Preprints [2865, 2869]. Preprocessor [653]. Presburger [977]. Prescriptive [1910, 1909]. Presence [1430, 987, 1294, 1208, 988, 105, 2530]. Present [1812, 1614, 1813]. presentation [2529]. presented [2860, 2863, 2870, 2877, 2885, 2888, 2892]. Preserve [2078]. preserves [1422]. preserving [414, 376]. Press [1489, 439, 440, 606, 525, 465, 451, 157, 450]. Pretenuring [1993]. Pretty [1644]. previous [543]. Price [1695, 286, 287]. Prime [26]. Primitive [87, 409]. Primitives [254, 1717, 1740, 1716, 2586, 2615, 1739]. Principal [2592, 2727]. Principality [2839]. Principals [2379]. principle [2539]. Principles [2853, 2854, 2856, 2857, 2860, 2863, 2866, 2868, 2870, 2877, 2885, 2888, 2890, 2892, 2897, 2901, 1410, 2302, 804, 2880, 1051, 2790]. Print [111]. Printing [1431, 1644]. Priority [1483, 1325]. Priority-Based [1483]. Privatization [1746, 650, 2608, 1216]. Probabilistic [520, 1467]. Problem [2157, 1316, 1317, 1004, 1467, 2390, 2231, 2553, 1271, 1031]. Problem-Oriented [1316, 1317, 2390]. Problems [1564, 1021, 1729, 1563, 932, 1802, 2725]. Procedure [1444, 1689, 285, 1273, 464, 2714, 1272, 102]. Procedures [896, 2612]. Proceedings [2852, 2853, 2854, 2855, 2856, 2857, 2861, 2862, 2864, 2867, 2868, 2871, 2886, 2887, 2889, 2893, 2894, 2895, 2896, 2897, 2898, 2899, 2900, 2901, 2902, 2879, 2859, 2878, 2884, 2890, 2858, 195]. Process [649, 890, 268, 906, 1485, 1712, 1510, 605, 565, 23, 70, 2519]. processes [167, 329, 2614, 2697, 2676, 2841, 2678, 2752, 2784, 148, 1257, 2825, 159]. Processing [2303, 2465, 1231, 2151, 2132, 2546, 134, 871, 615, 1332, 617]. Processor [2448, 1639, 248, 2135, 2154, 545, 243, 1640]. processor-network [545]. Processors [2282, 2328, 2133, 1514, 2137, 1536, 250, 316, 2502, 143, 555, 103, 1326, 1327, 2330]. Procs [789]. produce [1974]. producer [950]. produces [623]. Production [1035, 1421, 847, 321, 2228, 841, 842, 443]. production-quality [321]. productions [490]. products [2197]. Professional [1168, 1944, 1390, 1391]. Professionals [1050]. Proficiency [1183]. Profile [1993, 1285, 1286, 101]. Profile-Driven [1993]. Profile-Guided [1286, 1285]. Profiles [1983, 284]. Profiling [1682, 1754, 2581, 2807, 932, 1452, 2719, 1753]. Program [2864, 2871, 2894, 2900, 1995, 1374, 2269, 1612, 2142, 2252, 832, 2778, 1617, 1618, 649, 1432, 1693, 1423, 2130, 1517, 1043, 2036, 922, 2185, 2034, 981, 1180, 2293, 1621, 2358, 378, 1622, 2429, 365, 1623, 991, 1080, 1625, 1626, 2552, 932, 1629, 703, 1041, 2026, 983, 2161, 974, 1699, 1633, 2530, 725, 1642, 1136, 2902, 124, 2652, 453, 363, 726, 2639, 831, 328, 2740, 388, 1694, 2721, 543, 2707, 1518, 332, 123, 734, 735, 1471, 1124, 905, 2551, 2836, 2708, 495, 218, 2592, 289, 284, 581, 315, 2628, 1641, 125]. Programmable [653, 743, 1328, 1329]. Programmer [525, 2192]. Programmers [1397, 1394, 1357, 1392, 482, 1396, 1395, 1358, 1393]. Programming [2852, 2854, 2855, 2857, 2860, 2862, 2865, 2863, 2866, 2867, 2868, 2869, 2874, 2870, 2875, 2873, 2877, 2878, 2884, 2885, 2886, 2887, 2888, 2889, 2890, 2892, 2893, 2895, 2897, 2898, 2899, 669, 2117, 1410, 2252, 2302, 1964, 1583, 1161, 206, 49, 640, 1717, 2212, 1812, 2207, 1735, 816, 1939, 2074, 2084, 1490, 1591, 1355, 1408, 2157, 1608, 2457, 1720, 2085, 1659, 662, 1083, 715, 1432, 1038, 1581, 1167, 2015, 2880, 2168, 2374, 1726, 2463, 1973, 1026, 1001, 1168, 1615, 1882, 584, 1236, 2368, 1315, 1963, 803, 2210, 2211, 1841, 1137, 1940, 619, 1048, 1037, 1385, 1881, 1381, 607, 440, 1346, 135, 1068]. Programming [1084, 1267, 1295, 2158, 2201, 2437, 956, 274, 606, 885, 1561, 824, 582, 1169, 2444, 2169, 723, 1818, 2261, 815, 2308, 2041, 1955, 1146, 1674, 1628, 2054, 799, 2183, 668, 708, 645, 211, 1054, 1, 1728, 1912, 588, 1016, 1756, 51, 58, 1139, 1614, 2260, 1662, 1631, 1678, 2883, 1825, 724, 2882, 1922, 641, 948, 2379, 216, 1268, 1646, 2853, 2856, 2901, 1076, 766, 2323, 2763, 2022, 1582, 140, 1716, 85, 2213, 1813, 943, 154, 2637, 2631, 839, 57, 213, 487, 201, 486, 2668, 239, 193, 134, 1580, 1794, 1763, 340, 170, 470, 1027, 2771, 295, 296, 672]. programming [767, 415, 2764, 878, 879, 2155, 2824, 517, 855, 1235, 479, 345, 116, 2793, 2502, 2613, 1780, 168, 1232, 337, 227, 339, 2410, 2409, 697, 228, 1049, 397, 1795, 323, 610, 1817, 2562, 1657, 1266, 461, 83, 2436, 141, 199, 343, 618, 608, 676, 389, 1127, 2599, 176, 1496, 2632, 2309, 1930, 79, 2670, 1821, 53, 1147, 2357, 162, 460, 233, 2532, 2574, 2712, 2055, 1342, 1343, 62, 358, 2607, 1803, 218, 146, 1805, 1804, 1911, 1185, 1138, 96, 2259, 1505, 1632, 2576, 1934, 217, 2843, 469, 391, 945, 210, 964, 98, 1378, 1929, 1599, 1953]. programming-language [460]. Programs [1433, 994, 1904, 2385, 254, 438, 258, 1444, 1654, 1179, 2268, 1412, 1738, 1484, 2009, 1898, 428, 1570, 426, 642, 894, 427, 1523, 2384, 1696, 2177, 1252, 2021, 2030, 1860, 1274, 1705, 1893, 429, 2032, 995, 810, 984, 1740, 2277, 636, 710, 1159, 1856, 2035, 992, 2278, 720, 424, 2205, 1996, 631, 1375, 951, 1655, 316, 1341, 2581, 1241, 1413, 2682, 115, 1091, 291, 2609, 783, 2587, 2645, 357, 911, 1897, 2417, 2710, 117, 1032, 513, 2650, 2747, 1205, 60, 784, 2538, 2750, 2841, 454, 2529, 516, 452, 744, 745, 1253, 1449, 222, 898, 899, 2342, 927]. programs [928, 2020, 2124, 1214, 181, 110, 2341, 2634, 1859, 63, 512, 1892, 2635, 2669, 912, 913, 897, 2657, 1807, 2679, 2789, 778, 779, 2533, 1739, 1802, 768, 2711, 359, 2629, 360, 503, 130, 2398, 1192, 1193, 1493, 721, 938, 939, 937, 2658, 1595]. Progress [2298]. Project [1482, 2099, 1003, 14, 1236, 1070, 2014, 1235, 1280, 682, 2015, 1281, 8]. Project-Based [1236, 1235]. Projecting [894]. projection [2525, 2420, 38]. Projects [1278, 1283, 343]. PROLOG [239, 59, 2584, 1939, 205, 139, 671, 322, 2238, 814, 138, 997, 94, 227, 1815, 199, 1238, 1485, 695, 376, 2536, 65, 354, 668, 708, 47, 998]. Prominent [1614]. PROMIS [1189]. Promise [1848]. Promotes [1183]. Promotion [1978, 1705, 1977, 2063, 1979]. Proof [2369, 2766, 2379, 2798, 2575, 2790]. Proof-carrying [2766]. Proofs [890, 2518]. Propagation [949, 958, 1196, 1197, 736, 1201]. Proper [1183, 1994]. Properties [374, 480, 2812, 2826]. Proposal [889, 1059, 50, 79, 1597, 489]. Proposals [583]. Protection [2072, 587, 412]. Protein [1795]. Protocol [1303, 1311, 2336, 450, 1312, 2335, 1900, 180]. Protocols [1735, 1443, 1736, 1060, 1113, 548]. Prototype [2452, 583, 797, 582, 1599, 1054, 1281, 931, 283, 1280]. Prototype-Based [583, 1054, 1281, 1280]. Prototype-Instance [582]. Prototypes [1554]. Prototyping [2450, 502, 2661, 484]. provable [1469]. Provably [1631, 1632, 2749]. Provably-Secure [1631, 1632]. Provide [1240, 1239]. Providing [568]. Proving [2757, 1658, 2658]. Proxy [484]. pSather [637]. PSI [60]. PST [88]. psyche [62]. Pure [2732, 392, 878, 879, 1000]. purely [952, 479, 1803]. Purpose [1374, 1432, 1375, 849, 850, 2351, 851]. PUT [1105]. PUT/GET [1105]. Putting [2360, 2061, 2806, 2728]. Pygmy [2507].

QED [381]. QoS [2333, 2334]. Quadratic [2019, 1918, 1937, 2018]. Qualifiers [2287]. Quality [1990, 208, 321, 2228, 2389]. Quantification [813, 2662, 2723, 2602, 2703]. Quantitative [1522, 1333]. Quasi [2829, 2647, 2543]. Quasi-linear [2829]. Quasi-static [2647, 2543]. Queens [1031]. Queries [1177, 182]. Query [2159, 1859, 1860, 1546, 1545]. Query-Based [1860, 1859]. quest [448]. Question [2087, 2088]. queues [1764, 1765]. quick [486]. QuickBASIC [457]. Quickly [1431, 2672]. Quicksort [936].

R [381, 440, 2635]. Race [328, 327, 912, 913]. racehorses [208]. races [328]. Radioactive [1683]. Raising [2162]. Rambus [2119]. RAMpage [2119]. Random [2021, 464, 2020]. Range [990, 758, 759, 1220, 989, 1201]. rank [2839, 2370, 2376]. Rapid [2450, 1129, 484]. rate [1096]. Rationale [440]. Rationalized [1162]. Raw [2114]. RC5 [2464]. rdp [2063]. Re [1328, 1329, 2337, 1646, 836]. Re-Engineering [2337, 1646]. Re-usable [1328, 1329]. re-use [836]. Reachability [1896, 2412, 2693]. reachability-based [2412]. Reachable [427]. Reactive [2009, 1088, 2614, 1785, 2757]. Read [109, 270]. readable [227, 199]. reader [324]. reader-writer [324]. reading [2792, 2389]. Ready [26, 1577, 1576]. Real [1071, 482, 2176, 2881, 1845, 2332, 1005, 63, 1814, 1966, 1874, 751, 752, 2336, 11, 2337, 593, 2333, 588, 1678, 35, 2861, 463, 1341, 1120, 1072, 1333, 1330, 2447, 190, 2487, 70, 1339, 2489, 2575, 452, 744, 745, 189, 1334, 1337, 77, 2335, 2240, 1288, 2493, 2599, 2495, 1331, 1092, 1093, 1342, 1343, 2241, 1332, 915, 2334, 284, 1338, 98]. Real-Time [2881, 1845, 2332, 1005, 1874, 2336, 11, 2337, 593, 588, 35, 63, 1814, 1966, 751, 752, 2333, 463, 1341, 1120, 1333, 1330, 2447, 70, 1339, 2489, 2575, 744, 745, 189, 1334, 1337, 77, 2335, 2240, 2493, 2599, 2495, 1331, 1092, 1093, 1342, 1343, 2241, 1332, 2334, 1338, 98]. Real-Time-and-Distributed [1071, 1072]. realisation [1762]. Realistic [1440, 764, 345]. Reality [1585, 2316, 1584]. reals [110]. realtime [454]. Rearranging [247]. Reasoning [2696, 2193, 2097, 181, 1294, 313, 2623]. REC [959, 141]. Receiver [1286, 1285]. reclamation [2266]. recognition [128, 204]. Recognizing [2596, 434]. Recollections [683]. recompilation [2644]. Reconciling [1065, 861]. Reconfigurable [1026, 1027, 2502]. reconfiguration [388, 325]. Reconsidered [1058, 1187, 1194, 1195]. reconstruction [2571]. record [2860, 2863, 2870, 2877, 2885, 2888, 2892, 2646, 2556, 2589, 107, 108, 2590, 38]. Recorder [436]. Recovering [2031]. recovery [728]. recurrence [128]. recurrent [1247]. Recursion [1994, 1345, 2368, 2371, 2426]. Recursive [625, 2275, 1898, 1019, 2206, 1535, 2151, 1364, 2554, 2819, 2603, 1897, 490, 516, 1456, 2063, 2642, 1534, 1807, 2391, 2632, 84]. Recycling [2182]. Reduce [1864, 1863, 2750, 1257, 1113]. Reduced [2326, 1422, 217]. reducible [2673]. Reducing [1075, 1109, 2682, 2403, 540, 1539, 1259, 247, 2062, 2789, 434]. Reduction [2284, 1151, 2733, 2819, 88, 225, 2577, 2515, 1777, 965, 1216, 2422, 966]. reductions [87, 978, 136]. Redundancies [1609]. Redundancy [980, 1701, 1978, 2153, 2315, 1219, 1979]. Redundant [982, 1976, 538, 1220, 319, 380]. Reengineering [417, 1382]. Reexecution [995]. Refactoring [1565]. Referees [1265]. Reference [1029, 1720, 368, 2131, 1058, 2111, 713, 712, 733, 535, 2234, 458]. References [2461, 1780, 2626, 368]. Refill [2136]. Refinement [307, 2573, 2424, 2635]. Reflection [1142, 2145, 345, 1141, 2418, 1450, 929]. Reflections [2145, 2217, 2256, 1083]. Reflective [1300, 575, 1305, 2086, 600, 1304, 2481]. reflexive [2754]. Refutation [442]. Region [1702, 1212, 2738]. region-based [1212]. regional [1476]. Regions [2006, 981, 792, 2665, 2247]. Register [256, 1207, 299, 1978, 1979, 1705, 1704, 259, 991, 755, 519, 102, 1977, 1483, 1990, 2638, 104, 2741, 67, 758, 759, 2743, 1703, 103, 2610, 520]. Regular [826, 724, 2423, 467]. Reid [466]. Related [1056, 1044]. Relating [2538, 2370]. Relation [264]. Relational [1830, 28, 2791, 2622, 1853, 2549, 1829, 900]. Relations [630, 1699, 128, 2620, 2540]. relationship [2525]. relaxation [2731]. Relaxed [1749]. Relaxing [510, 511]. Release [1489, 1319, 1320, 814]. Release-to-Release [1319, 1320]. Relevant [2837]. reliable [537]. Reliance [349]. Relocating [1442]. Remarks [1131, 275]. Rematerialization [521]. remembered [2229]. Remembering [1355, 1356]. Remote [1671, 2093, 2481, 1631, 2356, 1632, 1672]. Removal [319, 1976, 2635]. Removing [2388]. Remy [38]. Rendering [1755]. Renvoise [1609]. Reordering [1989, 2635, 131, 509]. Reorganisation [2178]. Reorganization [2290, 169, 297, 298]. Reorganizing [1557, 1556]. Repair [1227]. Repetition [2113]. replace [218]. Replay [438, 422, 1445, 897]. Replication [975, 522, 1200, 751, 752]. replicator [118]. Reply [622, 2443]. Report [2059, 1233, 1488, 1376, 1929, 1719, 2298, 1866, 1497, 1879, 1935, 1950, 2148, 2219, 2257, 1477, 2435, 479, 1232, 2060, 1869, 2305, 1870, 1875, 620, 1500, 815, 1598, 1810, 2311, 633, 1605, 197, 2500, 1121, 2677, 1081, 1145, 276, 795, 2097, 619, 215, 2681, 1377, 581]. Reporting [6]. Representation [2269, 643, 125, 1156, 1157, 1165, 1561, 812, 782, 2738, 1759, 2678, 222, 1414, 1415, 1124, 2802, 811, 1767, 2674, 1010]. Representations [2876, 1158, 740, 741, 1467, 132, 1164]. Representing [1430, 2686, 2840, 105, 2378, 1238, 1080]. Reproduce [1884]. reproducing [328]. Request [2128, 2197]. Requirement [624]. Requirements [1156, 1608, 1573, 1523, 1070, 1621, 858, 2408, 2390, 1622]. Research [2251, 819, 1615, 1624, 888, 1842, 887, 235, 1123]. residual [81]. resilient [2810]. Resolution [561]. Resource [1156, 2068, 1907, 1908, 1096, 1357, 1358]. Resource-Bounded [1908, 1907]. Resources [1604, 2441, 1988, 2260, 314, 1786]. Respect [803]. Response [2224]. Responsiveness [1065]. restores [1207]. Restoring [433]. Restructuring [1751, 1565, 559]. Results [813, 2369, 943, 2798]. resyntaxed [1597]. Retargetable [1436, 1165, 303, 1111, 386, 496]. Retargeting [1680]. Rete [443]. Retiming [1748]. retrieval [395]. Retrieving [2198, 2199]. Return [2314, 1895, 1894, 2313]. Returned [706]. Reusability [1824]. Reusable [801, 591, 1567, 164, 2664]. Reuse [1550, 2276, 1547, 2106, 563, 564, 1567, 1825, 1553, 2107, 411, 1301, 406, 838]. reusing [179]. revenge [2374]. Reverse [2268, 1680, 761, 762, 2393]. Review [1953, 662, 1603, 1638, 1167, 1670, 466, 1168, 1882, 381, 439, 1166, 1226, 607, 440, 274, 481, 606, 525, 465, 451, 480, 885, 1227, 1169, 705, 1231, 1229, 1230, 157, 1170, 1228, 450, 796, 1171, 1669, 938, 939]. Reviewers [1543, 2138]. Revised [2060, 222]. Revised5 [2059]. Revisited [8, 175, 1452]. Revisiting [2747]. revival [2684]. Revolution [1868]. rewards [220]. Rewriting [1366, 2161, 1642, 2567, 932, 1641, 930]. RIGAL [206]. Right [1919]. rings [2696]. Rio [1520]. RISC [272, 553, 273, 2282, 2535, 2508]. RISCs [256]. Rivieres [450]. RMI [2455, 2456]. Robert [1167, 1229]. Robot [1669]. Robotics [1715, 1670, 2200, 1669]. Rochester [1497, 1498]. Role [1071, 576, 2410, 1712, 2077, 32, 1574, 1072, 1657, 1455]. Roles [876]. Rolling [2762]. ROOM [2447]. root [2492, 2765]. root-stable [2765]. Routine [711]. Routine-Texts [711]. routines [92]. Rover [1342, 1343]. Row [1289]. Roy [1230]. RPC [1900]. RSets [1486]. RT [452]. RT-MODULA2 [452]. RTsynchronizer [1331]. Rule [1346, 611]. Rule-based [1346]. Rules [18, 1369, 1828, 1297, 1943, 229, 1298, 376, 2790]. Run [1915, 2085, 2290, 1581, 2296, 996, 1434, 1906, 2295, 1998, 1594, 1098, 2736, 1580, 1905, 2297, 2243, 1216, 2542, 840]. Run-Time [1915, 1581, 1531, 996, 1434, 1906, 2295, 2085, 2296, 1998, 1594, 1098, 2736, 1580, 1905, 2297, 2243, 1216, 2542, 2350, 840]. Running [995, 2246, 218]. runs [543]. Runtime [422, 2123, 2160, 1262, 839, 774]. runtimes [1339]. Russia [1376, 1377]. Rüthing [722].

S [481, 348]. S-expressions [348]. Safe [2074, 2084, 828, 2631, 2542, 799, 2846, 515, 2795, 341, 2531]. Safety [1429, 1781, 1493]. Saint [1376, 1377]. Salesman [2049]. San [2853, 2862, 2873, 2877, 2892, 2901, 1612, 2858, 2868]. Sanford [274]. Save [2254, 2229, 2253]. saves [1207]. Scalability [1834, 1833]. Scalable [1915, 2005, 265, 2473, 324, 1832, 738, 907, 1831, 557, 2607]. Scalar [1424, 1977]. Scale [2469, 1516, 1756, 34, 343, 838]. Scaling [1900]. scan [1990]. Scanners [2307, 2306]. Scanning [318, 467, 2229]. scans [978]. Scenario [1051]. Scenario-Based [1051]. Scene [2211, 2054]. schedulable [744, 745]. schedular [771]. Schedule [2112]. Schedule-Independent [2112]. scheduler [772]. Schedulers [1692]. schedules [314]. Scheduling [1275, 1208, 256, 1087, 2133, 2332, 1325, 2114, 259, 2535, 1903, 1519, 812, 660, 2329, 728, 544, 1341, 304, 303, 1422, 1902, 300, 301, 756, 757, 302, 760, 143, 1209, 512, 555, 2669, 755, 305, 306, 811, 1263]. Schefstrom [1166, 1226]. Schema [1659, 169]. Schemata [1939]. Scheme [265, 266, 60, 1253, 2197, 791, 741, 2059, 943, 709, 2060, 2204, 1463, 2444, 369, 219, 1178]. SCHEMEXEROX [740]. School [807, 808]. Schreiber [1229]. Science [466, 477]. Sciences [381]. Scientific [640, 1653, 2474, 2437, 2449, 114, 2509, 2436, 1063]. Scientists [1614]. SCM [1501]. SCONE [1315]. Scope [1030, 2713, 2842]. scoped [617]. scopes [2647]. scoping [2647]. Screeners [2307, 2306]. SCSI [2116]. Seamless [1608, 2457]. Search [2763, 1953, 1882, 1881, 574, 573, 60, 2718, 1031]. Searching [237]. Seat [1397, 1394, 1392]. seatbelts [1396, 1395, 1393]. Seattle [2854]. Second [2854, 2869, 1696, 2817, 2420, 2858, 1968, 1969, 815]. Second-Order [1696, 2817]. secrecy [2824]. Section [1969, 1970]. Sections [428]. Secure [2336, 2823, 1631, 2335, 1632]. Securing [1867]. Security [861, 2826, 860, 858, 856, 857]. Segregating [2111]. SEIS [2446]. Seismic [2446]. SEL [872, 873]. Selected [1968, 1969, 1970]. Selecting [439]. selection [1467, 1221, 2845, 475, 2355]. Selective [1204, 2004, 2685, 2653]. Selector [852]. Self [1504, 1890, 2081, 1891, 2195, 1051, 1058, 1402, 379, 927, 928, 608, 137, 1065, 1279]. Self-4.0 [1279]. Self-Applicable [1891, 1890]. self-application [379].

Self-Guided [1051]. Self-Indexed [1504, 1402]. Self-Inflicted [2081]. self-modifying [608]. Self-Reference [1058]. Self-Similar [2195]. self-similarity [927, 928]. Semantic [1582, 1583, 719, 346, 18, 2394, 1295, 1079, 2585, 591, 123, 1802, 2574, 409, 289]. Semantic-Based [1583, 1582]. Semantics [2864, 2871, 2900, 625, 1179, 2320, 441, 2187, 2185, 2443, 1486, 2615, 1495, 2273, 376, 1628, 1293, 1630, 2902, 2586, 2698, 2546, 2583, 2707, 2606, 89, 1885, 2272, 313, 2619, 78, 1463, 373, 2783, 1496, 230, 488, 369, 2422, 2787]. Semantics-Based [2864, 2871, 2900, 2902]. Semi [975, 1060, 162]. Semi-Automatic [1060]. semi-dataflow [162]. Semi-static [975]. Sense [1854, 2487]. Sensing [1671, 1672]. Sensitive [1682, 1444, 1693, 987, 2814, 1196, 1197, 2627, 1694, 988, 756, 757, 1192, 1193]. Sensitivity [1895, 1894, 1797]. Sentinel [2025, 555]. Separate [968, 1886, 1287, 2659]. Separating [2617, 1086]. Separation [2711]. September [2898]. Sequence [1444, 1252, 1033]. Sequences [805, 806, 901, 531, 1255]. Sequential [2385, 2690, 2457, 329, 2697, 2715, 2689, 2711, 391]. sequentiality [2604]. Series [440]. Server [1577, 1073, 2246, 1576]. Servers [2128, 1541, 1087]. Service [1845, 1558]. Session [918, 920, 921, 1302, 923, 924, 925, 1301, 1283]. Sessions [2208]. Set [1958, 1717, 271, 1959, 1697, 1896, 1716, 2786, 1698, 1774, 1773, 217]. Set-Based [1697, 2786, 1698, 1774, 1773]. SetLog [230]. Sets [1995, 1177, 2041, 2229, 783, 1801, 44, 96]. Seven [873, 872]. seventeenth [2853]. Seventh [2884]. Several [2136, 1943]. Sh [2508]. Sh-BOOM [2508]. Shallow [350]. Shape [2760, 1085, 2724, 2725, 2835]. shape-analysis [2725]. share [2587].

Shared [1737, 1708, 1738, 599, 1531, 1533, 586, 637, 267, 1056, 2095, 1549, 1751, 1755, 2472, 658, 1733, 1514, 2137, 1445, 1530, 650, 2126, 1753, 1754, 59, 72, 1091, 1215, 2094, 1089, 557, 1259, 320, 74, 769, 770, 326, 324, 912, 913, 897, 792, 1114, 1115, 2350]. Shared-Memory [637, 658, 1514, 2137, 1445, 2126, 1754, 1753, 1091, 1215, 320, 769, 770, 324, 912, 913, 897]. Sharing [1737, 2163, 1472, 2126, 2121, 2646, 357, 2660, 1259, 2647, 2424, 2636]. Sharlit [501]. Shasha [1603, 1602]. Shasta [1530]. Shear [1755]. Shear-Warp [1755]. shell [229]. Shibin [818, 886]. Shoot [829]. Short [2383, 223]. Shortcut [1167]. Shot [1430]. Should [459, 1269, 2339, 4]. Showdown [1421, 2807]. Shrink [1858, 1857]. shuffling [1207]. Shuttle [2509]. sickness [463]. Side [1018, 1887, 2627, 732, 2689]. side-effects [1887]. Sifting [1076]. SIG [1712]. SIGACT [2860, 2863, 2870, 2877, 2885, 2888, 2892, 2901, 2302]. SIGADA [1946, 2434, 2435]. SIGForth [2858]. sigma [1451, 2655]. Signal [2465, 1536, 1210, 1211]. Signature [1010]. Signatures [2381, 2702, 2729, 2592]. Significant [715]. SIGPLAN [2865, 2874, 2886, 2887, 1378, 2142, 2252, 2222, 1874, 1361, 1359, 2448, 2852, 2854, 2855, 2857, 2860, 2862, 2865, 2863, 2866, 2864, 2867, 2868, 2869, 2870, 2871, 2876, 2877, 2885, 2888, 2889, 2892, 2893, 2894, 2895, 2896, 2897, 2898, 2899, 2900, 2901, 2902, 1379, 2304, 2302, 2010, 1023, 2208, 2221, 2880, 2881, 1164, 2300, 1944, 1390, 1391, 1362, 1810, 1635, 1664, 1650, 1649, 1648, 2050, 2140, 2883, 2248, 2249]. SIGPLAN-SIGACT [2860, 2863, 2870, 2877, 2885, 2888, 2892, 2901, 2302]. SIGPLAN-SIGSOFT [2142]. SIGPLAN/SIGSOFT [2894]. SIGSOFT [2142, 2894, 2140]. silicon [2502]. SIMD [510, 511]. Similar [273, 2195]. similarity [927, 928]. Simple [1429, 39, 1157, 2369, 2154, 2448, 1570, 1205, 2153, 523, 1675, 1437, 1802, 1113, 644, 2683, 823, 2832, 623, 2761]. Simplicity [1953, 1882, 1881]. Simplification [1982]. Simplifying [1461]. Simson [885]. Simulation [2460, 2135, 2482, 916, 342, 534, 697, 866, 676, 1099, 1100]. Simulations [2453, 2354, 2343, 64]. Single [1986, 2821, 1155, 2288, 1513, 550, 1512, 2629]. Single-Chip [1513, 1512]. SISAL [66]. Situated [1559]. Situation [1139, 1138]. Situation-Oriented [1139, 1138]. Six [631, 1466]. Sixteen [1614]. Sixth [2873]. Size [291, 1221, 1213]. sized [2425, 2757]. skeletons [1243]. SLAM [1720, 2824]. Slicing [994, 1688, 2030, 1560, 124, 2721, 289]. SLPX [449]. Small [2527, 1837, 802, 98]. Smalltalk [1078, 847, 696, 560, 820, 396, 2524, 1861, 673, 11, 186, 1057, 146, 24, 585]. Smalltalk-80 [560, 146]. Smart [459]. Smartest [2644]. SmartFiles [1322]. smartness [1493]. SML [2373]. SNOOPS [388]. Snowbird [276]. Soft [2663, 308, 1564, 1563]. SoftFLASH [1533]. Software [2894, 532, 595, 1737, 867, 2142, 1532, 10, 249, 1244, 2031, 1564, 276, 1846, 2851, 1916, 1645, 1531, 804, 1643, 801, 2106, 2107, 1014, 16, 1021, 1354, 381, 1311, 144, 2338, 1090, 18, 1189, 1384, 241, 621, 1621, 25, 1752, 1733, 2119, 54, 2075, 465, 1227, 2033, 539, 1421, 2340, 1530, 868, 1184, 31, 1332, 33, 1842, 2092, 2483, 1060, 1171, 597, 605, 1208, 404, 1328, 1329, 1563, 1847, 221, 2414, 164, 411, 1301, 1353, 557, 1312, 2730, 302, 384, 406, 1335, 1124, 1622, 2076, 1340, 2610, 519, 792, 1011, 838, 218, 1097, 2389, 1338, 275, 622]. software-based [1340]. Software-Only [1530]. Solution [2086, 932, 1271]. Solutions [977, 1032, 1073]. solved [1152, 2231]. solver [413]. Solving [2725]. SOM [1319, 1083, 1082, 1320]. Some [1609, 1483, 240, 943, 2417, 946, 947]. somewhere [2506]. Son [2374]. Sons [662, 607, 274, 481, 705, 796]. Sort [1504, 1402, 2643, 898, 899]. SOS [2790]. Sound [1291, 426, 2422, 2508]. Source [1424, 2031, 432, 1750, 1351, 1129]. Source-Level [1424, 1129]. Source-To-Source [1351]. South [2863, 1719]. Space [2279, 748, 1994, 2326, 2017, 1734, 1517, 2368, 1755, 253, 2114, 1752, 1730, 1731, 2357, 2402, 1469, 205, 2509, 2747, 2016, 1518, 550, 1452, 2719, 2509]. Space-Efficient [1731, 1730, 1452]. Space-Time [2114, 2357]. spaces [2242]. Spaniel [2223, 2155]. SPARC [271, 248]. Sparse [1490, 1591, 1491, 1592, 1978, 1158, 1159, 852, 2550, 1979]. spatial [416]. speaker [919]. SPEC [271]. Special [2886, 1374, 1375, 699, 1969, 2439, 1970, 2858, 2438, 168, 2190]. Special-Purpose [1374, 1375]. Specialisation [1693, 1694, 1779]. Specialization [1904, 2385, 1180, 1441, 1906, 2295, 1855, 1856, 2736, 1204, 1782, 1783, 1905, 370, 845, 846, 2522, 380, 2427]. specializing [2789]. Specific [425, 72, 395, 2742, 830]. Specification [719, 805, 1148, 1400, 1005, 2066, 2714, 165, 2698, 806, 2514, 602, 449, 1126, 444]. Specifications [560, 1036, 827, 1351, 312, 2526, 952, 99, 158, 2730, 931, 2321, 1022, 172, 1331, 91]. specify [1149]. Specifying [2079, 2618, 178, 313, 2391, 1011]. SpecInt95 [2339]. Spectra [2036]. Spectrum [1554]. Speculation [2133, 2115, 2749]. speculative [1216, 539, 617]. Speed [2193, 1990, 544, 330, 1170, 1228]. spell [2527]. spent [1091]. SPICE [2479]. Spill [1702]. Spilling [1702]. Spineless [2174]. SPLAW [1922]. Split [1557, 1555, 1556]. Splitting [990, 1770, 1769, 2403, 115, 989]. SPMD [659, 768]. Sponsored [2886]. Sponsorship [1649, 1648]. Spots [623]. Spreadsheets [819, 482]. Springer [885].

Springer-Verlag [885]. SQL [1323, 1382]. SQL-92 [1323]. Square [2364, 2492]. Squares [2019, 2018]. Squeak [1861]. SSA [1964, 1701, 730, 2805, 1977]. SSBUV [2509]. SSets [1486]. St [2885]. stable [2765]. Stack [1993, 1225, 2665]. StackThreads [2346]. StackThreads/MP [2346]. Stage [1912, 1805, 1804, 1911, 2422]. staged [2745, 2297, 2421, 2818]. stages [2617]. Staging [364, 2681]. Standard [1601, 803, 1510, 1511, 214, 222, 2568, 50, 530, 1406, 968, 2171, 789, 1437, 1206, 431, 1405, 2319, 817]. standardization [2513]. Standardized [1883]. Standards [1064, 1665, 1307, 2346]. Stanford [1089, 1112]. Starting [1359, 1360]. State [1047, 2294, 969, 2079, 1348, 1277, 2121, 2801, 2697, 1333, 1781, 1276, 145]. Statechart [1275]. stateful [1780]. statement [531]. Statements [889, 994, 1615, 940]. States [433, 427]. Static [1594, 400, 2569, 2025, 1572, 1179, 568, 2546, 649, 2417, 2029, 2380, 1425, 2750, 1986, 1334, 1155, 2030, 2288, 2103, 429, 1921, 3, 2816, 1821, 2629, 1042, 974, 1453, 1212, 334, 2804, 1202, 1203, 2390, 401, 1885, 975, 2647, 1325, 355, 2102, 78, 2844, 1201, 485, 569, 2543, 86, 297, 298, 1108]. static-graph [297, 298]. static/dynamic [1325]. Statically [1570, 1766, 2631, 518]. Statically-Typed [1570, 2631]. Statistical [2291, 1249, 1250]. Status [633, 833, 834, 1081]. steal [2073]. stealing [1324]. stealthy [2810]. Steele [1229]. Steering [1043, 1721]. Steffen [722]. Step [1920, 885, 867, 2511, 868]. sticky [2233]. Still [1825, 795]. Stitt [1227]. Stochastic [2841]. stock [791]. stop [2788]. Storage [1142, 1078, 1213, 2281, 1454, 2112, 1141, 2118, 2266, 1094, 1095]. store [758, 759]. Stores [1978, 1159, 1979]. story [1861]. Straitjacket [669]. Strategic [1615]. Strategies [1853, 1175, 2161, 882, 240]. Strategy [1756, 2671, 1241]. Stratified [2634]. Stream [246, 950]. Streaming [257]. Streamlining [1070]. Streams [2183, 1242]. Strength [1151, 527]. stress [180]. strict [479, 2124, 2719, 2711]. Strictness [286, 287, 2549, 2666, 2812, 2527, 2548]. String [2226, 1033, 87, 1796, 2225]. Strong [568, 175, 570, 569]. Strongly [1759, 295, 296, 571]. Strongtalk [847]. Structural [866, 2703, 1208, 416, 346, 1456, 2672]. Structure [2270, 2271, 1898, 981, 2021, 1528, 1897, 99, 2729, 2020, 390, 118, 2792, 905, 1198, 2836, 2636]. structure-sharing [2636]. Structured [2198, 640, 719, 2698, 1973, 639, 451, 824, 1674, 660, 174, 2199, 167, 536, 1243, 638]. Structures [640, 1029, 1967, 1242, 985, 1535, 1628, 2120, 1015, 2278, 129, 516, 76, 2368, 74, 2562, 1534, 2802, 1455, 2391, 1821]. Structuring [1784, 2718]. Stub [1018]. Student [1477, 1476]. students [1593, 1119]. studies [872, 873]. Study [10, 1720, 1432, 1038, 2469, 1849, 2134, 2007, 1548, 597, 1254, 736, 2242, 1476, 2588, 2409, 406, 333, 461, 2400, 945]. Style [1610, 2363, 1155, 51, 2775, 227, 2191, 2617, 199, 2589, 2817]. Styles [1016, 2687]. Subclassing [1277, 1294, 1276]. Subject [2408, 878, 879, 1297, 1298]. Subject-Oriented [1297, 2408, 878, 879, 1298]. Subjectivity [19]. Subobjects [1293]. Subpages [1539]. Subprograms [664, 663]. Subroutine [719]. Subroutines [625, 2832, 2808]. subscript [1220]. subscripted [104]. substitution [176]. substitutions [2516, 2655, 1801]. substrate [499]. Substring [279]. substrings [2596]. subsystem [2650]. Subtransitive [1700]. subtype [1462, 116, 142]. Subtypes [1290, 2704, 181, 2601, 827]. Subtyping [2554, 1012, 2179, 2180, 1187, 177, 2603, 2523, 1771, 2755, 2705, 2642, 2600, 2633, 1772, 2568, 1461, 2779]. successes [1600].

Successful [567]. Sufficient [1240, 1239]. SUIF [2344, 1123]. suitable [1133]. suite [205, 407]. Summary [718]. Summing [1364]. Sun [1900]. super [2652]. Superclass [1551]. Supercomputing [2143]. superoperators [2716]. superoptimizer [524]. superpages [1103]. Superscalar [970, 248, 250, 973, 304, 555, 243, 2330, 556]. Superscalar-Based [973]. Support [2873, 2884, 1188, 1443, 1708, 1725, 260, 1932, 2090, 1056, 2115, 1516, 1240, 1733, 2280, 1707, 1541, 2248, 1120, 654, 839, 1117, 2244, 518, 1931, 2595, 187, 1105, 550, 326, 1106, 2416, 912, 913, 1340, 1248, 1331, 539, 162, 1103, 2350, 1706, 1097, 1239, 2249]. Supporting [125, 1554, 1336, 74, 1530, 645, 1016, 2486, 388]. Supports [1944]. Suppression [1609]. Surpassing [1103]. Survey [820, 1417, 47, 196, 143, 1418]. Surviving [1520]. Swaps [580]. SwapsManager [579]. sweep [2244, 2580, 2245, 2235]. switch [544]. Switches [252]. Switching [11]. Symbol [2317]. Symbolic [416, 2430, 2638, 786]. symmetric [2556]. Symposium [2854, 2857, 2860, 2863, 2866, 2864, 2868, 2870, 2877, 2885, 2888, 2890, 2892, 2897, 2302, 276, 1866, 2880, 815, 2853, 2856, 12]. Synapse [98]. Synchronization [2001, 2772, 269, 1740, 1515, 2399, 2388, 57, 2750, 2245, 771, 772, 1214, 1088, 2773, 836, 324, 64, 1739, 549, 1256, 1464, 1261, 787, 788]. Synchronous [1161, 1470, 2183, 1210, 1211, 2777, 1505, 1808, 1806]. Synopsis [644]. Syntactic [2381, 2379, 363, 2842]. Syntax [441, 2317, 1943, 147, 462, 743, 476]. Syntax-Directed [2317]. Synthesis [1014, 2338, 1800, 193, 1253, 1807, 2664, 1332]. System [255, 49, 596, 263, 599, 1520, 422, 1554, 1531, 1436, 2096, 436, 2090, 1056, 1021, 439, 1315, 1400, 1819, 2338, 1005, 1607, 2048, 1035, 2103, 582, 2093, 1903, 1684, 11, 1279, 983, 2260, 1541, 2092, 335, 2227, 600, 1642, 1098, 2066, 1262, 303, 536, 322, 1596, 342, 534, 1111, 1902, 401, 2524, 158, 332, 121, 835, 2661, 326, 1106, 1797, 2102, 229, 1772, 2493, 2803, 2832, 2720, 623, 233, 1342, 1343, 1786, 2808, 182, 1103, 443, 549, 2259, 2498, 217, 2499, 1338, 1094, 1095, 410, 1641, 1820, 272, 469]. System-wide [1279]. System/6000 [272]. Systematic [1762, 1301, 716, 221, 317]. Systems [2875, 2873, 2878, 2884, 2896, 2899, 1915, 36, 1191, 1071, 2459, 603, 1974, 1432, 1581, 2339, 5, 2881, 423, 1053, 19, 30, 21, 18, 425, 2332, 2105, 1752, 1267, 9, 606, 480, 2317, 1624, 1514, 2336, 2331, 2137, 1079, 2337, 860, 660, 250, 2333, 31, 588, 33, 1052, 35, 2882, 1832, 1749, 537, 165, 1120, 1190, 1072, 1333, 1330, 864, 2656, 1580, 2374, 1831, 1339, 184, 2489, 2490, 2666, 374, 178, 2575, 793, 2757, 417, 310, 1334, 2370, 2352, 1593, 550, 1337, 2104, 1325, 2335, 1266, 2567, 866, 2599, 841, 842, 865, 172, 1331, 2834, 1086, 1332, 915, 900, 2334]. systems [2427, 2849, 297, 298, 347, 611, 1928, 2210, 2211, 1061, 1119, 1874, 1190]. systolic [130].

T [606, 451]. T3E [1515]. Table

[2136, 2333, 2334, 852, 523]. Table-Based [2333, 2334]. Tables [1066, 1289, 612, 1837, 1067]. tabling [2238]. Tag [295, 296, 1540]. Tag-free [295, 296]. Tagless [2174]. Tail [1994, 1364]. Tail-Recursive [1364]. Tailorable [1053]. Tailoring [1292, 15]. TAKE [976]. Takeuchi [1169]. Talk [888, 2417, 887, 2421, 2427]. Talking [1965]. Taming [2164]. Tapeworm [1099, 1100]. Target [1444]. Target-Sensitive [1444]. targeting [1116]. Task [117, 642, 812, 1728, 71, 811, 765]. task-level [71]. Tasking [1347]. tasks [1325, 1255]. Taste [817]. taxation [2674]. Taxonomy [583, 2553, 1045]. TB [339]. TC [1916, 1605]. TC-2 [1916, 1605]. tcc [1684]. Teach [4].

Teaching [1482, 1479, 1480, 1940, 1600, 964]. Teams [1085]. Teapot [1443]. Tears [2373]. Technical [1399, 1411, 1925, 2325, 2312, 1656, 1954, 2058, 2150, 2220, 2263, 1412, 1916, 2265, 2443, 2314, 634, 2151, 1649, 2226, 1380, 1101]. Technique [2369, 982, 1681, 2123, 2379, 1098, 483, 2683, 823, 302, 1116, 172, 1821, 96, 790]. Techniques [1284, 1372, 1388, 2067, 970, 1893, 361, 1227, 430, 2289, 404, 840, 1373, 1389, 1669, 657, 1892, 2230, 218, 58, 2389]. TECHNOLOGIES [1376, 2468, 1377, 872, 873]. Technology [1929, 36, 1589, 276, 1050, 580, 30, 37, 1278, 2261, 1665, 32, 867, 579, 419, 874, 875, 2262, 1930, 1600, 868, 856, 857, 882]. Telecommunication [11, 2340]. Template [626, 627, 1405, 2319, 1406]. templates [775]. Temporal [2193, 2575, 480, 2289]. temporary [963]. ten [2890, 292, 293]. Tenth [2882]. Term [2202, 2567, 2787]. Terminable [940]. Terminal [2219, 2488, 2218]. Terminals [2340]. termination [159]. Terminology [1646]. TErms [2677]. Terse [1606]. Test [985, 506, 1216]. tested [2394]. Testing [24, 2639, 911, 311, 281, 1288, 180, 1011]. Tests [1838]. Texas [2878, 2858]. textbook [150]. Texts [711]. textual [123]. Their [264, 583, 1602, 1603, 827, 1822, 1823, 926]. Themselves [1884]. theorem [2529]. Theoretic [1048, 2660, 1049]. theoretical [203]. Theoreticians [603]. Theory [1047, 2287, 621, 1748, 2166, 1136, 2827, 1468, 153, 241, 1471, 171, 2830, 2654, 2769, 2780, 2545, 2422, 275, 622, 161, 351, 2024]. there [375]. Thermodynamics [957]. Thin [2001]. Third [2857, 1376, 1065, 2858]. Thomas [1168]. Thoughts [1497, 1498, 2148, 343, 2149]. Thread [1519, 2121]. Threaded [260, 2030, 2457, 2004, 2823, 2350, 2345]. Threading [1142, 52, 1141]. Threads [632, 1517, 1852, 969, 188, 1518, 2711, 1453]. threat [532]. threatening [1224]. Three [1162, 1974, 1175]. Three-Tier [1974]. Tier [1974]. tightly [545]. tightly-coupled [545]. TIL [1438]. Tile [1221]. Tiling [1849, 2289, 1732, 2770]. Time [1915, 643, 2279, 1071, 2290, 1581, 1531, 2088, 1734, 2881, 1517, 1845, 1700, 996, 1843, 981, 1252, 2332, 1005, 1434, 2114, 1906, 2295, 429, 1874, 26, 2336, 11, 2337, 1895, 2719, 593, 588, 973, 35, 1004, 424, 2003, 1594, 1098, 2229, 463, 1341, 782, 2596, 1120, 1469, 1072, 97, 1333, 385, 2085, 1091, 2587, 1330, 2087, 2736, 2447, 2487, 1580, 70, 1205, 1339, 1905, 2296, 2297, 2749, 1518, 2489, 2575, 452, 2770, 744, 745, 1894, 1324, 366, 189, 1259, 1334, 2243, 1090, 1337, 77, 286, 287, 2335, 63, 1814, 1966, 2240, 2544, 912, 913, 2493, 751, 752, 1326, 1327]. time [2599, 2495, 2535, 1257, 2767, 305, 306, 2357, 1216, 1331, 2542, 1629, 1092, 1093, 1342, 1343, 2241, 2694, 971, 2726, 2333, 2402, 2350, 1332, 915, 2334, 1271, 2618, 1998, 1338, 98, 840]. time-critical [2535]. time-efficient [2749, 2402]. time-triggered [1337]. Time/Run [1531]. timed [300, 301, 2712]. timestamps [902]. Timethread [1071, 1072]. Timethread-Role [1071, 1072]. timing [189, 1334, 1336]. Tips [1479, 1357, 1405, 1924]. Titanium [2485]. Title [883, 1871, 2109]. TkGofer [1784]. TLB [2136, 1103]. TLB-Refill [2136]. TM [221]. TMS320C6x [2329]. Toaster [1965]. TOBEY [1160]. Together [1412, 1413]. Token [2147, 2146]. Too [2360]. Tool [1078, 2452, 1412, 1366, 2338, 1189, 2042, 1143, 1181, 206, 1120, 205, 1413, 57, 1333, 43, 2827, 767, 332, 1337, 1182, 2318, 230, 495, 501, 1338, 1166, 1226]. Toolbox [1036, 1932, 1479, 1917, 1877, 1494, 1357, 1405]. Toolkit [8, 1931, 415, 862, 863, 844, 1932]. Toolmaker [1633]. Tools [2894, 2896, 2142, 1958, 2339, 2881, 830, 1621, 1874, 1227, 983, 2446, 1959, 2514, 1339, 398, 1622, 915]. Top [244, 1789, 147, 490, 952]. Top-down [244, 147, 952]. Topics [1044]. topological [898, 899]. Topology [2131, 2657]. Torgerson [1168]. Toronto [2855]. Total [2708, 2538]. Tour [1604]. Towers [2252]. Trace [422, 2740]. Trace-and-Replay [422]. Trace-based [2740]. Tracing [995, 2365, 2581, 2266, 897, 2230]. Track [2303]. Tracking [1750]. Tractable [1274]. Trade [2119]. Trade-Offs [2119]. Tradeoffs [1737]. trailing [999]. Training [1050]. Traister [1167]. Trampolined [2363]. Transactions [18, 2366, 558, 409]. TransBase [339]. transfer [2403, 2124, 2429, 1086, 1107, 882]. Transform [2285]. Transformation [1617, 1903, 1626, 1136, 1330, 1902, 2684, 516, 2617, 376, 2708, 2530, 315]. Transformations [1688, 649, 2130, 635, 1625, 1980, 1258, 363, 87, 414, 385, 1215, 1759, 2772, 364, 510, 1259, 317, 378, 362, 1800, 509, 69, 511]. Transformers [1348, 2717]. Transforming [1898, 778, 779, 1897]. transforms [2773]. Transition [1400, 419]. translated [138]. Translating [1382, 2769]. Translation [1584, 2127, 1585, 1675, 92, 2375, 553, 547]. translations [2566]. translator [457, 2595, 1414, 1415]. Translators [2123, 1383, 2681]. Transparent [2349, 2381, 2702, 2701]. transparently [322]. Transport [1282]. Trap [1099, 1100, 2651]. trap-based [2651]. Trap-driven [1099, 1100]. Traveling [2049]. traversals [1775, 2391]. treadmill [463]. Treating [531, 145, 46, 1033]. treatment [2722]. Tree [2254, 1366, 981, 2178, 2680, 2253, 2594, 475, 2724, 2063, 2673]. Trees [1299, 1154, 100]. trellis [70, 571]. Trends [1354, 2299]. Tribute [2044]. triggered [1337]. Trip [276]. trivial [2422]. TRS [2360]. true [220]. Trust [2834]. Truth [564, 563]. Tse [451]. Tunable [603]. Tuning [6]. Tuples [2101]. Tupling [1775].

Turing [623]. Turtle [231]. Tutorial [2648, 1038]. Twentieth [2863]. Two [719, 263, 1967, 472, 1715, 800, 277, 151, 107, 108, 1942, 2054, 2055, 1695, 469, 123, 286, 287, 245]. Two-directional [107, 108]. Two-Level [719, 263, 800]. Two-valued [151]. TX [2901]. typability [2370, 2544]. Type [1284, 2002, 1761, 1760, 2781, 2383, 112, 113, 2593, 2382, 2187, 2744, 1987, 1779, 1047, 1291, 2287, 2096, 2377, 2846, 2163, 997, 996, 116, 1819, 1722, 1048, 1049, 2428, 2103, 2192, 2160, 666, 2641, 1624, 2376, 176, 2848, 1074, 2180, 1889, 341, 2371, 182, 1507, 1611, 1438, 998, 1838, 2375, 2227, 1136, 740, 741, 1826, 400, 2569, 2646, 1758, 72, 45, 941, 942, 828, 2631, 115, 802, 2393, 2827, 401, 2524, 2517, 2420, 2666, 2660, 2700, 2555, 2704, 2794, 2243, 2661, 2839, 2370, 1781, 2570, 2601, 2102, 492, 2795, 2832, 2728, 136, 402, 2720, 753, 2815, 2780, 2817]. type [1206, 1888, 2818, 2808, 2828, 1820, 800]. type-assignment [2593]. Type-Based [1987, 2375, 2428, 2848, 1758, 2393, 1206]. Type-dependent [112, 113]. Type-Directed [2160, 1889, 1438, 2744, 2420, 2376, 1888]. Type-driven [1761, 1760]. Type-Erasure [2187]. Type-Precision [2002]. Type-safe [2846, 341]. type-safe-definitely [2795]. type-specific [72]. Type-Theoretic [1048, 1049, 2660]. Typechecking [1046, 847]. Typed [2847, 1570, 2378, 2746, 2188, 2172, 2569, 2631, 115, 2683, 1759, 518, 2653, 295, 296, 848, 571, 2826, 2522, 2803, 365, 2623, 2825, 2665]. Types [2198, 2181, 2100, 2833, 2085, 2072, 2782, 1019, 2377, 2364, 2314, 808, 2151, 1230, 2000, 664, 2186, 807, 2756, 2663, 2554, 2199, 2768, 663, 2603, 226, 2762, 2425, 1793, 2393, 2760, 307, 470, 1460, 1459, 2757, 89, 2571, 160, 2839, 2624, 2829, 2659, 826, 2811, 2313, 2632, 926, 380, 2759, 336, 2392, 2843]. Typing [568, 2559, 2194, 2164, 845, 846, 3, 2590, 2780, 2663, 2646, 308, 2591, 175, 570, 2816, 2834, 2818, 569, 2543]. typings [2727, 2779]. Typographical [669].

ubiquitous [2399]. Ugly [1972]. UIMS [186]. UIST’90 [276]. UK [1969, 440, 1605]. Umbriel [1185]. UML [1846, 1847, 2390]. unattainable [209]. unbounded [1034]. Unboxed [2591]. Unboxing [2160]. uncaught [2848]. uncertain [760]. unconditional [522]. unCurrying [2796]. undecidable [2602]. Under-Appreciated [2184]. Understanding [1310, 2023, 2026, 492, 1338, 623]. Underused [1172, 1173]. Unfold [2184, 2185]. unfolding [374]. unfolding-based [374]. Unicode [934]. Unification [1367, 2710, 614, 2817]. Unified [1156, 2121, 2764, 2722, 317]. Uniform [640]. Uniformity [1279]. Unifying [1215, 451, 409]. union [2811]. Uniprep [934]. uniprocessors [554]. unique [2308, 2309, 2233]. Unitary [2202]. United [1968]. Units [2132, 2168, 2136, 2791, 1999]. Universal [209, 2486, 926, 441, 48]. universes [237]. University [1719, 439, 440, 606, 525, 465, 451, 122]. UNIX [1365, 443]. Unloading [1921]. unnecessary [2388]. unobtrusive [2656]. untyped [2422]. Update [2163, 1237]. updates [2547]. Updating [1029, 2725]. upon [2828]. USA [2874, 2872, 2875, 2873, 2869, 2896, 2899, 2900, 2902]. usable [1328, 1329]. Use [1013, 1141, 1142, 1555, 1412, 1573, 2088, 2501, 1548, 1413, 482, 2087, 733, 2707, 2805, 181, 827, 349, 1590, 836, 2608, 926, 1452, 1454, 96, 375]. Use-Once [1142, 1141]. Used [2339]. Useful [967, 1658, 922, 1821]. useless [2428, 2849]. User [254, 652, 276, 1176, 620, 1279, 1281, 1098, 919, 398, 1336, 1101, 1280, 1171]. user-friendly [1336]. Uses [2043]. Using [1995, 1066, 747, 1149, 1899, 1939, 832, 1075, 2455, 2459, 2232, 1974, 1432, 1032, 580, 426, 952, 2423, 1148, 1132, 1021, 1689, 1986, 394, 1476, 1315, 1539, 2729, 931, 1607, 771, 772, 1828, 1849, 9, 1305, 606, 430, 2691, 1740, 950, 380, 2477, 2135, 1058, 1280, 1281, 2226, 33, 1574, 2283, 1699, 992, 1067, 2354, 840, 1224, 2639, 1688, 916, 1207, 2403, 1221, 2326, 224, 914, 2650, 579, 1104, 2771, 877, 524, 972, 767, 415, 510, 2700, 551, 997, 2384, 2757, 1696, 566, 234, 2409, 371, 397, 2124, 2630, 1335, 1304, 362, 542, 2552, 506, 1801, 1739]. using [2759, 2349, 359, 786, 2225, 1447, 2353, 360, 2480, 2665, 2389, 998, 284, 2787, 100, 315, 1108, 511]. USSA [441]. Utah [276]. Utilization [271]. UTLB [2127].

v [2655, 2859]. Validation

[583, 804, 1143, 158, 349, 335]. Value [2133, 1527, 2674, 1887, 2814, 2676, 1218, 152, 1201, 1450, 2665, 46]. value-flow [2814]. value-passing [2676]. Valued [1172, 151, 2835, 1173]. Variability [1529]. Variable [2002, 2169, 2125, 246, 664, 663, 2428, 2647, 2849]. Variables [889, 1142, 2328, 2095, 1240, 341, 2638, 728, 128, 1141, 104, 367, 2094, 2606, 2564, 2622, 926, 508, 1239]. variation [722]. Varying [664, 663]. Vclusters [2089]. VCODE [1436]. VDM [1036, 1416, 931, 1274]. VDM-SL [1036, 530]. Vector [259, 1776, 149, 778, 779]. vector-based [1776]. Vegas [2889]. Venturing [1878]. Verbose [2194]. verifiable [99]. Verification [1163, 890, 2384, 1696, 2784, 2681]. verifier [2396]. Verifying [2697, 362, 480]. VeriSoft [2771]. Verlag [885]. Version [479, 534, 458, 347]. Versioned [586]. Versions [628, 123, 409]. Versus [634, 2807, 821, 2732, 1058, 82]. Verus [1333]. Very [1436, 2245]. VI [2873, 1543]. Via [1526, 553, 1702, 2738, 1249, 1250, 1109, 299, 540, 2029, 1529, 709, 932, 2693, 84, 2835, 335]. Vibrant [1643]. Vice [1371, 1647]. Vice-Chair [1647]. video [2488]. Vienna [652]. view [2573, 45, 919, 192, 340, 1384]. Viewing [166]. viewpoints [241, 275]. viewpoints-critical [275]. views [2416, 185]. VII [2884]. VIII [2138]. Virginia [2857]. Virtual [2002, 254, 1958, 1572, 1899, 1571, 1533, 1639, 1751, 1521, 1240, 2069, 1068, 2314, 2083, 1052, 2484, 1837, 2082, 1959, 830, 2394, 2351, 2313, 1640, 1239, 2831, 2116]. Virtuality [1725]. virtually [546]. VISA [2116]. Visibility [1369, 228]. visible [156]. VisiCLANG [156]. VisiCola [228, 1030]. Vision [1231, 1279, 71]. Visual [1720, 1581, 1580, 1168]. Visualization [2449, 1583, 1899, 2455, 2474, 1310, 2033, 1579, 1052, 1582, 922, 898, 899, 1578]. Visualizing [864, 865, 2035, 2092, 577, 184, 1862]. vizualization [1338]. VLIW [2329, 316, 970, 555, 246]. VM [2476]. void [1452]. volatile [537, 1094, 1095]. Volume [1755]. Volunteer [2477]. Vortex [1552]. VORX [61]. vs [1284, 1290, 1152, 1421, 2371]. vulnerability [558].

W [1168]. walk [231]. Warp [1755]. Washington [2854, 122]. Watchpoints [1140]. watermarking [2851]. wavefront [2353]. Way [1608, 507, 243, 623]. weak [1468, 2567]. Wearing [439]. Weather [66, 2197]. Web [1423, 2482, 125, 1884, 2409, 2444, 2483]. Web-based [2482, 2409]. Weeks [798]. well [2506]. We’re [1668, 2010, 1667]. WG21 [489]. Where [2010, 1091, 1290, 493, 1090]. Which [829, 391, 625, 80]. While [1840, 1839, 2789]. Whirlwind [1604]. White [2852]. Who [1657]. Whole [1517, 1518, 2293]. Whole-Program [1517, 1518]. Wide [2444, 2352, 1279]. Width [247]. Wiley [662, 607, 274, 481, 705, 796]. Williamsburg [2857]. window [558]. Windows [1167, 1170, 1228, 607]. WindView [1338]. Winter [481]. Wirth [1637, 1638]. Witches [1318]. Within [2367, 1555, 1451]. Without [2373, 582, 463, 962, 775, 1368, 2245, 1797, 135, 269, 2632, 2561, 999, 2674, 1837]. Woodger [440]. Words [1033]. Work [2360, 1056, 424, 2558, 2806, 2728]. Workbench [705]. Workflow [1078]. Working [2373, 525, 682, 1605]. Workloads [2137, 1101]. Workshop [2872, 2871, 2876, 2891, 2894, 2896, 2900, 1233, 1379, 2142, 1915, 1929, 2304, 1120, 2881, 1962, 1874, 1875, 620, 1599, 1499, 1500, 2861, 2902, 2500, 1164, 1232, 619, 1593, 1600, 1145, 1119]. Workshops [2300, 1870, 1648, 2858]. workstations [1465, 1116]. World [589, 1843, 1057, 1186, 2861, 590, 190, 2579, 1288, 2531, 2444]. Worlds [1899, 1282]. Worst [1326, 1327, 1324]. Worst-case [1326, 1327, 1324]. Wrap [1858, 1857]. WRAPGEN [1412, 1413]. wrapping [483]. Write [572, 489]. writer [324]. Writing [1443, 582, 206, 452]. written [1861]. www.cs.rutgers.edu [2252]. Wyoming [2506]. WYSIWYG [239].

X3J16/ISO [489]. XDP [782, 643].

XEROX [741]. XIL [1160]. XML [2375]. XSC [803].

YACC [1132]. YALE [2170, 2261, 2262]. Year [1870, 2827]. years [2890, 872, 873]. YIL [1160]. York [885, 705, 2852].

Zero [989, 990, 2327]. Zero-cost [989, 990]. ZF [138]. Zones [2414]. Zosel [1229]. Zuse [621].

References

Stroustrup:1995:WCJ

[1] B. Stroustrup. Why C++ is not just an object-oriented programming language. OOPS Messenger, 6(4):1–13, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Koenig:1995:ID

[2] A. Koenig. Idiomatic design. OOPS Messenger, 6(4):14–19, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Meyer:1995:ST

[3] B. Meyer. Static typing. OOPS Messenger, 6(4):20–29, October 1995. CO-DEN OOMEEO. ISSN 1055-6400.

Goldberg:1995:WSW

[4] A. Goldberg. What should we teach? OOPS Messenger, 6(4):30–45, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Fraser:1995:ADA

[5] Steven Fraser, Honna Segel, Jim Coplien, and Judith White. Application of domain analysis to objectoriented systems. OOPS Messenger, 6(4):46–49, October 1995. CODEN OOMEEO. ISSN 1055-

6400. URL http://www.acm.org/
pubs/citations/proceedings/oops/
260094/p46-fraser/.

Meadows:1995:BPT

[6] Debbie Meadows and Charles R. Gardner. Building and performance tuning a distributed reporting application for an OODBMS. OOPS Messenger, 6(4):50– 55, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Gardner:1995:POO

[7] Charles R. Gardner. Performance of an OODB in an online 7 × 24 × 365 manufacturing operation. OOPS Messenger, 6(4):56–60, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Strom:1995:NTP

[8] Stephen W. Strom, William F. Rich, and Matthew T. Verona. The Navigation Toolkit Project revisited. OOPS Messenger, 6(4):61–66, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Marney:1995:UOM

[9] Steven Marney and Mamdouh Ibrahim. Using objects to manage in-memory data intensive expert systems. OOPS Messenger, 6(4):67–71, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Burgett:1995:AOS

[10] Jeff L. Burgett and Sheryl J. Adam. Applying object-oriented software engineering methods to the development of call center software: Case study. OOPS Messenger, 6(4):72–76, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Radford:1995:IRE

[11] J. Radford. Implementing a real-time, embedded, telecommunication switching system in Smalltalk. OOPS Messenger, 6(4):77–82, October 1995. CO-DEN OOMEEO. ISSN 1055-6400.

Rosson:1995:ODS

[12] M. B. Rosson. OOPSLA ’95 Doctoral Symposium. OOPS Messenger, 6(4):83–84, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Fraser:1995:PCCa

[13] S. Fraser. Patterns: Cult to culture? OOPS Messenger, 6(4):85–88, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Hill:1995:OPM

[14] L. Hill. Object-oriented project management. OOPS Messenger, 6(4):89– 93, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Mancl:1995:TOA

[15] D. Mancl. Tailoring OO analysis and design methods. OOPS Messenger, 6(4):94–99, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Gangopadhyay:1995:FSD

[16] Dipayan Gangopadhyay, Wolfgang Pree, and Albert Schappert. Frameworkcentered software development. OOPS Messenger, 6(4):100–104, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Mili:1995:MO

[17] Hafedh Y. Mili, François Pachet, Ilham Benyahia, and Fred Eddy. Metamodeling in OO. OOPS Messenger, 6 (4):105–110, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Kilov:1995:SIC

[18] Haim Kilov, Bill Harvey, and Kevin Tyson. Semantic integration in complex systems: Collective behavior in business rules and software transactions. OOPS Messenger, 6(4):111–116, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Harrison:1995:SOS

[19] William Harrison, Harold Ossher, and Hafedh Mili. Subjectivity in objectoriented systems. OOPS Messenger, 6(4):117–122, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Cummins:1995:OEF

[20] Fred Cummins, Roman Cunis, and Gail Harris. The object engine: Foundation for next generation architectures. OOPS Messenger, 6(4):123–127, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Kafura:1995:DPC

[21] Dennis Kafura, Greg Lavender, and Doug Schmidt. Design patterns for concurrent, parallel, and distributed object-oriented systems. OOPS Messenger, 6(4):128–131, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Meszaros:1995:PDM

[22] G. Meszaros. Patterns for decision making in architectural design. OOPS Messenger, 6(4):132–137, October 1995. CODEN OOMEEO. ISSN 1055-6400.

deChampeaux:1995:OPM

[23] Dennis de Champeaux, Simon Horner, and Granville Miller. OO process and metrics for effort estimation. OOPS Messenger, 6(4):138–142, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Yates:1995:TSA

[24] B. Yates. Testing Smalltalk applications. OOPS Messenger, 6(4):143–148, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Lieberherr:1995:AAS

[25] K. Lieberherr. Adaptable and adaptive software. OOPS Messenger, 6 (4):149–154, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Narayanaswamy:1995:OCF

[26] K. Narayanaswamy and Adrian Blakey. Are object-oriented CASE frameworks ready for prime time? OOPS Messenger, 6(4):155–158, October 1995. CO-DEN OOMEEO. ISSN 1055-6400.

Zorn:1995:ODB

[27] Benjamin G. Zorn and Akmal B. Chaudhri. Object database behavior, benchmarks, and performance. OOPS Messenger, 6(4):159–163, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Chung:1995:ORD

[28] Jen-Yao Chung, Yi-Jing Lin, and Daniel T. Chang. Objects and relational databases. OOPS Messenger, 6 (4):164–169, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Sutherland:1995:BOD

[29] J. Sutherland. Business object design and implementation. OOPS Messenger, 6(4):170–175, October 1995. CO-DEN OOMEEO. ISSN 1055-6400.

Jang:1995:LSO

[30] Y. Jang. Legacy systems and object technology. OOPS Messenger, 6 (4):176–179, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Taivalsaari:1995:OLS

[31] Antero Taivalsaari, Roland Trauter, and Eduardo Casais. Object-oriented legacy systems and software evolution. OOPS Messenger, 6(4):180–185, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Vaishnavi:1995:RCO

[32] Vijay Vaishnavi and Timothy Korson. Role of a corporate object technology center. OOPS Messenger, 6 (4):186–190, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Tiwary:1995:BLD

[33] Ashutosh Tiwary, Rajendra K. Raj, Douglas S. Lea, and Christopher J. Bosch. Building large distributed software systems using objects. OOPS Messenger, 6(4):191–195, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Brown:1995:DCL

[34] G. Brown. Design and construction of large-scale components. OOPS Messenger, 6(4):196–202, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Wilhelm:1995:ORS

[35] Bob Wilhelm, John Gilbert, and Mohamed Fayad. Object-oriented realtime systems. OOPS Messenger, 6 (4):203–208, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Arora:1995:OTH

[36] S. Arora. Object-oriented technology for health care and medical information systems. OOPS Messenger, 6 (4):209–212, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Laffra:1995:OTL

[37] C. Laffra. OO technology in large financial institutions. OOPS Messenger, 6(4):213–217, October 1995. CODEN OOMEEO. ISSN 1055-6400.

Wong:1995:IRF

[38] Limsoon Wong. An introduction to Remy’s fast polymorphic record projection. ACM SIGMOD Record, 24 (3):34–39, September 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cichelli:1980:MPH

[39] R. J. Cichelli. Minimal perfect hash functions made simple. Communications of the ACM, 23(1):17–19, January 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks in [41], and the author’s response in [40]. A simple improvement giving dramatic speedups is described in [1175].

Cichelli:1980:CMP

[40] R. J. Cichelli. On Cichelli’s minimal perfect hash functions method. Communications of the ACM, 23(12): 728–729, December 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This is the author’s response to the comments in [41] about [39].

Jaeschke:1980:CMP

[41] G. Jaeschke and G. Osterburg. On Cichelli’s minimal perfect hash functions method. Communications of the ACM, 23(12):728–729, December 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This letter to the editor contains comments on [39], together with a response from R. J. Cichelli [40].

Cook:1982:LOM

[42] Curtis R. Cook and R. R. Oldehoeft. A letter oriented minimal perfect hashing function. ACM SIGPLAN Notices, 17(9):18–27, September 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [1175] for significant speedup.

Caron:1990:BTD

[43] J. M. Caron and P. A. Darnell. Bugfind: a tool for debugging optimizing compilers. ACM SIGPLAN Notices, 25(1):17–22, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srinivasan:1990:CDS

[44] S. Srinivasan. Compiler design for sets in Pascal. ACM SIGPLAN Notices, 25 (1):23–24, January 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boute:1990:HVT

[45] R. T. Boute. A heretical view on type embedding. ACM SIGPLAN Notices, 25(1):25–28, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wong:1990:TFV

[46] L. Wong and B. C. Ooi. Treating failure as value. ACM SIGPLAN Notices, 25 (1):29–32, January 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Talia:1990:SCP

[47] D. Talia. Survey and comparison of PARLOG and Concurrent Prolog. ACM SIGPLAN Notices, 25(1):33–42, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chiu:1990:GAU

[48] P. P. K. Chiu and S. T. K. Fu. A generative approach to Universal Cross Assembler design. ACM SIGPLAN Notices, 25(1):43–51, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avenarius:1990:FLP

[49] Adrian Avenarius and Siegfried Oppermann. FWEB: A literate programming system for Fortran 8X. ACM SIG-PLAN Notices, 25(1):52–58, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Payne:1990:PLC

[50] M. Payne, C. Schaffert, and B. Wichmann. Proposal for a language compatible arithmetic standard. ACM SIG-PLAN Notices, 25(1):59–86, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomas:1990:BPS

[51] Edward J. Thomas and Paul W. Oman. A bibliography of programming style. ACM SIGPLAN Notices, 25(2): 7–16, February 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nordstrom:1990:TL

[52] D. J. Nordstrom. Threading Lisp. ACM SIGPLAN Notices, 25(2):17– 24, February 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radensky:1990:TII

[53] A. Radensky. Toward integration of the imperative and logic programming paradigms: Horn-clause programming in the Pascal environment. ACM SIG-PLAN Notices, 25(2):25–34, February 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mehndiratta:1990:SME

[54] B. Mehndiratta and P. S. Grover. Software metrics — an experimental analysis. ACM SIGPLAN Notices, 25(2): 35–41, February 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dinning:1990:ECM

[55] A. Dinning and E. Schonberg. An empirical comparison of monitoring algorithms for access anomaly detection. ACM SIGPLAN Notices, 25(3):1–10, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hseush:1990:MCP

[56] W. Hseush and G. E. Kaiser. Modeling concurrency in parallel debugging. ACM SIGPLAN Notices, 25(3):11–20, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Callahan:1990:AES

[57] D. Callahan, K. Kennedy, and J. Subhlok. Analysis of event synchronization in a parallel programming tool. ACM SIGPLAN Notices, 25(3):

21–30, March 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tick:1990:PTE

[58] E. Tick and N. Ichiyoshi. Programming techniques for efficiently exploiting parallelism in logic programming languages. ACM SIGPLAN Notices, 25 (3):31–39, March 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ambriola:1990:DDI

[59] V. Ambriola, P. Ciancarini, and M. Danelutto. Design and distributed implementation of the parallel logic language shared Prolog. ACM SIG-PLAN Notices, 25(3):40–49, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furuichi:1990:MLL

[60] M. Furuichi, K. Taki, and N. Ichiyoshi. A multi-level load balancing scheme for OR-parallel exhaustive search programs on the multi-PSI. ACM SIG-PLAN Notices, 25(3):50–59, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Katseff:1990:EHV

[61] H. P. Katseff, R. D. Gaglianello, and B. S. Robinson. The evolution of HPC/ VORX. ACM SIGPLAN Notices, 25 (3):60–69, March 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scott:1990:MMP

[62] M. L. Scott, T. J. LeBlanc, and B. D. Marsh. Multi-model parallel programming in psyche. ACM SIGPLAN Notices, 25(3):70–78, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1990:RTC

[63] K. Li, J. F. Naughton, and J. S. Plank. Real-time, concurrent checkpoint for parallel programs. ACM SIGPLAN Notices, 25(3):79–88, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nicol:1990:ASM

[64] D. M. Nicol. Analysis of synchronization in massively parallel discrete-event simulations. ACM SIGPLAN Notices, 25(3):89–98, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramkumar:1990:CKI

[65] B. Ramkumar and L. V. Kale. A Chare kernel implementation of a parallel Prolog compiler. ACM SIG-PLAN Notices, 25(3):99–108, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chang:1990:IBN

[66] P. S. Chang and G. K. Egan. An implementation of a barotropic Numerical Weather Prediction model in the functional language SISAL. ACM SIG-PLAN Notices, 25(3):109–117, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:1990:ERC

[67] R. Gupta. Employing register channels for the exploitation of instruction level parallelism. ACM SIGPLAN Notices, 25(3):118–127, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hieb:1990:CC

[68] R. Hieb and R. Kent Dybvig. Continuations and concurrency. ACM SIG-PLAN Notices, 25(3):128–136, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whitfield:1990:AOO

[69] D. Whitfield and M. L. Soffa. An approach to ordering optimizing transformations. ACM SIGPLAN Notices, 25(3):137–146, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Factor:1990:PTA

[70] M. Factor. The process trellis architecture for real-time monitors. ACM SIG-PLAN Notices, 25(3):147–155, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harvey:1990:ETL

[71] W. Harvey, D. Kalp, M. Tambe, D. McKeown, and A. Newell. The effectiveness of task-level parallelism for high-level vision. ACM SIGPLAN Notices, 25(3):156–167, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bennett:1990:MDS

[72] J. K. Bennett, J. B. Carter, and W. Zwaenepoel. Munin: distributed shared memory based on type-specific memory coherence. ACM SIGPLAN Notices, 25(3):168–176, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chien:1990:CACa

[73] A. A. Chien and W. J. Dally. Concurrent aggregates (CA). ACM SIG-PLAN Notices, 25(3):177–186, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koelbel:1990:SSD

[74] C. Koelbel, P. Mehrotra, and J. Van Rosendale. Supporting shared data structures on distributed memory architectures. ACM SIGPLAN Notices, 25(3):177–186, March 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chien:1990:CACb

[75] Andrew A. Chien and William J. Dally. Concurrent aggregates (CA). ACM SIGPLAN Notices, 25(3):187– 196, March 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herlihy:1990:MIH

[76] M. Herlihy. A methodology for implementing highly concurrent data structures. ACM SIGPLAN Notices, 25(3): 197–206, March 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Labreche:1990:IRT

[77] Pierre Labrèche and Louis Lamarche. Interactors: a real-time executive with multiparty interactions in C++. ACM SIGPLAN Notices, 25(4):20–32, April 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1990:SSF

[78] Brian Meek. The static semantics file. ACM SIGPLAN Notices, 25(4):33–42, April 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prather:1990:PCP

[79] Ronald E. Prather. Proposal for a categorical programming language. ACM SIGPLAN Notices, 25(4):43–50, April 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snelting:1990:HBL

[80] Gregor Snelting. How to build LR parsers which accept incomplete input. ACM SIGPLAN Notices, 25(4):51–58, April 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pagan:1990:CEG

[81] Frank G. Pagan. Comparative efficiency of general and residual parsers. ACM SIGPLAN Notices, 25(4):59–68, April 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stansifer:1990:IVF

[82] Ryan Stansifer. Imperative versus functional. ACM SIGPLAN Notices, 25 (4):69–72, April 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mandl:1990:PCP

[83] Robert Mandl. On ‘powerloop’ constructs in programming languages. ACM SIGPLAN Notices, 25(4):73–82, April 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roberts:1990:RAR

[84] George H. Roberts. From recursive ascent to recursive descent: via compiler optimizations. ACM SIGPLAN Notices, 25(4):83–89, April 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bal:1990:OLD

[85] Henri E. Bal. Orca: a language for distributed programming. ACM SIG-PLAN Notices, 25(5):17–24, May 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:1990:ISN

[86] Mark B. Wells. Identifiers and static name abstraction. ACM SIGPLAN Notices, 25(5):25–28, May 1990. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barnett:1990:PST

[87] Michael P. Barnett. Primitive string transformations as reductions to normal form. ACM SIGPLAN Notices, 25 (5):29–33, May 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barnett:1990:IPC

[88] Michael P. Barnett and Xu Rui. Infix to prefix conversion as a PST reduction. ACM SIGPLAN Notices, 25 (5):34–38, May 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:1990:DES

[89] Xinjie Jiang and Yongsen Xu. Diverse executable semantics definitions in NUSL and an implementation of functional types. ACM SIGPLAN Notices, 25(5):39–52, May 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Louden:1990:PCC

[90] Kenneth Louden. P-code and compiler portability: experience with a Modula-2 optimizing compiler. ACM SIGPLAN Notices, 25(5):53–59, May 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanHorebeek:1990:CFS

[91] Ivo Van Horebeek and Johan Lewi. Are constructive formal specifications less abstract? ACM SIGPLAN Notices, 25(5):60–68, May 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sundaresan:1990:TNP

[92] Neelakantan Sundaresan. Translation of nested Pascal routines to C. ACM SIGPLAN Notices, 25(5):69–81, May 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dobler:1990:CNC

[93] H. Dobler and K. Pirklbauer. Coco-2: a new compiler compiler. ACM SIGPLAN Notices, 25(5):82–90, May 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaiser:1990:GMM

[94] Klaus Kaiser. Generating the minimal model in Prolog. ACM SIG-PLAN Notices, 25(5):91–94, May 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keller:1990:GNN

[95] Daniel Keller. A guide to natural naming. ACM SIGPLAN Notices, 25 (5):95–102, May 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Treadway:1990:USA

[96] Peter L. Treadway. The use of sets as an application programming technique. ACM SIGPLAN Notices, 25 (5):103–116, May 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burshteyn:1990:MFG

[97] Boris Burshteyn. On the modification of the formal grammar at parse time. ACM SIGPLAN Notices, 25(5): 117–123, May 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deChamplain:1990:SSE

[98] Michel de Champlain. Synapse: a small and expressible object-based real-time programming language. ACM SIG-PLAN Notices, 25(5):124–134, May 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Giegerich:1990:SVC

[99] Robert Giegerich. On the structure of verifiable code generator specifications. ACM SIGPLAN Notices, 25(6): 1–8, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wendt:1990:FCG

[100] Alan L. Wendt. Fast code generation using automatically-generated decision trees. ACM SIGPLAN Notices, 25(6):9–15, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p9-wendt/

.

Pettis:1990:PGC

[101] Karl Pettis and Robert C. Hansen. Profile guided code positioning. ACM SIGPLAN Notices, 25(6):16–27, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p16-pettis/

.

Santhanam:1990:RAA

[102] Vatsa Santhanam and Daryl Odnert. Register allocation across procedure and module boundaries. ACM SIG-PLAN Notices, 25(6):28–39, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p28-santhanam/

.

Nickerson:1990:GCR

[103] Brian R. Nickerson. Graph coloring register allocation for processors with multi-register operands. ACM SIGPLAN Notices, 25(6):40– 52, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p40-nickerson/

.

Callahan:1990:IRA

[104] David Callahan, Steve Carr, and Ken Kennedy. Improving register allocation for subscripted variables. ACM SIGPLAN Notices, 25 (6):53–65, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p53-callahan/

.

Hieb:1990:RCP

[105] Robert Hieb, R. Kent Dybvig, and Carl Bruggeman. Representing control in the presence of first-class continuations. ACM SIGPLAN Notices, 25 (6):66–77, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p66-hieb/

.

Kessler:1990:FBD

[106] Peter B. Kessler. Fast breakpoints. design and implementation. ACM SIGPLAN Notices, 25(6):78– 84, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p78-kessler/

.

Pugh:1990:TDR

[107] William Pugh and Grant Weddell. Two-directional record layout for multiple inheritance. ACM SIGPLAN Notices, 25(6):85–91, June 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pugh:1990:TRL

[108] William Pugh and Grant Weddell. Two-directional record layout for multiple inheritance. ACM SIGPLAN Notices, 25(6):85–91, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/93542/p85-pugh/

.

Clinger:1990:HRF

[109] William D. Clinger. How to read floating point numbers accurately. ACM SIGPLAN Notices, 25(6):92– 101, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/93542/p92-clinger/

. See also output algorithms in [?, 111, 1431, ?, ?].

Lee:1990:OPC

[110] Vernon A. Lee, Jr. and Hans-J. Boehm. Optimizing programs over the constructive reals. ACM SIG-PLAN Notices, 25(6):102–111, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/93542/p102-lee/

.

Steele:1990:HPF

[111] Guy L. Steele Jr. and Jon L. White. How to print floating-point numbers accurately. ACM SIGPLAN Notices, 25(6):112–126, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p112-steele/

. See also input algorithm in [109, ?], and a faster output algorithm in [1431] and [?], IBM S/360 algorithms in [?] for both IEEE 754 and S/360 formats, and a twenty-year retrospective in [?]. In electronic mail dated Wed, 27 Jun 1990 11:55:36 EDT, Guy Steele reported that an intrepid pre-SIGPLAN 90 conference implementation of what is stated in the paper revealed 3 mistakes:

1. Table 5 (page 124): insert k <-- 0 after assertion, and also delete k <-- 0 from Table 6.

2. Table 9 (page 125):

for  -1:USER!("");
substitute -1:USER!("0");

and delete the comment.

3. Table 10 (page 125):

for  fill(-k, "0")
substitute fill(-k-1, "0")

Cormack:1990:TDP

[112] Gordon V. Cormack and Andrew K. Wright. Type-dependent parameter inference. ACM SIGPLAN Notices, 25 (6):127–136, June 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cormack:1990:TPI

[113] Gordon V. Cormack and Andrew K. Wright. Type-dependent parameter inference. ACM SIGPLAN Notices, 25 (6):127–136, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p127-cormack/

.

Anderson:1990:CHA

[114] Steven Anderson and Paul Hudak. Compilation of Haskell array comprehensions for scientific computing. ACM SIGPLAN Notices, 25(6):137– 149, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p137-anderson/

.

Chambers:1990:ITA

[115] Craig Chambers and David Ungar. Iterative type analysis and extended message splitting: Optimizing dynamically-typed object-oriented programs. ACM SIGPLAN Notices, 25 (6):150–164, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p150-chambers/

.

Jacobs:1990:TDS

[116] Dean Jacobs. Type declarations as subtype constraints in logic programming. ACM SIGPLAN Notices, 25(6): 165–173, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p165-jacobs/

.

Debray:1990:TGA

[117] Saumya K. Debray, Nai-Wei Lin, and Manuel Hermenegildo. Task granularity analysis in logic programs. ACM SIGPLAN Notices, 25(6):174– 188, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p174-debray/

.

Mitsolides:1990:GRC

[118] Thanasis Mitsolides and Malcolm Harrison. Generators and the replicator control structure in the parallel environment of ALLOY. ACM SIGPLAN Notices, 25(6):189–196, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p189-mitsolides/

.

Teitelbaum:1990:HAG

[119] Tim Teitelbaum and Richard Chapman. Higher-order attribute grammars and editing environments. ACM SIGPLAN Notices, 25(6):197–208, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p197-teitelbaum/

.

Teitelbaum:1990:HOA

[120] Tim Teitelbaum and Richard Chapman. Higher-order attribute grammars and editing environments. ACM SIG-PLAN Notices, 25(6):197–208, June 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jourdan:1990:DIE

[121] Martin Jourdan, Didier Parigot, Catherine Julié, Olivier Durin, and Carole Le Bellec. Design, implementation and evaluation of the FNC-2 attribute grammar system. ACM SIGPLAN Notices, 25(6):209– 222, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p209-jourdan/

.

Henry:1990:UWI

[122] Robert R. Henry, Kenneth M. Whaley, and Bruce Forstall. The University of Washington illustrating compiler. ACM SIGPLAN Notices, 25(6): 223–233, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p223-henry/

.

Horwitz:1990:IST

[123] Susan Horwitz. Identifying the semantic and textual differences between two versions of a program. ACM SIGPLAN Notices, 25(6):234– 245, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

proceedings/pldi/93542/p234-horwitz/

.

Agrawal:1990:DPS

[124] Hiralal Agrawal and Joseph R. Horgan. Dynamic program slicing. ACM SIGPLAN Notices, 25(6):246– 256, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p246-agrawal/

.

Ballance:1990:PDW

[125] Robert A. Ballance, Arthur B. Maccabe, and Karl J. Ottenstein. The Program Dependence Web: A representation supporting control-, data- , and demand-driven interpretation of imperative languages. ACM SIGPLAN Notices, 25(6):257–271, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p257-ottenstein/

.

Gupta:1990:FLA

[126] Rajiv Gupta. A fresh look at optimizing array bound checking. ACM SIGPLAN Notices, 25(6):272– 282, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p272-gupta/

.

Gupta:1990:FLO

[127] Rajiv Gupta. A fresh look at optimizing array bound checking. ACM SIG-PLAN Notices, 25(6):272–282, June 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ammarguellat:1990:ARI

[128] Zahira Ammarguellat and W. L. Harrison, III. Automatic recognition of induction variables and recurrence relations by abstract interpretation. ACM SIGPLAN Notices, 25(6): 283–295, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/93542/p283-ammarguellat/

.

Chase:1990:APS

[129] David R. Chase, Mark Wegman, and F. Kenneth Zadeck. Analysis of pointers and structures. ACM SIGPLAN Notices, 25(6):296–310, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/93542/p296-chase/

.

Tseng:1990:CPL

[130] Ping-Sheng Tseng. Compiling programs for a linear systolic array. ACM SIGPLAN Notices, 25(6):311– 321, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p311-tseng/

.

Sarkar:1990:IRF

[131] Vivek Sarkar. Instruction reordering for fork-join parallelism. ACM SIGPLAN Notices, 25(6):322–336, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p322-sarkar/

.

Cytron:1990:CRC

[132] Ron Cytron, Jeanne Ferrante, and Vivek Sarkar. Compact representations for control dependence. ACM SIGPLAN Notices, 25(6):337– 351, June 1990. CODEN SIN-ODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/93542/p337-cytron/

.

Eriksson:1990:CEM

[133] Mikael Eriksson. A correct example of multiple inheritance. ACM SIG-PLAN Notices, 25(7):7–10, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Drashansky:1990:CFP

[134] Tsvetan T. Drashansky. A concept for functional programming and distributed data processing in a local area network and its implementation. ACM SIGPLAN Notices, 25(7):11–18, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lins:1990:PEO

[135] C. Lins. Programming without enumerations in Oberon. ACM SIG-PLAN Notices, 25(7):19–27, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paakki:1990:OTE

[136] Jukka Paakki, Anssi Karhinen, and Tomi Silander. Orthogonal type extensions and reductions. ACM SIG-PLAN Notices, 25(7):28–38, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1990:SIB

[137] C. K. Yuen and W. F. Wong. A self interpreter for BaLinda Lisp. ACM SIGPLAN Notices, 25(7):39–58, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenfield:1990:BZE

[138] Peter Greenfield. Bagof/3 to ZF expression: an example Prolog application manually translated to Miranda. ACM SIGPLAN Notices, 25(7):59–62, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyd:1990:PC

[139] Joanne L. Boyd and Gerald M. Karam. Prolog in ‘C’. ACM SIGPLAN Notices, 25(7):63–71, July 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Axford:1990:ELC

[140] Tom Axford. An elementary language construct for parallel programming. ACM SIGPLAN Notices, 25 (7):72–80, July 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McIntosh:1990:PLR

[141] Harold V. McIntosh and Gerardo Cisneros. The programming languages REC and convert. ACM SIGPLAN Notices, 25(7):81–94, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakkinen:1990:EBS

[142] Markku Sakkinen. On embedding Boolean as a subtype of integer. ACM SIGPLAN Notices, 25(7):95–96, July 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krishamurthy:1990:BSP

[143] Sanjay M. Krishamurthy. A brief survey of papers on scheduling for pipelined processors. ACM SIGPLAN Notices, 25(7):97–106, July 1990. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ince:1990:ABS

[144] Darrel Ince. An annotated bibliography of software metrics. ACM SIG-PLAN Notices, 25(8):15–23, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wong:1990:TFS

[145] L. Wong and B. C. Ooi. Treating failure as state. ACM SIGPLAN Notices, 25 (8):24–26, August 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steigerwald:1990:CPS

[146] R. Steigerwald and M. Nelson. Concurrent programming in Smalltalk-80. ACM SIGPLAN Notices, 25(8):27–36, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elsworth:1990:ACC

[147] W. F. Elsworth and M. B. A. Parkes. Automated compiler construction based on top-down syntax analysis and attribute evaluation. ACM SIGPLAN Notices, 25(8):37–42, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milanese:1990:CFG

[148] V. Milanese. Coupling FORTH and GKS in monitoring processes. ACM SIGPLAN Notices, 25(8):43–52, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dekeyser:1990:EEV

[149] J. Dekeyser, Ph. Marquet, and Ph. Pruex. EVA: an explicit vector language. ACM SIGPLAN Notices, 25 (8):53–71, August 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:1990:CTB

[150] V. Schneider. Compiler textbook bibliographies considered harmful. ACM SIGPLAN Notices, 25(8):72–74, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1990:TVD

[151] B. Meek. Two-valued datatypes. ACM SIGPLAN Notices, 25(8):75–79, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1990:FJO

[152] B. Meek. Failure is not just one value. ACM SIGPLAN Notices, 25(8):80–83, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Issendorff:1990:TOa

[153] H. Von Issendorff. A theory of organization. ACM SIGPLAN Notices, 25 (8):84–102, August 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1990:AP

[154] M. Berry. Adverbial programming. ACM SIGPLAN Notices, 25(8):103– 105, August 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jonsson:1990:FCN

[155] D. Jonsson. The flow of control notations Pancode and boxcharts. ACM SIGPLAN Notices, 25(8):106–119, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Resler:1990:VVC

[156] D. Resler. VisiCLANG: a visible compiler for CLANG. ACM SIGPLAN Notices, 25(8):120–123, August 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spackman:1990:BRI

[157] Stephen P. Spackman. Book review: Interpretation and Instruction Path Coprocessing by Eddy H. Debaere & Jan M. Van Campenhout: (MIT Press 1990). ACM SIGPLAN Notices, 25(9): 7–9, September 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gyorkos:1990:DSS

[158] J. Gyorkos, I. Rozman, and T. Welzer. Dynamical system specifications as validation media. ACM SIGPLAN Notices, 25(9):10–16, September 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Talla:1990:NTO

[159] D. Talla. Notes on termination of OCCAM processes. ACM SIGPLAN Notices, 25(9):17–24, September 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jung:1990:CT

[160] C. Jung and E. Sibert. Communication types. ACM SIGPLAN Notices, 25(9): 25–30, September 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonIssendorff:1990:TO

[161] H. von Issendorff. A theory of organization. ACM SIGPLAN Notices, 25(9): 31–38, September 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rulghaver:1990:LSS

[162] A. B. Rulghaver and T. T. E. Yeo. Language support for a semi-dataflow parallel programming environment. ACM SIGPLAN Notices, 25(9):39– 47, September 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Booch:1990:DCB

[163] Grady Booch and Michael Vilot. The design of the C++ Booch Components. ACM SIGPLAN Notices, 25 (10):1–11, October 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gossain:1990:IDM

[164] Sanjiv Gossain and Bruce Anderson. An iterative-design model for reusable object-oriented software. ACM SIG-PLAN Notices, 25(10):12–27, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bear:1990:GSO

[165] Stephen Bear, Phillip Allen, Derek Coleman, and Fiona Hayes. Graphical specification of object oriented systems. ACM SIGPLAN Notices, 25(10): 28–37, October 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nierstrasz:1990:VOP

[166] Oscar Nierstrasz and Michael Papathomas. Viewing object as patterns of communicating agents. ACM SIG-PLAN Notices, 25(10):38–43, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andreoli:1990:BCS

[167] Jean-Marc Andreoli and Remo Pareschi. LO and behold! Concurrent structured processes. ACM SIGPLAN Notices, 25(10):44–56, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahn:1990:ASC

[168] K. Kahn and Vijay A. Saraswat. Actors as a special case of concurrent constraint (logic) programming. ACM SIGPLAN Notices, 25(10):57–66, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lerner:1990:BSE

[169] Barbara Staudt Lerner and A. Nico Habermann. Beyond schema evolution to database reorganization. ACM SIG-PLAN Notices, 25(10):67–76, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freeman-Benson:1990:KMO

[170] Bjorn N. Freeman-Benson. Kaleidoscope: mixing objects, constraints, and imperative programming. ACM SIG-PLAN Notices, 25(10):77–88, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meseguer:1990:LTC

[171] José Meseguer. A logical theory of concurrent objects. ACM SIGPLAN Notices, 25(10):101–115, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Purchase:1990:MPS

[172] Jan A. Purchase and Russel L. Winder. Message pattern specifications: a new technique for handling errors in parallel object oriented systems. ACM SIG-PLAN Notices, 25(10):116–125, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kafura:1990:GCA

[173] Dennis Kafura, Doug Washabaugh, and Jeff Nelson. Garbage collection of actors. ACM SIGPLAN Notices, 25(10):126–134, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deChampeaux:1990:SAO

[174] Dennis de Champeaux, Larry Constantine, Ivar Jacobson, Stephen Mellor, Paul Ward, and Edward Yourdon. Structured analysis and object oriented analysis. ACM SIGPLAN Notices, 25(10):135–139, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Madsen:1990:STO

[175] Ole Lehrmann Madsen, Boris Magnusson, and Birger Mølier-Pedersen. Strong typing of object-oriented languages revisited. ACM SIGPLAN Notices, 25(10):140–150, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palsberg:1990:TSO

[176] Jens Palsberg and Michael I. Schwartzbach. Type substitution for objectoriented programming. ACM SIG-PLAN Notices, 25(10):151–160, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

America:1990:POO

[177] Pierre America and Frank van der Linden. A parallel object-oriented language with inheritance and subtyping. ACM SIGPLAN Notices, 25(10):161– 168, October 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Helm:1990:CSB

[178] Richard Helm, Ian M. Holland, and Dipayan Gangopadhyay. Contracts: specifying behavioral compositions in object-oriented systems. ACM SIG-PLAN Notices, 25(10):169–180, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berlin:1990:WOC

[179] Lucy M. Berlin. When objects collide experiences with reusing multiple class hierarchies. ACM SIGPLAN Notices, 25(10):181–193, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paepcke:1990:PST

[180] Andreas Paepcke. PCLOS: stress testing CLOS experiencing the metaobject protocol. ACM SIGPLAN Notices, 25(10):194–211, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavens:1990:RAO

[181] Gary T. Leavens and William E. Weihl. Reasoning about object-oriented programs that use subtypes. ACM SIG-PLAN Notices, 25(10):212–223, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Straube:1990:TCQ

[182] Dave D. Straube and M. Tamer Oezsu. Type consistency of queries in an object-oriented database system. ACM SIGPLAN Notices, 25(10):224– 233, October 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stein:1990:IOD

[183] Jacob Stein, Tim Andrews, Bill Kent, Kate Rotzell, and Dan Weinreb. Issues in object database management. ACM SIGPLAN Notices, 25(10):235– 236, October 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haarslev:1990:FVO

[184] Volker Haarslev and Ralf Möller. A framework for visualizing objectoriented systems. ACM SIGPLAN Notices, 25(10):237–244, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ECOOP ’90 Conference on Object-Oriented Programming: Systems, Languages, and Applications.

McDonald:1990:PMV

[185] John Alan McDonald, Werner Stuetzle, and Andreas Buja. Painting multiple views of complex objects. ACM SIG-PLAN Notices, 25(10):245–257, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shan:1990:MUS

[186] Yen-Ping Shan. MoDE: a UIMS for Smalltalk. ACM SIGPLAN Notices, 25(10):258–268, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Habert:1990:CKS

[187] Sabine Habert and Laurence Mosseri. COOL: kernel support for objectoriented environments. ACM SIG-PLAN Notices, 25(10):269–275, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Faust:1990:POO

[188] John E. Faust and Henry M. Levy. The performance of an object-oriented threads package. ACM SIGPLAN Notices, 25(10):278–288, October 1990.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ishikawa:1990:OOR

[189] Yutaka Ishikawa and Hideyuki Tokuda. Object-oriented real-time language design: constructs for timing constraints. ACM SIGPLAN Notices, 25(10):289– 298, October 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeNatale:1990:ORW

[190] Rick DeNatale, Charles Irby, John LaLonde, Burton Leathers, and Reed Phillips. OOP in the real world. ACM SIGPLAN Notices, 25(10):299– 302, October 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bracha:1990:MBI

[191] Gilad Bracha and William Cook. Mixin-based inheritance. ACM SIG-PLAN Notices, 25(10):303–311, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carre:1990:PVN

[192] Bernard Carré and Jean-Marc Geib. The point of view notion for multiple inheritance. ACM SIGPLAN Notices, 25(10):312–321, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dony:1990:EHO

[193] Christophe Dony. Exception handling and object-oriented programming: towards a synthesis. ACM SIGPLAN Notices, 25(10):322–330, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rymer:1990:ODO

[194] John R. Rymer, Richard Mark Soley, William Stephen Andreas, Ian Fuller, Neal Jacobson, and Richard A. Demers. OOPSLA distributed object management. ACM SIGPLAN Notices, 25(10):331–333, October 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

staff:1990:AIC

[195] SIGPLAN Notices staff. Abstracts of the 1990 International Conference on Computer Languages Proceedings. ACM SIGPLAN Notices, 25(11):26– 34, November 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christiansen:1990:SAG

[196] H. Christiansen. A survey of adaptable grammars. ACM SIGPLAN Notices, 25(11):35–44, November 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andersen:1990:ELD

[197] B. Andersen. Ellie language definition report. ACM SIGPLAN Notices, 25(11):45–64, November 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ivanovic:1990:ICI

[198] M. Ivanovic and Z. Budimac. Involving coroutines in interaction between functional and conventional language. ACM SIGPLAN Notices, 25(11):65– 74, November 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McLaughlin:1990:PRP

[199] R. McLaughlin. A plea for a readable Prolog programming style. ACM SIG-PLAN Notices, 25(11):75–79, November 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kovats:1990:CAP

[200] T. Kovats. A conservative alternative to Pancode. ACM SIGPLAN Notices, 25(11):80–84, November 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:1990:FPP

[201] T. Christopher and D. Ameiss. Functional programming in a parallel environment: the implementation of FP in MDC. ACM SIGPLAN Notices, 25(11):85–94, November 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dulli:1990:DDG

[202] Susi Dulli and Vitaliano Milanese. A database design for graphical models. ACM SIGPLAN Notices, 25(12):30– 36, December 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lieu:1990:LTB

[203] Choun Tong Lieu and Djamil Sarni. O.LISP theoretical basis and description. ACM SIGPLAN Notices, 25(12): 37–44, December 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burshteyn:1990:GRF

[204] Boris Burshteyn. Generation and recognition of formal languages by modifiable grammars. ACM SIGPLAN Notices, 25(12):45–53, December 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bothe:1990:PSB

[205] K. Bothe. A Prolog space benchmark suite: a new tool to compare Prolog implementations. ACM SIGPLAN Notices, 25(12):54–60, December 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Auguston:1990:PLR

[206] M. I. Auguston. Programming language RIGAL as a compiler writing tool. ACM SIGPLAN Notices, 25(12): 61–69, December 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jellinghaus:1990:ELO

[207] Robert Jellinghaus. Eiffel Linda: an object-oriented Linda dialect. ACM SIGPLAN Notices, 25(12):70–84, December 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dotzel:1990:DMG

[208] Günter Dotzel. Does Modula-2 generate racehorses?: comparison of compiler generated code quality for floating point arithmetic. ACM SIGPLAN Notices, 25(12):85–88, December 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yovev:1990:UEU

[209] C. S. Yovev. Universal editor unattainable. ACM SIGPLAN Notices, 25(12): 89–92, December 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhai:1990:PIC

[210] Cheng xiang Zhai. Preliminary ideas of a conceptual programming language. ACM SIGPLAN Notices, 25(12):93– 100, December 1990. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1991:ABL

[211] Lisa M. C. Smith and Mansur H. Samadzadeh. An annotated bibliography of literate programming. ACM SIGPLAN Notices, 26(1):14–20, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pancake:1991:BPD

[212] Cherri M. Pancake and Sue Utter. A bibliography of parallel debuggers, 1990 edition. ACM SIGPLAN Notices, 26(1):21–37, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calliss:1991:CMC

[213] Frank W. Calliss. A comparison of module constructs in programming languages. ACM SIGPLAN Notices, 26 (1):38–46, January 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coward:1991:PFL

[214] P. David Coward. Path feasibility, linear optimizers and the evaluate standard form. ACM SIGPLAN Notices, 26 (1):47–56, January 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lujun:1991:CIR

[215] Shang Lujun. Cluster: an informal report. ACM SIGPLAN Notices, 26(1): 57–74, January 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhai:1991:PIC

[216] Cheng xiang Zhai. Preliminary Ideas of a Conceptual Programming Language. ACM SIGPLAN Notices, 26(1):77–84, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Watanabe:1991:ELS

[217] Tan Watanabe, Kazuko Sakuma, Hideyuki Arai, and Kohbun Umetani. Essential language el(α) — a reduced expression set language for system programming. ACM SIGPLAN Notices, 26(1):85–98, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stadel:1991:OOP

[218] Manfred Stadel. Object oriented programming techniques to replace software components on the fly in a running program. ACM SIGPLAN Notices, 26(1):99–108, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Teodosiu:1991:HOP

[219] Dan Teodosiu. HARE: an optimizing portable compiler for Scheme. ACM SIGPLAN Notices, 26(1):109–120, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Franz:1991:RGT

[220] Michael Franz. The rewards of generating true 32-bit code. ACM SIG-PLAN Notices, 26(1):121–123, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ejiogu:1991:TSM

[221] Lem O. Ejiogu. TM: a systematic methodology of software metrics. ACM SIGPLAN Notices, 26(1):124–132, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joy:1991:GRS

[222] Mike Joy and Tom Axford. GCODE: a revised standard for a graph representation for functional programs. ACM SIGPLAN Notices, 26(1):133–139, January 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pemberton:1991:SIA

[223] Steven Pemberton. A short introduction to the ABC language. ACM SIG-PLAN Notices, 26(2):11–16, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coyle:1991:BAI

[224] Christopher Coyle and Peter Crogono. Building abstract iterators using continuations. ACM SIGPLAN Notices, 26(2):17–24, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gold:1991:IRN

[225] E. Mark Gold. Incremental reduction with nested constraints. ACM SIG-PLAN Notices, 26(2):25–34, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cashman:1991:BET

[226] Mark Cashman. The benefits of enumerated types in Modula-2. ACM SIG-PLAN Notices, 26(2):35–39, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaplan:1991:PRP

[227] Randy M. Kaplan. A plea for readable pleas for a readable Prolog programming style. ACM SIGPLAN Notices, 26(2):41–50, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klug:1991:VML

[228] Michael Klug. VisiCola, a model and a language for visibility control in programming languages. ACM SIGPLAN Notices, 26(2):51–63, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Madhav:1991:ICF

[229] B. I. B. Madhav and Narayan Hegde. Implementing C function calls in rules for an expert system shell. ACM SIG-PLAN Notices, 26(2):64–66, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pellegrini:1991:STE

[230] Marco Pellegrini and Raimondo Sepe. SetLog, a tool for experimenting with new semantics. ACM SIGPLAN Notices, 26(2):67–74, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zoran:1991:TWT

[231] Putnik Zoran, Budimac Zoran, and Ivanovic Mirjana. Turtle walk through functional language. ACM SIGPLAN Notices, 26(2):75–82, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bernecky:1991:FA

[232] Robert Bernecky. Fortran 90 arrays. ACM SIGPLAN Notices, 26(2): 83–98, February 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Santo:1991:ESC

[233] Michele Di Santo and Wilma Russo. The ensemble system: concurrent programming on a personal computer. ACM SIGPLAN Notices, 26(2):99– 108, February 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keller:1991:AGC

[234] Wolfgang Keller. Automated generation of code using backtracking parsers for attribute grammars. ACM SIG-PLAN Notices, 26(2):109–117, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Placer:1991:MRN

[235] John Placer. Multiparadigm research: a new direction of language design. ACM SIGPLAN Notices, 26(3):9–17, March 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roberts:1991:NMG

[236] George H. Roberts. A note on modifiable grammars. ACM SIGPLAN Notices, 26(3):18, March 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roberts:1991:SDU

[237] George H. Roberts. Searching in discrete universes. ACM SIGPLAN Notices, 26(3):19–24, March 1991. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fantechi:1991:EFD

[238] A. Fantechi, S. Gnesi, F. Leggio, and P. Talini. Executing the formal definition of Ada. ACM SIG-PLAN Notices, 26(3):25–34, March 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Delrieux:1991:TII

[239] Claudio Delrieux, Pablo Azero, and Fernando Tohme. Toward integrating imperative and logic programming paradigms: WYSIWYG approach to PROLOG programming. ACM SIG-PLAN Notices, 26(3):35–44, March 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1991:SIS

[240] Paul R. Wilson. Some issues and strategies in heap management and memory hierarchies. ACM SIGPLAN Notices, 26(3):45–52, March 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Konrad:1991:SMM

[241] Erhard Konrad. Software metrics, measurement theory, and viewpoints. ACM SIGPLAN Notices, 26(3):53–62, March 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:1991:OOC

[242] Chung-Shyan Liu. On the objectorientedness of C++. ACM SIG-PLAN Notices, 26(3):63–67, March 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oyang:1991:EMW

[243] Yen-Jen Oyang. Exploiting multiway branching to boost superscalar processor performance. ACM SIG-PLAN Notices, 26(3):68–78, March 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dobler:1991:TPC

[244] H. Dobler. Top-down parsing in Coco-2. ACM SIGPLAN Notices, 26(3): 79–87, March 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shepherd:1991:WTP

[245] Janice C. Shepherd. Why a two pass front end? ACM SIGPLAN Notices, 26 (3):88–94, March 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolfe:1991:VIS

[246] Andrew Wolfe and John P. Shen. A variable instruction stream extension to the VLIW architecture. ACM SIG-PLAN Notices, 26(4):2–14, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Katevenis:1991:RBP

[247] Manolis Katevenis and Nestoras Tzartzanis. Reducing the branch penalty by rearranging instructions in double-width memory. ACM SIG-PLAN Notices, 26(4):15–27, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1991:FPP

[248] Roland L. Lee, Alex Y. Kwok, and Fayé A. Briggs. The floating-point performance of a superscalar SPARC processor. ACM SIGPLAN Notices, 26 (4):28–37, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Callahan:1991:SP

[249] David Callahan, Ken Kennedy, and Allan Porterfield. Software prefetching. ACM SIGPLAN Notices, 26(4):40–52, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sohi:1991:HBD

[250] Gurindar S. Sohi and Manoj Franklin. High-bandwidth data memory systems for superscalar processors. ACM SIGPLAN Notices, 26(4):53–62, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lam:1991:CPO

[251] Monica S. Lam, Edward E. Rothberg, and Michael E. Wolf. The cache performance and optimization of blocked algorithms. ACM SIGPLAN Notices, 26(4):63–74, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mogul:1991:ECS

[252] Jeffrey C. Mogul and Anita Borg. The effect of context switches on cache performance. ACM SIGPLAN Notices, 26 (4):75–84, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keppel:1991:PIF

[253] David Keppel. A portable interface for on-the-fly instruction space modification. ACM SIGPLAN Notices, 26 (4):86–95, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Appel:1991:VMP

[254] Andrew W. Appel and Kai Li. Virtual memory primitives for user programs. ACM SIGPLAN Notices, 26 (4):96–107, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1991:IAO

[255] Thomas E. Anderson, Henry M. Levy, Brian N. Bershad, and Edward D. Lazowska. The interaction of architecture and operating system design. ACM SIGPLAN Notices, 26(4): 108–120, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bradlee:1991:IRA

[256] David G. Bradlee, Susan J. Eggers, and Robert R. Henry. Integrating register allocation and instruction scheduling for RISCs. ACM SIGPLAN Notices, 26 (4):122–131, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benitez:1991:CGS

[257] Manuel E. Benitez and Jack W. Davidson. Code generation for streaming: An access/execute mechanism. ACM SIG-PLAN Notices, 26(4):132–141, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bagrodia:1991:EIH

[258] Rajiv Bagrodia and Sharad Mathur. Efficient implementation of high level parallel programs. ACM SIGPLAN Notices, 26(4):142–151, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mangione-Smith:1991:VRD

[259] William Mangione-Smith, Santosh G. Abraham, and Edward S. Davidson. Vector register design for polycyclic vector scheduling. ACM SIGPLAN Notices, 26(4):154–163, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culler:1991:FGP

[260] David E. Culler, Anurag Sah, Klaus E. Schauser, Thorsten von Eicken, and John Wawrzynek. Fine-grain parallelism with minimal hardware support: A compiler-controlled threaded abstract machine. ACM SIGPLAN Notices, 26(4):164–175, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wall:1991:LIL

[261] David W. Wall. Limits of instructionlevel parallelism. ACM SIGPLAN Notices, 26(4):176–188, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1991:PCP

[262] Edward K. Lee and Randy H. Katz. Performance consequences of parity placement in disk arrays. ACM SIG-PLAN Notices, 26(4):190–199, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cate:1991:ICC

[263] Vincent Cate and Thomas Gross. Integration of compression and caching for a two-level file system. ACM SIG-PLAN Notices, 26(4):200–211, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bolosky:1991:NPT

[264] William J. Bolosky, Michael L. Scott, Robert P. Fitzgerald, Robert J. Fowler, and Alan L. Cox. NUMA policies and their relation to memory architecture. ACM SIGPLAN Notices, 26(4): 212–221, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaiken:1991:LDS

[265] David Chaiken, John Kubiatowicz, and Anant Agarwal. LimitLESS directories: A scalable cache coherence scheme. ACM SIGPLAN Notices, 26 (4):224–234, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Min:1991:ECB

[266] Sang Lyul Min and Jong-Deok Choi. An efficient cache-based access anomaly detection scheme. ACM SIG-PLAN Notices, 26(4):235–244, April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gharachorloo:1991:PEM

[267] Kourosh Gharachorloo, Anoop Gupta, and John L. Hennessy. Performance evaluation of memory consistency models for shared memory multiprocessors. ACM SIGPLAN Notices, 26(4): 245–257, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freudenthal:1991:PCF

[268] Eric Freudenthal and Allan Gottlieb. Process coordination with fetch-and-increment. ACM SIGPLAN Notices, 26 (4):260–268, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mellor-Crummey:1991:SC

[269] John M. Mellor-Crummey and Michael L. Scott. Synchronization without contention. ACM SIGPLAN Notices, 26 (4):269–278, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1991:CRB

[270] Douglas Johnson. The case for a read barrier. ACM SIGPLAN Notices, 26 (4):279–287, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cmelik:1991:AMS

[271] Robert F. Cmelik, Shing I. Kong, David R. Ditzel, and Edmund J. Kelly.

An analysis of MIPS and SPARC instruction set utilization on the SPEC benchmarks. ACM SIGPLAN Notices, 26(4):290–301 (or 290–302??), April 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hall:1991:PCA

[272] C. Brian Hall and Kevin O’Brien. Performance characteristics of architectural features of the IBM RISC System/6000. ACM SIGPLAN Notices, 26 (4):303–309, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bhandarkar:1991:PAC

[273] Dileep Bhandarkar and Douglas W. Clark. Performance from architecture: Comparing a RISC and a CISC with similar hardware organization. ACM SIGPLAN Notices, 26(4): 310–319, April 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meeson:1991:BRD

[274] Reginald Meeson. Book review: Data Abstraction and Object-Oriented Programming in C++ by Keith Gorlen, Sanford Orlow, and Perry Plexico: (John Wiley & Sons, 1990). ACM SIGPLAN Notices, 26(5):25–26, May 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zuse:1991:EKS

[275] Horst Zuse and Peter Bollmann. Erhard Konrad: software metrics, measurement theory, and viewpointscritical remarks on a new approach.

ACM SIGPLAN Notices, 26(5):27–36, May 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1991:TRU

[276] Ellis S. Cohen. Trip report: UIST’90 The Annual Symposium on User Interface Software and Technology Snowbird, Utah, October 3–5, 1990. ACM SIGPLAN Notices, 26(5):37–42, May 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McNally:1991:TMI

[277] David J. McNally and Antony J. T. Davie. Two models for integrating persistence and lazy functional languages. ACM SIGPLAN Notices, 26(5):43–52, May 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sicherman:1991:ALD

[278] G. L. Sicherman. An Algorithmic Language for Database Operations. ACM SIGPLAN Notices, 26(5):53–58, May 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rekers:1991:SPA

[279] Jan Rekers and Wilco Koorn. Substring parsing for arbitrary context-free grammars. ACM SIGPLAN Notices, 26(5):59–66, May 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maydan:1991:EED

[280] Dror E. Maydan, John L. Hennessy, and Monica S. Lam. Efficient and exact data dependence analysis. ACM SIGPLAN Notices, 26(6):1–14, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p1-maydan/

.

Goff:1991:PDT

[281] Gina Goff, Ken Kennedy, and Chau-Wen Tseng. Practical dependence testing. ACM SIGPLAN Notices, 26(6):15–29, June 1991. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p15-goff/

.

Wolf:1991:DLO

[282] Michael E. Wolf and Monica S. Lam. A data locality optimizing algorithm. ACM SIGPLAN Notices, 26(6):30–44, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p30-wolf/

.

Morris:1991:CPC

[283] W. G. Morris. CCG: a prototype coagulating code generator. ACM SIGPLAN Notices, 26(6):45–58, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p45-morris/

.

Wall:1991:PPB

[284] David W. Wall. Predicting program behavior using real or estimated profiles. ACM SIGPLAN Notices, 26(6):59–70, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p59-wall/

.

McFarling:1991:PMI

[285] Scott McFarling. Procedure merging with instruction caches. ACM SIGPLAN Notices, 26(6):71–79, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p71-mcfarling/

.

Launchbury:1991:SBA

[286] John Launchbury. Strictness and binding-time analyses: two for the price of one. ACM SIGPLAN Notices, 26(6):80–91, June 1991. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p80-launchbury/

.

Launchbury:1991:SBT

[287] John Launchbury. Strictness and binding-time analyses: two for the price of one. ACM SIGPLAN Notices, 26(6):80–91, June 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Consel:1991:PPE

[288] Charles Consel and Siau Cheng Khoo. Parameterized partial evaluation. ACM SIGPLAN Notices, 26(6):92–106, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p92-consel/

.

Venkatesh:1991:SAP

[289] G. A. Venkatesh. The semantic approach to program slicing. ACM SIG-PLAN Notices, 26(6):107–119, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p107-venkatesh/

.

Whitfield:1991:AGG

[290] Deborah Whitfield and Mary Lou Soffa. Automatic generation of global optimizers. ACM SIGPLAN Notices, 26(6):120–129, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p120-whitfield/

.

Chatterjee:1991:SAI

[291] Siddhartha Chatterjee, Guy E. Blelloch, and Allan L. Fisher. Size and access inference for data-parallel programs. ACM SIGPLAN Notices, 26(6):130–144, June 1991. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p130-chatterjee/

.

Bromley:1991:FAT

[292] Mark Bromley, Steven Heller, Tim Mc-Nerney, and Guy L. Steele, Jr. Fortran at ten gigaflops: the Connection Machine convolution compiler. ACM SIG-PLAN Notices, 26(6):145–156, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p145-bromley/

.

Bromley:1991:FTG

[293] Mark Bromley, Steven Heller, Tim Mc-Nerney, and Guy L. Steele, Jr. Fortran at ten gigaflops: the Connection Machine convolution compiler. ACM SIG-PLAN Notices, 26(6):145–156, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:1991:MPG

[294] Hans-J. Boehm, Alan J. Demers, and Scott Shenker. Mostly parallel garbage collection. ACM SIG-PLAN Notices, 26(6):157–164, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p157-boehm/

.

Goldberg:1991:TFG

[295] Benjamin Goldberg. Tag-free garbage collection for strongly typed programming languages. ACM SIGPLAN Notices, 26(6):165–176, June 1991.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldberg:1991:TGC

[296] Benjamin Goldberg. Tag-free garbage collection for strongly typed programming languages. ACM SIG-PLAN Notices, 26(6):165–176, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/113445/p165-goldberg/

.

Wilson:1991:ESG

[297] Paul R. Wilson, Michael S. Lam, and Thomas G. Moher. Effective “static-graph” reorganization to improve locality in garbage-collected systems. ACM SIGPLAN Notices, 26(6): 177–191, June 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1991:ESR

[298] Paul R. Wilson, Michael S. Lam, and Thomas G. Moher. Effective “staticgraph” reorganization to improve locality in garbage-collected systems. ACM SIGPLAN Notices, 26(6):177–191, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/113445/p177-wilson/

.

Callahan:1991:RAH

[299] David Callahan and Brian Koblenz. Register allocation via hierarchical graph coloring. ACM SIGPLAN Notices, 26(6):192–203, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p192-callahan/

.

Gao:1991:TPM

[300] Guang R. Gao, Yue-Bong Wong, and Qi Ning. A timed Petri-net model for fine-grain loop scheduling. ACM SIG-PLAN Notices, 26(6):204–218, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p204-gao/

.

Gao:1991:TPN

[301] Guang R. Gao, Yue-Bong Wong, and Qi Ning. A timed Petri-net model for fine-grain loop scheduling. ACM SIG-PLAN Notices, 26(6):204–218, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jain:1991:CSN

[302] Suneel Jain. Circular scheduling: a new technique to perform software pipelining. ACM SIGPLAN Notices, 26(6):219–228, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p219-jain/

.

Bradlee:1991:MSR

[303] David G. Bradlee, Robert R. Henry, and Susan J. Eggers. The Marion system for retargetable instruction scheduling. ACM SIGPLAN Notices, 26(6):229–240, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p229-bradlee/

.

Bernstein:1991:GIS

[304] David Bernstein and Michael Rodeh. Global instruction scheduling for superscalar machines. ACM SIG-PLAN Notices, 26(6):241–255, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p241-bernstein/

.

Proebsting:1991:LOC

[305] Todd A. Proebsting and Charles N. Fischer. Linear-time, optimal code scheduling for delayed-load architectures. ACM SIGPLAN Notices, 26(6):256–267, June 1991. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p256-proebsting/

.

Proebsting:1991:LTO

[306] Todd A. Proebsting and Charles N. Fischer. Linear-time, optimal code scheduling for delayed-load architectures. ACM SIGPLAN Notices, 26(6): 256–267, June 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freeman:1991:RTM

[307] Tim Freeman and Frank Pfenning. Refinement types for ML. ACM SIG-PLAN Notices, 26(6):268–277, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p268-freeman/

.

Cartwright:1991:ST

[308] Robert Cartwright and Mike Fagan. Soft typing. ACM SIG-PLAN Notices, 26(6):278–292, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p278-cartwright/

.

Reppy:1991:CHC

[309] John H. Reppy. CML: A higher concurrent language. ACM SIG-PLAN Notices, 26(6):293–305, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/113445/p293-reppy/

.

Jaffar:1991:MMH

[310] Joxan Jaffar, Spiro Michaylov, and Roland H. C. Yap. A methodology for managing hard constraints in CLP systems. ACM SIGPLAN Notices, 26(6):306–316, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

                                        proceedings/pldi/113445/p306-jaffar/

.

Fritzson:1991:GAD

[311] Peter Fritzson, Tibor Gyimothy, Mariam Kamkar, and Nahid Shahmehri. Generalized algorithmic debugging and testing. ACM SIG-PLAN Notices, 26(6):317–326, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/113445/p317-fritzson/

.

Bertot:1991:ODS

[312] Yves Bertot. Occurrences in debugger specifications. ACM SIG-PLAN Notices, 26(6):327–337, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/113445/p327-bertot/

.

Kishon:1991:MSF

[313] Amir Kishon, Paul Hudak, and Charles Consel. Monitoring semantics: a formal framework for specifying, implementing, and reasoning about execution monitors. ACM SIG-PLAN Notices, 26(6):338–352, June 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/113445/p338-kishon/

.

Nicolau:1991:OSP

[314] Alexandru Nicolau and Haigeng Wang. Optimal schedules for parallel prefix computation with bounded resources. ACM SIGPLAN Notices, 26(7):1–10, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:1991:PPT

[315] J. Allan Yang and Young il Choo. Parallel-program transformation using a metalanguage. ACM SIGPLAN Notices, 26(7):11–20, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bakewell:1991:MCP

[316] Hester Bakewell, Donna J. Quammen, and Pearl Y. Wang. Mapping concurrent programs to VLIW processors. ACM SIGPLAN Notices, 26(7):21–27, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:1991:UFS

[317] Lee-Chung Lu. A unified framework for systematic loop transformations. ACM SIGPLAN Notices, 26(7):28–38, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ancourt:1991:SPD

[318] Corinne Ancourt and François Irigoin. Scanning polyhedra with DO loops. ACM SIGPLAN Notices, 26(7):39–50, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krothapalli:1991:RRD

[319] V. P. Krothapalli and P. Sadayappan. Removal of redundant dependences in DOACROSS loops with constant dependences. ACM SIGPLAN Notices, 26(7):51–60, July 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ju:1991:EAD

[320] Dz-Ching Ju and Wai-Mee Ching. Exploitation of APL data parallelism on a shared-memory MIMD machine. ACM SIGPLAN Notices, 26(7):61–72, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hatcher:1991:PQC

[321] Philip J. Hatcher, Anthony J. Lapadula, Robert R. Jones, Michael J. Quinn, and Ray J. Anderson. A production-quality C* compiler for Hypercube multicomputers. ACM SIG-PLAN Notices, 26(7):73–82, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Costa:1991:APP

[322] Vítor Santos Costa, David H. D. Warren, and Rong Yang. Andorra I: a parallel Prolog system that transparently exploits both And-and or-parallelism. ACM SIGPLAN Notices, 26(7):83–93, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lam:1991:CGP

[323] Monica S. Lam and Martin C. Rinard. Coarse-grain parallel programming in Jade. ACM SIGPLAN Notices, 26 (7):94–105, July 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mellor-Crummey:1991:SRW

[324] John M. Mellor-Crummey and Michael L. Scott. Scalable reader-writer synchronization for shared-memory multiprocessors. ACM SIGPLAN Notices, 26 (7):106–113, July 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feeley:1991:DNR

[325] Michael J. Feeley, Brian N. Bershad, Jeffrey S. Chase, and Henry M. Levy. Dynamic node reconfiguration in a parallel-distributed environment. ACM SIGPLAN Notices, 26(7):114–121, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaRowe:1991:EOS

[326] Richard P. LaRowe, Jr., James T. Wilkes, and Carla S. Ellis. Exploiting operating system support for dynamic page placement on a NUMA shared memory multiprocessor. ACM SIGPLAN Notices, 26(7):122–132, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Netzer:1991:IAD

[327] Robert H. B. Netzer and Barton P. Miller. Improving the accuracy of data race detection. ACM SIG-PLAN Notices, 26(7):133–144, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choi:1991:RFR

[328] Jong-Deok Choi and Sang Lyul Min. Race Frontier: reproducing data races in parallel-program debugging. ACM SIGPLAN Notices, 26(7):145–154, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:1991:OPC

[329] David F. Bacon and Robert E. Strom. Optimistic parallelization of communicating sequential processes. ACM SIGPLAN Notices, 26(7):155–166, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kung:1991:PNC

[330] H. T. Kung, Peter Steenkiste, Marco Gubitoso, and Manpreet Khaira. Parallelizing a new class of large applications over high-speed networks. ACM SIGPLAN Notices, 26(7):167–177, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meier:1991:PPC

[331] Ulrike Meier and Rudolf Eigenmann. Parallelization and performance of Conjugate Gradient algorithms on the Cedar hierarchical-memory multiprocessor. ACM SIGPLAN Notices, 26 (7):178–188, July 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hollingsworth:1991:IAS

[332] Jeffrey K. Hollingsworth, R. Bruce Irvin, and Barton P. Miller. The integration of application and system based metrics in a parallel program performance tool. ACM SIG-PLAN Notices, 26(7):189–200, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malony:1991:EBP

[333] Allen D. Malony. Event-based performance perturbation: a case study. ACM SIGPLAN Notices, 26(7):201– 212, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balasundaram:1991:SPE

[334] Vasanth Balasundaram, Geoffrey Fox, Ken Kennedy, and Ulrich Kremer. A static performance estimator to guide data partitioning decisions. ACM SIG-PLAN Notices, 26(7):213–223, July 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Waters:1991:SVC

[335] Richard C. Waters. System validation via constraint modeling. ACM SIGPLAN Notices, 26(8):27–36, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakkinen:1991:ADE

[336] Markku Sakkinen. Another defense of enumerated types. ACM SIG-PLAN Notices, 26(8):37–41, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaplan:1991:PL

[337] Randy M. Kaplan. The IL programming language. ACM SIGPLAN Notices, 26(8):42–49, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heymann:1991:CAM

[338] Jürgen Heymann. A comprehensive analytical model for garbage collection algorithms. ACM SIGPLAN Notices, 26 (8):50–59, August 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kempe:1991:CTA

[339] J. Kempe, T. Lenz, B. Freitag, H. Shütz, and G. Specht. CL/TB an allegro common Lisp programming interface for TransBase. ACM SIG-PLAN Notices, 26(8):60–69, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fiterman:1991:MVP

[340] Charles Fiterman. Multiple view programming languages. ACM SIGPLAN Notices, 26(8):70–73, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:1991:TSL

[341] Diomidis Spinellis. Type-safe linkage for variables and functions. ACM SIGPLAN Notices, 26(8):74– 79, August 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://kerkis.math.
                                        aegean.gr/~dspin/pubs/jrnl/1991-
                                        SIGPLAN-CType/html/tsl.html.

Eckart:1991:CAS

[342] J. Dana Eckart. A cellular automata simulation system. ACM SIGPLAN Notices, 26(8):80–85, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McLaughlin:1991:TLS

[343] Robert McLaughlin. Thoughts on large scale programming projects. ACM SIGPLAN Notices, 26(8):86–89, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butcher:1991:LO

[344] Paul Butcher and Hussein Zedan. Lucinda — an overview. ACM SIG-PLAN Notices, 26(8):90–100, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ishikawa:1991:RFR

[345] Yutaka Ishikawa. Reflection facilities and realistic programming. ACM SIG-PLAN Notices, 26(8):101–110, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gruzlewski:1991:SCS

[346] Tadeusz Gruźlewski and Zbigniew Weiss. Semantic correctness of structural editing. ACM SIGPLAN Notices, 26(8):111–120, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wong:1991:DVF

[347] K. C. Wong. Detection of version features in distributed systems. ACM SIGPLAN Notices, 26(8):121–127, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuasa:1991:BIE

[348] Kei Yuasa. A browsing interface for S-expressions. ACM SIGPLAN Notices, 26(8):128–136, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacDonell:1991:RCD

[349] Stephen MacDonell. Reliance on correlation data for complexity metric use and validation. ACM SIGPLAN Notices, 26(8):137–144, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1991:SBM

[350] Henry G. Baker. Shallow binding makes functional arrays fast. ACM SIGPLAN Notices, 26(8):145–147, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonIssendorff:1991:TO

[351] Hermann von Issendorff. A theory of organization. ACM SIGPLAN Notices, 26(8):148–150, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Takano:1991:GPC

[352] Akihiko Takano. Generalized partial computation for a lazy functional language. ACM SIGPLAN Notices, 26(9): 1–11, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Romanenko:1991:IM

[353] Alexander Romanenko. Inversion and metacomputation. ACM SIGPLAN Notices, 26(9):12–22, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sahlin:1991:DAF

[354] Dan Sahlin. Determinacy analysis for full Prolog. ACM SIGPLAN Notices, 26(9):23–30, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lisper:1991:DSA

[355] Björn Lisper. Detecting static algorithms by partial evaluation. ACM SIGPLAN Notices, 26(9):31– 42, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hickey:1991:TPE

[356] Timothy J. Hickey and Donald A. Smith. Toward the partial evaluation of CLP languages. ACM SIGPLAN Notices, 26(9):43–51, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cortesi:1991:AIL

[357] Agostino Cortesi and Gilbert Filé. Abstract interpretation of logic programs: an abstract domain for groundness, sharing, freeness and compoundness analysis. ACM SIGPLAN Notices, 26(9):52–61, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1991:PEP

[358] Donald A. Smith. Partial evaluation of pattern matching in constraint logic programming languages. ACM SIG-PLAN Notices, 26(9):62–71, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sherman:1991:OEP

[359] David Sherman, Robert Strandh, and Irène Durand. Optimization of equational programs using partial evaluation. ACM SIGPLAN Notices, 26(9): 72–82, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sundaresh:1991:BIP

[360] R. S. Sundaresh. Building incremental programs using partial evaluation. ACM SIGPLAN Notices, 26(9): 83–93, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meyer:1991:TPE

[361] Uwe Meyer. Techniques for partial evaluation of imperative languages. ACM SIGPLAN Notices, 26(9):94– 105, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McNerney:1991:VCC

[362] Timothy S. McNerney. Verifying the correctness of compiler transformations on basic blocks using abstract interpretation. ACM SIGPLAN Notices, 26(9):106–115, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ariola:1991:SAP

[363] Zena M. Ariola. A syntactic approach to program transformations. ACM SIGPLAN Notices, 26(9):116– 129, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hannan:1991:STA

[364] John Hannan. Staging transformations for abstract machines. ACM SIG-PLAN Notices, 26(9):130–141, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Niel:1991:PBP

[365] Anne De Niel, Eddy Bevers, and Karel De Vlaminck. Program bifurcation for a polymorphically typed functional language. ACM SIGPLAN Notices, 26(9):142–153, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunt:1991:BTA

[366] Sebastian Hunt and David Sands. Binding time analysis: a new PERspective. ACM SIGPLAN Notices, 26 (9):154–165, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gomard:1991:GLV

[367] Carsten K. Gomard and Peter Sestoft. Globalization and live variables. ACM SIGPLAN Notices, 26(9):166– 177, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:1991:REA

[368] Young Gil Park and Benjamin Goldberg. Reference escape analysis: optimizing reference counting based on the lifetime of references. ACM SIG-PLAN Notices, 26(9):178–189, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shivers:1991:SSC

[369] Olin Shivers. The semantics of Scheme control-flow analysis. ACM SIG-PLAN Notices, 26(9):190–198, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klimov:1991:DSE

[370] Andrei V. Klimov. Dynamic specialization in extended functional language with monotone objects. ACM SIG-PLAN Notices, 26(9):199–210, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Khoo:1991:CIU

[371] Siau Cheng Khoo and R. S. Sundaresh. Compiling inheritance using partial evaluation. ACM SIG-PLAN Notices, 26(9):211–222, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holst:1991:PEF

[372] Carsten Kehler Holst and Darsten Krogh Gomard. Partial evaluation is fuller laziness. ACM SIGPLAN Notices, 26 (9):223–233, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:1991:MLN

[373] Robert Muller. M-LISP: its natural semantics and equational logic. ACM SIGPLAN Notices, 26(9):234– 242, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1991:PUB

[374] Torben Amtoft Hansen. Properties of unfolding-based meta-level systems. ACM SIGPLAN Notices, 26(9):243– 254, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:1991:TUL

[375] Philip Wadler. Is there a use for linear logic? ACM SIGPLAN Notices, 26(9):255–273, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Proietti:1991:SPT

[376] Maurizio Proietti and Alberto Pettorossi. Semantics preserving transformation rules for Prolog. ACM SIG-PLAN Notices, 26(9):274–284, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michaylov:1991:CPS

[377] Spiro Michaylov and Frank Pfenning. Compiling the polymorphic λ-calculus. ACM SIGPLAN Notices, 26(9):285– 296, September 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mason:1991:PTC

[378] Ian A. Mason and Carolyn L. Talcott. Program transformations for configuring components. ACM SIGPLAN Notices, 26(9):297–308, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gluck:1991:TMS

[379] Robert Glück. Towards multiple selfapplication. ACM SIGPLAN Notices, 26(9):309–320, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruf:1991:UTA

[380] Erik Ruf and Daniel Weise. Using types to avoid redundant specialization. ACM SIGPLAN Notices, 26(9): 321–333, September 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hook:1991:BRS

[381] Audrey A. Hook. Book review: The Software Factory: Managing Software Development and Maintenance (2nd. Ed) by James R. Johnson: (QED Information Sciences, Inc. 1991). ACM SIGPLAN Notices, 26(10):8, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Creak:1991:GFI

[382] G. Alan Creak. Garbage-further investigations. ACM SIGPLAN Notices, 26 (10):9–10, October 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kilov:1991:OCB

[383] Haim Kilov. Objects concepts and bibliography. ACM SIGPLAN Notices, 26(10):11–12, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeJacq:1991:FPO

[384] Jean Pierre LeJacq. Function preconditions in object oriented software. ACM SIGPLAN Notices, 26(10):13–18, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carpentieri:1991:CTT

[385] B. Carpentieri and G. Mou. Compiletime transformations and optimization of parallel Divide-and-Conquer algorithms. ACM SIGPLAN Notices, 26 (10):19–28, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1991:RCA

[386] Christopher W. Fraser. A retargetable compiler for ANSI C. ACM SIG-PLAN Notices, 26(10):29–43, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//storage.webhop.net/documents/
overview.pdf.

Tang:1991:CDC

[387] L. S. Tang. C++’s destructors can be destructive. ACM SIGPLAN Notices, 26(10):44–52, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Craske:1991:SOO

[388] Noel Craske. SNOOPS: An objectoriented language enhancement supporting dynamic program reconfiguration. ACM SIGPLAN Notices, 26(10): 53–63, October 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nelson:1991:COO

[389] Michael L. Nelson. Concurrency & object-oriented programming. ACM SIGPLAN Notices, 26(10):63–72, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miloucheva:1991:CSE

[390] Ilka Miloucheva and Hans Loeper. Compiler structure engineering with attribute grammars. ACM SIGPLAN Notices, 26(10):73–82, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1991:WMP

[391] C. K. Yuen. Which model of programming for LISP: sequential, functional or mixed? ACM SIGPLAN Notices, 26(10):83–92, October 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chambers:1991:MPO

[392] Craig Chambers and David Ungar. Making pure object-oriented languages practical. ACM SIGPLAN Notices, 26 (11):1–15, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sharma:1991:PGG

[393] Ravi Sharma and Mary Lou Soffa. Parallel generational garbage collection. ACM SIGPLAN Notices, 26(11):16– 32, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hayes:1991:UKO

[394] Barry Hayes. Using key object opportunism to collect old objects. ACM SIGPLAN Notices, 26(11):33– 46, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Helm:1991:IIR

[395] Richard Helm and Yoëlle S. Maarek. Integrating information retrieval and domain specific approaches for browsing and retrieval in object-oriented class libraries. ACM SIGPLAN Notices, 26(11):47–61, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gold:1991:PIC

[396] Eric Gold and Mary Beth Rosson. Portia: an instance-centered environment for Smalltalk. ACM SIGPLAN Notices, 26(11):62–74, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Konstan:1991:DGU

[397] Joseph A. Konstan and Lawrence A. Rowe. Developing a GUIDE using object-oriented programming. ACM SIGPLAN Notices, 26(11):75– 88, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guimaraes:1991:BGU

[398] Nuno Guimarães. Building generic user interface tools: an experience with multiple inheritance. ACM SIGPLAN Notices, 26(11):89–96, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gibbs:1991:CMA

[399] Simon Gibbs. Composite multimedia and active objects. ACM SIG-PLAN Notices, 26(11):97–112, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agrawal:1991:STC

[400] Rakesh Agrawal, Linda G. Demichiel, and Bruce G. Lindsay. Static type checking of multi-methods. ACM SIGPLAN Notices, 26(11):113–128, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghelli:1991:STS

[401] Giorgio Ghelli. A static type system for message passing. ACM SIGPLAN Notices, 26(11):129–145, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palsberg:1991:OOT

[402] Jens Palsberg and Michael I. Schwartzbach. Object-oriented type inference. ACM SIGPLAN Notices, 26(11):146– 161, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reed:1991:IMC

[403] David R. Reed, Marty Cagan, Ted Goldstein, and Barbara Moo. Issues in moving from C to C++. ACM SIGPLAN Notices, 26(11):163– 165, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

America:1991:FTO

[404] Pierre America, Derek Coleman, Roger Duke, Doug Lea, and Gary Leavens. Formal techniques for OO software development. ACM SIGPLAN Notices, 26(11):166–170, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hayes:1991:CMO

[405] Fiona Hayes and Derek Coleman. Coherent models for object-oriented analysis. ACM SIGPLAN Notices, 26(11): 171–183, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewis:1991:ESO

[406] John A. Lewis, Sallie M. Henry, Dennis G. Kafura, and Robert S. Schulman. An empirical study of the objectoriented paradigm and software reuse. ACM SIGPLAN Notices, 26(11):184– 196, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chidamber:1991:TMS

[407] Shyam R. Chidamber and Chris F. Kemerer. Towards a metrics suite for object oriented design. ACM SIGPLAN Notices, 26(11):197–211, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andreoli:1991:CFD

[408] Jean-Marc Andreoli and Remo Pareschi. Communication as fair distribution of knowledge. ACM SIGPLAN Notices, 26(11):212–229, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simmel:1991:KBS

[409] Sergui S. Simmel and Ivan Godard. The Kala basket: a semantic primitive unifying object transactions, access control, versions, and configurations. ACM SIGPLAN Notices, 26(11):230– 246, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yaseen:1991:EKO

[410] Rahim Yaseen, Stanley Y. W. Su, and Herman Lam. An extensible kernel object management system. ACM SIGPLAN Notices, 26(11):247– 263, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griss:1991:ESR

[411] Martin Griss, Sam S. Adams, Howard Baetjer, Jr., Brad J. Cox, and Adele Goldberg. The economics of software reuse. ACM SIGPLAN Notices, 26(11):264–270, November 1991.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hogg:1991:IAP

[412] John Hogg. Islands: aliasing protection in object-oriented languages. ACM SIGPLAN Notices, 26(11):271– 285, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilk:1991:EOO

[413] Michael R. Wilk. Equate: an objectoriented constraint solver. ACM SIGPLAN Notices, 26(11):286–298, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergstein:1991:OPC

[414] Paul L. Bergstein. Object-preserving class transformations. ACM SIGPLAN Notices, 26(11):299–313, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanson:1991:EDI

[415] Eric N. Hanson, Tina M. Harvey, and Mark A. Roth. Experiences in DBMS implementation using an object-oriented persistent programming language and a database toolkit. ACM SIGPLAN Notices, 26(11):314– 328, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1991:SSD

[416] Dan Benson and Greg Zick. Symbolic and spatial database for structural biology. ACM SIGPLAN Notices, 26(11):329–339, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jacobson:1991:ROS

[417] Ivar Jacobson and Fredrik Lindström. Reengineering of old systems to an object-oriented architecture. ACM SIGPLAN Notices, 26(11):340–350, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bobrow:1991:OA

[418] Daniel Bobrow, Carl Hewitt, Jean-François Perror, Reid Smith, and Howard Shrobe. OOP and AI. ACM SIGPLAN Notices, 26(11):351– 354, November 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hazeltine:1991:MTO

[419] Nelson Hazeltine, Tim Higenberg, Reed Philip, and David Taylor. Managing the transition to object-oriented technology. ACM SIGPLAN Notices, 26(11):355–358, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snyder:1991:HGY

[420] Alan Snyder. How to get your paper accepted at OOPSLA. ACM SIGPLAN Notices, 26(11):359–363, November 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malony:1991:MPP

[421] Allen D. Malony and Daniel A. Reed. Models for performance perturbation analysis. ACM SIGPLAN Notices, 26 (12):15–25, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choi:1991:BRR

[422] Jong-Deok Choi and Janice M. Stone. Balancing runtime and replay costs in a trace-and-replay system. ACM SIG-PLAN Notices, 26(12):26–35, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1991:CAA

[423] Victor Jon Griswold. Core algorithms for autonomous monitoring of distributed systems. ACM SIGPLAN Notices, 26(12):36–45, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zernik:1991:AWT

[424] Dror Zernik and Larry Rudolph. Animating work and time for debugging parallel programs foundation and experience. ACM SIGPLAN Notices, 26 (12):46–56, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kilpatrick:1991:CAS

[425] Carol Kilpatrick and Karsten Schwan. ChaosMON — application-specific monitoring and display of performance information for parallel and distributed systems. ACM SIGPLAN Notices, 26 (12):57–67, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Francioni:1991:DPP

[426] Joan M. Francioni, Larry Albright, and Jay Alan Jackson. Debugging parallel programs using sound. ACM SIG-PLAN Notices, 26(12):68–75, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Helmbold:1991:CRS

[427] David P. Helmbold and Charles E. Mc-Dowell. Computing reachable states of parallel programs. ACM SIGPLAN Notices, 26(12):76–84, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dinning:1991:DAA

[428] Anne Dinning and Edith Schonberg. Detecting access anomalies in programs with critical sections. ACM SIGPLAN Notices, 26(12):85–96, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Masticola:1991:MAP

[429] Stephen P. Masticola and Barbara G. Ryder. A model of Ada programs for static deadlock detection in polynomial time. ACM SIGPLAN Notices, 26(12): 97–107, December 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pineo:1991:DPC

[430] Patricia Prather Pineo and Mary Lou Soffa. Debugging parallelized code using code liberation techniques. ACM SIGPLAN Notices, 26(12):108–119, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tolmach:1991:DCE

[431] Andrew P. Tolmach and Andrew W. Appel. Debuggable concurrency extensions for Standard ML. ACM SIG-PLAN Notices, 26(12):120–131, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohn:1991:SLD

[432] Robert Cohn. Source level debugging of automatically parallelized code. ACM SIGPLAN Notices, 26(12):132– 143, December 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldberg:1991:RCG

[433] Arthur P. Goldberg, Ajei Gopal, Andy Lowry, and Rob Strom. Restoring consistent global states of distributed computations. ACM SIGPLAN Notices, 26(12):144–154, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spezialetti:1991:ARD

[434] Madalene Spezialetti. An approach to reducing delays in recognizing distributed event occurrences. ACM SIG-PLAN Notices, 26(12):155–166, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:1991:CDG

[435] Robert Cooper and Keith Marzullo. Consistent detection of global predicates. ACM SIGPLAN Notices, 26 (12):167–174, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gorlick:1991:FRA

[436] Michael M. Gorlick. The flight recorder: An architectural aid for system monitoring. ACM SIGPLAN Notices, 26(12):175–183, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruegge:1991:PPD

[437] Bernd Bruegge. A portable platform for distributed event environments. ACM SIGPLAN Notices, 26(12):184– 193, December 1991. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:1991:HAR

[438] David F. Bacon and Seth Copen Goldstein. Hardware-assisted replay of multiprocessor programs. ACM SIG-PLAN Notices, 26(12):194–206, December 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hook:1992:BRS

[439] Audrey A. Hook. Book review: Selecting an Ada Compilation System Edited by J. Dawes, M. J. Pickett, and A. Wearing: (Cambridge University Press 1990). ACM SIGPLAN Notices, 27 (1):13, January 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1992:BRR

[440] J. A. N. Lee. Book review: The Rationale for the Design of the Ada Programming Language, The Ada Companion Series by J. ichbiah, J. Barnes, R. Firth, and M. Woodger.: (Cambridge University Press, Cambridge, UK). ACM SIGPLAN Notices, 27 (1):14, January 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burshteyn:1992:UUS

[441] Boris Burshteyn. USSA: Universal Syntax and Semantics Analyzer. ACM SIGPLAN Notices, 27(1):42–60, January 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahan:1992:ARL

[442] W. Kahan. Analysis and refutation of the LCAS. ACM SIGPLAN Notices, 27 (1):61–74, January 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thirion:1992:EUC

[443] Bernard Thirion. Edison, a Unix and C friendly Rete based production system. ACM SIGPLAN Notices, 27(1): 75–84, January 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:1992:OOS

[444] Pei-Chi Wu and Feng-Jian Wang. An object-oriented specification for compiler. ACM SIGPLAN Notices, 27(1): 85–94, January 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1992:CAB

[445] H. G. Baker. Computing A ∗ B (mod N) efficiently in ANSI C. ACM SIGPLAN Notices, 27(1):95–98, January 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1992:CME

[446] Henry G. Baker. Computing A*B (mod N) efficiently in ANSI C. ACM SIG-PLAN Notices, 27(1):95–98, January 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ince:1992:APC

[447] D. C. Ince. Arrays and pointers considered harmful. ACM SIGPLAN Notices, 27(1):99–104, January 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Furtado:1992:AGQ

[448] Antonio L. Furtado. Analogy by generalization — and the quest of the grail. ACM SIGPLAN Notices, 27(1): 105–113, January 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nourani:1992:PMS

[449] Cyrus F. Nourani. Parallel module specification on SLPX. ACM SIG-PLAN Notices, 27(1):114–115, January 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1992:BRA

[450] Alan Wexelblat. Book review: The Art of the Meta-Object Protocol by Gregor Kiczales, Jim des Rivieres, and Daniel G. Bobrow: (MIT Press, 1991). ACM SIGPLAN Notices, 27(2):9, February 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michael:1992:BRU

[451] Bret Michael. Book review: A Unifying Framework for Structured Analysis and Design Models by T. H. Tse: (Cambridge University Press, 1991). ACM SIGPLAN Notices, 27(2):10–11, February 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hernandez:1992:RME

[452] Juan Hernández and Juan Antonio Sanchez. RT-MODULA2: an embedded in MODULA2 language for writing concurrent and real time programs. ACM SIGPLAN Notices, 27(2): 26–36, February 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arab:1992:EPC

[453] Mouloud Arab. Enhancing program comprehension: formatting and documenting. ACM SIGPLAN Notices, 27(2):37–46, February 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gwinn:1992:OOP

[454] J. M. Gwinn. Object-oriented programs in realtime. ACM SIGPLAN Notices, 27(2):47–56, February 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryan:1992:LCF

[455] Steven Ryan. Linear control flow analysis. ACM SIGPLAN Notices, 27(2): 57–64, February 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Placer:1992:IDA

[456] John Placer. Integrating destructive assignment and lazy evaluation in the multiparadigm language G-2. ACM SIGPLAN Notices, 27(2):65–74, February 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caringal:1992:FIQ

[457] Rizaldo B. Caringal and Phan Minh Dung. A FORTRAN IV to Quick-BASIC translator. ACM SIGPLAN Notices, 27(2):75–87, February 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parr:1992:PRM

[458] T. J. Parr, H. G. Dietz, and W. E. Cohen. PCCTS reference manual: version 1.00. ACM SIGPLAN Notices, 27(2): 88–165, February 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1992:CSC

[459] Henry G. Baker, Jr. CONS should not CONS its arguments, or, A lazy alloc is a smart alloc. ACM SIG-PLAN Notices, 27(3):24–34, March 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Submitted to Communications of the ACM.

Salomon:1992:FDP

[460] Daniel J. Salomon. Four Dimensions of programming-language independence. ACM SIGPLAN Notices, 27 (3):35–53, March 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mamone:1992:ESM

[461] Salvatore Mamone. Empirical study of motivation in a entry level programming course. ACM SIGPLAN Notices, 27(3):54–60, March 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Watt:1992:CSC

[462] J. M. Watt. A compact syntax chart for Ada. ACM SIGPLAN Notices, 27 (3):61–65, March 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1992:TRT

[463] Henry G. Baker. The treadmill: realtime garbage collection without motion sickness. ACM SIGPLAN Notices, 27 (3):66–70, March 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anastasio:1992:OCL

[464] Thomas A. Anastasio and William W. Carlson. An observation on the C library procedure random(). ACM SIGPLAN Notices, 27(3):71–74, March 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michael:1992:BRD

[465] James Bret Michael. Book review: Discrete Mathematics: An Introduction for Software Engineers, by Mike Piff: (Cambridge University Press, 1991). ACM SIGPLAN Notices, 27 (4):10, April 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gladwell:1992:BRF

[466] Ian Gladwell. Book review: Fortran 90 Explained by M. Metcalf and J. Reid: (Oxford Science). ACM SIG-PLAN Notices, 27(4):11, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:1992:SRL

[467] Pei-Chi Wu, Feng-Jian Wang, and Kai-Ru Young. Scanning regular languages by dual finite automata. ACM SIGPLAN Notices, 27(4):12–16, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Angele:1992:MAC

[468] Jürgen Angele and Detlef Küpper. Modula-2 an alternative to C? ACM SIGPLAN Notices, 27(4):17–26, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Young:1992:TMO

[469] Norman Young. Two models of objectoriented programming and the Common Lisp Object System. ACM SIGPLAN Notices, 27(4):27–36, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gauthier:1992:NFP

[470] Michel Gauthier. Noised or filtered programming? (about enumerated types). ACM SIGPLAN Notices, 27(4):37–40, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ipser:1992:ELD

[471] Edward A. Ipser, Jr. Exploratory language design. ACM SIGPLAN Notices, 27(4):41–50, April 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fischer:1992:TFC

[472] Ronald Fischer. Two flaws of the current C++ language definition. ACM SIGPLAN Notices, 27(4):51–54, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryan:1992:LEB

[473] Steven Ryan. Linear expression bounding. ACM SIGPLAN Notices, 27(4): 55–58, April 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryan:1992:LDF

[474] Steven Ryan. Linear data flow analysis. ACM SIGPLAN Notices, 27(4):59–67, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1992:BFO

[475] Christopher W. Fraser, Robert R. Henry, and Todd A. Proebsting. BURG: fast optimal instruction selection and tree parsing. ACM SIG-PLAN Notices, 27(4):68–76, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vandenBosch:1992:BSE

[476] Peter N. van den Bosch. A bibliography on syntax error handling in context free languages. ACM SIGPLAN Notices, 27 (4):77–86, April 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Appel:1992:PMS

[477] Andrew W. Appel. Is POPL mathematics or science? ACM SIGPLAN Notices, 27(4):87–89, April 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hudak:1992:GIH

[478] Paul Hudak and Joseph H. Fasel. A gentle introduction to Haskell. ACM SIGPLAN Notices, 27(5):1–52, May 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hudak:1992:RPL

[479] Paul Hudak, Simon Peyton Jones, Philip Wadler, Brian Boutel, Jon Fairbairn, Joseph Fasel, María M. Guzmán, Kevin Hammond, John Hughes, Thomas Johnsson, Dick Kieburtz, Rishiyur Nikhil, Will Partain, and John Peterson. Report on the programming language Haskell: a non-strict, purely functional language (version 1.2). ACM SIGPLAN Notices, 27(5):Ri–Rx, R1–R163, May 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michael:1992:BRV

[480] James Bret Michael. Book review: Verifying Temporal Properties of Systems, by Julian Charles Bradfield: (Birkhäuser Boston, Cambridge, Massachusetts, 1992). ACM SIGPLAN Notices, 27(6):8, June 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meeson:1992:BRIa

[481] Reginald Meeson. Book review: Implementation of Functional Multiprocessors by O. Boudillet, J. P. Gupta, and S. C. Winter: (John Wiley & Sons Inc. 1991). ACM SIGPLAN Notices, 27(6): 9, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casimir:1992:RPD

[482] Rommert J. Casimir. Real programmers don’t use spreadsheets. ACM SIGPLAN Notices, 27(6):10–16, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cargille:1992:BWT

[483] Jon Cargille and Barton P. Miller. Binary wrapping: a technique for instrumenting object code. ACM SIG-PLAN Notices, 27(6):17–18, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavenworth:1992:PIR

[484] Burt Leavenworth. Proxy: an interpreter for rapid prototyping. ACM SIGPLAN Notices, 27(6):19–28, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sabatella:1992:LEC

[485] Marc Sabatella. Lazy evaluation of C++ static constructors. ACM SIG-PLAN Notices, 27(6):29–36, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coolsaet:1992:QIP

[486] K. Coolsaet. A quick introduction to the programming language MIKE. ACM SIGPLAN Notices, 27(6):37–46, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carrasco:1992:NLM

[487] José M. García Carrasco. A new language for multicomputer programming. ACM SIGPLAN Notices, 27(6):47–53, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prather:1992:SC

[488] Ronald E. Prather. The semantics of CAT. ACM SIGPLAN Notices, 27 (6):54–63, June 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

onExtensions:1992:HWC

[489] Corporate X3J16 Working Group on Extensions. How to write a C++ language extension proposal for ANSI-X3J16/ISO-WG21. ACM SIG-PLAN Notices, 27(6):64–71, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frost:1992:GAG

[490] Richard A. Frost. Guarded attribute grammars: top down parsing and left recursive productions. ACM SIG-PLAN Notices, 27(6):72–75, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ponder:1992:PCH

[491] Carl Ponder and Bill Bush. Polymorphism considered harmful. ACM SIG-PLAN Notices, 27(6):76–79, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mody:1992:UTD

[492] R. P. Mody. On understanding type declarations in C. ACM SIGPLAN Notices, 27(6):80–83, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakkinen:1992:WEA

[493] Markku Sakkinen. Where is the evidence against arrays and pointers? ACM SIGPLAN Notices, 27(6):84–85, June 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brooks:1992:NAD

[494] Gary Brooks, Gilbert J. Hansen, and Steve Simmons. A new approach to debugging optimized code. ACM SIGPLAN Notices, 27(7):1– 11, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p1-brooks/

.

Sosic:1992:DTP

[495] Rok Sosič. Dynascope. A tool for program directing. ACM SIGPLAN Notices, 27(7):12–21, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p12-sosic/

.

Ramsey:1992:RD

[496] Norman Ramsey and David R. Hanson. A retargetable debugger. ACM SIGPLAN Notices, 27 (7):22–31, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//storage.webhop.net/documents/
retargetable.pdf;   http://www.

acm.org:80/pubs/citations/proceedings/
pldi/143095/p22-ramsey/.

Holzle:1992:DOC

[497] Urs Hölzle, Craig Chambers, and David Ungar. Debugging optimized code with dynamic deoptimization. ACM SIGPLAN Notices, 27 (7):32–43, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p32-holzle/

.

Ogata:1992:DIH

[498] Kazuhiro Ogata, Satoshi Kurihara, Mikio Inari, and Norihisa Doi. The design and implementation of HoME. ACM SIGPLAN Notices, 27(7):44– 54, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p44-ogata/

.

Jagannathan:1992:CSC

[499] Suresh Jagannathan and Jim Philbin. A customizable substrate for concurrent languages. ACM SIGPLAN Notices, 27(7):55–67, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p55-jagannathan/

.

Wortman:1992:CCM

[500] David B. Wortman and Michael D. Junkin. A concurrent compiler for Modula-2+. ACM SIGPLAN Notices, 27(7):68–81, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p68-wortman/

.

Tjiang:1992:STB

[501] Steven W. K. Tjiang and John L. Hennessy. Sharlit — a tool for building optimizers. ACM SIGPLAN Notices, 27(7):82–93, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p82-tjiang/

.

Chen:1992:PFC

[502] Marina Chen and James Cowie. Prototyping Fortran-90 compilers for massively parallel machines. ACM SIG-PLAN Notices, 27(7):94–105, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p94-chen/

.

Tan:1992:CDA

[503] Jichang Tan and I-Peng Lin. Compiling dataflow analysis of logic programs. ACM SIGPLAN Notices, 27 (7):106–115, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

proceedings/pldi/143095/p106-tan/

.

Park:1992:EAL

[504] Young Gil Park and Benjamin Goldberg. Escape analysis on lists. ACM SIGPLAN Notices, 27(7):116– 127, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p116-park/

.

Jaffar:1992:AMC

[505] Joxan Jaffar, Peter J. Stuckey, Spiro Michaylov, and Roland H. C. Yap. An abstract machine for CLP(R). ACM SIGPLAN Notices, 27(7):128– 139, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p128-jaffar/

.

Pugh:1992:EFD

[506] William Pugh and David Wonnacott. Eliminating false data dependences using the Omega test. ACM SIGPLAN Notices, 27(7):140– 151, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p140-pugh/

.

Maslov:1992:DEW

[507] Vadim Maslov. Delinearization: an efficient way to break multiloop dependence equations. ACM SIG-PLAN Notices, 27(7):152–161, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p152-maslov/

.

Wolfe:1992:BIV

[508] Michael Wolfe. Beyond induction variables. ACM SIGPLAN Notices, 27 (7):162–174, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p162-wolfe/

.

Sarkar:1992:GFI

[509] Vivek Sarkar and Radhika Thekkath. A general framework for iterationreordering loop transformations. ACM SIGPLAN Notices, 27(7):175–187, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p175-sarkar/

.

Hanxleden:1992:RSC

[510] Reinhard v. Hanxleden and Ken Kennedy. Relaxing SIMD control flow constraints using loop transformations. ACM SIGPLAN Notices, 27(7):188– 199, July 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonHanxleden:1992:RSC

[511] Reinhard von Hanxleden and Ken Kennedy. Relaxing SIMD control flow constraints using loop transformations. ACM SIGPLAN Notices, 27 (7):188–199, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p188-v__
hanxleden/.

Lucco:1992:DSM

[512] Steven Lucco. A dynamic scheduling method for irregular parallel programs. ACM SIGPLAN Notices, 27 (7):200–211, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p200-lucco/

.

Dhamdhere:1992:HAL

[513] Dhananjay M. Dhamdhere, Barry K. Rosen, and F. Kenneth Zadeck. How to analyze large programs efficiently and informatively. ACM SIGPLAN Notices, 27(7):212–223, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p212-dhamdhere/

.

Knoop:1992:LCM

[514] Jens Knoop, Oliver Rüthing, and Bernhard Steffen. Lazy code motion. ACM SIGPLAN Notices, 27(7):

224–234, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p224-knoop/

.

Landi:1992:SAA

[515] William Landi and Barbara G. Ryder. A safe approximate algorithm for interprocedural aliasing. ACM SIGPLAN Notices, 27(7):235– 248, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p235-landi/

.

Hendren:1992:ARP

[516] Laurie J. Hendren, Joseph Hummel, and Alexandru Nicolau. Abstractions for recursive pointer data structures. improving the analysis and transformation of imperative programs. ACM SIGPLAN Notices, 27 (7):249–260, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p249-hendren/

.

Hoover:1992:AIC

[517] Roger Hoover. Alphonse: Incremental computation as a programming abstraction. ACM SIGPLAN Notices, 27 (7):261–272, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p261-hoover/

.

Diwan:1992:CSG

[518] Amer Diwan, Eliot Moss, and Richard Hudson. Compiler support for garbage collection in a statically typed language. ACM SIGPLAN Notices, 27 (7):273–282, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p273-diwan/

.

Rau:1992:RAS

[519] B. R. Rau, M. Lee, P. P. Tirumalai, and M. S. Schlansker. Register allocation for software pipelined loops. ACM SIGPLAN Notices, 27 (7):283–299, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p283-rau/

.

Proebsting:1992:PRA

[520] Todd A. Proebsting and Charles N. Fischer. Probabilistic register allocation. ACM SIGPLAN Notices, 27(7): 300–310, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p300-proebsting/

.

Briggs:1992:R

[521] Preston Briggs, Keith D. Cooper, and Linda Torczon. Rematerialization. ACM SIGPLAN Notices, 27(7): 311–321, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/143095/p311-briggs/

.

Mueller:1992:AUJ

[522] Frank Mueller and David B. Whalley. Avoiding unconditional jumps by code replication. ACM SIGPLAN Notices, 27(7):322–330, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/143095/p322-mueller/

.

Proebsting:1992:SEB

[523] Todd A. Proebsting. Simple and efficient BURS table generation. ACM SIGPLAN Notices, 27(7):331– 340, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/143095/p331-proebsting/

.

Granlund:1992:EBU

[524] Torbjörn Granlund and Richard Kenner. Eliminating branches using a superoptimizer and the GNU C compiler. ACM SIGPLAN Notices, 27(7): 341–352, July 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/143095/p341-granlund/

.

Meeson:1992:BRM

[525] Reginald Meeson. Book review: ML for the Working Programmer by Laurence C. Paulson: (Cambridge University Press, 429 pp., 1991). ACM SIG-PLAN Notices, 27(8):14, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cardelli:1992:MLD

[526] Luca Cardelli, James Donahue, Lucille Glassman, Mick Jordan, Bill Kalsow, and Greg Nelson. Modula-3 language definition. ACM SIGPLAN Notices, 27 (8):15–42, August 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamel:1992:ISC

[527] Lutz H. Hamel. Industrial strength compiler construction with equations. ACM SIGPLAN Notices, 27(8):43–50, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lowry:1992:HLO

[528] Andy Lowry. The Hermes language in outline form. ACM SIGPLAN Notices, 27(8):51–70, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1992:HNL

[529] C. K. Yuen. Hamming numbers, lazy evaluation, and eager disposal. ACM SIGPLAN Notices, 27(8):71–75, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plat:1992:OIV

[530] Nico Plat and Peter Gorm Larsen. An overview of the ISO/VDM-SL standard. ACM SIGPLAN Notices, 27(8): 76–82, August 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mossenbock:1992:TSS

[531] Hanspeter Mössenböck. Treating statement sequences as block objects. ACM SIGPLAN Notices, 27(8):83–86, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1992:SPE

[532] Paul W. Abrahams. Software patents: an example of the threat. ACM SIG-PLAN Notices, 27(8):87–88, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1992:LLL

[533] Henry G. Baker. Lively linear Lisp: “look ma, no garbage!”. ACM SIG-PLAN Notices, 27(8):89–98, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eckart:1992:CAS

[534] J. Dana Eckart. A cellular automata simulation system: version 2.0. ACM SIGPLAN Notices, 27(8):99–106, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eckart:1992:CLR

[535] J. Dana Eckart. Cellang 2.0: language reference manual. ACM SIG-PLAN Notices, 27(8):107–112, August 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burrows:1992:LDC

[536] Michael Burrows, Charles Jerian, Butler Lampson, and Timothy Mann. On-line data compression in a logstructured file system. ACM SIG-PLAN Notices, 27(9):2–9, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p2-burrows/

.

Baker:1992:NVM

[537] Mary Baker, Satoshi Asami, Etienne Deprit, John Ousterhout, and Margo Seltzer. Non-volatile memory for fast, reliable file systems. ACM SIG-PLAN Notices, 27(9):10–22, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p10-baker/

.

Holland:1992:PDC

[538] Mark Holland and Garth A. Gibson. Parity declustering for continuous operation in redundant disk arrays. ACM SIGPLAN Notices, 27 (9):23–35, September 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p23-holland/

.

Rogers:1992:SSS

[539] Anne Rogers and Kai Li. Software support for speculative loads. ACM SIG-PLAN Notices, 27(9):38–50, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p38-rogers/

.

Chen:1992:RML

[540] Tien-Fu Chen and Jean-Loup Baer. Reducing memory latency via nonblocking and prefetching caches. ACM SIGPLAN Notices, 27(9): 51–61, September 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p51-chen/

.

Mowry:1992:DEC

[541] Todd C. Mowry, Monica S. Lam, and Anoop Gupta. Design and evaluation of a compiler algorithm for prefetching. ACM SIGPLAN Notices, 27(9):62–73, September 1992.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p62-mowry/

.

Pan:1992:IAD

[542] Shien-Tai Pan, Kimming So, and Joseph T. Rahmeh. Improving the accuracy of dynamic branch prediction using branch correlation. ACM SIG-PLAN Notices, 27(9):76–84, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p76-pan/

.

Fisher:1992:PCB

[543] Joseph A. Fisher and Stefan M. Freudenberger. Predicting conditional branch directions from previous runs of a program. ACM SIG-PLAN Notices, 27(9):85–95, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p85-fisher/

.

Anderson:1992:HSS

[544] Thomas E. Anderson, Susan S. Owicki, James B. Saxe, and Charles P. Thacker. High speed switch scheduling for local area networks. ACM SIG-PLAN Notices, 27(9):98–110, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/asplos/143365/p98-anderson/

.

Henry:1992:TCP

[545] Dana S. Henry and Christopher F. Joerg. A tightly-coupled processornetwork interface. ACM SIGPLAN Notices, 27(9):111–122, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p111-
                                        henry/.

Wheeler:1992:CMV

[546] Bob Wheeler and Brian N. Bershad. Consistency management for virtually indexed caches. ACM SIG-PLAN Notices, 27(9):124–136, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p124-
                                        wheeler/.

Chiueh:1992:EAT

[547] Tzi cker Chiueh and Randy H. Katz. Eliminating the address translation bottleneck for physical address cache. ACM SIGPLAN Notices, 27(9):137–148, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p137-
                                        chiueh/.

Veenstra:1992:PEO

[548] Jack E. Veenstra and Robert J. Fowler. A performance evaluation of optimal hybrid cache coherency protocols. ACM SIGPLAN Notices, 27(9):149–160, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p149-
veenstra/.

Torrellas:1992:CCS

[549] Josep Torrellas, Anoop Gupta, and John Hennessy. Characterizing the caching and synchronization performance of a multiprocessor operating system. ACM SIGPLAN Notices, 27(9):162–174, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p162-
torrellas/.

Koldinger:1992:ASS

[550] Eric J. Koldinger, Jeffrey S. Chase, and Susan J. Eggers. Architecture support for single address space operating systems. ACM SIGPLAN Notices, 27(9):175–186, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p175-
koldinger/.

Harty:1992:ACP

[551] Kieran Harty and David R. Cheriton. Application-controlled physical memory using external pagecache management. ACM SIGPLAN Notices, 27(9):187–197, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p187-
                                        harty/.

Wahbe:1992:EDB

[552] Robert Wahbe. Efficient data breakpoints. ACM SIGPLAN Notices, 27 (9):200–212, September 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p200-
                                        wahbe/.

Andrews:1992:MCC

[553] Kristy Andrews and Duane Sand. Migrating a CISC computer family onto RISC via object code translation. ACM SIGPLAN Notices, 27 (9):213–222, September 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p213-
                                        andrews/.

Bershad:1992:FME

[554] Brian N. Bershad, David D. Redell, and John R. Ellis. Fast mutual exclusion for uniprocessors. ACM SIG-PLAN Notices, 27(9):223–233, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p223-
                                        bershad/.

Mahlke:1992:SSV

[555] Scott A. Mahlke, William Y. Chen, Wen mei W. Hwu, B. Ramakrishna Rau, and Michael S. Schlansker. Sentinel scheduling for VLIW and superscalar processors. ACM SIG-PLAN Notices, 27(9):238–247, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p238-
mahlke/.

Smith:1992:ESP

[556] Michael D. Smith, Mark Horowitz, and Monica S. Lam. Efficient superscalar performance through boosting. ACM SIGPLAN Notices, 27 (9):248–259, September 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p248-
smith/.

Hill:1992:CSM

[557] Mark D. Hill, James R. Larus, Steven K. Reinhardt, and David A. Wood. Cooperative shared memory: software and hardware for scalable multiprocessor. ACM SIG-PLAN Notices, 27(9):262–273, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/143365/p262-
hill/.

Kubiatowicz:1992:CWV

[558] John Kubiatowicz, David Chaiken, and Anant Agarwal. Closing the window of vulnerability in multiphase memory transactions. ACM SIG-PLAN Notices, 27(9):274–284, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p274-
                                        kubiatowicz/.

Li:1992:ANL

[559] Wei Li and Keshav Pingali. Access normalization: loop restructuring for NUMA compilers. ACM SIG-PLAN Notices, 27(9):285–295, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/143365/p285-
                                        li/.

Cook:1992:ISS

[560] William R. Cook. Interfaces and specifications for the Smalltalk-80 collection classes. ACM SIGPLAN Notices, 27 (10):1–15, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ducournau:1992:MCR

[561] R. Ducournau, M. Habib, M. Huchard, and Marie-Laure Mugnier. Monotonic conflict resolution mechanisms for inheritance. ACM SIGPLAN Notices, 27(10):16–24, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ossher:1992:CIH

[562] Harold Ossher and William H. Harrison. Combination of inheritance hierarchies. ACM SIGPLAN Notices, 27(10): 25–40, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCollough:1992:RPT

[563] Paul McCollough, Bob Atkinson, Adele Goldberg, Martin Griss, and John Morrison. Reuse (panel): truth or fiction. ACM SIGPLAN Notices, 27(10): 41–44, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCullough:1992:RTF

[564] Paul McCullough, Robert G. Atkinson, Adele Goldberg, Martin L. Griss, and John Morrison. Reuse: Truth or fiction (panel). ACM SIGPLAN Notices, 27(10):41–44, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deChampeaux:1992:POO

[565] Dennis de Champeaux, Doug Lea, and Penelope Faure. The process of object-oriented design. ACM SIG-PLAN Notices, 27(10):45–62, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1992:DFU

[566] Ralph E. Johnson. Documenting frameworks using patterns. ACM SIG-PLAN Notices, 27(10):63–76, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:1992:WCS

[567] Chamond Liu, Stephen Goetze, and Bill Glynn. What contributes to successful object-oriented learning? ACM SIGPLAN Notices, 27(10):77–86, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bulman:1992:OLP

[568] David Bulman, S. Tucker Taft, Bertrand Meyer, Greg Nelson, and Mike Kilian. OOP in languages providing strong, static typing (panel). ACM SIGPLAN Notices, 27(10):87, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taft:1992:PPO

[569] S. Tucker Taft. Position paper for OOPSLA ’92 panel on OOP for languages based on strong, static typing. ACM SIGPLAN Notices, 27 (10):88, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meyer:1992:EST

[570] Bertrand Meyer. Ensuring strong typing in an object-oriented language (abstract). ACM SIGPLAN Notices, 27 (10):89–90, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kilian:1992:TWW

[571] Michael F. Kilian. Trellis: what we have learned from a strongly typed language. ACM SIGPLAN Notices, 27 (10):91, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hosking:1992:CPE

[572] Antony L. Hosking, J. Eliot B. Moss, and Darko Stefanovic. A comparative performance evaluation of write barrier implementations. ACM SIGPLAN Notices, 27(10):92–109, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andre:1992:OMS

[573] Pascal André and Jean-Claude Royer. Optimizing method search with lookup caches and incremental coloring. ACM SIGPLAN Notices, 27(10):110–126, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Royer:1992:OMS

[574] Jean-Claude Royer. Optimizing method search with lookup caches and incremental coloring. ACM SIGPLAN Notices, 27(10):110–126, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Masuhara:1992:OOC

[575] Hidehiko Masuhara, Satoshi Matsuoka, Takuo Watanabe, and Akinori Yonezawa. Object-oriented concurrent reflective languages can be implemented efficiently. ACM SIGPLAN Notices, 27(10):127–144, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeNatale:1992:RMC

[576] Rick DeNatale, Grady Booch, Peter Coad, Dave A. Thomas, and John Tibbets. The role of methods and CASE in OO development (panel). ACM SIGPLAN Notices, 27(10):145, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wangler:1992:VOM

[577] Michael F. Wangler and Peeter Hansen. Visualizing objects: Methods for exploring human computer interaction concepts. ACM SIGPLAN Notices, 27 (10):146–153, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calder:1992:OOI

[578] Paul R. Calder and Mark A. Linton. The object-oriented implementation of a document editor. ACM SIGPLAN Notices, 27(10):154–165, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eggenschwiler:1992:EUO

[579] Thomas Eggenschwiler and Erich Gamma. ET++SwapsManager: using object technology in the financial engineering domain. ACM SIGPLAN Notices, 27(10):166–177, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eggenschwiler:1992:SMU

[580] Thomas Eggenschwiler and Erich Gamma. ET++ swaps manager: Using object technology in the financial engineering domain. ACM SIGPLAN Notices, 27(10):166–177, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walsh:1992:PDD

[581] James F. Walsh. Preliminary defect data from the iterative development of a large C++ program (experience report). ACM SIGPLAN Notices, 27(10): 178–183, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Myers:1992:DPP

[582] Brad A. Myers, Dario A. Giuse, and Brad T. Vander Zanden. Declarative programming in a prototype-instance system: Object-oriented programming without writing methods. ACM SIG-PLAN Notices, 27(10):184–200, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dony:1992:PBL

[583] Christophe Dony, Jacques Malenfant, and Pierre Cointe. Prototype-based languages: From a new taxonomy to constructive proposals and their validation. ACM SIGPLAN Notices, 27(10): 201–217, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horn:1992:CPB

[584] Bruce Horn. Constraint patterns as a basis for object-oriented programming. ACM SIGPLAN Notices, 27(10): 218–233, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yelland:1992:ECF

[585] Phillip M. Yelland. Experimental classification facilities for Smalltalk. ACM SIGPLAN Notices, 27(10):235– 246, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feeley:1992:DSM

[586] Michael J. Feeley and Henry M. Levy. Distributed shared memory with versioned objects. ACM SIGPLAN Notices, 27(10):247–262, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Richardson:1992:CEF

[587] Joel Richardson, Peter Schwarz, and Luis-Felipe Cabrera. CACL: Efficient fine-grained protection for objects. ACM SIGPLAN Notices, 27(10): 263–275, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Takashio:1992:DOO

[588] Kazunori Takashio and Mario Tokoro. DROL: An object-oriented programming language for distributed realtime systems. ACM SIGPLAN Notices, 27(10):276–294, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Constantine:1992:EOH

[589] Larry L. Constantine, Ivar Jacobson, Meilir Page-Jones, John Palmer, and Steven Weiss. From events to objects: The heresy of event-orientation in a world of objects (panel). ACM SIG-PLAN Notices, 27(10):295–297, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Constantine:1992:EOP

[590] Larry L. Constantine, Ivar Jacobson, Meilir Page-Jones, John Palmer, and Steven Weiss. From events to objects (panel): the heresy of event-orientation in a world of objects. ACM SIGPLAN Notices, 27(10):295–297, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stacy:1992:ESI

[591] Webb Stacy, Richard Helm, Gail E. Kaiser, and Bertrand Meyer. Ensuring semantic integrity of reusable objects (panel). ACM SIGPLAN Notices, 27(10):298–302, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ishikawa:1992:CMA

[592] Yutaka Ishikawa. Communication mechanism on autonomous objects. ACM SIGPLAN Notices, 27(10):303– 314, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Satoh:1992:FRT

[593] Ichiro Satoh and Mario Tokoro. A formalism for real-time concurrent objectoriented computing. ACM SIGPLAN Notices, 27(10):315–326, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lohr:1992:CA

[594] Klaus-Peter Löhr. Concurrency annotations. ACM SIGPLAN Notices, 27 (10):327–340, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aksit:1992:OOO

[595] Mehmet Aksit and Lodewijk Bergmans. Obstacles in object-oriented software development. ACM SIGPLAN Notices, 27(10):341–358, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruegge:1992:OOS

[596] Bernd Bruegge, Jim Blythe, Jeffrey Jackson, and Jeff Shufelt. Objectoriented system modeling with OMT. ACM SIGPLAN Notices, 27(10):359– 376, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deChampeaux:1992:CSO

[597] Dennis de Champeaux, Al Anderson, and Ed Feldhousen. Case study of object-oriented software development. ACM SIGPLAN Notices, 27(10):377– 391, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1992:OOM

[598] Peter Wegner, William L. Scherlis, James M. Purtilo, David C. Luckham, and Ralph E. Johnson. Object-oriented megaprogramming (panel). ACM SIG-PLAN Notices, 27(10):392–396, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chase:1992:LSO

[599] Jeffrey S. Chase, Henry M. Levy, Edward D. Lazowska, and Miche Baker-Harvey. Lightweight shared objects in a 64-bit operating system. ACM SIG-PLAN Notices, 27(10):397–413, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yokote:1992:ARO

[600] Yasuhiko Yokote. The Apertos reflective operating system: The concept and its implementation. ACM SIG-PLAN Notices, 27(10):414–434, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiczales:1992:IDD

[601] Gregor Kiczales and John Lamping. Issues in the design and documentation of class libraries. ACM SIGPLAN Notices, 27(10):435–451, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiczales:1992:IDS

[602] Gregor Kiczales and John Lamping. Issues in the design and specification of class libraries. ACM SIGPLAN Notices, 27(10):435–451, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clyde:1992:TFO

[603] Stephen W. Clyde, David E. Embley, and Scott N. Woodfield. Tunable formalism in object-oriented systems analysis: Meeting the needs of both theoreticians and practitioners. ACM SIG-PLAN Notices, 27(10):452–465, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buhr:1992:AP

[604] Raymond J. A. Buhr and Ronald S. Casselman. Architectures with pictures. ACM SIGPLAN Notices, 27(10): 466–483, October 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deChampeaux:1992:OSD

[605] Dennis de Champeaux, Robert Balzer, Dave Bulman, Kathleen Culver-Lozo, Ivar Jacobson, and Stephen J. Mellor. The OO software development process (panel). ACM SIGPLAN Notices, 27(10):484–489, October 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meeson:1992:BRIb

[606] Reginald Meeson. Book review: An Introduction to Functional Programming Systems Using Haskell by Antony J. T. Davie: (Cambridge University Press, 1992). ACM SIGPLAN Notices, 27 (11):11, November 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lawyer:1992:BRW

[607] Steve Lawyer. Book review: Windows Graphics Programming with Borland C++ by Loren Heiny: (John Wiley & Sons, Inc., 1992). ACM SIG-PLAN Notices, 27(11):12, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mody:1992:FPS

[608] R. P. Mody. Functional programming is not self-modifying code. ACM SIG-PLAN Notices, 27(11):13–14, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1992:LCE

[609] Henry G. Baker. Less complex elementary functions. ACM SIGPLAN Notices, 27(11):15–16, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:1992:COO

[610] Bing Liu and Yuen-Wah Ku. ConstraintLisp: an object-oriented constraint programming language. ACM SIGPLAN Notices, 27(11):17–26, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yulin:1992:LLE

[611] Feng Yulin, Huang Tao, and Li Jing. LogC: a language and environment for embedded rule based systems. ACM SIGPLAN Notices, 27(11):27– 32, November 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Panti:1992:MOH

[612] Maurizio Panti and Salvatore Valenti. A modulus oriented hash function for the construction of minimal perfect tables. ACM SIGPLAN Notices, 27(11):

33–38, November 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cabasino:1992:DPE

[613] S. Cabasino, P. S. Paolucci, and G. M. Todesco. Dynamic parsers and evolving grammars. ACM SIGPLAN Notices, 27(11):39–48, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fickert:1992:UBF

[614] Chris Fickert and Thomas Sudkamp. Unification based FP interpreters. ACM SIGPLAN Notices, 27(11):49– 58, November 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakharov:1992:MPH

[615] Alexander Sakharov. Macro processing in high-level languages. ACM SIG-PLAN Notices, 27(11):59–66, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weisert:1992:MDC

[616] Conrad Weisert. Macros for defining C++ classes. ACM SIGPLAN Notices, 27(11):67–76, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1992:AMD

[617] C. K. Yuen. An abstract machine design for lexically scoped parallel Lisp with speculative processing. ACM SIG-PLAN Notices, 27(11):77–84, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mijanovic:1992:MNA

[618] Srdjan Mijanovic. MANEVRO: a new approach to class based programming. ACM SIGPLAN Notices, 27(11):85– 89, November 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klint:1992:PER

[619] Paul Klint, Thomas Reps, and Gregor Snelting. Programming environments: report on an international workshop at Dagstuhl Castle. ACM SIGPLAN Notices, 27(11):90–96, November 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Myers:1992:RCW

[620] Brad A. Myers. Report on the CHI’91 Workshop on Languages for Developing User Interfaces. ACM SIGPLAN Notices, 27(12):8–12, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Konrad:1992:AMT

[621] Erhard Konrad. Application of measurement theory to software metrics — comments on the Bollmann-Zuse approach. ACM SIGPLAN Notices, 27 (12):13–19, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See reply [622].

Zuse:1992:REK

[622] Horst Zuse and Peter Bollmann-Sdorra. Reply to: Erhard Konrad:

“Application of measurement theory to software metrics. Comments on the Bollmann-Zuse approach”. ACM SIG-PLAN Notices, 27(12):20–27, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [621].

Pool:1992:DTD

[623] Robert Pool. Did Turing Discover How the Leopard Got Its Spots?: Understanding the way a simple chemical system produces patterns may offer insights into animal development. ACM SIGPLAN Notices, 27(12):28, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dumke:1992:MBC

[624] Reiner R. Dumke, Kerstin Neumann, and Kerstin Stoeffler. The metric based compiler — A current requirement. ACM SIGPLAN Notices, 27(12): 29–38, December 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1992:ISS

[625] Henry G. Baker. Inlining semantics for subroutines which are recursive. ACM SIGPLAN Notices, 27(12):39–46, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCluskey:1992:TI

[626] Glen McCluskey and Robert B. Murray. Template instantiation for C++. ACM SIGPLAN Notices, 27(12):47– 56, December 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCluskey:1992:TIC

[627] Glen McCluskey and Robert B. Murray. Template instantiation for C++. ACM SIGPLAN Notices, 27(12):47– 56, December 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wong:1992:FIP

[628] K. C. Wong. Feature inversion: A practice on language versions determination. ACM SIGPLAN Notices, 27(12): 57–60, December 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stansifer:1992:CE

[629] Ryan Stansifer. The calculation of Easter. ACM SIGPLAN Notices, 27 (12):61–65, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See also [1023].

Razek:1992:COR

[630] Gottfried Razek. Combining objects and relations. ACM SIGPLAN Notices, 27(12):66–70, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zorn:1992:EMS

[631] Benjamin G. Zorn and Dirk Grunwald. Empirical measurements of six allocation-intensive C programs. ACM SIGPLAN Notices, 27(12):71–80, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cattaneo:1992:ACT

[632] G. Cattaneo, G. Di Giore, and M. Ruotolo. Another C threads library. ACM SIGPLAN Notices, 27(12):81–90, December 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steele:1993:HPF

[633] Guy L. Steele, Jr. High Performance Fortran: Status report. ACM SIGPLAN Notices, 28(1):1–4, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pratt:1993:KCP

[634] Terrence W. Pratt. Kernel-control parallel versus data parallel: A technical comparison. ACM SIGPLAN Notices, 28(1):5–8, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1993:LTN

[635] Wei Li and Keshav Pingali. Loop transformations for NUMA machines. ACM SIGPLAN Notices, 28(1):9–12, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sharma:1993:DDA

[636] S. Sharma, Chua-Huang Huang, and P. Sadayappan. On data dependence analysis for compiling programs on distributed-memory machines (extended abstract). ACM SIGPLAN Notices, 28(1):13–16, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1993:SML

[637] Jerome A. Feldman, Chu-Cheow Lim, and Thomas Rauber. The shared-memory language pSather on a distributed-memory multiprocessor. ACM SIGPLAN Notices, 28(1):17–20, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lemke:1993:LAI

[638] Max Lemke and Daniel Quinlan. Library for architecture-independent development of structured grid applications. ACM SIGPLAN Notices, 28(1): 21–23, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lemke:1993:PPC

[639] Max Lemke and Daniel Quinlan. P++, a parallel C++ array class library for architecture-independent development of structured grid applications. ACM SIGPLAN Notices, 28(1):21–23, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baden:1993:LPP

[640] Scott B. Baden and Scott R. Kohn. Lattice parallelism: A parallel programming model for manipulating non-uniform structured scientific data structures. ACM SIGPLAN Notices, 28 (1):24–27, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yelick:1993:PMI

[641] Katherine A. Yelick. Programming models for irregular applications. ACM SIGPLAN Notices, 28(1):28–31, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gross:1993:CTD

[642] T. Gross, S. Hinrichs, G. Lueh, D. O’Hallaron, J. Stichnoch, and J. Subhlok. Compiling task and data parallel programs for iWarp (extended abstract). ACM SIGPLAN Notices, 28 (1):32–35, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bala:1993:EDPa

[643] Vasanth Bala and Jeanne Ferrante. Explicit data placement (XDP): A methodology for explicit compile-time representation and optimization of data movement (extended abstract). ACM SIGPLAN Notices, 28(1):36–39, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snyder:1993:SOS

[644] Lawrence Snyder. Synopsis of Orca, A simple language implementation of phase abstractions. ACM SIGPLAN Notices, 28(1):40–43, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seevers:1993:PPE

[645] Bradley K. Seevers, Michael J. Quinn, and Philip J. Hatcher. A parallel programming environment supporting multiple data-parallel modules. ACM SIGPLAN Notices, 28(1):44–47, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maniatty:1993:EDA

[646] Bill Maniatty, Boleslaw K. Szymanski, and Balaram Sinharoy. Efficiency of data alignment on Maspar. ACM SIG-PLAN Notices, 28(1):48–51, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ponnusamy:1993:EDM

[647] Ravi Ponnusamy, Joel Saltz, Raja Das, Charles Koelbel, and Alok N. Choudhary. Embedding data mappers with distributed memory machine compilers. ACM SIGPLAN Notices, 28(1): 52–55, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:1993:IMD

[648] Peter R. Bailey and Malcolm C. Newey. Implementing ML on distributed memory multiprocessors. ACM SIGPLAN Notices, 28(1):56–59, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cox:1993:PTS

[649] Stuart Cox, Shell-Ying Huang, Paul Kelly, Junxian Liu, and Frank Taylor. Program transformations for static process networks. ACM SIGPLAN Notices, 28(1):60–63, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tu:1993:APS

[650] Peng Tu and David A. Padua. Array privatization for shared and distributed memory machines (extended abstract). ACM SIGPLAN Notices, 28 (1):64–67, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chatterjee:1993:OEA

[651] Siddhartha Chatterjee, John R. Gilbert, Robert Schreiber, and Shang-Hua Teng. Optimal evaluation of array expressions on massively parallel machines (extended abstract). ACM SIG-PLAN Notices, 28(1):68–71, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chapman:1993:UDM

[652] Barbara M. Chapman, Piyush Mehrotra, and Hans P. Zima. User defined mappings in Vienna Fortran. ACM SIGPLAN Notices, 28(1):72–75, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosing:1993:PPA

[653] Matthew Rosing, Robert B. Schnabel, and Robert P. Weaver. A programmable preprocessor approach to efficient parallel language design. ACM SIGPLAN Notices, 28(1):76–79, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brewer:1993:PLS

[654] Eric A. Brewer, Adrian Colbrook, Wilson C. Hsieh, Paul Wang, and William E. Weihl. Pipes: linguistic support for ordered asynchronous invocations. ACM SIGPLAN Notices, 28 (1):80, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carr:1993:DC

[655] Harold Carr, Robert R. Kessler, and Mark R. Swanson. Distributed C++. ACM SIGPLAN Notices, 28(1):81, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:1993:AGD

[656] S. K. S. Gupta, S. D. Kaushik, C.-H. Huang, J. R. Johnson, R. W. Johnson, and P. Sadayappan. On the automatic generation of data distributions. ACM SIGPLAN Notices, 28(1):82, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ikei:1993:FCT

[657] Mitsuru Ikei and Michael Wolfe. Functional compiler techniques for an imperative language. ACM SIGPLAN Notices, 28(1):83, January 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kranz:1993:IMPa

[658] David Kranz, Beng-Hong Lim, Kirk L. Johnson, John Kubiatowicz, and Anant Agarwal. Integrating message-passing and shared-memory: Early experience. ACM SIGPLAN Notices, 28 (1):84, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Majumdar:1993:ESC

[659] Aloke Majumdar and Marina C. Chen. Efficient SPMD constructs for asynchronous message passing architectures. ACM SIGPLAN Notices, 28 (1):85, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1993:DSI

[660] Sharon L. Smith. Dynamic scheduling of irregularly structured parallel computations in heterogeneous distributed systems. ACM SIGPLAN Notices, 28 (1):86, January 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zenith:1993:EMI

[661] Steven Ericsson Zenith. Ease: the model and its implementation. ACM SIGPLAN Notices, 28(1):87, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1993:BRD

[662] Jacques Cohen. Book review: DLP — A Language for Distributed Logic Programming by A. Eliens: (John Wiley and Sons, 1992). ACM SIGPLAN Notices, 28(2):13, February 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baowen:1993:SVN

[663] Xu Baowen. On subprograms with a variable number of parameters of varying types. ACM SIGPLAN Notices, 28(2):14–20, February 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:1993:SVN

[664] Baowen Xu. On subprograms with A variable number of parameters of varying types. ACM SIGPLAN Notices, 28(2):14–20, February 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Waite:1993:ELD

[665] W. M. Waite. An executable language definition. ACM SIGPLAN Notices, 28(2):21–40, February 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper describes a subset of C, called C–, defined as a literate program using the Eli and FunnelWeb systems.

Nigro:1993:TEO

[666] Libero Nigro. On the type extensions of Oberon-2. ACM SIGPLAN Notices, 28(2):41–44, February 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trehan:1993:COO

[667] Rajiv Trehan, Nobuyuki Sawashima, Akira Morishita, Ichiro Tomoda, Toru Imai, and Ken-Ichi Maeda. Concurrent Object Oriented ‘C’ (cooC). ACM SIG-PLAN Notices, 28(2):45–52, February 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schreiweis:1993:IPPa

[668] Uwe Schreiweis, A. Keune, and Horst Langendörfer. An integrated Prolog programming environment. ACM SIG-PLAN Notices, 28(2):53–60, February 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1993:TEP

[669] Paul W. Abrahams. Typographical extensions for programming languages: Breaking out of the ASCII straitjacket. ACM SIGPLAN Notices, 28(2): 61–68, February 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1993:MCP

[670] Per Brinch Hansen. Monitors and Concurrent Pascal: a personal history. ACM SIGPLAN Notices, 28(3):1–35, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p1-hansen/

.

Colmerauer:1993:BP

[671] Alain Colmerauer and Philippe Roussel. The birth of Prolog. ACM SIG-PLAN Notices, 28(3):37–52, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p37-colmerauer/

.

Griswold:1993:HIP

[672] Ralph E. Griswold and Madge T. Griswold. History of the Icon programming language. ACM SIG-PLAN Notices, 28(3):53–68, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p53-griswold/

.

Kay:1993:EHS

[673] Alan C. Kay. The early history of Smalltalk. ACM SIG-PLAN Notices, 28(3):69–95, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p69-kay/

.

Lindsey:1993:HA

[674] C. H. Lindsey. A history of ALGOL 68. ACM SIGPLAN Notices, 28(3):97–132, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p97-lindsey/

.

Liskov:1993:HC

[675] Barbara Liskov. A history of CLU. ACM SIGPLAN Notices, 28(3):133–147, March 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/154766/p133-liskov/

.

Nance:1993:HDE

[676] Richard E. Nance. A history of discrete event simulation programming languages. ACM SIGPLAN Notices, 28(3):149–175, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/154766/p149-nance/

.

Rather:1993:EF

[677] Elizabeth D. Rather, Donald R. Colburn, and Charles H. Moore. The evolution of Forth. ACM SIG-PLAN Notices, 28(3):177–199, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p177-rather/

.

Ritchie:1993:DCL

[678] Dennis M. Ritchie. The development of the C language. ACM SIG-PLAN Notices, 28(3):201–208, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p201-ritchie/

.

Sammet:1993:BDF

[679] Jean E. Sammet. The beginning and development of FORMAC (FORmula MAnipulation Compiler). ACM SIG-PLAN Notices, 28(3):209–230, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p209-sammet/

.

Steele:1993:EL

[680] Guy L. Steele, Jr. and Richard P. Gabriel. The evolution of Lisp. ACM SIGPLAN Notices, 28(3):231–270, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p231-steele/

.

Stroustrup:1993:HC

[681] Bjarne Stroustrup. A history of C++: 1979–1991. ACM SIG-PLAN Notices, 28(3):271–297, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p271-stroustrup/

.

Whitaker:1993:APD

[682] William A. Whitaker. Ada—the project: the DoD High Order Language Working Group. ACM SIG-PLAN Notices, 28(3):299–331, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p299-whitaker/

.

Wirth:1993:RAD

[683] N. Wirth. Recollections about the development of Pascal. ACM SIG-PLAN Notices, 28(3):333–342, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

proceedings/plan/154766/p333-wirth/

.

Gargaro:1993:BIA

[684] Anthony Gargaro. A brief introduction to Ada. ACM SIG-PLAN Notices, 28(3):343–344, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p343-gargaro/

.

Bowlden:1993:IA

[685] Henry J. Bowlden. An introduction to ALGOL 68. ACM SIG-PLAN Notices, 28(3):345–346, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p345-bowlden/

.

McKeeman:1993:BIC

[686] W. M. McKeeman. A brief introduction to C. ACM SIGPLAN Notices, 28(3):347–348, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/154766/p347-mckeeman/

.

Schwarz:1993:BIC

[687] Jerry Schwarz. A brief introduction to C++. ACM SIGPLAN Notices, 28(3):349–350, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

proceedings/plan/154766/p349-schwarz/

.

Guttag:1993:BIC

[688] John Guttag. A brief introduction to CLU. ACM SIGPLAN Notices, 28(3):351–352, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/154766/p351-guttag/

.

Hayden:1993:BIC

[689] Charles Hayden. A brief introduction to Concurrent Pascal. ACM SIG-PLAN Notices, 28(3):353–354, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p353-hayden/

.

Griesmer:1993:BIF

[690] James H. Griesmer. A brief introduction to FORMAC. ACM SIG-PLAN Notices, 28(3):355–356, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p355-griesmer/

.

Koopman:1993:BIF

[691] Philip J. Koopman, Jr. A brief introduction to Forth. ACM SIG-PLAN Notices, 28(3):357–358, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/plan/154766/p357-koopman/

.

Hanson:1993:BII

[692] David R. Hanson. A brief introduction to Icon. ACM SIG-PLAN Notices, 28(3):359–360, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/154766/p359-hanson/

.

Sussman:1993:BIL

[693] G. J. Sussman, G. L. Steele, Jr., and R. P. Gabriel. A brief introduction to Lisp. ACM SIG-PLAN Notices, 28(3):361–362, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/154766/p361-lee/

.

Lee:1993:BIP

[694] Peter Lee. A brief introduction to Pascal. ACM SIGPLAN Notices, 28(3):363–364, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/154766/p363-lee/

.

Pereira:1993:BIP

[695] Fernando Pereira. A brief introduction to Prolog. ACM SIG-PLAN Notices, 28(3):365–366, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p365-pereira/

.

Budd:1993:BIS

[696] Tim Budd. A brief introduction to Smalltalk. ACM SIGPLAN Notices, 28(3):367–368, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/154766/p367-budd/

.

Kiviat:1993:BID

[697] Philip Kiviat. A brief introduction to discrete-event simulation programming languages. ACM SIG-PLAN Notices, 28(3):369–370, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/154766/p369-kiviat/

.

Anonymous:1993:CME

[698] Anonymous. Computer museum ephemera announcement. ACM SIG-PLAN Notices, 28(3):viii–??, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1993:SOH

[699] Anonymous. Special offer for HOPL books. ACM SIGPLAN Notices, 28 (3):xiii–??, March 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1993:A

[700] John A. N. Lee. Acknowledgements. ACM SIGPLAN Notices, 28(3):vii–??, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1993:GI

[701] John A. N. Lee and Jean E. Sammet. General introduction. ACM SIG-PLAN Notices, 28(3):i–??, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mahoney:1993:MH

[702] Michael S. Mahoney. ‘what makes history?’. ACM SIGPLAN Notices, 28 (3):x–??, March 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sammet:1993:PD

[703] Jean E. Sammet. Program development. ACM SIGPLAN Notices, 28 (3):iii–??, March 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1993:EI

[704] Richard L. Wexelblat. Editor’s introduction. ACM SIGPLAN Notices, 28 (3):ix–??, March 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Norris:1993:BRM

[705] John W. Norris. Book review: Micro Focus Workbench: Developing Mainframe Applications on the PC by Alida Jatich and Phil Nowak: (John Wiley & Sons, New York, NY 1992). ACM SIGPLAN Notices, 28(4):7, April 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weisert:1993:KR

[706] Conrad Weisert. Has the King returned? ACM SIGPLAN Notices, 28 (4):9–10, April 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hernandez:1993:APL

[707] J. Hernández, Pedro de Miguel, M. Barrena, J. M. Martínez, A. Polo, and M. Nieto. ALBA: A parallel language based on actors. ACM SIGPLAN Notices, 28(4):11–20, April 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schreiweis:1993:IPPb

[708] Uwe Schreiweis, A. Keune, and Horst Langendörfer. An integrated Prolog programming environment. ACM SIGPLAN Notices, 28(4):21–28, April 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ivanovic:1993:DIL

[709] Mirjana Ivanovic and Zoran Budimac. A definition of an ISWIM-like language via Scheme. ACM SIGPLAN Notices, 28(4):29–38, April 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shu:1993:ADO

[710] William S. Shu. Adapting a debugger for optimized programs. ACM SIGPLAN Notices, 28(4):39–44, April 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stolz:1993:ART

[711] Otto Stolz. Anonymous routine-texts: an orthogonal approach to block objects. ACM SIGPLAN Notices, 28 (4):45–48, April 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baowen:1993:CPP

[712] Xu Baowen. CRL/Pascal: a Pascaloriented cross reference language and its applications. ACM SIGPLAN Notices, 28(4):49–54, April 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:1993:CPP

[713] Baowen Xu. CRL/Pascal: A Pascaloriented cross reference language and its applications. ACM SIGPLAN Notices, 28(4):49–54, April 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1993:IMG

[714] Henry G. Baker. ‘infant mortality’ and generational garbage collection. ACM SIGPLAN Notices, 28(4):55–57, April 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davison:1993:JSM

[715] Andrew Davison. 15th July 1972: A significant moment for logic programming. ACM SIGPLAN Notices, 28 (4):58–60, April 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Templ:1993:SAM

[716] Josef Templ. A systematic approach to multiple inheritance implementation. ACM SIGPLAN Notices, 28(4):61–66, April 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hall:1993:PCC

[717] Philip W. Hall, IV. Parsing with C++ constructors. ACM SIGPLAN Notices, 28(4):67–68, April 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hu:1993:CCK

[718] Shibin Hu. Comment on Cooper and Kennedy’s flow-insensitive interprocedural summary information computation algorithm. ACM SIGPLAN Notices, 28(5):3–8, May 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [818, 886].

Badii:1993:DSS

[719] M. Badii and F. Abdollahzadeh. Dynamic semantic specification by twolevel grammars for a block structured language with subroutine parameters. ACM SIGPLAN Notices, 28(5):9–18, May 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuhua:1993:APE

[720] Zheng Yuhua, Tu Honglei, and Xie Li. And/or parallel execution of logic programs: Exploiting dependent andparallelism. ACM SIGPLAN Notices, 28(5):19–28, May 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuhua:1993:PEL

[721] Zheng Yuhua, Tu Honglei, and Xie Li. And/or parallel execution of logic programs: exploiting dependent Andparallelism. ACM SIGPLAN Notices, 28(5):19–28, May 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Drechsler:1993:VKR

[722] Karl-Heinz Drechsler and Manfred P. Stadel. A variation of Knoop, Rüthing, and Steffen’s Lazy Code Motion. ACM SIGPLAN Notices, 28(5):29–38, May 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Passia:1993:FFI

[723] Jacek Passia and Klaus-Peter Löhr. Fips: A functional-imperative language for explorative programming. ACM SIGPLAN Notices, 28(5):39–48, May 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wentworth:1993:GRE

[724] E. P. Wentworth. Generalized regular expressions — a programming exercise in Haskell. ACM SIGPLAN Notices, 28(5):49–54, May 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:1993:PSP

[725] Baowen Xu. PIDL / Pascal: A Pascaloriented program interface description language and its applications. ACM SIGPLAN Notices, 28(5):53–62, May 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baowen:1993:PPP

[726] Xu Baowen. PIDL/Pascal: a Pascaloriented program interface description language and its applications. ACM SIGPLAN Notices, 28(5):55–62, May 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wahbe:1993:PDB

[727] Robert Wahbe, Steven Lucco, and Susan L. Graham. Practical data breakpoints: design and implementation. ACM SIGPLAN Notices, 28 (6):1–12, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p1-wahbe/

.

Adl-Tabatabai:1993:DRE

[728] Ali-Reza Adl-Tabatabai and Thomas Gross. Detection and recovery of endangered variables caused by instruction scheduling. ACM SIGPLAN Notices, 28(6):13–25, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p13-adl-
tabatabai/.

Boyd:1993:IAO

[729] Mickey R. Boyd and David B. Whalley. Isolation and analysis of optimization errors. ACM SIGPLAN Notices, 28(6):26–35, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p26-boyd/

.

Cytron:1993:EAM

[730] Ron Cytron and Reid Gershbein. Efficient accommodation of may-alias information in SSA form. ACM SIGPLAN Notices, 28(6):36–45, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p36-cytron/

.

Bourdoncle:1993:ADH

[731] François Bourdoncle. Abstract debugging of higher-order imperative languages. ACM SIGPLAN Notices, 28 (6):46–55, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p46-bourdoncle/

.

Landi:1993:IMS

[732] William Landi, Barbara G. Ryder, and Sean Zhang. Interprocedural modification side effect analysis with pointer aliasing. ACM SIGPLAN Notices, 28 (6):56–67, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p56-landi/

.

Duesterwald:1993:PDF

[733] Evelyn Duesterwald, Rajiv Gupta, and Mary Lou Soffa. A practical data flow framework for array reference analysis and its use in optimizations. ACM SIGPLAN Notices, 28 (6):68–77, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p68-duesterwald/

.

Johnson:1993:DBP

[734] Richard Johnson and Keshav Pingali. Dependence-based program analysis. ACM SIGPLAN Notices, 28(6):78–89, June 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1993:DPA

[735] Richard Johnson and Keshav Pingali. Dependence-based program analysis. ACM SIGPLAN Notices, 28 (6):78–89, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p78-johnson/

.

Grove:1993:ICP

[736] Dan Grove and Linda Torczon. Interprocedural constant propagation: a study of jump function implementations. ACM SIGPLAN Notices, 28 (6):90–99, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p90-grove/

.

Graham:1993:OIA

[737] Susan L. Graham, Steven Lucco, and Oliver Sharp. Orchestrating interactions among parallel computations. ACM SIGPLAN Notices, 28(6): 100–111, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p100-graham/

.

Anderson:1993:GOP

[738] Jennifer M. Anderson and Monica S. Lam. Global optimizations for parallelism and locality on scalable parallel machines. ACM SIGPLAN Notices, 28(6):112–125, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p112-anderson/

.

Amarasinghe:1993:COC

[739] Saman P. Amarasinghe and Monica S. Lam. Communication optimization and code generation for distributed memory machines. ACM SIGPLAN Notices, 28(6):126–138, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p126-amarasinghe/

.

Adams:1993:FCD

[740] Norman Adams, Pavel Curtis, and Mike Spreitzer. First-class datatype representations in SCHEMEXE-ROX. ACM SIGPLAN Notices, 28(6): 139–146, June 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:1993:FDR

[741] Norman Adams, Pavel Curtis, and Mike Spreitzer. First-class datatype representations in SCHEME XE-ROX. ACM SIGPLAN Notices, 28(6): 139–146, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p139-adams/

.

Sitaram:1993:HC

[742] Dorai Sitaram. Handling control. ACM SIGPLAN Notices, 28(6):147– 155, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p147-sitaram/

.

Weise:1993:PSM

[743] Daniel Weise and Roger F. Crew. Programmable syntax macros. ACM SIGPLAN Notices, 28(6):156–165, June 1993. CODEN SINODQ.

ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p156-weise/

.

Hong:1993:CRP

[744] Seongsoo Hong and Richard Gerber. Compiling real-time programs into schedulable code. ACM SIG-PLAN Notices, 28(6):166–176, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p166-hong/

.

Hong:1993:CRT

[745] Seongsoo Hong and Richard Gerber. Compiling real-time programs into schedulable code. ACM SIG-PLAN Notices, 28(6):166–176, June 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grunwald:1993:ICL

[746] Dirk Grunwald, Benjamin Zorn, and Robert Henderson. Improving the cache locality of memory allocation. ACM SIGPLAN Notices, 28(6):177– 186, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p177-grunwald/

.

Barrett:1993:ULP

[747] David A. Barrett and Benjamin G. Zorn. Using lifetime predictors to improve memory allocation performance. ACM SIGPLAN Notices, 28 (6):187–196, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p187-barrett/

.

Boehm:1993:SEC

[748] Hans-Juergen Boehm. Space efficient conservative garbage collection. ACM SIGPLAN Notices, 28(6): 197–206, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p197-boehm/

.

Dybvig:1993:GGB

[749] R. Kent Dybvig, Carl Bruggeman, and David Eby. Guardians in a generationbased garbage collector. ACM SIG-PLAN Notices, 28(6):207–216, June 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dybvig:1993:GGG

[750] R. Kent Dybvig, Carl Bruggeman, and David Eby. Guardians in a generation-based garbage collector. ACM SIGPLAN Notices, 28(6):207– 216, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p207-dybvig/

.

Nettles:1993:RRG

[751] Scott Nettles and James O’Toole. Real-time replication garbage collection. ACM SIGPLAN Notices, 28(6): 217–226, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p217-nettles/

.

Nettles:1993:RTR

[752] Scott Nettles and James O’Toole. Real-time replication garbage collection. ACM SIGPLAN Notices, 28(6): 217–226, June 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Peterson:1993:ITC

[753] John Peterson and Mark P. Jones. Implementing type classes. ACM SIGPLAN Notices, 28(6):227–236, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p227-peterson/

.

Flanagan:1993:ECC

[754] Cormac Flanagan, Amr Sabry, Bruce F. Duba, and Matthias Felleisen. The essence of compiling with continuations. ACM SIGPLAN Notices, 28(6): 237–247, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p237-flanagan/

.

Pinter:1993:RAI

[755] Shlomit S. Pinter. Register allocation with instruction scheduling: a new approach. ACM SIGPLAN Notices, 28 (6):248–257, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p248-pinter/

.

Huff:1993:LMS

[756] Richard A. Huff. Lifetime-sensitive modulo scheduling. ACM SIG-PLAN Notices, 28(6):258–267, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p258-huff/

.

Huff:1993:LSM

[757] Richard A. Huff. Lifetime-sensitive modulo scheduling. ACM SIGPLAN Notices, 28(6):258–267, June 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kolte:1993:LRA

[758] Priyadarshan Kolte and Mary Jean Harrold. Load/store range analysis for global register allocation. ACM SIGPLAN Notices, 28(6):268– 277, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p268-kolte/

.

Kolte:1993:LSR

[759] Priyadarshan Kolte and Mary Jean Harrold. Load/store range analysis for global register allocation. ACM SIG-PLAN Notices, 28(6):268–277, June 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kerns:1993:BSI

[760] Daniel R. Kerns and Susan J. Eggers. Balanced scheduling: instruction scheduling when memory latency is uncertain. ACM SIGPLAN Notices, 28(6):278–289, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p278-kerns/

.

Warter:1993:RI

[761] Nancy J. Warter, Scott A. Mahlke, Wen-Mei W. Hwu, and B. Ramakrishna Rau. Reverse If-Conversion. ACM SIGPLAN Notices, 28(6):290– 299, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/155090/p290-warter/

.

Warter:1993:RIC

[762] Nancy J. Warter, Scott A. Mahlke, Wen-Mei W. Hwu, and B. Ramakrishna Rau. Reverse If-Conversion. ACM SIGPLAN Notices, 28(6):290– 299, June 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:1993:BPF

[763] Thomas Ball and James R. Larus. Branch prediction for free. ACM SIGPLAN Notices, 28(6):300–313, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/155090/p300-ball/

.

Culler:1993:LTR

[764] David E. Culler, Richard M. Karp, David A. Patterson, Abhijit Sahay, Klaus E. Schauser, Eunice Santos, Ramesh Subramonian, and Thorsten von Eicken. LogP: towards a realistic model of parallel computation. ACM SIGPLAN Notices, 28(7):1–12, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Subhlok:1993:ETD

[765] Jaspal Subhlok, James M. Stichnoth, David R. O’Hallaron, and Thomas Gross. Exploiting task and data parallelism on a multicomputer. ACM SIGPLAN Notices, 28(7):13–22, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agha:1993:AOD

[766] Gul Agha and Christian J. Callsen. ActorSpace: an open distributed programming paradigm. ACM SIG-PLAN Notices, 28(7):23–32, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hall:1993:EUP

[767] Mary W. Hall, Timothy J. Harvey, Ken Kennedy, Nathaniel McIntosh, Kathryn S. McKinley, Jeffrey D. Oldham, Michael H. Paleczny, and Gerald Roth. Experiences using the ParaScope Editor: an interactive parallel programming tool. ACM SIG-PLAN Notices, 28(7):33–43, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarukkai:1993:PAH

[768] Sekhar R. Sarukkai and Allen D. Malony. Perturbation analysis of high level instrumentation for SPMD programs. ACM SIGPLAN Notices, 28(7):44–53, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kranz:1993:IMPb

[769] David Kranz, Kirk Johnson, Anant Agarwal, John Kubiatowicz, and Beng-Hong Lim. Integrating message-passing and shared-memory: early experience. ACM SIGPLAN Notices, 28(7):54–63, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kranz:1993:IMS

[770] David Kranz, Kirk L. Johnson, Anant Agarwal, John Kubiatowicz, and Beng-Hong Lim. Integrating message-passing and shared-memory: early experience. ACM SIGPLAN Notices, 28(7):54–63, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kontothanasiss:1993:USI

[771] Leonidas Kontothanasiss and Robert W. Wisniewski. Using schedular information to achieve optimal barrier synchronization performance. ACM SIG-PLAN Notices, 28(7):64–72, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kontothanassis:1993:USI

[772] Leonidas Kontothanassis and Robert W. Wisniewski. Using scheduler information to achieve optimal barrier synchronization performance. ACM SIG-PLAN Notices, 28(7):64–72, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huelsbergen:1993:CCG

[773] Lorenz Huelsbergen and James R. Larus. A concurrent copying garbage collector for languages that distinguish (im)mutable data. ACM SIG-PLAN Notices, 28(7):73–82, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leung:1993:IPR

[774] Shun-Tak Leung and John Zahorjan. Improving the performance of runtime parallelization. ACM SIG-PLAN Notices, 28(7):83–91, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chapman:1993:HPF

[775] Barbara M. Chapman, Piyush Mehrotra, and Hans P. Zima. High performance Fortran without templates: an alternative model for distribution and alignment. ACM SIGPLAN Notices, 28 (7):92–101, July 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blelloch:1993:IPN

[776] Guy E. Blelloch, Siddhartha Chatterjee, Jonathan C. Hardwick, Jay Sipelstein, and Marco Zagha. Implementation of a portable nested data-parallel language. ACM SIGPLAN Notices, 28 (7):102–111, July 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rao:1993:ELD

[777] Pushpa Rao and Clifford Walinsky. An equational language for dataparallelism. ACM SIGPLAN Notices, 28(7):112–118, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prins:1993:THD

[778] Jan F. Prins and Daniel W. Palmer. Transforming high-level data-parallel programs into vector operations. ACM SIGPLAN Notices, 28(7):119–128, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prins:1993:THL

[779] Jan F. Prins and Daniel W. Palmer. Transforming high-level data-parallel programs into vector operations. ACM SIGPLAN Notices, 28(7):119–128, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Masticola:1993:NA

[780] Stephen P. Masticola and Barbara G. Ryder. Non-concurrency analysis. ACM SIGPLAN Notices, 28(7):129– 138, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Masticola:1993:NCA

[781] Stephen P. Masticola and Barbara G. Ryder. Non-concurrency analysis. ACM SIGPLAN Notices, 28(7):129– 138, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bala:1993:EDPb

[782] Vasanth Bala, Jeanne Ferrante, and Larry Carter. Explicit data placement (XDP): a methodology for explicit compile-time representation and optimization of data movement. ACM SIGPLAN Notices, 28(7):139–148, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chatterjee:1993:GLA

[783] Siddhartha Chatterjee, John R. Gilbert, Fred J. E. Long, Robert Schreiber, and Shang-Hua Teng. Generating local addresses and communication sets for data-parallel programs. ACM SIGPLAN Notices, 28 (7):149–158, July 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grunwald:1993:DFE

[784] Dirk Grunwald and Harini Srinivasan. Data flow equations for explicitly parallel programs. ACM SIGPLAN Notices, 28(7):159–168, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chakrabarti:1993:IIA

[785] Soumen Chakrabarti and Katherine A. Yelick. Implementing an irregular application on a distributed memory multiprocessor. ACM SIGPLAN Notices, 28(7):169–178, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siegl:1993:PAS

[786] Kurt Siegl. Parallelizing algorithms for symbolic computation using ∥Maple∥. ACM SIGPLAN Notices, 28(7):179– 186, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yeung:1993:EFG

[787] Donald Yeung and Anant Agarwal. Experience with fine-grain synchronization in MIMD machines for preconditioned conjugate gradient. ACM SIGPLAN Notices, 28(7):187–192, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yeung:1993:EFS

[788] Donald Yeung and Anant Agarwal. Experience with fine-grain synchronization in MIMD machines for preconditioned conjugate gradient. ACM SIGPLAN Notices, 28(7):187–197, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morrisett:1993:PLP

[789] J. Gregory Morrisett and Andrew P. Tolmach. Procs and locks: a portable multiprocessing platform for Standard ML of New Jersey. ACM SIG-PLAN Notices, 28(7):198–207, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wagner:1993:LPT

[790] David B. Wagner and Bradley G. Calder. Leapfrogging: a portable technique for implementing efficient futures. ACM SIGPLAN Notices, 28 (7):208–217, July 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taura:1993:EIS

[791] Kenjiro Taura, Satoshi Matsuoka, and Akinori Yonezawa. An efficient implementation scheme of concurrent objectoriented languages on stock multicomputers. ACM SIGPLAN Notices, 28 (7):218–228, July 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sandhu:1993:SRA

[792] Harjinder S. Sandhu, Benjamin Gamsa, and Songnian Zhou. The shared regions approach to software cache coherence on multiprocessors. ACM SIGPLAN Notices, 28(7):229–238, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hsieh:1993:CME

[793] Wilson C. Hsieh, Paul Wang, and William E. Weihl. Computation migration: enhancing locality for distributed-memory parallel systems. ACM SIGPLAN Notices, 28(7):239– 248, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chandra:1993:DLL

[794] Rohit Chandra, Anoop Gupta, and John L. Hennessy. Data locality and load balancing in COOL. ACM SIG-PLAN Notices, 28(7):249–259, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guthery:1993:WSH

[795] Scott Guthery. Are we still having fun?: a minority report from HOPL-II. ACM SIGPLAN Notices, 28(8):1, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1993:BRO

[796] Alan Wexelblat. Book review: Obfuscated C and Other Mysteries by Don Libes: (John Wiley & Sons, Inc.). ACM SIGPLAN Notices, 28(8):2, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hesse:1993:PIB

[797] Jan Hesse, Rainer König, Filippo Logi, and Jens Herder. A prototype of an interface builder for the Common Lisp Interface Manager — CLIB. ACM SIGPLAN Notices, 28(8):19–28, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chandra:1993:CEW

[798] Kailash Chandra. C++ in eight weeks. ACM SIGPLAN Notices, 28(8):29–38, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schilling:1993:FSP

[799] Jonathan L. Schilling. Fail-safe programming in compiler optimization. ACM SIGPLAN Notices, 28(8):39–42, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grune:1993:TLG

[800] Dick Grune. Two-level grammars are more expressive than Type 0 grammars: or are they? ACM SIG-PLAN Notices, 28(8):43–45, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fei:1993:ATM

[801] Zongming Fei and Jian Lu. An approach toward mechanization of acquiring reusable software components. ACM SIGPLAN Notices, 28(8):46–52, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Czeck:1993:DSL

[802] Edward W. Czeck and James M. Feldman. On defusing a small landmine in the type casting of pointers in the “C” language. ACM SIGPLAN Notices, 28 (8):53–56, August 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Janssens:1993:HDP

[803] Pascal Janssens and Annie Cuyt. How does PASCAL-XSC compare to other programming languages with respect to the IEEE standard? ACM SIG-PLAN Notices, 28(8):57–66, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ejiogu:1993:FPF

[804] Lem O. Ejiogu. Five principles for the formal validation of models of software metrics. ACM SIGPLAN Notices, 28 (8):67–76, August 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cashman:1993:ESC

[805] Mark Cashman. Edicates — A specification of calling sequences. ACM SIGPLAN Notices, 28(8):77–80, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caslunan:1993:ESC

[806] Mark Caslunan. Edicates — a specification of calling sequences. ACM SIGPLAN Notices, 28(8):77–80, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deLaRocqueRodriguez:1993:TS

[807] Noemi de La Rocque Rodriguez, Roberto Ierusalimschy, and José Lucas Rangel. Types in school. ACM SIGPLAN Notices, 28(8):81–89, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodriguez:1993:TS

[808] Noemi de la Rocque Rodriguez, Roberto Ierusalimschy, and José Lucas Rangel. Types in school. ACM SIGPLAN Notices, 28(8):81–89, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shute:1993:AAB

[809] Malcolm J. Shute. Abotec: an automatic back-of-the-envelope calculator. ACM SIGPLAN Notices, 28(8):90–98, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Philippsen:1993:CMI

[810] Michael Philippsen, Ernst A. Heinz, and Paul Lukowicz. Compiling machine-independent parallel programs. ACM SIGPLAN Notices, 28(8): 99–108, August 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rost:1993:DDD

[811] Johann Rost. “D2R”: a dynamic dataflow representation for task scheduling. ACM SIGPLAN Notices, 28(8):109–116, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rost:1993:DDR

[812] Johann Rost. D^{2}R: A dynamic dataflow representation for task scheduling. ACM SIGPLAN Notices, 28(8):109–116, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arro:1993:PBQ

[813] Henrik Arro, Jonas Barklund, and Johan Bevemyr. Parallel bounded quantification — preliminary results. ACM SIGPLAN Notices, 28(8):117–124, August 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Getzinger:1993:ARA

[814] Thomas W. Getzinger, Ralph Clarke Haygood, and Peter Van Roy. Announcing the release of Aquarius Prolog. ACM SIGPLAN Notices, 28(9): 5, September 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pfeiffer:1993:RSA

[815] Phil Pfeiffer. Report on the Second Annual Alan J. Perlis Symposium on Programming Languages. ACM SIG-PLAN Notices, 28(9):6–12, September 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bielak:1993:OOP

[816] Richie Bielak. Object oriented programming: The fundamentals. ACM SIGPLAN Notices, 28(9):13– 14, September 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woodman:1993:TMS

[817] Mark Woodman. A taste of the Modula-2 Standard. ACM SIGPLAN Notices, 28(9):15–24, September 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Banning:1993:CCSa

[818] John Banning. Comment on a comment by Shibin Hu. ACM SIG-PLAN Notices, 28(9):25, September 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [718, 818].

Bell:1993:SRA

[819] Doug Bell and Mike Parr. Spreadsheets: A research agenda. ACM SIG-PLAN Notices, 28(9):26–28, September 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gao:1993:SIC

[820] Yaoqing Gao and Chung Kwong Yuen. A survey of implementations of concurrent, parallel and distributed Smalltalk. ACM SIGPLAN Notices, 28(9): 29–35, September 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liffers:1993:IVC

[821] Rainer H. Liffers. Inheritance versus containment. ACM SIGPLAN Notices, 28(9):36–38, September 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heymann:1993:PID

[822] Jurgen Heymann. A 100% portable inline-debugger. ACM SIGPLAN Notices, 28(9):39–46, September 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldsby:1993:SIT

[823] Michael E. Goldsby. A simple implementation technique for mixin inheritance. ACM SIGPLAN Notices, 28(9): 47–56, September 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:1993:OOP

[824] Bernd Müller. Is object-oriented programming structured programming? ACM SIGPLAN Notices, 28(9):57– 66, September 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marlowe:1993:PIA

[825] Thomas J. Marlowe, William G. Landi, Barbara G. Ryder, Jong-Deok Choi, Michael G. Burke, and Paul Carini. Pointer-induced aliasing: A clarification. ACM SIGPLAN Notices, 28(9): 67–70, September 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nierstrasz:1993:RTA

[826] Oscar Nierstrasz. Regular types for active objects. ACM SIGPLAN Notices, 28(10):1–15, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liskov:1993:STU

[827] Barbara Liskov and Jeannette M. Wing. Specifications and their use in defining subtypes. ACM SIGPLAN Notices, 28(10):16–28, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruce:1993:SDT

[828] Kim B. Bruce, Jon Crabtree, Thomas P. Murtagh, Robert van Gent, Allyn Dimock, and Robert Muller. Safe and decidable type checking in an objectoriented language. ACM SIGPLAN Notices, 28(10):29–46, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mendez:1993:WMB

[829] Rich Mendez, Rebecca Wirfs-Brock, James Rumbaugh, Stephen J. Mellor, Ivar Jacobson, and Grady Booch. Which method is best? shoot out at the OO corral. ACM SIGPLAN Notices, 28(10):46–47 (or 47–47??), October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krueger:1993:TDA

[830] Keith Krueger, David Loftesness, Amin Vahdat, and Thomas Anderson. Tools for the development of applicationspecific virtual memory management. ACM SIGPLAN Notices, 28(10):48– 64, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruegge:1993:FDP

[831] Bernd Bruegge, Tim Gottschalk, and Bin Luo. A framework for dynamic program analyzers. ACM SIGPLAN Notices, 28(10):65–82, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bingham:1993:EDU

[832] Tim Bingham, Nancy Hobbs, and Dave Husson. Experiences developing and using an object-oriented library for program manipulation. ACM SIGPLAN Notices, 28(10):83–89, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanStee:1993:SOO

[833] J. G. Van Stee, Dan Clarke, David Filani, Dmitry Lenkov, and Raymond Obin. Status of object-oriented COBOL (panel). ACM SIGPLAN Notices, 28(10):90, October 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanStee:1993:SOC

[834] J. G. van Stee, Megan Adams, Dan Clarke, Dimitri Lenkov, and Raymond Obin. Status of object-oriented COBOL. ACM SIGPLAN Notices, 28 (10):90, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kale:1993:CPC

[835] Laxmikant V. Kalé and Sanjeev Krishnan. CHARM++: A portable concurrent object oriented system based on C++. ACM SIGPLAN Notices, 28(10): 91–108, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matsuoka:1993:HEE

[836] Satoshi Matsuoka, Kenjiro Taura, and Akinori Yonezawa. Highly efficient and encapsulated re-use of synchronization code in concurrent objectoriented languages. ACM SIGPLAN Notices, 28(10):109–126, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taft:1993:AAO

[837] S. Tucker Taft. Ada 9X. from abstraction-oriented to object-oriented. ACM SIGPLAN Notices, 28(10):127– 136, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seidewitz:1993:DSL

[838] Ed Seidewitz, Brad Balfour, Sam S. Adams, David M. Wade, and Brad Cox. Developing software for largescale reuse (panel). ACM SIGPLAN Notices, 28(10):137–143, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cahill:1993:AGG

[839] Vinny Cahill, Seán Baker, Chris Horn, and Gradimir Starovic. The Amadeus GRT: generic runtime support for distributed persistent programming. ACM SIGPLAN Notices, 28(10):144– 161, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bal:1993:ODO

[840] Henri E. Bal and M. Frans Kaashoek. Object distribution in Orca using Compile-Time and Run-Time techniques. ACM SIGPLAN Notices, 28 (10):162–177, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odeh:1993:OEO

[841] Mohammed H. Odeh and Julian A. Padget. Object-oriented execution of OPS5 production systems. ACM SIG-PLAN Notices, 28(10):178–190, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odeh:1993:OOE

[842] Mohammed H. Odeh and Julian A. Padget. Object-oriented execution of OPS5 production systems. ACM SIG-PLAN Notices, 28(10):178–190, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coppola:1993:DBG

[843] Angela Jo Coppola and William Ruth. Doing business with governments. ACM SIGPLAN Notices, 28(10):191, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strauss:1993:IIG

[844] Paul S. Strauss. IRIS Inventor, A 3D graphics toolkit. ACM SIGPLAN Notices, 28(10):192–200, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lamping:1993:TSI

[845] John Lamping. Typing the specialization interface. ACM SIGPLAN Notices, 28(10):201–214, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Larnping:1993:TSI

[846] John Larnping. Typing the specialization interface. ACM SIGPLAN Notices, 28(10):201–214, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bracha:1993:STS

[847] Gilad Bracha and David Griswold. Strongtalk: typechecking Smalltalk in a production environment. ACM SIG-PLAN Notices, 28(10):215–230, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hauck:1993:IME

[848] Franz J. Hauck. Inheritance modeled with explicit bindings: An approach to typed inheritance. ACM SIGPLAN Notices, 28(10):231–239, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoydalsvik:1993:POA

[849] Geir Magne Hoydalsvik and Guttorm Sindre. On the purpose of objectoriented analysis. ACM SIGPLAN Notices, 28(10):240–255, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoydalsvik:1993:POO

[850] Geir Magne Høydalsvik and Guttorm Sindre. On the purpose of objectoriented analysis. ACM SIGPLAN Notices, 28(10):240–255, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomas:1993:DPO

[851] Dave Thomas, Adele Goldberg, James Coplien, Peter Coad, and Geir Magne Høydalsvik. A discussion of On the Purpose of Object-Oriented Analysis. ACM SIGPLAN Notices, 28(10):256– 258, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Driesen:1993:STI

[852] Karel Driesen. Selector table indexing & sparse arrays. ACM SIGPLAN Notices, 28(10):259–270, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casean:1993:EHM

[853] Yves Casean. Efficient handling of multiple inheritance hierarchies. ACM SIGPLAN Notices, 28(10):271– 287, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caseau:1993:EHM

[854] Yves Caseau. Efficient handling of multiple inheritance hierarchies. ACM SIGPLAN Notices, 28(10):271– 287, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hosking:1993:OFH

[855] Antony L. Hosking and J. Eliot B. Moss. Object fault handling for persistent programming languages: A performance evaluation. ACM SIGPLAN Notices, 28(10):288–303, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thuraisingham:1993:IOO

[856] Bhavani Thuraisingham. Integrating object-oriented technology and security technology. ACM SIGPLAN Notices, 28(10):304, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thuraisingham:1993:IOT

[857] Bhavani M. Thuraisingham. Integrating object technology and security technology: A panel discussion. ACM SIGPLAN Notices, 28(10):304, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ting:1993:MSR

[858] T. C. Ting. Modeling security requirements for applications (panel). ACM SIGPLAN Notices, 28(10):305, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sell:1993:OOA

[859] Peter J. Sell. Object oriented approach to MLS database application design (panel). ACM SIGPLAN Notices, 28 (10):306, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sandhu:1993:SOS

[860] Ravi S. Sandhu. Security for OODBMS (or systems) (panel). ACM SIG-PLAN Notices, 28(10):307, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keefe:1993:ROM

[861] Thomas F. Keefe. Reconciling objects and multilevel security (panel). ACM SIGPLAN Notices, 28(10):308, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koved:1993:GOO

[862] Larry Koved and Wayne L. Wooten. GROOP: an object-oriented toolkit for animated 3D graphics. ACM SIG-PLAN Notices, 28(10):309–325, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koved:1993:GOT

[863] Larry Koved and Wayne L. Wooten. GROOP: An object-oriented toolkit for animated 3D graphics. ACM SIG-PLAN Notices, 28(10):309–325, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DePauw:1993:VBO

[864] Wim De Pauw, Richard Helm, Doug Kimelman, and John Vlissides. Visualizing the behavior of object-oriented systems. ACM SIGPLAN Notices, 28 (10):326–337, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pauw:1993:VBO

[865] Wim De Pauw, Richard Helm, Doug Kimelman, and John Vlissides. Visualizing the behavior of object-oriented systems. ACM SIGPLAN Notices, 28 (10):326–337, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minoura:1993:SAO

[866] Toshimi Minoura, Shirish S. Pargaonkar, and Kurt Rehfuss. Structural active object systems for simulation. ACM SIGPLAN Notices, 28(10): 338–355, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1993:SAN

[867] Bruce Anderson, Mary Shaw, Larry Best, and Kent Beck. Software architecture: the next step for object technology (panel). ACM SIGPLAN Notices, 28(10):356–359, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1993:SAN

[868] Mary Shaw, Larry Best, and Kent Beck. Software architecture: The next step for object technology (panel). ACM SIGPLAN Notices, 28(10):356– 362, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shan:1993:MIE

[869] Yen-Ping Shan, Tom Cargill, Brad Cox, William Cook, Mary Loomis, and Alan Snyder. Is multiple inheritance essential to OOP? (panel). ACM SIG-PLAN Notices, 28(10):360–363, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shan:1993:PMI

[870] Yeng-Ping Shan, Tom Cargill, Brad Cox, William Cook, Mary Loomis, and Alan Snyder. Panel: Is multiple inheritance essential to OOP? ACM SIGPLAN Notices, 28(10):363, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nierstrasz:1993:PDP

[871] Oscar Nierstrasz. Panel: Distributed processing. ACM SIGPLAN Notices, 28(10):364, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stark:1993:IOO

[872] Mike Stark. Impacts of object-oriented technologies: seven years of SEL studies. ACM SIGPLAN Notices, 28(10): 365–373, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stark:1993:IOT

[873] Mike Stark. Impacts of object-oriented technologies: Seven years of SEL studies. ACM SIGPLAN Notices, 28(10): 365–373, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Litvintchonk:1993:PET

[874] Steven D. Litvintchonk, Mike Stark, Brad Balfour, Mohamed Fayad, and Bernie Rosenfeld. Panel: Evolving toward object-oriented technology in large organizations. ACM SIGPLAN Notices, 28(10):374–375, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Litvintchouk:1993:ETO

[875] Steven D. Litvintchouk, Mike Stark, Brad Balfour, Mohamed Fayad, and Bernie Rosenfeld. Evolving toward object-oriented technology in large organizations (panel). ACM SIGPLAN Notices, 28(10):374–375, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Civello:1993:RCO

[876] Franco Civello. Roles for composite objects in object-oriented analysis and design. ACM SIGPLAN Notices, 28(10): 376–393, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Godin:1993:BMA

[877] Robert Godin and Hafedh Mili. Building and maintaining analysis-level class hierarchies using Galois Lattices. ACM SIGPLAN Notices, 28(10):394–410, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1993:SOP

[878] William Harrison and Harold Ossher. Subject-oriented programming: a critique of pure objects. ACM SIGPLAN Notices, 28(10):411–428, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1993:SPC

[879] William Harrison and Harold Ossher. Subject-oriented programming (a critique of pure objects). ACM SIGPLAN Notices, 28(10):411–428, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beck:1993:HGP

[880] Kent Beck, Grady Booch, William Cook, Richard Gabriel, and Rebecca Wirfs-Brock. How to get a paper accepted at OOPSLA (panel). ACM SIG-PLAN Notices, 28(10):429–436, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1993:HGP

[881] Ralph E. Johnson, Kent Beck, Grady Booch, William Cook, Richard Gabriel, and Rebecca Wirfs-Brock. How to get a paper accepted at OOPSLA (panel). ACM SIGPLAN Notices, 28(10):429– 436, October 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deChampeaux:1993:SOO

[882] Dennis de Champeaux, Andrew J. Baer, Brian Bernsen, Alan R. Korncoff, Tim Korson, and Daniel S. Tkach. Strategies for object-oriented technology transfer (panel). ACM SIGPLAN Notices, 28(10):437–447, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1993:TI

[883] Anonymous. Title index. ACM SIG-PLAN Notices, 28(10):448, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1993:AI

[884] Anonymous. Author index. ACM SIG-PLAN Notices, 28(10):449–??, October 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Michael:1993:BRN

[885] J. Bret Michael. Book review: NeXTStep Programming: Step One, Object-Oriented Applications, by Simson L. Garfinkel and Michael K. Mahoney: (Springer-Verlag, New York, 1993). ACM SIGPLAN Notices, 28 (11):2, November 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Banning:1993:CCSb

[886] John Banning. Comment on a comment by Shibin Hu. ACM SIG-PLAN Notices, 28(11):3, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [718, 886].

Jones:1993:HGG

[887] Simon L. Peyton Jones, John Hughes, and John Launchbury. How to give a good research talk. ACM SIGPLAN Notices, 28(11):9–12, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

PeytonJones:1993:HGG

[888] Simon L. Peyton Jones, John Hughes, and John Launchbury. How to give a good research talk. ACM SIGPLAN Notices, 28(11):9–12, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adrian:1993:MPC

[889] Frank A. Adrian. A modest proposal concerning variables and assignment statements. ACM SIGPLAN Notices, 28(11):13–15, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dehbonei:1993:VPB

[890] Babak Dehbonei and Fernando Mejia. Verification of proofs for the B formal development process. ACM SIGPLAN Notices, 28(11):16–21, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1993:CGI

[891] Henry G. Baker. Complex Gaussian integers for ‘Gaussian graphics’. ACM SIGPLAN Notices, 28(11):22– 27, November 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deOliveiraGuimaraes:1993:PM

[892] José de Oliveira Guimarães. Parametrized methods. ACM SIGPLAN Notices, 28 (11):28–32, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guimaraes:1993:PM

[893] José de Oliveira Guimarães. Parametrized methods. ACM SIGPLAN Notices, 28 (11):28–32, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harman:1993:PFM

[894] Mark Harman and Sebastian Danicic. Projecting functional models of imperative programs. ACM SIGPLAN Notices, 28(11):33–41, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hetherington:1993:IEP

[895] Tony Hetherington. An introduction to the Extended Pascal language. ACM SIGPLAN Notices, 28(11):42– 51, November 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiessling:1993:BP

[896] Heiko Kießling and Uwe Krüger. Blocks and procedures. ACM SIGPLAN Notices, 28(11):52–61, November 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Netzer:1993:OTR

[897] Robert H. B. Netzer. Optimal tracing and replay for debugging sharedmemory parallel programs. ACM SIG-PLAN Notices, 28(12):1–11, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kimelman:1993:FTS

[898] Doug Kimelman and Dror Zernik. Onthe-fly topological sort — a basis for interactive debugging and live visualization of parallel programs. ACM SIG-PLAN Notices, 28(12):12–20, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kimelman:1993:OTS

[899] Doug Kimelman and Dror Zernik. Onthe-fly topological sort: A basis for interactive debugging and live visualization of parallel programs. ACM SIG-PLAN Notices, 28(12):12–20, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tomlinson:1993:DRG

[900] Alexander I. Tomlinson and Vijay K. Garg. Detecting relational global predicates in distributed systems. ACM SIGPLAN Notices, 28(12):21–31, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hurfin:1993:DAS

[901] Michel Hurfin, Noël Plouzeau, and Michel Raynal. Detecting atomic sequences of predicates in distributed computations. ACM SIGPLAN Notices, 28(12):32–42, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Valot:1993:CAD

[902] Céline Valot. Characterizing the accuracy of distributed timestamps. ACM SIGPLAN Notices, 28(12):43–52, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rover:1993:MAM

[903] Diane T. Rover and Abdul Waheed. Multiple-domain analysis methods. ACM SIGPLAN Notices, 28(12): 53–63, December 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rover:1993:MDA

[904] Diane T. Rover and Abdul Waheed. Multiple-domain analysis methods. ACM SIGPLAN Notices, 28(12): 53–63, December 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pancake:1993:CPP

[905] Cherri M. Pancake. Customizable portrayals of program structure. ACM SIGPLAN Notices, 28(12):64–74, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kunz:1993:PCD

[906] Thomas Kunz. Process clustering for distributed debugging. ACM SIG-PLAN Notices, 28(12):75–84, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cuny:1993:ADS

[907] Janice E. Cuny, George Forman, Alfred Hough, Joydip Kundu, Calvin Lin, Lawrence Snyder, and David W. Stemple. The Ariadne debugger: scalable application of event-based abstraction. ACM SIGPLAN Notices, 28(12): 85–95, December 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

May:1993:PPE

[908] John May and Francine Berman. Panorama: A portable, extensible parallel debugger. ACM SIGPLAN Notices, 28(12):96–106, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jun:1993:FDA

[909] Yong-Kee Jun and Kern Koh. Onthe-fly detection of access anomalies in nested parallel loops. ACM SIG-PLAN Notices, 28(12):107–117, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jun:1993:ODA

[910] Yong-Kee Jun and Kern Koh. Onthe-fly detection of access anomolies in nested parallel loops. ACM SIG-PLAN Notices, 28(12):107–117, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Damodaran-Kamal:1993:NTD

[911] Suresh K. Damodaran-Kamal and Joan M. Francioni. Nondeterminancy: testing and debugging in message passing parallel programs. ACM SIG-PLAN Notices, 28(12):118–128, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mellor-Crummey:1993:CSE

[912] John Mellor-Crummey. Compile-time support for efficient data race detection in shared-memory parallel programs. ACM SIGPLAN Notices, 28(12):129– 139, December 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mellor-Crummey:1993:CTS

[913] John Mellor-Crummey. Compile-time support for efficient data race detection in shared-memory parallel programs. ACM SIGPLAN Notices, 28(12):129– 139, December 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crovella:1993:PDU

[914] Mark E. Crovella and Thomas J. LeBlanc. Performance debugging using parallel performance predicates. ACM SIGPLAN Notices, 28(12):140– 150, December 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Timmerman:1993:HLT

[915] M. Timmerman, Frank Gielen, and P. Lambrix. High level tools for the debugging of real time multiprocessor systems. ACM SIGPLAN Notices, 28(12):151–157, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brewer:1993:DPA

[916] Eric A. Brewer and William E. Weihl. Developing parallel applications using high-performance simulation. ACM SIGPLAN Notices, 28(12):158–168, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pancake:1993:BPD

[917] Cherri M. Pancake and Robert H. B. Netzer. A bibliography of parallel debuggers, 1993 edition. ACM SIG-PLAN Notices, 28(12):169–186, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ACM/ONR Workshop on Parallel and Distributed Debugging.

Brown:1993:SC

[918] Jeff Brown, B. Irvin, and G. Forman. Session 6. chair. ACM SIGPLAN Notices, 28(12):xvi–??, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cargille:1993:OFM

[919] J. Cargille, A. Tomlinson, G. Forman, and F. Gielen. Open forum. moderator, Bart Miller, speaker, Pat Helland. the users’ view of debugging. ACM SIG-PLAN Notices, 28(12):xviii–??, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choi:1993:SC

[920] Jong Choi, J. Lumpp, and F. Gielen. Session 5. chair. ACM SIGPLAN Notices, 28(12):xv–??, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Francioni:1993:SC

[921] Joan Francioni, S. Damodaran-Kamal, J. Hollingsworth, and C. McDowell. Session 3. chair. ACM SIGPLAN Notices, 28(12):ix–??, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hollingsworth:1993:PVY

[922] J. Hollingsworth, S. Damodaran-Kamal, B. Irvin, and J. Lumpp. Program visualization: Are you drawing useful pictures? ACM SIGPLAN Notices, 28(12):xi–??, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marzullo:1993:SC

[923] Keith Marzullo, G. Forman, and J. Lumpp. Session 2. chair. ACM SIG-PLAN Notices, 28(12):viii–??, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1993:SC

[924] Bart Miller, F. Gielen, and Suresh K. Damodaran-Kamal. Session 1. chair. ACM SIGPLAN Notices, 28(12):vii– ??, December 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Recd:1993:SC

[925] Dan Recd, A. Tomlinson, and J. Cargille. Session 4. chair. ACM SIG-PLAN Notices, 28(12):xiii–??, December 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radenski:1994:MTM

[926] Atanas Radenski. Module types, module variables, and their use as a universal encapsulation mechanism. ACM SIGPLAN Notices, 29(1):3–8, January 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kokol:1994:SCP

[927] P. Kokol. The self-similarity and computer programs. ACM SIGPLAN Notices, 29(1):9–12, January 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kokol:1994:SSC

[928] Peter Kokol. The self-similarity and computer programs. ACM SIGPLAN Notices, 29(1):9–12, January 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:1994:CRP

[929] Ming-Yuan Zhu. Computational reflection in Power Epsilon. ACM SIG-PLAN Notices, 29(1):13–19, January 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zeng:1994:EPA

[930] Li-Xin Zeng and Ming-Yuan Zhu. Extending PowerEpsilon with algebraic rewriting. ACM SIGPLAN Notices, 29 (1):20–26, January 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kans:1994:UAP

[931] Aaron Kans and Clive Hayton. Using ABC to prototype VDM specifications. ACM SIGPLAN Notices, 29(1): 27–36, January 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reiser:1994:PPP

[932] John F. Reiser and Joseph P. Skudlarek. Program profiling problems, and a solution via machine language rewriting. ACM SIGPLAN Notices, 29(1): 37–45, January 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Conway:1994:PCC

[933] Damian Conway. Parsing with C++ classes. ACM SIGPLAN Notices, 29 (1):46–52, January 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Durst:1994:UPC

[934] Martin J. Dürst. Uniprep: preparing a C/C++ compiler for Unicode. ACM SIGPLAN Notices, 29(1):53, January 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alblas:1994:BPP

[935] Henk Alblas, Rieks op den Akker, Paul Oude Luttighuis, and Klaas Sikkel. A bibliography on parallel parsing. ACM SIGPLAN Notices, 29(1): 54–65, January 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1994:LLQ

[936] Henry G. Baker, Jr. A “linear logic” Quicksort. ACM SIGPLAN Notices, 29(2):13–18, February 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:1994:PBM

[937] Genjiang Zhu, Li Xie, and Zhongxiu Sun. A path-based method of parallelizing C++ programs. ACM SIG-PLAN Notices, 29(2):19–24, February 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:1994:REA

[938] K. Zhang. A review of exploitation of AND-parallelism and combined AND/OR-parallelism in logic programs. ACM SIGPLAN Notices, 29(2):25–32, February 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:1994:REP

[939] Kang Zhang. A review of exploitation of AND-parallelism and combined AND/OR-parallelism in logic programs. ACM SIGPLAN Notices, 29(2):25–32, February 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bantchev:1994:TSD

[940] Boyko B. Bantchev. Terminable statements and destructive computation. ACM SIGPLAN Notices, 29(2):33– 38, February 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brown:1994:NLT

[941] Bob Brown. Non-linear type extensions. ACM SIGPLAN Notices, 29(2): 39–43, February 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brown:1994:NTE

[942] B. Brown. Nonlinear type extensions. ACM SIGPLAN Notices, 29(2): 39–43, February 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1994:DSE

[943] A. Michael Berman. Does Scheme enhance an introductory programming course?: some preliminary empirical results. ACM SIGPLAN Notices, 29(2): 44–48, February 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Black:1994:FOO

[944] Andrew Black and Jens Palsberg. Foundations of object-oriented languages. ACM SIGPLAN Notices, 29 (3):3–11, March 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1994:PMC

[945] C. K. Yuen and M. D. Feng. Parallel multiplication: a case study in parallel programming. ACM SIG-PLAN Notices, 29(3):12–17, March 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srinivasan:1994:CLA

[946] S. Srinivasan. A critical look at some Ada features. ACM SIGPLAN Notices, 29(3):18–22, March 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srinivasan:1994:CLS

[947] S. Srinivasan. A critical look at some Ada features. ACM SIGPLAN Notices, 29(3):18–22, March 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1994:PPL

[948] C. K. Yuen. Programming the premature loop exit: from functional to navigational. ACM SIGPLAN Notices, 29 (3):23–27, March 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakharov:1994:PCAa

[949] Alexander Sakharov. Propagation of Constants and Assertions. ACM SIG-PLAN Notices, 29(3):28–32, March 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rizzo:1994:UPC

[950] Michael Rizzo. Using producer and consumer manipulators to extend stream I/O formatting in C++. ACM SIGPLAN Notices, 29(3):33–35, March 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anjaneyulu:1994:BAP

[951] K. S. R. Anjaneyulu. Bug analysis of Pascal programs. ACM SIGPLAN Notices, 29(4):15–22, April 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frost:1994:UMA

[952] R. A. Frost. Using memoization to achieve polynomial complexity of purely functional executable specifications of non-deterministic top-down parsers. ACM SIGPLAN Notices, 29 (4):23–30, April 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dekeyser:1994:GDL

[953] J.-L. Dekeyser, D. Lazure, and P. Marquet. A geometrical data-parallel language. ACM SIGPLAN Notices, 29 (4):31–40, April 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dekeyser:1994:GDP

[954] Jean-Luc Dekeyser, Dominique Lazure, and Philippe Marquet. A geometrical data-parallel language. ACM SIGPLAN Notices, 29(4):31–40, April 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Loeliger:1994:DIO

[955] Jon Loeliger and R. Metzger. Developing an interprocedural optimizing compiler. ACM SIGPLAN Notices, 29 (4):41–48, April 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1994:PLT

[956] Brian L. Meek. Programming languages: towards greater commonality. ACM SIGPLAN Notices, 29(4):49–57, April 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1994:TGC

[957] Henry G. Baker. Thermodynamics and garbage collection. ACM SIGPLAN Notices, 29(4):58–63, April 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakharov:1994:PCAb

[958] Alexander Sakharov. Propagation of constants and assertions. ACM SIG-PLAN Notices, 29(5):3–6, 6, May 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cisneros:1994:CR

[959] Gerardo Cisneros. Configurable REC. ACM SIGPLAN Notices, 29(5):7–16, May 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andersen:1994:GFG

[960] Birger Andersen. A general, finegrained, machine independent, objectoriented language. ACM SIGPLAN Notices, 29(5):17–26, May 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andersen:1994:GFM

[961] B. Andersen. A general, fine-grained, machine independent, object-oriented language. ACM SIGPLAN Notices, 29 (5):17–26, May 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burson:1994:CC

[962] Scott L. Burson. Continuations without copying. ACM SIGPLAN Notices, 29(5):27–30, May 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sharafuddin:1994:PLD

[963] Ahsan J. Sharafuddin, Nathan Ida, and James E. Grover. Passing large dynamic objects out of function frame boundaries: the temporary linked list method. ACM SIGPLAN Notices, 29 (5):31–36, May 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zivkova:1994:ETO

[964] Emilia Zivkova. Experience in teaching object-oriented programming with C++. ACM SIGPLAN Notices, 29 (5):37–40, May 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manwu:1994:ALN

[965] Xu Manwu, Lu Jianfeng, Zeng Fancong, and Dai Jingwen. Agent language NUML and its reduction implementation model based on HOπ. ACM SIGPLAN Notices, 29(5):41–48, May 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:1994:ALN

[966] Manwu Xu, Jianfeng Lu, Fancong Zeng, and Jingwen Dai. Agent language NUML and its reduction implementation model based on HO pi. ACM SIGPLAN Notices, 29(5):41–48, May 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sabry:1994:CPU

[967] Amr Sabry and Matthias Felleisen. Is continuation-passing useful for data flow analysis? ACM SIGPLAN Notices, 29(6):1–12, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p1-sabry/

.

Appel:1994:SCS

[968] Andrew W. Appel and David B. Mac-Queen. Separate compilation for Standard ML. ACM SIGPLAN Notices, 29 (6):13–23, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p13-appel/

.

Launchbury:1994:LFS

[969] John Launchbury and Simon L. Peyton Jones. Lazy functional state threads. ACM SIGPLAN Notices, 29 (6):24–35, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

proceedings/pldi/178243/p24-launchbury/

.

Ebcioglu:1994:VCT

[970] Kemal Ebcioğlu, Randy D. Groves, Ki-Chang Kim, Gabriel M. Silberman, and Isaac Ziv. VLIW compilation techniques in a superscalar environment. ACM SIGPLAN Notices, 29 (6):36–48, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p36-ebcioglu/

.

Srivastava:1994:LTO

[971] Amitabh Srivastava and David W. Wall. Link-time optimization of address calculation on a 64–bit architecture. ACM SIGPLAN Notices, 29 (6):49–60, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p49-srivastava/

.

Granlund:1994:DII

[972] Torbjörn Granlund and Peter L. Montgomery. Division by invariant integers using multiplication. ACM SIGPLAN Notices, 29(6):61– 72, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL ftp://

ftp.cwi.nl/pub/pmontgom/divcnst.
psa4.gz;   ftp://ftp.cwi.nl/pub/
pmontgom/divcnst.psl.gz;   http:
//www.acm.org:80/pubs/citations/

                                        proceedings/pldi/178243/p61-granlund/

.

Wang:1994:PCT

[973] Ko-Yang Wang. Precise compile-time performance prediction for superscalarbased computers. ACM SIGPLAN Notices, 29(6):73–84, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p73-wang/

.

Wagner:1994:ASE

[974] Tim A. Wagner, Vance Maverick, Susan L. Graham, and Michael A. Harrison. Accurate static estimators for program optimization. ACM SIGPLAN Notices, 29(6):85– 96, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p85-wagner/

.

Krall:1994:ISS

[975] Andreas Krall. Improving semi-static branch prediction by code replication. ACM SIGPLAN Notices, 29(6): 97–106, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p97-krall/

.

vonHanxleden:1994:GTB

[976] Reinhard von Hanxleden and Ken Kennedy. GIVE-N-TAKE — a balanced code placement framework. ACM SIGPLAN Notices, 29(6):107– 120, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p107-von_
hanxleden/.

Pugh:1994:CSP

[977] William Pugh. Counting solutions to Presburger formulas: How and why. ACM SIGPLAN Notices, 29(6): 121–134, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p121-pugh/

.

Fisher:1994:PCS

[978] Allan L. Fisher and Anwar Mohammed Ghuloum. Parallelizing complex scans and reductions. ACM SIGPLAN Notices, 29(6):135–146, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p135-fisher/

.

Knoop:1994:PDC

[979] Jens Knoop, Oliver Rüthing, and Bernhard Steffen. Partial dead code elimination. ACM SIGPLAN Notices, 29 (6):147–158, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/pldi/178243/p147-knoop/

.

Briggs:1994:EPR

[980] Preston Briggs and Keith D. Cooper. Effective partial redundancy elimination. ACM SIGPLAN Notices, 29(6): 159–170, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p159-briggs/

.

Johnson:1994:PST

[981] Richard Johnson, David Pearson, and Keshav Pingali. The program structure tree: Computing control regions in linear time. ACM SIGPLAN Notices, 29(6):171–185, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p171-johnson/

.

Davidson:1994:MAC

[982] Jack W. Davidson and Sanjay Jinturkar. Memory access coalescing: A technique for eliminating redundant memory accesses. ACM SIGPLAN Notices, 29(6):186–195, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p186-davidson/

.

Srivastava:1994:ASB

[983] Amitabh Srivastava and Alan Eustace. ATOM: A system for building customized program analysis tools. ACM SIGPLAN Notices, 29(6):196– 205, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p196-srivastava/

.

Reinhold:1994:CPG

[984] Mark B. Reinhold. Cache performance of garbage-collected programs. ACM SIGPLAN Notices, 29(6):206– 217, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p206-reinhold/

.

Hummel:1994:GDD

[985] Joseph Hummel, Laurie J. Hendren, and Alexandru Nicolau. A general data dependence test for dynamic, pointer-based data structures. ACM SIGPLAN Notices, 29(6):218– 229, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p218-hummel/

.

Deutsch:1994:IMA

[986] Alain Deutsch. Interprocedural mayalias analysis for pointers: beyond klimiting. ACM SIGPLAN Notices, 29 (6):230–241, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p230-deutsch/

.

Emami:1994:CIP

[987] Maryam Emami, Rakesh Ghiya, and Laurie J. Hendren. Context-sensitive interprocedural points-to analysis in the presence of function pointers. ACM SIGPLAN Notices, 29(6):242– 256, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p242-emami/

.

Emami:1994:CSI

[988] Maryam Emami, Rakesh Ghiya, and Laurie J. Hendren. Context-sensitive interprocedural points-to analysis in the presence of function pointers. ACM SIGPLAN Notices, 29(6):242– 256, June 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kurlander:1994:ZCR

[989] Steven M. Kurlander and Charles N. Fischer. Zero-cost range splitting. ACM SIGPLAN Notices, 29(6):257– 265, June 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kurlander:1994:ZRS

[990] Steven M. Kurlander and Charles N. Fischer. Zero-cost range splitting.

ACM SIGPLAN Notices, 29(6):257– 265, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p257-kurlander/

.

Norris:1994:RAP

[991] Cindy Norris and Lori L. Pollock. Register allocation over the program dependence graph. ACM SIGPLAN Notices, 29(6):266–277, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p266-norris/

.

Wismuller:1994:DGO

[992] Roland Wismüller. Debugging of globally optimized programs using data flow analysis. ACM SIGPLAN Notices, 29(6):278–289, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p278-wismuller/

.

Austin:1994:EDA

[993] Todd M. Austin, Scott E. Breach, and Gurindar S. Sohi. Efficient detection of all pointer and array access errors. ACM SIGPLAN Notices, 29(6):290–301, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/pldi/178243/p290-austin/

.

Agrawal:1994:SPJ

[994] Hiralal Agrawal. On slicing programs with jump statements. ACM SIGPLAN Notices, 29(6):302–312, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p302-agrawal/

.

Netzer:1994:OTI

[995] Robert H. B. Netzer and Mark H. Weaver. Optimal tracing and incremental reexecution for debugging long-running programs. ACM SIGPLAN Notices, 29(6):313–325, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p313-netzer/

.

Holzle:1994:ODD

[996] Urs Hölzle and David Ungar. Optimizing dynamically-dispatched calls with run-time type feedback. ACM SIGPLAN Notices, 29(6):326–336, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/178243/p326-holzle/

.

Hentenryck:1994:TAP

[997] P. Van Hentenryck, A. Cortesi, and B. Le Charlier. Type analysis of Prolog using type graphs. ACM SIG-PLAN Notices, 29(6):337–348, June 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanHentenryck:1994:TAP

[998] P. Van Hentenryck, A. Cortesi, and B. Le Charlier. Type analysis of Prolog using type graphs. ACM SIGPLAN Notices, 29(6):337– 348, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p337-van_
hentenryck/.

VanHentenryck:1994:BTC

[999] Pascal Van Hentenryck and Viswanath Ramachandran. Backtracking without trailing in CLP(R_{Lin}). ACM SIGPLAN Notices, 29(6):349–360, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/178243/p349-van_
hentenryck/.

Kuhne:1994:HOO

[1000] Thomas Kühne. Higher order objects in pure object-oriented languages. ACM SIGPLAN Notices, 29(7):15–20, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grogono:1994:GMO

[1001] Peter Grogono and Mark Gargul. A graph model for object oriented programming. ACM SIGPLAN Notices, 29(7):21–28, July 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderLinden:1994:FMO

[1002] Frank J. van der Linden. Formal methods: from object-based to objectoriented. ACM SIGPLAN Notices, 29 (7):29–38, July 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davison:1994:KED

[1003] Andrew Davison. KL1 and the early days of the FGCS project. ACM SIG-PLAN Notices, 29(7):39–41, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:1994:CPA

[1004] Pei-Chi Wu and Feng-Jian Wang. Is circularity problem for attribute grammars exponential-time complete? ACM SIGPLAN Notices, 29(7):42, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kirner:1994:DCR

[1005] Tereza G. Kirner. Detection of cycle in real-time system specification. ACM SIGPLAN Notices, 29(7):43–50, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolfe:1994:CFG

[1006] M. Wolfe. J^{+} = J (control flow graphs). ACM SIGPLAN Notices, 29 (7):51–53, July 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolfe:1994:J

[1007] Michael Wolfe. J+=J. ACM SIG-PLAN Notices, 29(7):51–53, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dyadkin:1994:MP

[1008] Lev J. Dyadkin. Multibox parsers.

ACM SIGPLAN Notices, 29(7):54–60, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baquero:1994:CAC

[1009] Carlos Baquero and Francisco Moura.

Concurrency annotations in C++. ACM SIGPLAN Notices, 29(7):61–67, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Auerbach:1994:CSR

[1010] Joshua S. Auerbach and James R. Russell. The Concert Signature Representation: IDL as intermediate language. ACM SIGPLAN Notices, 29(8):1–12, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/185084/p1-auerbach/

.

Sankar:1994:AID

[1011] Sriram Sankar and Roger Hayes. ADL — an interface definition language for specifying and testing software. ACM SIGPLAN Notices, 29(8):13–21, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/185084/p13-sankar/

.

Katiyar:1994:PSI

[1012] Dinesh Katiyar, David Luckham, John Mitchell, and Sigurd Meldal. Polymorphism and subtyping in interfaces. ACM SIGPLAN Notices, 29(8):22–34, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/185084/p22-katiyar/

.

Allen:1994:BDU

[1013] Robert Allen and David Garlan. Beyond definition/use: Architectural interconnection. ACM SIGPLAN Notices, 29(8):35–45, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/185084/p35-allen/

.

Freitag:1994:PAS

[1014] Burkhard Freitag, Tiziana Margaria, and Bernhard Steffen. A pragmatic approach to software synthesis. ACM SIGPLAN Notices, 29(8):46–58, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/185084/p46-freitag/

.

Russell:1994:CIL

[1015] James R. Russell, Robert E. Strom, and Daniel M. Yellin. A checkable interface language for pointerbased structures. ACM SIGPLAN Notices, 29(8):59–73, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/185084/p59-russell/

.

Tan:1994:ILS

[1016] Yang Meng Tan. Interface language for supporting programming styles. ACM SIGPLAN Notices, 29(8):74–83, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/185084/p74-tan/

.

Cohen:1994:API

[1017] Don Cohen, Neil Goldman, and K. Narayanaswamy. Adding performance information to ADT interfaces: Why and how. ACM SIG-PLAN Notices, 29(8):84–93, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/185084/p84-cohen/

.

Kessler:1994:CSI

[1018] Peter B. Kessler. A client-side stub interpreter. ACM SIGPLAN Notices, 29(8):94–100, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/185084/p94-kessler/

.

Gay:1994:IDL

[1019] David E. Gay. Interface definition language conversions: Recursive types. ACM SIGPLAN Notices, 29(8):101–110, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/185084/p101-gay/

.

Kahrs:1994:IEM

[1020] Stefan Kahrs, Donald Sannella, and Andrzej Tarlecki. Interfaces and extended ML. ACM SIGPLAN Notices, 29(8):111–118, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/185084/p111-kahrs/

.

Hamilton:1994:UII

[1021] Graham Hamilton and Sanjay Radia. Using interface inheritance to address problems in system software evolution. ACM SIGPLAN Notices, 29(8):119–128, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/185084/p119-hamilton/

.

Leavens:1994:IIS

[1022] Gary T. Leavens. Inheritance of interface specifications (extended abstract). ACM SIGPLAN Notices, 29(8):129–138, August 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/185084/p129-leavens/

.

Bevan:1994:LEC

[1023] Stephen J. Bevan. Letter to the editor: SIGPLAN Notices 27(12) 1992: “The Calculation of Easter”. ACM SIGPLAN Notices, 29(9):8, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [629].

Conway:1994:PCD

[1024] Damian Conway. Parsing with C++ deferred expressions. ACM SIGPLAN Notices, 29(9):9–16, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Romanovsky:1994:ACC

[1025] A. Romanovsky. About conversations for concurrent OO languages. ACM SIGPLAN Notices, 29(9):17– 21, September 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Giraud-Carrier:1994:RDF

[1026] Christophe Giraud-Carrier. A reconfigurable data flow machine for implementing functional programming languages. ACM SIGPLAN Notices, 29 (9):22–28, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Giraud-Carrier:1994:RDM

[1027] Christophe Giraud-Carrier. A reconfigurable dataflow machine for implementing functional programming languages. ACM SIGPLAN Notices, 29 (9):22–28, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Justice:1994:MAC

[1028] Timothy P. Justice, Rajeev K. Pandey, and Timothy A. Budd. A multiparadigm approach to compiler construction. ACM SIGPLAN Notices, 29 (9):29–37, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1994:MRC

[1029] Henry G. Baker. Minimizing reference count updating with deferred and anchored pointers for functional data structures. ACM SIGPLAN Notices, 29(9):38–43, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klug:1994:BOV

[1030] Michael Klug. Basic operations of the VisiCola scope model. ACM SIGPLAN Notices, 29(9):44–50, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1994:BFS

[1031] C. K. Yuen and M. D. Feng. Breadthfirst search in the Eight Queens Problem. ACM SIGPLAN Notices, 29(9): 51–55, September 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Desainte-Catherine:1994:UAG

[1032] M. Desainte-Catherine and K. Barbar. Using attribute grammars to find solutions for musical equational programs. ACM SIGPLAN Notices, 29(9): 56–63, September 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:1994:PST

[1033] Pei-Chi Wu and Feng-Jian Wang.

Padded string: Treating string as sequence of machine words. ACM SIG-PLAN Notices, 29(9):64–67, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:1994:PUA

[1034] Pei-Chi Wu and Feng-Jian Wang. Pool:

an unbounded array. ACM SIGPLAN Notices, 29(9):68–71, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1994:PSL

[1035] Jing Li and Yulin Feng. A production system language KDOPS. ACM SIG-PLAN Notices, 29(9):72–76, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elmstrom:1994:IVS

[1036] René Elmstrøm, Peter Gorm Larsen, and Poul Bøgh Lassen. The IFAD VDM-SL toolbox: A practical approach to formal specifications. ACM SIGPLAN Notices, 29(9):77– 80, September 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristensen:1994:CMP

[1037] Bent Bruun Kristensen and Kasper Østerbye. Conceptual modeling and programming languages. ACM SIG-PLAN Notices, 29(9):81–90, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dekker:1994:GLC

[1038] Anthony H. Dekker. The game of Life: A CLEAN programming tutorial and case study. ACM SIGPLAN Notices, 29(9):91–104, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maruyama:1994:COO

[1039] Katsumi Maruyama and Nicholas Raguideau. Concurrent object-oriented language “COOL”. ACM SIGPLAN Notices, 29(9):105–114, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neto:1994:AAC

[1040] João José Neto. Adaptive automata for context-dependent languages. ACM SIGPLAN Notices, 29(9):115–124, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shelekhov:1994:OAP

[1041] Vladimir I. Shelekhov and Sergey V.

Kuksenko. Object analysis of program. ACM SIGPLAN Notices, 29(9):125– 134, September 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Umrigar:1994:FSD

[1042] Zerksis D. Umrigar. Fully static dimensional analysis with C++. ACM SIG-PLAN Notices, 29(9):135–139, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gu:1994:ABI

[1043] Weiming Gu, Jeffrey Vetter, and Karsten Schwan. An annotated bibliography of interactive program steering. ACM SIGPLAN Notices, 29(9):140– 148, September 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sankaran:1994:BGC

[1044] Nandakumar Sankaran. A bibliography on garbage collection and related topics. ACM SIGPLAN Notices, 29(9):149–158, September 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1994:TD

[1045] Brian Meek. A taxonomy of datatypes. ACM SIGPLAN Notices, 29(9):159– 167, September 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chambers:1994:TMM

[1046] Craig Chambers and Gary T. Leavens. Typechecking and modules for multimethods. ACM SIGPLAN Notices, 29 (10):1–15, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eifrig:1994:AOT

[1047] Jonathan Eifrig, Scott Smith, Valery Trifonov, and Amy Zwarico. Application of OOP type theory: State, decidability, integration. ACM SIGPLAN Notices, 29(10):16–30, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kobayashi:1994:TFC

[1048] N. Kobayashi and A. Yonezawa. Typetheoretic foundations for concurrent object-oriented programming. ACM SIGPLAN Notices, 29(10):31–45, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kobayashi:1994:TTF

[1049] Naoki Kobayashi and Akinori Yonezawa. Type-theoretic foundations for concurrent object-oriented programming. ACM SIGPLAN Notices, 29(10):31– 45, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1994:TPO

[1050] Judy Cohen, Mary Lynn Manns, Susan Lilly, Richard P. Gabriel, Janet Conway, and Desmond D’Souza. Training professionals in object technology (panel). ACM SIGPLAN Notices, 29 (10):46–50, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robertson:1994:OSG

[1051] Scott P. Robertson, John M. Carroll, Robert L. Mack, Mary Beth Rosson, Sherman R. Alpert, and Jürgen Koenenmann-Belliveau. ODE: A self-guided, scenario-based learning environment for object-oriented design principles. ACM SIGPLAN Notices, 29(10):51–64, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vion-Dury:1994:VII

[1052] Jean-Yves Vion-Dury and Miguel Santana. Virtual images: Interactive visualization of distributed object-oriented systems. ACM SIGPLAN Notices, 29 (10):65–84, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gronbaek:1994:BTH

[1053] Kaj Grønbæk and Jawahar Malhotra. Building tailorable hypermedia systems: the embedded interpreter approach. ACM SIGPLAN Notices, 29 (10):85–101, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1994:PBL

[1054] Randall B. Smith, Mark Lentczner, Walter R. Smith, Antero Taivalsaari, and David Ungar. Prototype-based languages: Object lessons from class-free programming (panel). ACM SIGPLAN Notices, 29(10):102–112, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Puaut:1994:DGC

[1055] Isabelle Puaut. A distributed garbage collector for active objects. ACM SIG-PLAN Notices, 29(10):113–128, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hagimont:1994:PSO

[1056] Daniel Hagimont, P.-Y. Chevalier, A. Freyssinet, S. Krakowiak, S. Lacourte, J. Mossière, and X. Rousset de Pina. Persistent shared object support in the Guide system: Evaluation and related work. ACM SIGPLAN Notices, 29(10):129–144, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shan:1994:SBW

[1057] Yen-Ping Shan, Ken Auer, Andrew J. Bear, Jim Adamczyk, Adele Goldberg, Tom Love, and Dave Thomas. Smalltalk in the business world: the good, the bad, and the future (panel). ACM SIGPLAN Notices, 29(10):145– 152, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seidewitz:1994:GVI

[1058] Ed Seidewitz. Genericity versus inheritance reconsidered: Self-reference using generics. ACM SIGPLAN Notices, 29(10):153–163, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ducournau:1994:PMM

[1059] R. Ducournau, M. Habib, M. Huchard, and M. L. Mugnier. Proposal for a monotonic multiple inheritance linearization. ACM SIGPLAN Notices, 29 (10):164–175, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yellin:1994:IPS

[1060] Daniel M. Yellin and Robert E.

Strom. Interfaces, protocols, and the semi-automatic construction of software adaptors. ACM SIGPLAN Notices, 29(10):176–190, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karpovich:1994:EFS

[1061] John F. Karpovich, Andrew S. Grimshaw, and James C. French. ExtensibLe File Systems (ELFS): An object-oriented approach to high performance file I/O. ACM SIGPLAN Notices, 29(10):191–204, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rothering:1994:DOI

[1062] Darryl James Rothering. Development of an OO infrastructure for mainframe database applications. ACM SIGPLAN Notices, 29(10):205–211, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinelli:1994:MOO

[1063] Andrea Spinelli, Paolo Salvaneschi, Mauro Cadei, and Marino Rocca. MI: an object oriented environment for integration of scientific applications. ACM SIGPLAN Notices, 29(10):212– 222, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Monarchi:1994:MSH

[1064] David Monarchi, Grady Booch, Brian Henderson-Sellers, Ivar Jacobson, Steve Mellor, James Rumbaugh, and Rebecca Wirfs-Brock. Methodology standards: Help or hindrance? ACM SIGPLAN Notices, 29(10):223–228, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holzle:1994:TGS

[1065] Urs Hölzle and David Ungar. A third generation SELF implementation: Reconciling responsiveness with performance. ACM SIGPLAN Notices, 29(10):229–243, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amiel:1994:OMD

[1066] E. Amiel, O. Gruber, and E. Simon. Optimizing multi-method dispatch using compressed dispatch tables. ACM SIGPLAN Notices, 29(10):244, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amiel:1994:OMM

[1067] Eric Amiel, Olivier Gruber, and Eric Simon. Optimizing multi-method dispatch using compressed dispatch tables. ACM SIGPLAN Notices, 29(10): 244–258, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lopez:1994:ICI

[1068] Gus Lopez, Bjorn Freeman-Benson, and Alan Borning. Implementing constraint imperative programming languages: The Kaleidoscope ’93 virtual machine. ACM SIGPLAN Notices, 29(10):259–271, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristensen:1994:CAA

[1069] Bent Bruun Kristensen. Complex associations: Abstractions in objectoriented modeling. ACM SIGPLAN Notices, 29(10):272–286, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jordan:1994:SPC

[1070] Richard Jordan, Ruth Smilan, and Alex Wilkinson. Streamlining the project cycle with object-oriented requirements. ACM SIGPLAN Notices, 29(10):287–300, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buhr:1994:TMO

[1071] R. J. A. Buhr and R. S. Casselman. Timethread-role maps for object-oriented design of real-time-and-distributed systems. ACM SIGPLAN Notices, 29(10):301, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buhr:1994:TRM

[1072] R. J. A. Buhr and R. S. Casselman. Timethread-role maps for object-oriented design of real-time-and-distributed systems. ACM SIGPLAN Notices, 29(10):301–316, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heuser:1994:DDC

[1073] Lutz Heuser, John Dilley, Hari Madduri, Steven Rabin, and Shawn Woods. Development of distributed and client/ server object-oriented applications: industry solutions. ACM SIGPLAN Notices, 29(10):317–323, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plevyak:1994:PCT

[1074] John Plevyak and Andrew A. Chien. Precise concrete type inference for object-oriented languages. ACM SIG-PLAN Notices, 29(10):324–340, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bogle:1994:RCD

[1075] Phillip Bogle and Barbara Liskov. Reducing cross domain call overhead using batched futures. ACM SIGPLAN Notices, 29(10):341–354, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agesen:1994:SGD

[1076] Ole Agesen and David Ungar. Sifting out the gold. Delivering compact applications from an exploratory object-oriented programming environment. ACM SIGPLAN Notices, 29(10): 355–370, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Atwood:1994:ODP

[1077] Thomas Atwood, Jnan Dash, Jacob Stein, Michael Stonebraker, and Mary Loomis. Objects and databases (panel). ACM SIGPLAN Notices, 29 (10):371–372, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beck:1994:PSW

[1078] Bob Beck and Steve Hartley. Persistent storage for a workflow tool implemented in Smalltalk. ACM SIGPLAN Notices, 29(10):373–387, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Resende:1994:SLO

[1079] Rodolfo F. Resende, Divyakant Agrawal, and Amr El Abbadi. Semantic locking in object-oriented database systems. ACM SIGPLAN Notices, 29(10): 388–402, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Onodera:1994:ERC

[1080] Tamiya Onodera. Experience with representing C++ program information in an object-oriented database. ACM SIGPLAN Notices, 29(10):403– 413, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carey:1994:SRO

[1081] Michael J. Carey, David J. De-Witt, Chander Kant, and Jeffrey F. Naughton. A status report on the OO7 OODBMS benchmarking effort. ACM SIGPLAN Notices, 29(10):414– 426, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Forman:1994:CBA

[1082] Ira R. Forman, Scott Danforth, and Hari Madduri. Composition of Before/ After metaclasses in SOM. ACM SIG-PLAN Notices, 29(10):427–439, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danforth:1994:RMP

[1083] Scott Danforth and Ira R. Forman. Reflections on metaclass programming in SOM. ACM SIGPLAN Notices, 29(10): 440–452, October 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lortz:1994:CCE

[1084] Victor B. Lortz and Kang G. Shin. Combining contracts and exemplarbased programming for class hiding and customization. ACM SIGPLAN Notices, 29(10):453–467, October 1994.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1994:HDT

[1085] Steven Fraser, Kent Beck, Grady Booch, Derek Coleman, Jim Coplien, Richard Helm, and Kenny Rubin. How do teams shape objects? — how do objects shape teams? ACM SIGPLAN Notices, 29(10):468–473, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thekkath:1994:SDC

[1086] Chandramohan A. Thekkath, Henry M.

Levy, and Edward D. Lazowska. Separating data and control transfer in distributed operating systems. ACM SIG-PLAN Notices, 29(11):2–11, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p2-thekkath/

.

Chandra:1994:SPM

[1087] Rohit Chandra, Scott Devine, Ben Verghese, Anoop Gupta, and Mendel Rosenblum. Scheduling and page migration for multiprocessor compute servers. ACM SIGPLAN Notices, 29(11):12–24, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p12-chandra/

.

Lim:1994:RSA

[1088] Beng-Hong Lim and Anant Agarwal. Reactive synchronization algorithms for multiprocessors. ACM SIG-PLAN Notices, 29(11):25–35, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p25-lim/

.

Heinlein:1994:IMP

[1089] John Heinlein, Kourosh Gharachorloo, Scott Dresser, and Anoop Gupta. Integration of message passing and shared memory in the Stanford FLASH multiprocessor. ACM SIGPLAN Notices, 29(11):38–50, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p38-heinlein/

.

Karamcheti:1994:SOM

[1090] Vijay Karamcheti and Andrew A. Chien. Software overhead in messaging layers: where does the time go? ACM SIGPLAN Notices, 29 (11):51–60, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p51-karamcheti/

.

Chandra:1994:WTS

[1091] Satish Chandra, James R. Larus, and Anne Rogers. Where is time spent in message-passing and sharedmemory programs? ACM SIG-PLAN Notices, 29(11):61–73, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p61-chandra/

.

Schmidt:1994:PHA

[1092] William J. Schmidt and Kelvin D.

Nilsen. Performance of a hardwareassisted real-time garbage collector. ACM SIGPLAN Notices, 29(11):76– 85, November 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:1994:PHR

[1093] William J. Schmidt and Kelvin D.

Nilsen. Performance of a hardwareassisted real-time garbage collector. ACM SIGPLAN Notices, 29 (11):76–85, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p76-schmidt/

.

Wu:1994:ENM

[1094] Michael Wu and Willy Zwaenepoel.

eNVy: a non-volatile, main memory storage system. ACM SIG-PLAN Notices, 29(11):86–97, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p86-wu/

.

Wu:1994:ENV

[1095] Michael Wu and Willy Zwaenepoel. eNVy: a non-volatile, main memory storage system. ACM SIGPLAN Notices, 29(11):86–97, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Upton:1994:RAH

[1096] Michael Upton, Thomas Huff, Trevor Mudge, and Richard Brown. Resource allocation in a high clock rate microprocessor. ACM SIG-PLAN Notices, 29(11):98–109, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p98-upton/

.

Thekkath:1994:HSS

[1097] Chandramohan A. Thekkath and Henry M. Levy. Hardware and software support for efficient exception handling. ACM SIGPLAN Notices, 29(11):110–119, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p110-
                                        thekkath/.

Argade:1994:TMR

[1098] Pramod V. Argade, David K. Charles, and Craig Taylor. A technique for monitoring run-time dynamics of an operating system and a microprocessor executing user applications. ACM SIGPLAN Notices, 29 (11):122–131, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p122-
argade/.

Uhlig:1994:TDS

[1099] Richard Uhlig, David Nagle, Trevor Mudge, and Stuart Sechrest. Trapdriven simulation with Tapeworm II. ACM SIGPLAN Notices, 29(11):132– 144, November 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Uhlig:1994:TST

[1100] Richard Uhlig, David Nagle, Trevor Mudge, and Stuart Sechrest. Trapdriven simulation with Tapeworm II. ACM SIGPLAN Notices, 29 (11):132–144, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p132-
uhlig/.

Maynard:1994:CCC

[1101] Ann Marie Grizzaffi Maynard, Colette M. Donnelly, and Bret R. Olszewski. Contrasting characteristics and cache performance of technical and multi-user commercial workloads. ACM SIGPLAN Notices, 29 (11):145–156, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

proceedings/asplos/195473/p145-
maynard/.

Bershad:1994:ACM

[1102] Brian N. Bershad, Dennis Lee, Theodore H. Romer, and J. Bradley Chen. Avoiding conflict misses dynamically in large direct-mapped caches. ACM SIGPLAN Notices, 29 (11):158–170, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p158-
bershad/.

Talluri:1994:STP

[1103] Madhusudhan Talluri and Mark D. Hill. Surpassing the TLB performance of superpages with less operating system support. ACM SIGPLAN Notices, 29(11):171–182, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p171-
talluri/.

Gallagher:1994:DMD

[1104] David M. Gallagher, William Y. Chen, Scott A. Mahlke, John C. Gyllenhaal, and Wen mei W. Hwu. Dynamic memory disambiguation using the memory conflict buffer. ACM SIGPLAN Notices, 29(11):183–193, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p183-
gallagher/.

Hayashi:1994:AAS

[1105] Kenichi Hayashi, Tsunehisa Doi, Takeshi Horie, Yoichi Koyanagi, Osamu Shiraki, Nobutaka Imamura, Toshiyuki Shimizu, Hiroaki Ishihata, and Tatsuya Shindo. AP1000+: architectural support of PUT/GET interface for parallelizing compiler. ACM SIGPLAN Notices, 29(11): 196–207, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p196-
hayashi/.

Larus:1994:LMS

[1106] James R. Larus, Brad Richards, and Guhan Viswanathan. LCM: memory system support for parallel language implementation. ACM SIGPLAN Notices, 29(11):208–218, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p208-
larus/.

Woo:1994:PAI

[1107] Steven Cameron Woo, Jaswinder Pal Singh, and John L. Hennessy. The performance advantages of integrating block data transfer in cache-coherent multiprocessors. ACM SIGPLAN Notices, 29(11):219–229, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p219-
woo/.

Young:1994:IAS

[1108] Cliff Young and Michael D. Smith. Improving the accuracy of static branch prediction using branch correlation. ACM SIGPLAN Notices, 29 (11):232–241, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p232-
                                        young/.

Calder:1994:RBC

[1109] Brad Calder and Dirk Grunwald. Reducing branch costs via branch alignment. ACM SIGPLAN Notices, 29(11):242–251, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p242-
                                        calder/.

Carr:1994:COI

[1110] Steve Carr, Kathryn S. McKinley, and Chau-Wen Tseng. Compiler optimizations for improving data locality. ACM SIGPLAN Notices, 29 (11):252–262, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p252-
                                        carr/.

Engler:1994:DER

[1111] Dawson R. Engler and Todd A. Proebsting. DCG: an efficient, retargetable dynamic code generation system. ACM SIGPLAN Notices, 29 (11):263–272, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p263-
engler/.

Heinrich:1994:PIF

[1112] Mark Heinrich, Jeffrey Kuskin, David Ofelt, John Heinlein, Joel Baxter, Jaswinder Pal Singh, Richard Simoni, Kourosh Gharachorloo, David Nakahira, Mark Horowitz, Anoop Gupta, Mendel Rosenblum, and John Hennessy. The performance impact of flexibility in the Stanford FLASH multiprocessor. ACM SIGPLAN Notices, 29(11):274–285, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p274-
heinrich/.

Skeppstedt:1994:SCA

[1113] Jonas Skeppstedt and Per Stenström.

Simple compiler algorithms to reduce ownership overhead in cache coherence protocols. ACM SIGPLAN Notices, 29(11):286–296, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p286-
skeppstedt/.

Schoinas:1994:FAC

[1114] Ioannis Schoinas, Babak Falsafi, Alvin R. Lebeck, Steven K. Reinhardt, James R. Larus, and David A.

Wood. Fine-grain access control for distributed shared memory. ACM SIGPLAN Notices, 29 (11):297–306, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p297-
                                        schoinas/.

Schoinas:1994:FGA

[1115] Ioannis Schoinas, Babak Falsafi, Alvin R. Lebeck, Steven K. Reinhardt, James R. Larus, and David A. Wood. Fine-grain access control for distributed shared memory. ACM SIGPLAN Notices, 29(11):297–306, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laudon:1994:IMT

[1116] James Laudon, Anoop Gupta, and Mark Horowitz. Interleaving: a multithreading technique targeting multiprocessors and workstations. ACM SIGPLAN Notices, 29(11): 308–318, November 1994. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/195473/p308-
                                        laudon/.

Carter:1994:HSF

[1117] Nicholas P. Carter, Stephen W. Keckler, and William J. Dally. Hardware support for fast capabilitybased addressing. ACM SIGPLAN Notices, 29(11):319–327, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p319-
carter/.

Thekkath:1994:EMH

[1118] Radhika Thekkath and Susan J. Eggers. The effectiveness of multiple hardware contexts. ACM SIGPLAN Notices, 29(11):328–337, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/195473/p328-
thekkath/.

Kielmann:1994:WDS

[1119] Thilo Kielmann, Mira Mezini, and Arjan Loeffen. 4th Workshop for Doctoral Students in Object-Oriented Systems. ACM SIGPLAN Notices, 29(12): 7–10, December 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bestavros:1994:WLC

[1120] Azer Bestavros, Richard Gerber, and Stephen P. Masticola. Workshop on language, compiler, and tool support for real-time systems. ACM SIGPLAN Notices, 29(12):11–17, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:1994:ROO

[1121] Joseph Bergin. A report on objectoriented extensions to Pascal. ACM SIGPLAN Notices, 29(12):18–24, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ganesh:1994:FLB

[1122] Amit Ganesh. Fusing loops with backward inter loop data dependence. ACM SIGPLAN Notices, 29(12):25–30, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1994:SIR

[1123] Robert P. Wilson, Robert S. French, Christopher S. Wilson, Saman P. Amarasinghe, Jennifer M. Anderson, Steve W. K. Tjiang, Shih-Wei Liao, Chau-Wen Tseng, Mary W. Hall, Monica S. Lam, and John L. Hennessy. SUIF: an infrastructure for research on parallelizing and optimizing compilers. ACM SIGPLAN Notices, 29(12):31–37, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malloy:1994:EPR

[1124] Brian A. Malloy, John D. McGregor, Anand Krishnaswamy, and Murali Medikonda. An extensible program representation for object-oriented software. ACM SIGPLAN Notices, 29(12): 38–47, December 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Codenie:1994:AMP

[1125] Wim Codenie, Koen De Hondt, Theo D’Hondt, and Patrick Steyaert. Agora: message passing as a foundation for exploring OO language concepts. ACM SIGPLAN Notices, 29(12):48–57, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shang:1994:CS

[1126] David L. Shang. Covariant specification. ACM SIGPLAN Notices, 29(12): 58–65, December 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nelson:1994:CCC

[1127] Michael L. Nelson. Considerations in choosing a concurrent/distributed object-oriented programming language. ACM SIGPLAN Notices, 29 (12):66–71, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weisert:1994:CPP

[1128] Conrad Weisert. Comment on poor practice in coding examples. ACM SIGPLAN Notices, 29(12):72, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butt:1994:RDS

[1129] Farooq Butt. Rapid development of a source-level debugger for PowerPC microprocessors. ACM SIGPLAN Notices, 29(12):73–77, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agarwal:1994:CIO

[1130] Rakesh Agarwal. The C++ interface in objectivity. ACM SIGPLAN Notices, 29(12):78–86, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Skudlarek:1994:RMI

[1131] Joseph P. Skudlarek. Remarks on “A methodology for implementing highly concurrent data”. ACM SIGPLAN Notices, 29(12):87–93, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hahne:1994:UYL

[1132] Bruce Hahne and Hiroyuki Satō. Using YACC and Lex with C++. ACM SIGPLAN Notices, 29(12):94–103, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:1994:BCO

[1133] Thomas Wang. Better C: an Object-Oriented C language with automatic memory manager suitable for interactive applications. ACM SIGPLAN Notices, 29(12):104–111, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:1994:EMF

[1134] Thomas Wang. Eliminate memory fragmentation through holes in the heap. ACM SIGPLAN Notices, 29 (12):112–113, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:1995:LOC

[1135] Dianxiang Xu and Guoliang Zheng. Logical objects with constraints. ACM SIGPLAN Notices, 30(1):5–10, January 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhu:1995:PTC

[1136] Ming-Yuan Zhu. Program transformation in constructive type theory. ACM SIGPLAN Notices, 30(1):11–19, January 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:1995:NIM

[1137] Myung Ho Kim. A new iteration mechanism for the C++ programming language. ACM SIGPLAN Notices, 30(1): 20–26, January 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tin:1995:TSO

[1138] Erkan Tin, Varol Akman, and Murat Ersan. Towards situation-oriented programming languages. ACM SIG-PLAN Notices, 30(1):27–36, January 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tin:1995:TSP

[1139] E. Tin, V. Akman, and M. Ersan. Towards situation-oriented programming languages. ACM SIGPLAN Notices, 30 (1):27–??, January 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Copperman:1995:PMW

[1140] Max Copperman and Jeff Thomas.

Poor man’s watchpoints. ACM SIG-PLAN Notices, 30(1):37–44, January 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1995:UOV

[1141] Henry G. Baker. “use-once” variables and linear objects: storage management, reflection and multi-threading. ACM SIGPLAN Notices, 30(1):45–52, January 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1995:VLO

[1142] H. G. Baker. ‘use-once’ variables and linear objects-storage management, reflection and multi-threading. ACM SIGPLAN Notices, 30(1):45–??, January 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mellier:1995:OTD

[1143] Pierre Mellier and François Grize. OVIDE: A tool for data acquisition and validation. ACM SIGPLAN Notices, 30 (1):53–61, January 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maruyama:1995:NCS

[1144] Katsumi Maruyama. A name change from “COOL” to “ACOOL”. ACM SIGPLAN Notices, 30(2):2, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Castagna:1995:FOO

[1145] Giuseppe Castagna and Gary T. Leavens. Foundations of Object-Oriented Languages: 2nd Workshop report. ACM SIGPLAN Notices, 30(2):5–11, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raiha:1995:BLA

[1146] L. Raiha. A brief look at extension programming before and now. ACM SIG-PLAN Notices, 30(2):12–20, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raiha:1995:BLE

[1147] Liisa Räihä. A brief look at extension programming before and now. ACM SIGPLAN Notices, 30(2):12–20, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gallud:1995:SGM

[1148] J. A. Gallud and J. M. García. The specification of a generic multicomputer using Lotos. ACM SIGPLAN Notices, 30(2):21–24, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beaty:1995:PUO

[1149] Steven J. Beaty. ParsesraP: Using one grammar to specify both input and output. ACM SIGPLAN Notices, 30(2): 25–32, February 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boulanger:1995:OOM

[1150] Jean-Louis Boulanger. Object oriented method for Axiom. ACM SIGPLAN Notices, 30(2):33–41, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wu:1995:SRM

[1151] Youfeng Wu. Strength reduction of multiplications by integer constants. ACM SIGPLAN Notices, 30(2):42– 48, February 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferenczi:1995:GMV

[1152] Szabolcs Ferenczi. Guarded methods vs. inheritance anomaly: inheritance anomaly solved by nested guarded method calls. ACM SIGPLAN Notices, 30(2):49–58, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nourani:1995:IMA

[1153] Cyrus F. Nourani. Intelligent and multi agent object level computing: the preliminary overview. ACM SIGPLAN Notices, 30(2):59–64, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sreedhar:1995:ICD

[1154] Vugranam C. Sreedhar, Guang R. Gao, and Yong fong Lee. Incremental computation of dominator trees. ACM SIGPLAN Notices, 30(3):1–12, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/202529/p1-sreedhar/

.

Kelsey:1995:CBC

[1155] Richard A. Kelsey. A correspondence between continuation passing style and static single assignment form. ACM SIGPLAN Notices, 30(3):13–22, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/202529/p13-kelsey/

.

Berson:1995:GGU

[1156] David A. Berson, Rajiv Gupta, and Mary Lou Soffa. GURRR: A global unified resource requirements representation. ACM SIG-PLAN Notices, 30(3):23–34, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/202529/p23-berson/

.

Click:1995:SGB

[1157] Cliff Click and Michael Paleczny. A simple graph-based intermediate representation. ACM SIGPLAN Notices, 30(3):35–49, March 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/202529/p35-click/

.

Ruf:1995:OSR

[1158] Erik Ruf. Optimizing sparse representations for dataflow analysis. ACM SIGPLAN Notices, 30(3):50–61, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/202529/p50-ruf/

.

Steensgaard:1995:SFS

[1159] Bjarne Steensgaard. Sparse functional stores for imperative programs. ACM SIGPLAN Notices, 30(3):62–70, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/202529/p62-steensgaard/

.

OBrien:1995:XYI

[1160] Kevin O’Brien, Kathryn M. O’Brien, Martin Hopkins, Arvin Shepherd, and Ron Unrau. XIL and YIL: The intermediate languages of TOBEY. ACM SIGPLAN Notices, 30(3):71–82, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/202529/p71-o_brien/

.

Aubry:1995:GDF

[1161] Pascal Aubry and Thierry Gautier. GC: The data-flow graph format of synchronous programming. ACM SIG-PLAN Notices, 30(3):83–93, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/202529/p83-aubry/

.

Chitnis:1995:RTI

[1162] Sachin V. Chitnis, Manoranjan Satpathy, and Sundeep Oberoi. Rationalized three instruction machine. ACM SIGPLAN Notices, 30(3):94–102, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/202529/p94-chitnis/

.

Broustaut:1995:VAC

[1163] Frédéric Broustaut, Christian Fabre, François de Ferrière, Éric Ivanov, and Mauro Fiorentini. Verification of ANDF components. ACM SIG-PLAN Notices, 30(3):103–110, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/202529/p103-broustaut/

.

Gosling:1995:JIB

[1164] James Gosling. Java intermediate bytecodes: ACM SIGPLAN workshop on intermediate representations (IR’95). ACM SIGPLAN Notices, 30(3):111–118, March 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/202529/p111-gosling/

.

Lewis:1995:CMR

[1165] Brian T. Lewis, L. Peter Deutsch, and Theodore C. Goldstein. Clarity MCode: A retargetable intermediate representation for compilation. ACM SIGPLAN Notices, 30(3):119–128, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/202529/p119-lewis/

.

Krauter:1995:BRTa

[1166] Klaus Krauter. Book review: Tool Integration: Environments and Frameworks, Edited by: D. Schefstrom and G. van den Broek. ACM SIG-PLAN Notices, 30(4):7, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Falanga:1995:BRO

[1167] Rod Falanga. Book review: ObjectWindows for C++: A Shortcut to Windows Programming by Robert J. Traister. ACM SIGPLAN Notices, 30(4):8, April 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grogono:1995:BRV

[1168] Peter Grogono. Book review: Visual Basic Professional 3.9 Programming by Thomas W. Torgerson. ACM SIGPLAN Notices, 30(4):9, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noble:1995:BRP

[1169] James Noble. Book review: Parallel Logic Programming by Akikazu Takeuchi. ACM SIGPLAN Notices, 30(4):10, April 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:1995:BRHa

[1170] Diomidis Spinellis. Book review: High-Speed Windows Applications: Multitasking Design Methods by Dr. Bruce E. Krell. ACM SIGPLAN Notices, 30(4):11, April 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Youngblut:1995:BRU

[1171] Christine Youngblut. Book review:

User Interface Software, edited by Len Bass and Prasun Dewan. ACM SIG-PLAN Notices, 30(4):12, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schilling:1995:DCU

[1172] Jonathan Schilling. Dynamicallyvalued constants: An underused language feature. ACM SIGPLAN Notices, 30(4):13–20, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schilling:1995:DVC

[1173] Jonathan L. Schilling. Dynamicallyvalued constants: an underused language feature. ACM SIGPLAN Notices, 30(4):13–20, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

English:1995:MC

[1174] John English. Multithreading in C++.

ACM SIGPLAN Notices, 30(4):21–28, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trono:1995:CTS

[1175] John A. Trono. A comparison of three strategies for computing letter oriented, minimal perfect hashing functions. ACM SIGPLAN Notices, 30(4):

29–35, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper introduces a simple modification that produces dramatic speedups to the algorithms of Cichelli [39] and Cook and Oldehoeft [42].

Kokol:1995:UIM

[1176] Peter Kokol, Ivan Rozman, and Vlado Venuti. User interface metrics. ACM SIGPLAN Notices, 30(4):36–38, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wong:1995:PQA

[1177] Limsoon Wong. Polymorphic queries across sets, bags, and lists. ACM SIGPLAN Notices, 30(4):39–44, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tessman:1995:AGF

[1178] Thant Tessman. Adding generic functions to Scheme. ACM SIGPLAN Notices, 30(4):45–50, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Becker:1995:EEP

[1179] George Becker and Neil V. Murray. Efficient execution of programs with static semantics. ACM SIGPLAN Notices, 30(4):51–60, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kleinrubatscher:1995:FPS

[1180] Paul Kleinrubatscher, Albert Kriegshaber, Robert Zöchling, and Robert Glück. Fortran program specialization. ACM SIGPLAN Notices, 30(4):61–70, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mernik:1995:LAT

[1181] Marjan Mernik, Nikolay Korbar, and Viljem Žumer. LISA: A tool for automatic language implementation. ACM SIGPLAN Notices, 30(4):71–79, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mernik:1995:LTA

[1182] Marjan Mernik, Nikolaj Korbar, and Viljem Žumer. LISA: a tool for automatic language implementation. ACM SIGPLAN Notices, 30(4):71–79, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cigas:1995:PPP

[1183] John F. Čigas. Proper packaging promotes parameter passing proficiency. ACM SIGPLAN Notices, 30(4):80, April 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stiglic:1995:HEO

[1184] Bruno Stiglic, Marjan Heričko, and Ivan Rozman. How to evaluate objectoriented software development? ACM SIGPLAN Notices, 30(5):3–10, May 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Terry:1995:UAM

[1185] P. D. Terry. Umbriel — another minimal programming language. ACM SIG-PLAN Notices, 30(5):11–17, May 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Syropoulos:1995:BIW

[1186] Apostolos Syropoulos and Alexandros Karakos. Bottom in the imperative world. ACM SIGPLAN Notices, 30 (5):18–20, May 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shang:1995:CDS

[1187] David L. Shang. Covariant deep subtyping reconsidered. ACM SIG-PLAN Notices, 30(5):21–28, May 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butt:1995:IFS

[1188] Farooq Butt. Implementing FOR-TRAN77 support in the GNU gdb debugger. ACM SIGPLAN Notices, 30 (5):29–36, May 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kokol:1995:PSM

[1189] Peter Kokol, Viljem Žumer, Janez Brest, and Marjan Mernik. PROMIS: A software metrics tool generator. ACM SIGPLAN Notices, 30(5):37–42, May 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borovikov:1995:SIO

[1190] Igor A. Borovikov. L-systems with inheritance: an object-oriented extension of L-systems. ACM SIGPLAN Notices, 30(5):43–60, May 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Borovikov:1995:LIO

[1191] Igor A. Borovikov. L-systems with inheritance: An object-oriented extension of L-systems. ACM SIG-PLAN Notices, 30(5):43–??, May 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilson:1995:ECP

[1192] Robert P. Wilson and Monica S.

Lam. Efficient context-sensitive pointer analysis for C programs. ACM SIGPLAN Notices, 30(6):1–12, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p1-wilson/

.

Wilson:1995:ECS

[1193] Robert P. Wilson and Monica S. Lam.

Efficient context-sensitive pointer analysis for C programs. ACM SIG-PLAN Notices, 30(6):1–12, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruf:1995:CAA

[1194] Erik Ruf. Context-insensitive alias analysis reconsidered. ACM SIG-PLAN Notices, 30(6):13–22, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p13-ruf/

.

Ruf:1995:CIA

[1195] Erik Ruf. Context-insensitive alias analysis reconsidered. ACM SIG-PLAN Notices, 30(6):13–22, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carini:1995:FIC

[1196] Paul R. Carini and Michael Hind. Flow-sensitive interprocedural constant propagation. ACM SIG-PLAN Notices, 30(6):23–31, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p23-carini/

.

Carini:1995:FSI

[1197] Paul R. Carini and Michael Hind. Flow-sensitive interprocedural constant propagation. ACM SIGPLAN Notices, 30(6):23–31, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pingali:1995:ADS

[1198] Keshav Pingali and Gianfranco Bilardi. APT: a data structure for optimal control dependence computation. ACM SIGPLAN Notices, 30(6):32–46, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p32-pingali/

.

Tu:1995:EBP

[1199] Peng Tu and David Padua. Efficient building and placing of gating functions. ACM SIGPLAN Notices, 30(6):47–55, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p47-tu/.

Mueller:1995:ACB

[1200] Frank Mueller and David B. Whalley. Avoiding conditional branches by code replication. ACM SIG-PLAN Notices, 30(6):56–66, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p56-mueller/

.

Patterson:1995:ASB

[1201] Jason R. C. Patterson. Accurate static branch prediction by value range propagation. ACM SIG-PLAN Notices, 30(6):67–78, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p67-patterson/

.

Calder:1995:CBS

[1202] Brad Calder, Dirk Grunwald, Donald Lindsay, James Martin, Michael Mozer, and Benjamin Zorn. Corpusbased static branch prediction. ACM SIGPLAN Notices, 30(6):79–92, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calder:1995:CSB

[1203] Brad Calder, Dirk Grunwald, Donald Lindsay, James Martin, Michael Mozer, and Benjamin G. Zorn. Corpusbased static branch prediction. ACM SIGPLAN Notices, 30(6):79–92, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p79-calder/

.

Dean:1995:SSO

[1204] Jeffrey Dean, Craig Chambers, and David Grove. Selective specialization for object-oriented languages. ACM SIGPLAN Notices, 30(6):93–102, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p93-dean/

.

Fernandez:1995:SEL

[1205] Mary F. Fernández. Simple and effective link-time optimization of Modula-3 programs. ACM SIG-PLAN Notices, 30(6):103–115, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p103-fernandez/

.

Shao:1995:TBC

[1206] Zhong Shao and Andrew W. Appel. A type-based compiler for Standard ML. ACM SIGPLAN Notices, 30(6):116–129, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p116-shao/

.

Burger:1995:RAU

[1207] Robert G. Burger, Oscar Waddell, and R. Kent Dybvig. Register allocation using lazy saves, eager restores, and greedy shuffling. ACM SIG-PLAN Notices, 30(6):130–138, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p130-burger/

.

Altman:1995:SMS

[1208] Erik R. Altman, R. Govindarajan, and Guang R. Gao. Scheduling and mapping: software pipelining in the presence of structural hazards. ACM SIG-PLAN Notices, 30(6):139–150, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p139-altman/

.

Lo:1995:IBS

[1209] Jack L. Lo and Susan J. Eggers.

Improving balanced scheduling with compiler optimizations that increase instruction-level parallelism. ACM SIGPLAN Notices, 30(6):151–162, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p151-lo/

.

Amagbegnon:1995:IDF

[1210] Pascalin Amagbégnon, Loïc Besnard, and Paul Le Guernic. Implementation of the data-flow synchronous language SIGNAL. ACM SIGPLAN Notices, 30 (6):163–173, June 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amagbegnon:1995:IDS

[1211] Pascalin Amagbégnon, Loïc Besnard, and Paul Le Guernic. Implementation of the data-flow synchronous language SIGNAL. ACM SIG-PLAN Notices, 30(6):163–173, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p163-amagbegnon/

.

Aiken:1995:BSM

[1212] Alexander Aiken, Manuel Fähndrich, and Raph Levien. Better static memory management: improving regionbased analysis of higher-order languages. ACM SIGPLAN Notices, 30(6):174–185, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p174-aiken/

.

Liao:1995:SAD

[1213] Stan Liao, Srinivas Devadas, Kurt Keutzer, Steve Tjiang, and Albert Wang. Storage assignment to decrease code size. ACM SIGPLAN Notices, 30(6):186–195, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p186-liao/

.

Krishnamurthy:1995:OPP

[1214] Arvind Krishnamurthy and Katherine Yelick. Optimizing parallel programs with explicit synchronization. ACM SIGPLAN Notices, 30(6):196–204, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p196-krishnamurthy/

.

Cierniak:1995:UDC

[1215] Michał Cierniak and Wei Li. Unifying data and control transformations for distributed shared-memory machines. ACM SIGPLAN Notices, 30(6):205–217, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p205-cierniak/

.

Rauchwerger:1995:LTS

[1216] Lawrence Rauchwerger and David Padua. The LRPD test: speculative run-time parallelization of loops with privatization and reduction parallelization. ACM SIG-PLAN Notices, 30(6):218–232, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p218-rauchwerger/

.

Knoop:1995:PAM

[1217] Jens Knoop, Oliver Rüthing, and Bernhard Steffen. The power of assignment motion. ACM SIG-PLAN Notices, 30(6):233–245, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p233-knoop/

.

Click:1995:GCM

[1218] Cliff Click. Global code motion: global value numbering. ACM SIG-PLAN Notices, 30(6):246–257, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p246-click/

.

Agrawal:1995:IPR

[1219] Gagan Agrawal, Joel Saltz, and Raja Das. Interprocedural partial redundancy elimination and its application to distributed memory compilation. ACM SIGPLAN Notices, 30(6):258–269, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

proceedings/pldi/207110/p258-agrawal/

.

Kolte:1995:ERA

[1220] Priyadarshan Kolte and Michael Wolfe.

Elimination of redundant array subscript range checks. ACM SIG-PLAN Notices, 30(6):270–278, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p270-kolte/

.

Coleman:1995:TSS

[1221] Stephanie Coleman and Kathryn S.

McKinley. Tile size selection using cache organization and data layout. ACM SIGPLAN Notices, 30(6):279–290, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p279-coleman/

.

Larus:1995:EME

[1222] James R. Larus and Eric Schnarr.

EEL: machine-independent executable editing. ACM SIGPLAN Notices, 30(6):291–300, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/207110/p291-larus/

.

Larus:1995:EMI

[1223] James R. Larus and Eric Schnarr. EEL:

machine-independent executable editing. ACM SIGPLAN Notices, 30(6):

291–300, June 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barrett:1995:GCU

[1224] David A. Barrett and Benjamin G. Zorn. Garbage collection using a dynamic threatening boundary. ACM SIGPLAN Notices, 30(6):301–314, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/207110/p301-barrett/

.

Ertl:1995:SCI

[1225] M. Anton Ertl. Stack caching for interpreters. ACM SIGPLAN Notices, 30(6):315–327, June 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/207110/p315-ertl/

.

Krauter:1995:BRTb

[1226] Klaus Krauter. Book review: Tool Integration: Environments and Frameworks, Edited by: D. Schefstrom and G. van den Broek. ACM SIGPLAN Notices, 30(7):2, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Na:1995:BRD

[1227] Piaw Na. Book review: Debugging: Creative Techniques and Tools for Software Repair, by Martin Stitt. ACM SIGPLAN Notices, 30(7):3, July 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:1995:BRHb

[1228] Diomidis Spinellis. Book review: High-Speed Windows Applications: Multitasking Design Methods, by Dr. Bruce E. Krell. ACM SIGPLAN Notices, 30(7):4, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raghavachari:1995:BRH

[1229] Mukund Raghavachari. Book review:

The High Performance Fortran Handbook by Charles Koelbel, David Loveman, Robert Schreiber, Guy Steele Jr., and Mary Zosel. ACM SIGPLAN Notices, 30(7):5, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sankaran:1995:BRC

[1230] Nandakumar Sankaran. Book review:

Categories for Types by Roy L. Crole. ACM SIGPLAN Notices, 30(7):6, July 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parsons:1995:BRP

[1231] Peter Parsons. Book review: Parallel Algorithms for Digital image Processing, Computer Vision and Neural Networks, by Ed. I. Pitas. ACM SIGPLAN Notices, 30(7):7, July 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kamin:1995:RWF

[1232] Samuel Kamin and Eric Golin. Report of a workshop on future directions in programming languages and compilers. ACM SIGPLAN Notices, 30(7):9– 28, July 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1995:RFD

[1233] Anonymous. Report: Future directions workshop. ACM SIGPLAN Notices, 30(7):9–??, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deVivo:1995:BEC

[1234] Marco de Vivo, Gabriela O. de Vivo, and Luis Gonzalez. A brief ESSAY on capabilities. ACM SIGPLAN Notices, 30(7):29–36, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Howatt:1995:PBA

[1235] James Howatt. A project-based approach to programming language evaluation. ACM SIGPLAN Notices, 30 (7):37–40, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Howatt:1995:PAP

[1236] James Howatt. A project-based approach to programming language evaluation. ACM SIGPLAN Notices, 30 (7):37–??, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hallett:1995:FUM

[1237] Jon Hallett. Formulating update messages. ACM SIGPLAN Notices, 30 (7):41–44, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Navrat:1995:RCA

[1238] Pavol Návrat and Maria Bieliková.

Representing calendrical algorithms and data in Prolog and Prolog III languages. ACM SIGPLAN Notices, 30 (7):45–51, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

XueDong:1995:IVI

[1239] Li XueDong and Zheng Guoliang. Introducing virtual instance variables in classes to provide sufficient support for encapsulation. ACM SIG-PLAN Notices, 30(7):52–56, July 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1995:IVI

[1240] XueDong Li and GuoLiang Zheng.

Introducing virtual instance variables in classes to provide sufficient support for encapsulation. ACM SIG-PLAN Notices, 30(7):52–??, July 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bordawekar:1995:MCS

[1241] Rajesh Bordawekar, Alok Choudhary, Ken Kennedy, Charles Koelbel, and Michael Paleczny. A model and compilation strategy for out-of-core data parallel programs. ACM SIGPLAN Notices, 30(8):1–10, August 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gotwals:1995:PSL

[1242] Jacob Gotwals, Suresh Srinivas, and Dennis Gannon. pC++/streams: A library for I/O on complex distributed data structures. ACM SIGPLAN Notices, 30(8):11–18, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Darlington:1995:PSS

[1243] John Darlington, Yi ke Guo, Hing Wing To, and Jin Yang. Parallel skeletons for structured composition. ACM SIG-PLAN Notices, 30(8):19–28, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carlisle:1995:SCC

[1244] Martin C. Carlisle and Anne Rogers. Software caching and computation migration in Olden. ACM SIGPLAN Notices, 30(8):29–38, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carlson:1995:DDA

[1245] William W. Carlson and Jesse M. Draper. Distributed data access in AC. ACM SIGPLAN Notices, 30(8):39–47, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andersson:1995:GPC

[1246] Niclas Andersson and Peter Fritzson. Generating parallel code from object oriented mathematical models. ACM SIGPLAN Notices, 30(8):48–57, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghuloum:1995:FPI

[1247] Anwar M. Ghuloum and Allan L.

Fisher. Flattening and parallelizing irregular, recurrent loop nests. ACM SIGPLAN Notices, 30(8):58–67, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mukherjee:1995:ESI

[1248] Shubhendu S. Mukherjee, Shamik D.

Sharma, Mark D. Hill, James R. Larus, Anne Rogers, and Joel Saltz. Efficient support for irregular applications on distributed-memory machines. ACM SIGPLAN Notices, 30(8):68–79, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brewer:1995:HLO

[1249] Eric A. Brewer. High-level optimization via automated statistical modeling. ACM SIGPLAN Notices, 30(8): 80–91, August 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brewer:1995:HOA

[1250] E. A. Brewer. High-level optimization via automated statistical modeling. ACM SIGPLAN Notices, 30(8): 80–91, August 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Collard:1995:FAD

[1251] Jean-François Collard, Denis Barthou, and Paul Feautrier. Fuzzy array dataflow analysis. ACM SIGPLAN Notices, 30(8):92–101, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kennedy:1995:LTA

[1252] Ken Kennedy, Nenad Nedeljkovic, and Ajay Sethi. A linear-time algorithm for computing the memory access sequence in data-parallel programs. ACM SIG-PLAN Notices, 30(8):102–111, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hwang:1995:AOS

[1253] Gwan-Hwan Hwang, Jenq Kuen Lee, and Dz-Ching Ju. An array operation synthesis scheme to optimize Fortran 90 programs. ACM SIGPLAN Notices, 30(8):112–122, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bader:1995:PAI

[1254] David A. Bader and Joseph JáJá. Parallel algorithms for image histogramming and connected components with an experimental study (extended abstract). ACM SIGPLAN Notices, 30 (8):123–133, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Subhlok:1995:OMS

[1255] Jaspal Subhlok and Gary Vondran. Optimal mapping of sequences of data parallel tasks. ACM SIGPLAN Notices, 30(8):134–143, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tseng:1995:COE

[1256] Chau-Wen Tseng. Compiler optimizations for eliminating barrier synchronization. ACM SIGPLAN Notices, 30 (8):144–155, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Philippsen:1995:AAA

[1257] Michael Philippsen. Automatic alignment of array data and processes to reduce communication time on DMPPs. ACM SIGPLAN Notices, 30(8):156– 165, August 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1995:DCT

[1258] Jennifer M. Anderson, Saman P. Amarasinghe, and Monica S. Lam. Data and computation transformations for multiprocessors. ACM SIGPLAN Notices, 30(8):166–178, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jeremiassen:1995:RFS

[1259] Tor E. Jeremiassen and Susan J. Eggers. Reducing false sharing on shared memory multiprocessors through compile time data transformations. ACM SIGPLAN Notices, 30(8):179–188, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spertus:1995:ELB

[1260] Ellen Spertus and William J. Dally.

Evaluating the locality benefits of active messages. ACM SIGPLAN Notices, 30(8):189–198, August 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wisniewski:1995:HPS

[1261] Robert W. Wisniewski, Leonidas I. Kontothanassis, and Michael L. Scott. High performance synchronization algorithms for multiprogrammed multiprocessors. ACM SIGPLAN Notices, 30(8):199–206, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blumofe:1995:CEM

[1262] Robert D. Blumofe, Christopher F. Joerg, Bradley C. Kuszmaul, Charles E. Leiserson, Keith H. Randall, and Yuli Zhou. Cilk: an efficient multithreaded runtime system. ACM SIGPLAN Notices, 30(8):207–216, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wallach:1995:OAM

[1263] Deborah A. Wallach, Wilson C. Hsieh, Kirk L. Johnson, M. Frans Kaashoek, and William E. Weihl. Optimistic active messages: a mechanism for scheduling communication with computation. ACM SIGPLAN Notices, 30 (8):217–226, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1995:CO

[1264] Anonymous. Conference organization. ACM SIGPLAN Notices, 30(8):III–??, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1995:ER

[1265] Anonymous. External referees. ACM SIGPLAN Notices, 30(8):IV–??, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maclean:1995:OOP

[1266] Stuart Maclean and Sean Smith.

Object-oriented programming for embedded systems. ACM SIGPLAN Notices, 30(9):3–8, September 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLean:1995:OPE

[1267] Stuart MacLean and Sean Smith.

Object-oriented programming for embedded systems. ACM SIGPLAN Notices, 30(9):3–??, September 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deVivo:1995:PAC

[1268] Gabriela O. de Vivo and Marco de Vivo. A pragmatic approach to C++, Eiffel and Ada 9X programming. ACM SIGPLAN Notices, 30(9): 9–16, September 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1995:CSC

[1269] Henry G. Baker. CONS should not CONS its arguments, Part II: Cheney on the M.T.A. ACM SIGPLAN Notices, 30(9):17–20, September 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1995:FL

[1270] C. K. Yuen. A functional FOR loop. ACM SIGPLAN Notices, 30(9): 21–24, September 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Traunmuller:1995:CPS

[1271] Karl Traunmüller. The checkers problem — a solution with linear time complexity. ACM SIGPLAN Notices, 30(9):25–32, September 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1995:WPC

[1272] Brian L. Meek. What is a procedure call? ACM SIGPLAN Notices, 30(9): 33–40, September 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meek:1995:WIP

[1273] Brian L. Meek. What is a procedure call? ACM SIGPLAN Notices, 30(9): 33–??, September 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:1995:IDD

[1274] Jian Lu. Introducing data decomposition into VDM for tractable development of programs. ACM SIGPLAN Notices, 30(9):41–50, September 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:1995:ESF

[1275] Arthur Allen and Dennis de Champeaux. Extending the statechart formalism: Event scheduling and disposition. ACM SIGPLAN Notices, 30 (10):1–16, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sane:1995:OOS

[1276] Aamod Sane and Roy Campbell.

Object-oriented state machines: subclassing, composition, delegation, and genericity. ACM SIGPLAN Notices, 30(10):17–32, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sane:1995:OSM

[1277] A. Sane and R. Campbell. Objectoriented state machines: Subclassing, composition, delegation and genericity. ACM SIGPLAN Notices, 30(10): 17–32, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malan:1995:LEL

[1278] Ruth Malan, Derek Coleman, and Reed Letsinger. Lessons from the experiences of leading-edge object technology projects in Hewlett–Packard. ACM SIGPLAN Notices, 30(10):33–46, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1995:SUI

[1279] Randall B. Smith, John Maloney, and David Ungar. The Self-4.0 user interface: Manifesting a system-wide vision of concreteness, uniformity, and flexibility. ACM SIGPLAN Notices, 30(10): 47–60, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1995:UPB

[1280] Walter R. Smith. Using a prototypebased language for user interface: the Newton project’s experience. ACM SIGPLAN Notices, 30(10):61–72, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1995:UPL

[1281] W. R. Smith. Using a prototype-based language for user interface: The Newton Project’s experience. ACM SIG-PLAN Notices, 30(10):61–72, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ungar:1995:AOT

[1282] David Ungar. Annotating objects for transport to other worlds. ACM SIG-PLAN Notices, 30(10):73–87, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1995:MOO

[1283] Laura Hill, Kenny Rubin, John Daniels, Charles Berman, James Coplien, and Doug Johnson. Managing object oriented projects (panel session). ACM SIGPLAN Notices, 30(10): 88–90, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agesen:1995:TFV

[1284] Ole Agesen and Urs Hölzle. Type feedback vs. concrete type inference: A comparison of optimization techniques for object-oriented languages. ACM SIGPLAN Notices, 30(10):91–107, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grove:1995:PGR

[1285] David Grove, Jeffrey Dean, Charles Garrett, and Craig Chambers. Profileguided receiver class prediction. ACM SIGPLAN Notices, 30(10):108–123, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grove:1995:PRC

[1286] D. Grove, J. Dean, C. Garrett, and C. Chambers. Profile-guided receiver class prediction. ACM SIGPLAN Notices, 30(10):108–123, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Myers:1995:BOL

[1287] Andrew C. Myers. Bidirectional object layout for separate compilation. ACM SIGPLAN Notices, 30(10):124– 139, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McGregor:1995:OTR

[1288] John D. McGregor, Ed Berard, Don Firesmith, Brian Marick, and Dave Thomson. OO testing in the real world (panel): lessons for all. ACM SIGPLAN Notices, 30(10):140, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Driesen:1995:MRD

[1289] Karel Driesen and Urs Hölzle. Minimizing row displacement dispatch tables. ACM SIGPLAN Notices, 30(10): 141–155, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Day:1995:SVW

[1290] Mark Day, Robert Gruber, Barbara Liskov, and Andrew C. Myers. Subtypes vs. where clauses: Constraining parametric polymorphism. ACM SIG-PLAN Notices, 30(10):156–168, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eifrig:1995:SPT

[1291] Jonathan Eifrig, Scott Smith, and Valery Trifonov. Sound polymorphic type inference for objects. ACM SIG-PLAN Notices, 30(10):169–184, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Armour:1995:TOA

[1292] Frank Armour, Todd Cotton, Geoff Hambrick, Barbara Moo, and Dennis Mancl. Tailoring OO analysis and design methods (panel). ACM SIGPLAN Notices, 30(10):185–186, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rossie:1995:ASS

[1293] Jonathan G. Rossie, Jr. and Daniel P.

Friedman. An algebraic semantics of subobjects. ACM SIGPLAN Notices, 30(10):187–199, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stata:1995:MRP

[1294] Raymie Stata and John V. Guttag.

Modular reasoning in the presence of subclassing. ACM SIGPLAN Notices, 30(10):200–214, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malenfant:1995:SDD

[1295] J. Malenfant. On the semantic diversity of delegation-based programming languages. ACM SIGPLAN Notices, 30(10):215–230, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1995:PCCb

[1296] Steven Fraser, Grady Booch, Frank Buschmann, Jim Coplien, Norm Kerth, Ivar Jacobson, and Mary Beth Rosson. Patterns: cult to culture? ACM SIG-PLAN Notices, 30(10):231–234, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ossher:1995:SCR

[1297] H. Ossher, M. Kaplan, W. Harrison, and A. Katz. Subject-oriented composition rules. ACM SIGPLAN Notices, 30(10):235–250, October 1995.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ossher:1995:SOC

[1298] Harold Ossher, Matthew Kaplan, William Harrison, Alexander Katz, and Vincent Kruskal. Subject-oriented composition rules. ACM SIGPLAN Notices, 30(10):235–250, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Riehle:1995:HWE

[1299] Dirk Riehle. How and why to encapsulate class trees. ACM SIGPLAN Notices, 30(10):251–264, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ducasse:1995:RMF

[1300] S. Ducasse, M. Blay-Fornarino, and A. M. Pinna-Dery. A reflective model for first class dependencies. ACM SIG-PLAN Notices, 30(10):265–280, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griss:1995:SSR

[1301] Martin Griss, Ted Biggerstaff, Sallie Henry, Ivar Jacobson, Doug Lea, and Will Tracz. Systematic software reuse (panel session): objects and frameworks are not enough. ACM SIGPLAN Notices, 30(10):281–282, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marcus:1995:FDO

[1302] Bob Marcus, Bob Atkinson, Chris Horn, and Hari Madduri. The future of distributed object computing (panel session). ACM SIGPLAN Notices, 30(10):283–284, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chiba:1995:MPC

[1303] Shigeru Chiba. A metaobject protocol for C++. ACM SIGPLAN Notices, 30(10):285–299, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Masuhara:1995:CML

[1304] Hidehiko Masuhara, Satoshi Matsuoka, Kenichi Asai, and Akinori Yonezawa. Compiling away the meta-level in object-oriented concurrent reflective languages using partial evaluation. ACM SIGPLAN Notices, 30(10):300– 315, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Masuhara:1995:CMO

[1305] H. Masuhara, S. Matsuoka, K. Asai, Yonezawa, and A. Compiling away the meta-level in object-oriented concurrent reflective languages using partial evaluation. ACM SIGPLAN Notices, 30(10):300–315, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mulet:1995:TME

[1306] Philippe Mulet, Jacques Malenfant, and Pierre Cointe. Towards a methodology for explicit composition of MetaObjects. ACM SIGPLAN Notices, 30(10):316–330, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cattell:1995:ODS

[1307] Rick Cattell, Frank Manola, Richard Soley, Jeff Sutherland, and Mary Loomis. Objects and database standards (panel). ACM SIGPLAN Notices, 30(10):331–332, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gossain:1995:ODE

[1308] Sanjiv Gossain, Don Batory, Hassan Gomaa, Mitch Lubars, Christopher Pidgeon, and Ed Seidewitz. Objects and domain engineering (panel). ACM SIGPLAN Notices, 30(10):333– 336, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Helm:1995:PP

[1309] Richard Helm. “patterns in practice”. ACM SIGPLAN Notices, 30(10):337– 341, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lange:1995:IVD

[1310] Danny B. Lange and Yuichi Nakamura. Interactive visualization of design patterns can help in framework understanding. ACM SIGPLAN Notices, 30(10):342–357, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hueni:1995:FNP

[1311] H. Hueni, R. Johnson, and R. Engel.

A framework for network protocol software. ACM SIGPLAN Notices, 30(10): 358–369, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huni:1995:FNP

[1312] Hermann Hüni, Ralph Johnson, and Robert Engel. A framework for network protocol software. ACM SIG-PLAN Notices, 30(10):358–369, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

AlbrechtSchmid:1995:CAM

[1313] H. Albrecht Schmid. Creating the architecture of a manufacturing framework by design patterns. ACM SIG-PLAN Notices, 30(10):370–384, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmid:1995:CAM

[1314] Hans Albrecht Schmid. Creating the architecture of a manufacturing framework by design patterns. ACM SIG-PLAN Notices, 30(10):370–384, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Itoh:1995:SUC

[1315] Jun ichiro Itoh, Yasuhiko Yokote, and Mario Tokoro. SCONE: Using concurrent objects for low-level operating system programming. ACM SIGPLAN Notices, 30(10):385–398, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristensen:1995:POM

[1316] A. Kristensen and C. Low. Problemoriented object memory: Customizing consistency. ACM SIGPLAN Notices, 30(10):399–413, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristensen:1995:POO

[1317] Anders Kristensen and Colin Low. Problem-oriented object memory: customizing consistency. ACM SIGPLAN Notices, 30(10):399–413, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barton:1995:EOW

[1318] Don Barton and Tom Arnold. Evolving to objects — the Witches’ Brew. ACM SIGPLAN Notices, 30(10):414– 425, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Forman:1995:IFR

[1319] I. R. Forman, M. H. Conner, S. H. Danforth, and L. K. Raper. Ira Forman, “Release-to-Release Binary Compatibility in SOM”. ACM SIGPLAN Notices, 30(10):426–438, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Forman:1995:RRB

[1320] Ira R. Forman, Michael H. Conner, Scott H. Danforth, and Larry K. Raper.

Release-to-release binary compatibility in SOM. ACM SIGPLAN Notices, 30(10):426–438, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vayda:1995:LB

[1321] Thomas P. Vayda. “lessons from the battlefield”. ACM SIGPLAN Notices, 30(10):439–452, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haines:1995:SOA

[1322] Matthew Haines, Piyush Mehrotra, and John Van Rosendale. Smart-Files: An OO approach to data file interoperability. ACM SIGPLAN Notices, 30(10):453–466, October 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiernan:1995:ESO

[1323] Jerry Kiernan and Michael J. Carey.

Extending SQL-92 for OODB access: Design and implementation experience. ACM SIGPLAN Notices, 30(10):467– 480, October 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huang:1995:PWE

[1324] T.-Y. Huang and W.-S. Liu. Predicting the worst-case execution time of the concurrent execution of instructions and cycle-stealing DMA I/O operations. ACM SIGPLAN Notices, 30(11): 1–6, November 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1995:SHA

[1325] Jongwon Lee, Sungyoung Lee, and Hyungill Kim. Scheduling of hard aperiodic tasks in hybrid static/ dynamic priority systems. ACM SIG-PLAN Notices, 30(11):7–19, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/216636/p7-lee/.

Nilsen:1995:WCE

[1326] Kelvin D. Nilsen and Bernt Rygg. Worst-case execution time analysis on modern processors. ACM SIGPLAN Notices, 30(11):20–30, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nilsen:1995:WET

[1327] Kelvin D. Nilsen and Bernt Rygg. Worst-case execution time analysis on modern processors. ACM SIG-PLAN Notices, 30(11):20–30, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/216636/p20-nilsen/

.

Bonfatti:1995:RSD

[1328] Flavio Bonfatti, Gianni Gadda, and Paola Daniela Monari. Re-usable software design for programmable logic controllers. ACM SIGPLAN Notices, 30(11):31–40, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

proceedings/plan/216636/p31-bonfatti/

.

Bonfatti:1995:RUS

[1329] Flavio Bonfatti, Gianni Gadda, and Paola Daniela Monari. Re-usable software design for programmable logic controllers. ACM SIGPLAN Notices, 30(11):31–40, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chung:1995:LCT

[1330] Tai M. Chung and Hank G. Dietz. Language constructs and transformation for hard real-time systems. ACM SIG-PLAN Notices, 30(11):41–49, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/216636/p41-chung/

.

Ren:1995:RLS

[1331] Shangping Ren and Gul A. Agha.

RTsynchronizer: language support for real-time specifications in distributed systems. ACM SIGPLAN Notices, 30(11):50–59, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/216636/p50-ren/

.

Thoen:1995:SSR

[1332] Filip Thoen, Marco Cornero, Gert Goossens, and Hugo De Man. Software synthesis for real-time information processing systems. ACM SIG-PLAN Notices, 30(11):60–69, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/216636/p60-thoen/

.

Campos:1995:VTQ

[1333] S. Campos, E. Clarke, W. Marrero, and M. Minea. Verus: a tool for quantitative analysis of finitestate real-time systems. ACM SIG-PLAN Notices, 30(11):70–78, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/216636/p70-campos/

.

Jourdan:1995:STA

[1334] M. Jourdan and F. Maraninchi. Static timing analysis of real-time systems. ACM SIGPLAN Notices, 30 (11):79–87, November 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/216636/p79-jourdan/

.

Li:1995:PAE

[1335] Yau-Tsun Steven Li and Sharad Malik. Performance analysis of embedded software using implicit path enumeration. ACM SIGPLAN Notices, 30(11):88–98, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/216636/p88-li/.

Ko:1995:SUF

[1336] Lo Ko, David B. Whalley, and Marion G. Harmon. Supporting user-friendly analysis of timing constraints. ACM SIGPLAN Notices, 30 (11):99–107, November 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/216636/p99-ko/.

Kopetz:1995:CCT

[1337] Hermann Kopetz and Roman Nossal. The Cluster Compiler — a tool for the design of time-triggered real-time systems. ACM SIGPLAN Notices, 30(11):108–116, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/216636/p108-kopetz/

.

Wilner:1995:WTU

[1338] David Wilner. WindView: a tool for understanding real-time embedded software through system vizualization. ACM SIGPLAN Notices, 30 (11):117–123, November 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/216636/p117-wilner/

.

Gerber:1995:AIB

[1339] Rich Gerber, Steve Tjiang, David Whalley, David Wilner, and Mike Wolfe. Appropriate interfaces between design tools, languages, compilers and runtimes in real-time systems (panel). ACM SIGPLAN Notices, 30(11):124, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/216636/p124-gerber/

.

Mueller:1995:CSS

[1340] Frank Mueller. Compiler support for software-based cache partitioning. ACM SIGPLAN Notices, 30 (11):125–133, November 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/216636/p125-mueller/

.

Bakkers:1995:AGS

[1341] André Bakkers, Johan Sunter, and Evert Ploeg. Automatic generation of scheduling and communication code in real-time parallel programs. ACM SIGPLAN Notices, 30 (11):134–145, November 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/216636/p134-bakkers/

.

Schneider:1995:CCB

[1342] Stan Schneider, Vincent Chen, Jay Steele, and Gerardo Pardo-Castellote. The ControlShell component-based real-time programming system, and its application to the Marsokhod Martian Rover. ACM SIGPLAN Notices, 30(11):146–155, November 1995.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:1995:CCR

[1343] Stan Schneider, Vincent Chen, Jay Steele, and Gerardo Pardo-Castellote. The ControlShell component-based real-time programming system, and its application to the Marsokhod Martian Rover. ACM SIGPLAN Notices, 30(11):146–155, November 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/216636/p146-schneider/

.

Weisert:1995:MCP

[1344] Conrad Weisert. Making C++ practical for business applications: 3 essential classes. ACM SIGPLAN Notices, 30(12):4–8, December 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1995:EPR

[1345] Per Brinch Hansen. Efficient parallel recursion. ACM SIGPLAN Notices, 30(12):9–16, December 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1995:RBO

[1346] Tao Li. A rule-based and objectoriented AI programming language. ACM SIGPLAN Notices, 30(12):17– 24, December 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1995:NAE

[1347] Gansheng Li. A new approach for efficient implementation of Ada multitasking. ACM SIGPLAN Notices, 30 (12):25–31, December 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramsdell:1995:CCS

[1348] John D. Ramsdell. CST: C state transformers. ACM SIGPLAN Notices, 30(12):32–36, December 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Genjiang:1995:NMD

[1349] Zhu Genjiang, Xie Li, and Sun Zhongxiu. A new methodology of data dependence analysis for parallelizing C++. ACM SIGPLAN Notices, 30(12):37– 42, December 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noyelle:1995:DC

[1350] Yves L. Noyelle. Disciplined C. ACM SIGPLAN Notices, 30(12):43–50, December 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xinger:1995:AES

[1351] Zhang Xinger. An approach to executable specifications, based on formal source-to-source conversion. ACM SIGPLAN Notices, 30(12):51–58, December 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:1996:NEN

[1352] Ron K. Cytron. Notices: Editorial: A new look for an old friend. ACM SIG-PLAN Notices, 31(1):1, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Helm:1996:PAS

[1353] Richard Helm. Patterns, architecture and software. ACM SIGPLAN Notices, 31(1):2–3, January 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Helm:1996:PLT

[1354] Richard Helm. Patterns: Language trends: Patterns, architecture and software. ACM SIGPLAN Notices, 31 (1):2–3, January 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:PLPa

[1355] Preston Briggs. Programming: Literate programming: Remembering. ACM SIGPLAN Notices, 31(1):4–5, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:R

[1356] Preston Briggs. Remembering. ACM SIGPLAN Notices, 31(1):4–5, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1996:LTCa

[1357] G. Bowden Wise. Language tips: C++ toolbox: A new resource for C++ programmers and an invitation for participation. ACM SIGPLAN Notices, 31 (1):6–8, January 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1996:NRC

[1358] G. Bowden Wise. A new resource for C++ programmers and an invitation for participation. ACM SIG-PLAN Notices, 31(1):6–8, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stewart:1996:OHS

[1359] Kate Stewart. Online: ../SIG-PLAN/Hotlist: Starting points. ACM SIGPLAN Notices, 31(1):9–11, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stewart:1996:SP

[1360] Kate Stewart. Starting points. ACM SIGPLAN Notices, 31(1):9–11, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1996:AEC

[1361] Barbara Ryder. Activities: Executive committee news: Letter from the Chair: Changes to ACM SIGPLAN Bylaws. ACM SIGPLAN Notices, 31 (1):12–17, January 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1996:LCC

[1362] Barbara Ryder. Letter from the Chair: changes to ACM SIGPLAN bylaws. ACM SIGPLAN Notices, 31(1):12–17, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CAM

[1363] Anonymous. Conference announcements — MASPLAS’96. ACM SIG-PLAN Notices, 31(1):18, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuen:1996:TRP

[1364] C. K. Yuen and M. D. Feng. Tailrecursive parallel summing of a multilist. ACM SIGPLAN Notices, 31(1): 20–21, January 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zeller:1996:DFG

[1365] Andreas Zeller and Dorothea Lütkehaus.

DDD—A free graphical front-end for UNIX debuggers. ACM SIGPLAN Notices, 31(1):22–27, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gough:1996:BTR

[1366] K. John Gough. Bottom-up tree rewriting tool MBURG. ACM SIGPLAN Notices, 31(1):28–31, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barrett:1996:NCC

[1367] Myles F. Barrett and Marshall E.

Giguere. A note on covariance and contravariance unification. ACM SIG-PLAN Notices, 31(1):32–35, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamilton:1996:MIO

[1368] Jennifer Hamilton. A model for implementing an object-oriented design without language extensions. ACM SIGPLAN Notices, 31(1):36–43, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klug:1996:TCV

[1369] Michael Klug. Towards a classification of visibility rules. ACM SIG-PLAN Notices, 31(1):44–50, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:1996:CCCa

[1370] Ron K. Cytron. Conference corner: Call for participation: MAS-PLAS ’96. ACM SIGPLAN Notices, 31(2):1, February 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:1996:AEC

[1371] Mary Lou Soffa. Activities: Executive committee news: Letter from the Vice Chair for Conferences. ACM SIGPLAN Notices, 31(2):2–3, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1996:CPOa

[1372] A. Michael Berman. Curricular patterns: OO techniques in the classroom: Column constructor. ACM SIGPLAN Notices, 31(2):4–5, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1996:OTCa

[1373] A. Michael Berman. OO techniques in the classroom: column constructor. ACM SIGPLAN Notices, 31(2): 4–5, February 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agarwal:1996:LSL

[1374] Pankaj Agarwal. Languages: Specialpurpose languages: Describing biological cells by computer program. ACM SIGPLAN Notices, 31(2):6–7, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Software for modelling of Dictyostelium discoideum is available at ftp:/

/ibc.wustl.edu/pub/agarwal/cpl/
cpl.tar.Z and at http://ibc.wustl.
edu/~agarwal/cpl.

Agarwal:1996:SPL

[1375] Pankaj Agarwal. Special-purpose languages: describing biological cells by computer programs. ACM SIGPLAN Notices, 31(2):6–7, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CCa

[1376] Gabby Silberman. Conference corner: Conference report: The Third International Conference on PAR-ALLEL COMPUTING TECHNOLO-GIES (PaCT-95), Saint Petersburg, Russia. ACM SIGPLAN Notices, 31 (2):8–9, February 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silberman:1996:CRI

[1377] Gabby Silberman. Conference report: The International Conference on Parallel Computing Technologies (PaCT-95), Saint Petersburg, Russia. ACM SIGPLAN Notices, 31(2):8–9, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CPAa

[1378] Anonymous. Call for participation: ACM SIGPLAN International Conference on Functional Programming. ACM SIGPLAN Notices, 31(2):10, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CPAb

[1379] Anonymous. Call for participation: ACM SIGPLAN workshop on functional languages in introductory computing. ACM SIGPLAN Notices, 31 (2):11, February 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1996:TC

[1380] Richard L. Wexelblat. Technical correspondence. ACM SIGPLAN Notices, 31(2):12, February 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lastovetsky:1996:MMP

[1381] Alexey L. Lastovetsky. mpC: A multiparadigm programming language for massively parallel computers. ACM SIGPLAN Notices, 31(2):13–20, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tuovinen:1996:TSD

[1382] Antti-Pekka Tuovinen and Jukka Paakki. Translating SQL for database reengineering. ACM SIGPLAN Notices, 31(2):21–26, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parr:1996:LLT

[1383] Terence J. Parr and Russell W. Quong.

LL and LR translators need k > 1 lookahead. ACM SIGPLAN Notices, 31(2):27–34, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kokol:1996:SCA

[1384] Peter Kokol, Janez Brest, and Viljem Žumer. Software complexity — an alternative view. ACM SIGPLAN Notices, 31(2):35–41, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kristensen:1996:CPC

[1385] Bent Bruun Kristensen and Kasper Østerbye. A conceptual perspective on the comparison of object-oriented programming languages. ACM SIG-PLAN Notices, 31(2):42–54, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Appel:1996:IEC

[1386] Andrew W. Appel. Intensional equality ;=) for continuations. ACM SIG-PLAN Notices, 31(2):55–57, February 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hailpern:1996:AEC

[1387] Brent Hailpern. Activities: Executive committee news: Awards announcements. ACM SIGPLAN Notices, 31 (3):1, March 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1996:CPOb

[1388] A. Michael Berman. Curricular patterns: OO techniques in the classroom: Philadelphia dreaming. ACM SIGPLAN Notices, 31(3):2–3, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1996:OTCb

[1389] A. Michael Berman. OO techniques in the classroom: Philadelphia dreaming. ACM SIGPLAN Notices, 31(3):2–3, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rossen:1996:AEC

[1390] Mary Beth Rossen. Activities: Executive committee news: The SIG-PLAN Professional Activities Committee. ACM SIGPLAN Notices, 31(3):4, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rossen:1996:SPA

[1391] Mary Beth Rossen. The SIG-PLAN Professional Activities Committee. ACM SIGPLAN Notices, 31(3):4, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yellin:1996:DPC

[1392] Dan Yellin. Debates: Point/ counterpoint: Do programmers need seat belts? ACM SIGPLAN Notices, 31(3):5, March 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yellin:1996:PCD

[1393] Dan Yellin. Point/counterpoint: do programmers need seatbelts? ACM SIGPLAN Notices, 31(3):5, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strom:1996:DPD

[1394] Rob Strom. Debates: Point: Do programmers need seat belts? ACM SIG-PLAN Notices, 31(3):6–7, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strom:1996:PDP

[1395] Rob Strom. Point: do programmers need seatbelts? ACM SIG-PLAN Notices, 31(3):6–7, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabriel:1996:CDP

[1396] Richard Gabriel. Counterpoint: do programmers need seatbelts? ACM SIGPLAN Notices, 31(3):8–10, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabriel:1996:DCD

[1397] Richard Gabriel. Debates: Counterpoint: Do programmers need seat belts? ACM SIGPLAN Notices, 31 (3):8–10, March 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:1996:CCCb

[1398] Ron K. Cytron. Conference corner: Calendar. ACM SIGPLAN Notices, 31(3):11, March 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:TCa

[1399] Anonymous. Technical correspondence. ACM SIGPLAN Notices, 31 (3):13, March 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jia:1996:FTS

[1400] Guoping Jia and Guoliang Zheng. Fair transition system specification: An integrated approach. ACM SIGPLAN Notices, 31(3):14–21, March 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vowels:1996:P

[1401] Robin A. Vowels. PL/I for OS/2. ACM SIGPLAN Notices, 31(3):22–27, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:1996:NSA

[1402] Sunny Y. Wang. A new sort algorithm: Self-indexed sort. ACM SIGPLAN Notices, 31(3):28–36, March 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comment [1504].

Berman:1996:BO

[1403] A. Michael Berman. On beyond OOP.

ACM SIGPLAN Notices, 31(4):1–3, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1996:CPB

[1404] A. Michael Berman. Curricular patterns: On beyond OOP. ACM SIG-PLAN Notices, 31(4):1–3, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1996:LTCb

[1405] G. Bowden Wise. Language tips: C++ toolbox: An overview of the Standard Template Library. ACM SIG-PLAN Notices, 31(4):4–10, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1996:OST

[1406] G. Bowden Wise. An overview of the standard template library. ACM SIG-PLAN Notices, 31(4):4–10, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:AP

[1407] Preston Briggs. Automatic parallelization. ACM SIGPLAN Notices, 31 (4):11–15, April 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:PLPb

[1408] Preston Briggs. Programming: Literate programming: Automatic parallelization. ACM SIGPLAN Notices, 31 (4):11–15, April 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CCC

[1409] Anonymous. Conference corner: Calendar. ACM SIGPLAN Notices, 31 (4):16–17, April 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CPPa

[1410] Anonymous. Call for papers: Principles of programming languages. ACM SIG-PLAN Notices, 31(4):18, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:TCb

[1411] Anonymous. Technical correspondence. ACM SIGPLAN Notices, 31 (4):19, March 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruccoleri:1996:TCW

[1412] Robert E. Bruccoleri. Technical correspondence: WRAPGEN — A tool for the use of Fortran and C together in portable programs. ACM SIGPLAN Notices, 31(4):20–27, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruccoleri:1996:WTU

[1413] Robert E. Bruccoleri. WRAPGEN: a tool for the use of Fortran and C together in portable programs. ACM SIGPLAN Notices, 31(4):20–27, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karr:1996:TCC

[1414] Michael Karr. A translator from C to a λ-calculus representation. ACM SIGPLAN Notices, 31(4):28–31, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karr:1996:TCS

[1415] Michael Karr. A translator from C to a λ-calculus representation. ACM SIGPLAN Notices, 31(4):28–31, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldsack:1996:ADD

[1416] S. J. Goldsack, K. Lano, and E. H.

Dürr. Annealing and data decomposition in VDM^{++}. ACM SIGPLAN Notices, 31(4):32–38, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodriguez:1996:SOO

[1417] Bradford J. Rodriguez and W. F. S.

Poehlman. A survey of object-oriented Forths. ACM SIGPLAN Notices, 31 (4):39–42, April 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodriquez:1996:SOO

[1418] Bradford J. Rodriquez and W. F. S. Poehlman. A survey of object-oriented Forths. ACM SIGPLAN Notices, 31 (4):39–42, April 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zettel:1996:TOO

[1419] Leonard Zettel. Toward an objectoriented Forth. ACM SIGPLAN Notices, 31(4):43–44, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whitty:1996:OOM

[1420] Robin Whitty. Object-oriented metrics: An annotated bibliography. ACM SIGPLAN Notices, 31(4):45–75, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruttenberg:1996:SPS

[1421] John Ruttenberg, G. R. Gao, A. Stoutchinin, and W. Lichtenstein. Software pipelining showdown: Optimal vs. heuristic methods in a production compiler. ACM SIGPLAN Notices, 31(5):1–11, May 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p1-ruttenberg/

.

Eichenberger:1996:RMM

[1422] Alexandre E. Eichenberger and Edward S. Davidson. A reduced multipipeline machine description that preserves scheduling constraints. ACM SIGPLAN Notices, 31(5):12–22, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p12-eichenberger/

.

Flanagan:1996:CBW

[1423] Cormac Flanagan, Matthew Flatt, Shriram Krishnamurthi, Stephanie Weirich, and Matthias Felleisen. Catching bugs in the web of program invariants. ACM SIGPLAN Notices, 31(5):23–32, May 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p23-flanagan/

.

Adl-Tabatabai:1996:SLD

[1424] Ali-Reza Adl-Tabatabai and Thomas Gross. Source-level debugging of scalar optimized code. ACM SIG-PLAN Notices, 31(5):33–43, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p33-adl-
tabatabai/.

Evans:1996:SDD

[1425] David Evans. Static detection of dynamic memory errors. ACM SIG-PLAN Notices, 31(5):44–53, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p44-evans/

.

Rinard:1996:CAN

[1426] Martin C. Rinard and Pedro C. Diniz. Commutativity analysis: A new analysis framework for parallelizing compilers. ACM SIGPLAN Notices, 31(5):54– 67, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p54-rinard/

.

Chakrabarti:1996:GCA

[1427] Soumen Chakrabarti, Manish Gupta, and Jong-Deok Choi. Global communication analysis and optimization. ACM SIGPLAN Notices, 31(5):68–78, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p68-chakrabarti/

.

Trinder:1996:GPP

[1428] P. W. Trinder, K. Hammond, J. S. Mattson, Jr., A. S. Partridge, and S. L. Peyton Jones. GUM: A portable parallel implementation of Haskell. ACM SIGPLAN Notices, 31(5):79–88, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p79-trinder/

.

Boehm:1996:SGC

[1429] Hans-J. Boehm. Simple garbagecollector-safety. ACM SIGPLAN Notices, 31(5):89–98, May 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p89-boehm/

.

Bruggeman:1996:RCP

[1430] Carl Bruggeman, Oscar Waddell, and R. Kent Dybvig. Representing control in the presence of oneshot continuations. ACM SIG-PLAN Notices, 31(5):99–107, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p99-bruggeman/

.

Burger:1996:PFP

[1431] Robert G. Burger and R. Kent Dybvig. Printing floating-point numbers quickly and accurately. ACM SIG-PLAN Notices, 31(5):108–116, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p108-burger/

. This paper offers a significantly faster algorithm than that of [111], together with a correctness proof and an implementation in Scheme. See also [109, ?, ?, ?].

Dawson:1996:PPA

[1432] Steven Dawson, C. R. Ramakrishnan, and David S. Warren. Practical program analysis using general purpose logic programming systems — A case study. ACM SIG-PLAN Notices, 31(5):117–126, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p117-dawson/

.

Adl-Tabatabai:1996:ELI

[1433] Ali-Reza Adl-Tabatabai, Geoff Langdale, Steven Lucco, and Robert Wahbe. Efficient and language-independent mobile programs. ACM SIGPLAN Notices, 31(5):127–136, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p127-adl-
tabatabai/.

Lee:1996:OMR

[1434] Peter Lee and Mark Leone. Optimizing ML with run-time code generation. ACM SIGPLAN Notices, 31(5):137–148, May 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p137-lee/

.

Auslander:1996:FED

[1435] Joel Auslander, Matthai Philipose, Craig Chambers, Susan J. Eggers, and Brian N. Bershad. Fast, effective dynamic compilation. ACM SIG-PLAN Notices, 31(5):149–159, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p149-auslander/

.

Engler:1996:VRE

[1436] Dawson R. Engler. VCODE: A retargetable, extensible, very fast dynamic code generation system. ACM SIG-PLAN Notices, 31(5):160–170, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p160-engler/

.

Reppy:1996:SOS

[1437] John Reppy and Jon Riecke. Simple objects for Standard ML. ACM SIG-PLAN Notices, 31(5):171–180, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p171-reppy/

.

Tarditi:1996:TTD

[1438] D. Tarditi, G. Morrisett, P. Cheng, C. Stone, R. Harper, and P. Lee. TIL: A type-directed optimizing compiler for ML. ACM SIGPLAN Notices, 31(5):181–192, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p181-tarditi/

.

Jagannathan:1996:FDI

[1439] Suresh Jagannathan and Andrew Wright. Flow-directed inlining. ACM SIGPLAN Notices, 31(5):193–205, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/231379/p193-jagannathan/

.

Sperber:1996:RCP

[1440] Michael Sperber and Peter Thiemann. Realistic compilation by partial evaluation. ACM SIGPLAN Notices, 31(5):206–214, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/231379/p206-sperber/

.

Knoblock:1996:DS

[1441] Todd B. Knoblock and Erik Ruf. Data specialization. ACM SIG-PLAN Notices, 31(5):215–225, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/231379/p215-knoblock/

.

Ramsey:1996:RMI

[1442] Norman Ramsey. Relocating machine instructions by currying. ACM SIG-PLAN Notices, 31(5):226–236, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/231379/p226-ramsey/

.

Chandra:1996:TLS

[1443] Satish Chandra, Brad Richards, and James R. Larus. Teapot: Language support for writing memory coherence protocols. ACM SIG-PLAN Notices, 31(5):237–248, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p237-chandra/

.

Bailey:1996:TSC

[1444] Mark W. Bailey and Jack W. Davidson. Target-sensitive construction of diagnostic programs for procedure calling sequence generators. ACM SIG-PLAN Notices, 31(5):249–257, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p249-bailey/

.

Russinovich:1996:RCN

[1445] Mark Russinovich and Bryce Cogswell.

Replay for concurrent non-deterministic shared-memory applications. ACM SIGPLAN Notices, 31(5):258–266, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/231379/p258-russinovich/

.

Ramalingam:1996:DFF

[1446] G. Ramalingam. Data flow frequency analysis. ACM SIGPLAN Notices, 31(5):267–277, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/231379/p267-ramalingam/

.

Sreedhar:1996:NFE

[1447] Vugranam C. Sreedhar, Guang R. Gao, and Yong-Fong Lee. A new framework for exhaustive and incremental data flow analysis using DJ graphs. ACM SIGPLAN Notices, 31(5):278–290, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/231379/p278-sreedhar/

.

Bilardi:1996:FGC

[1448] Gianfranco Bilardi and Keshav Pingali. A framework for generalized control dependence. ACM SIG-PLAN Notices, 31(5):291–300, May 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/231379/p291-bilardi/

.

Jones:1996:LFM

[1449] Simon Peyton Jones, Will Partain, and André Santos. Let-floating: Moving bindings to give faster programs. ACM SIGPLAN Notices, 31(6):1–12, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sabry:1996:RCV

[1450] Amr Sabry and Philip Wadler. A reflection on call-by-value. ACM SIG-PLAN Notices, 31(6):13–24, June 1996.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hardin:1996:FBE

[1451] Thérèse Hardin, Luc Maranget, and Bruno Pagano. Functional backends within the lambda-sigma calculus. ACM SIGPLAN Notices, 31(6):25–33, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rojemo:1996:LDV

[1452] Niklas Röjemo and Colin Runciman.

Lag, drag, void and use — heap profiling and space-efficient compilation revisited. ACM SIGPLAN Notices, 31 (6):34–41, June 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1996:SDP

[1453] David S. Wise and Joshua Walgenbach. Static and dynamic partitioning of pointers as links and threads. ACM SIGPLAN Notices, 31(6):42–49, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Serrano:1996:SUA

[1454] Manuel Serrano and Marc Feeley. Storage use analysis and its applications. ACM SIGPLAN Notices, 31(6):50–61, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Okasaki:1996:RLE

[1455] Chris Okasaki. The role of lazy evaluation in amortized data structures. ACM SIGPLAN Notices, 31(6):62–72, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hu:1996:DSH

[1456] Zhenjiang Hu, Hideya Iwasaki, and Masato Takeichi. Deriving structural hylomorphisms from recursive definitions. ACM SIGPLAN Notices, 31 (6):73–82, June 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abadi:1996:ACD

[1457] Martín Abadi, Butler Lampson, and Jean-Jacques Lévy. Analysis and caching of dependencies. ACM SIG-PLAN Notices, 31(6):83–91, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lawall:1996:OIW

[1458] Julia L. Lawall and Harry G. Mairson. Optimality and inefficiency: What isn’t a cost model of the lambda calculus? ACM SIGPLAN Notices, 31 (6):92–101, June 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Howard:1996:ICP

[1459] Brian T. Howard. Inductive, coinductive, and pointed types. ACM SIG-PLAN Notices, 31(6):102–109, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gostanza:1996:NLP

[1460] Pedro Palao Gostanza, Ricardo Peña, and Manuel Núñez. A new look at pattern matching in abstract data types. ACM SIGPLAN Notices, 31 (6):110–121, June 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pottier:1996:SSC

[1461] François Pottier. Simplifying subtyping constraints. ACM SIGPLAN Notices, 31(6):122–133, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghelli:1996:CKF

[1462] Giorgio Ghelli. Complexity of kernel Fun subtype checking. ACM SIG-PLAN Notices, 31(6):134–145, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moreau:1996:SSF

[1463] Luc Moreau. The semantics of Scheme with future. ACM SIGPLAN Notices, 31(6):146–156, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turbak:1996:FCS

[1464] Franklyn Turbak. First-class synchronization barriers. ACM SIGPLAN Notices, 31(6):157–168, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flanagan:1996:PDP

[1465] Cormac Flanagan and Rishiyur S.

Nikhil. pHluid: the design of a parallel functional language implementation on workstations. ACM SIG-PLAN Notices, 31(6):169–179, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thiemann:1996:CSL

[1466] Peter J. Thiemann. Cogen in six lines. ACM SIGPLAN Notices, 31(6): 180–189, June 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chuang:1996:PAP

[1467] Tyng-Ruey Chuang and Wen L. Hwang. A probabilistic approach to the problem of automatic selection of data representations. ACM SIG-PLAN Notices, 31(6):190–200, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferreira:1996:TWB

[1468] William Ferreira, Matthew Hennessy, and Alan Jeffrey. A theory of weak bisimulation for core CML. ACM SIG-PLAN Notices, 31(6):201–212, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blelloch:1996:PTS

[1469] Guy E. Blelloch and John Greiner. A provable time and space efficient implementation of NESL. ACM SIG-PLAN Notices, 31(6):213–225, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caspi:1996:SKN

[1470] Paul Caspi and Marc Pouzet. Synchronous Kahn networks. ACM SIG-PLAN Notices, 31(6):226–238, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1996:ELC

[1471] Shinn-Der Lee and Daniel P. Friedman.

Enriching the lambda calculus with contexts: Toward a theory of incremental program construction. ACM SIG-PLAN Notices, 31(6):239–250, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Queinnec:1996:SCT

[1472] Christian Queinnec and David De Roure. Sharing code through firstclass environments. ACM SIGPLAN Notices, 31(6):251–261, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duggan:1996:MM

[1473] Dominic Duggan and Constantinos Sourelis. Mixin modules. ACM SIG-PLAN Notices, 31(6):262–273, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:ECN

[1474] Anonymous. Executive Committee news. ACM SIGPLAN Notices, 31 (7):1–??, July 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CPPb

[1475] Anonymous. Call for papers: PLDI ’97. ACM SIGPLAN Notices, 31(7):2– ??, July 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hind:1996:URC

[1476] Michael Hind and Phil Pfeiffer. Using regional conferences to mentor student development: a case study. ACM SIG-PLAN Notices, 31(7):4–7, July 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hind:1996:CRM

[1477] Michael Hind and Phil Pfeiffer. Conference report: Mentoring student development. ACM SIGPLAN Notices, 31(7):4–??, July 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CCb

[1478] Anonymous. Conference calendar. ACM SIGPLAN Notices, 31(7):8–??, July 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gibbon:1996:LTC

[1479] C. Gibbon and C. Higgins. Language tips: C++ toolbox: Teaching objectoriented design with heuristics. ACM SIGPLAN Notices, 31(7):12–16, July 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gibbon:1996:TOO

[1480] C. Gibbon and C. Higgins. Teaching object-oriented design with heuristics. ACM SIGPLAN Notices, 31(7):12–16, July 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CPJ

[1481] Anonymous. Call for papers: JMLC ’97. ACM SIGPLAN Notices, 31(7):17– ??, July 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aiken:1996:CPP

[1482] Alexander Aiken. Cool: A portable project for teaching compiler construction. ACM SIGPLAN Notices, 31 (7):19–24, July 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sorkin:1996:SCP

[1483] Arthur Sorkin. Some comments on ‘The Priority-Based Coloring Approach to Register Allocation’. ACM SIGPLAN Notices, 31(7):25–29, July 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:1996:EOO

[1484] Jiun-Liang Chen and Feng-Jian Wang.

Encapsulation in object-oriented programs. ACM SIGPLAN Notices, 31 (7):30–32, July 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Panti:1996:PPC

[1485] M. Panti, A. Cucchiarelli, M. Mattiucci, and S. Valenti. Process to process communication in Prolog. ACM SIGPLAN Notices, 31(7):33–39, July 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milanese:1996:IRS

[1486] Vitaliano Milanese. Interpreting RSets, SSets and features by denotational semantics. ACM SIGPLAN Notices, 31 (7):40–47, July 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:LC

[1487] Anonymous. Letter from the Chair. ACM SIGPLAN Notices, 31(8):1, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:AR

[1488] Anonymous. 1995–96 annual report. ACM SIGPLAN Notices, 31(8):2–3, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:PR

[1489] Anonymous. Press release. ACM SIG-PLAN Notices, 31(8):4, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:LPSa

[1490] Preston Briggs. Literate programming: Sparse matrix manipulation. ACM SIGPLAN Notices, 31(8):5–7, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:SMMa

[1491] Preston Briggs. Sparse matrix manipulation. ACM SIGPLAN Notices, 31 (8):5–7, August 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Loui:1996:WGA

[1492] Ronald P. Loui. Why GAWK for AI?

ACM SIGPLAN Notices, 31(8):8–9, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1996:CCB

[1493] G. Bowden Wise. Casting in C++:

bringing safety and smartness to your programs. ACM SIGPLAN Notices, 31 (8):10–15, August 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1996:CT

[1494] G. Bowden Wise. C++ toolbox. ACM SIGPLAN Notices, 31(8):10–15, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Papaspynou:1996:DSC

[1495] Nikolaos S. Papaspynou. Denotational semantics in C++. ACM SIG-PLAN Notices, 31(8):16–25, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Papaspyrou:1996:FPD

[1496] Nikolaos S. Papaspyrou. A framework for programming denotational semantics in C++. ACM SIGPLAN Notices, 31(8):16–25, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1996:FRT

[1497] Paul Frenger. Forth report: Thoughts on the 1996 Rochester FORTH conference. ACM SIGPLAN Notices, 31 (8):26–27, August 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1996:TRF

[1498] Paul Frenger. Thoughts on the 1996 Rochester FORTH Conference. ACM SIGPLAN Notices, 31(8):26–27, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Norton:1996:IWP

[1499] Charles D. Norton. The International Workshop on Parallel C++ (IWPC++) Kanazawa, Ishikawa prefecture, Japan. ACM SIGPLAN Notices, 31(8):28–30, August 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Norton:1996:WR

[1500] Charles D. Norton. Workshop report: The International Workshop on Parallel C++ (IWPC++), Kanazawa, Ishikawa Prefecturem, Japan. ACM SIGPLAN Notices, 31(8):28–30, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CPS

[1501] Anonymous. Call for papers: SCM 7. ACM SIGPLAN Notices, 31(8):31, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CPPc

[1502] Anonymous. Call for papers: PPoPP 97. ACM SIGPLAN Notices, 31(8): 32–33, August 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:CCd

[1503] Anonymous. Conference calendar.

ACM SIGPLAN Notices, 31(8):34–39, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andersson:1996:CSI

[1504] Arne Andersson and Andrej Brodnik.

Comment on “self-indexed sort”. ACM SIGPLAN Notices, 31(8):40–42, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [1402].

Vialle:1996:PPP

[1505] Stéphane Vialle, Thierry Cornu, and Yannick Lallement. ParCeL-1: a parallel programming language based on autonomous and synchronous actors. ACM SIGPLAN Notices, 31(8):43–51, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abdullah:1996:HPA

[1506] Mohd Hanafiah Abdullah. hcc: a portable ANSI C compiler (with a code generator for the PowerPCs). ACM SIGPLAN Notices, 31(8):52–59, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Syropoulos:1996:NTCa

[1507] Apostolos Syropoulos. A note on type checking linear functional languages. ACM SIGPLAN Notices, 31(8):60–63, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butt:1996:PMC

[1508] Farooq Butt. Porting the mcc compiler into an interactive development environment. ACM SIGPLAN Notices, 31 (8):64–73, August 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butt:1996:PTP

[1509] Farooq Butt. Porting the mcc PowerPC C/C++ compiler into an interactive development environment. ACM SIGPLAN Notices, 31(8):64–73, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pronk:1996:IIS

[1510] C. Pronk and M. Schönhacker. ISO/ IEC 10514-1, the Standard for Modula-2: Process aspects. ACM SIGPLAN Notices, 31(8):74–83, August 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schonhacker:1996:IIS

[1511] M. Schönhacker and C. Pronk. ISO/ IEC 10514-1, the Standard for Modula-2: Changes, clarifications and additions. ACM SIGPLAN Notices, 31(8): 84–95, August 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olukotun:1996:CSC

[1512] Kunle Olukotun, Basem A. Nayfeh, Lance Hammond, Ken Wilson, and Kunyung Chang. The case for a singlechip multiprocessor. ACM SIGPLAN Notices, 31(9):2–11, September 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olukotun:1996:CSM

[1513] Kunle Olukotun, Basem A. Nayfeh, Lance Hammond, Ken Wilson, and Kunyung Chang. The case for a single-chip multiprocessor. ACM SIG-PLAN Notices, 31(9):2–11, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p2-olukotun/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Pai:1996:EMC

[1514] Vijay S. Pai, Parthasarathy Ranganathan, Sarita V. Adve, and Tracy Harton. An evaluation of memory consistency models for sharedmemory systems with ILP processors. ACM SIGPLAN Notices, 31(9):12– 23, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p12-pai/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Scott:1996:SCT

[1515] Steven L. Scott. Synchronization and communication in the T3E multiprocessor. ACM SIGPLAN Notices, 31 (9):26–36, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p26-scott/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Krishnamurthy:1996:EAS

[1516] Arvind Krishnamurthy, Klaus E. Schauser, Chris J. Scheiman, Randolph Y. Wang, David E. Culler, and Katherine Yelick. Evaluation of architectural support for global address-based communication in largescale parallel machines. ACM SIG-PLAN Notices, 31(9):37–48, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p37-krishnamurthy/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Grunwald:1996:WOT

[1517] Dirk Grunwald and Richard Neves. Whole-program optimization for time and space efficient threads. ACM SIG-PLAN Notices, 31(9):50–59, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p50-grunwald/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Grunwald:1996:WPO

[1518] Dirk Grunwald and Richard Neves.

Whole-program optimization for time and space efficient threads. ACM SIG-PLAN Notices, 31(9):50–59, September 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Philbin:1996:TSC

[1519] James Philbin, Jan Edler, Otto J.

Anshus, Craig C. Douglas, and Kai Li. Thread scheduling for cache locality. ACM SIGPLAN Notices, 31 (9):60–71, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p60-philbin/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Chen:1996:RFC

[1520] Peter M. Chen, Wee Teck Ng, Subhachandra Chandra, Christopher Aycock, Gurushankar Rajamani, and David Lowell. The Rio file cache: Surviving operating system crashes. ACM SIGPLAN Notices, 31(9):74– 83, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p74-chen/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Lee:1996:PDV

[1521] Edward K. Lee and Chandramohan A. Thekkath. Petal: Distributed virtual disks. ACM SIGPLAN Notices, 31 (9):84–92, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p84-lee/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

McKinley:1996:QAL

[1522] Kathryn S. McKinley and Olivier Temam. A quantitative analysis of loop nest locality. ACM SIG-PLAN Notices, 31(9):94–104, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p94-mckinley/

. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Huang:1996:IBR

[1523] Andrew S. Huang and John Paul Shen. The intrinsic bandwidth requirements of ordinary programs. ACM SIGPLAN Notices, 31(9):105– 114, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p105-

huang/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Seznec:1996:MAB

[1524] André Seznec, Stéphan Jourdan, Pascal Sainrat, and Pierre Michaud. Multiple-block ahead branch predictors. ACM SIGPLAN Notices, 31(9): 116–127, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p116-

seznec/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Seznec:1996:MBA

[1525] André Seznec, Stéphan Jourdan, Pascal Sainrat, and Pierre Michaud. Multiple-block ahead branch predictors. ACM SIGPLAN Notices, 31(9): 116–127, September 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:1996:ABP

[1526] I-Cheng K. Chen, John T. Coffey, and Trevor N. Mudge. Analysis of branch prediction via data compression. ACM SIGPLAN Notices, 31(9): 128–137, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p128-

chen/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Lipasti:1996:VLL

[1527] Mikko H. Lipasti, Christopher B. Wilkerson, and John Paul Shen. Value locality and load value prediction. ACM SIGPLAN Notices, 31(9):138– 147, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p138-

lipasti/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Romer:1996:SPI

[1528] Theodore H. Romer, Dennis Lee, Geoffrey M. Voelker, Alec Wolman, Wayne A. Wong, Jean-Loup Baer, Brian N. Bershad, and Henry M. Levy.

The structure and performance of interpreters. ACM SIGPLAN Notices, 31 (9):150–159, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p150-

romer/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Fox:1996:ANC

[1529] Armando Fox, Steven D. Gribble, Eric A. Brewer, and Elan Amir. Adapting to network and client variability via on-demand dynamic distillation. ACM SIGPLAN Notices, 31(9):160– 170, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p160-

fox/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Scales:1996:SLO

[1530] Daniel J. Scales, Kourosh Gharachorloo, and Chandramohan A. Thekkath. Shasta: A low overhead, software-only approach for supporting fine-grain shared memory. ACM SIGPLAN Notices, 31(9):174– 185, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/asplos/237090/p174-

scales/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Dwarkadas:1996:ICT

[1531] Sandhya Dwarkadas, Alan L. Cox, and Willy Zwaenepoel. An integrated compile-time/run-time software distributed shared memory system. ACM SIGPLAN Notices, 31(9):186– 197, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p186-

dwarkadas/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Bianchini:1996:HCL

[1532] R. Bianchini, L. I. Kontothanassis, R. Pinto, M. De Maria, M. Abud, and C. L. Amorim. Hiding communication latency and coherence overhead in software DSMs. ACM SIGPLAN Notices, 31(9):198–209, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p198-

bianchini/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Erlichson:1996:SAP

[1533] Andrew Erlichson, Neal Nuckolls, Greg Chesson, and John Hennessy. Soft-FLASH: Analyzing the performance of clustered distributed virtual shared memory. ACM SIGPLAN Notices, 31 (9):210–220, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p210-

erlichson/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Luk:1996:CBP

[1534] Chi-Keung Luk and Todd C. Mowry.

Compiler-based prefetching for recursive data structures. ACM SIG-PLAN Notices, 31(9):222–233, September 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Luk:1996:CPR

[1535] Chi-Keung Luk and Todd C. Mowry.

Compiler-based prefetching for recursive data structures. ACM SIGPLAN Notices, 31(9):222–233, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p222-

luk/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Saghir:1996:EDD

[1536] Mazen A. R. Saghir, Paul Chow, and Corinna G. Lee. Exploiting dual datamemory banks in digital signal processors. ACM SIGPLAN Notices, 31 (9):234–243, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p234-

saghir/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Bugnion:1996:CDP

[1537] Edouard Bugnion, Jennifer M. Anderson, Todd C. Mowry, Mendel Rosenblum, and Monica S. Lam. Compilerdirected page coloring for multiprocessors. ACM SIGPLAN Notices, 31(9): 244–255, September 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bugnion:1996:CPC

[1538] Edouard Bugnion, Jennifer M. Anderson, Todd C. Mowry, Mendel Rosenblum, and Monica S. Lam. Compilerdirected page coloring for multiprocessors. ACM SIGPLAN Notices, 31(9): 244–255, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p244-

bugnion/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Jamrozik:1996:RNL

[1539] Hervé A. Jamrozik, Michael J. Feeley, Geoffrey M. Voelker, James Evans, II, Anna R. Karlin, Henry M. Levy, and Mary K. Vernon. Reducing network latency using subpages in a global memory environment. ACM SIGPLAN Notices, 31(9):258– 267, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p258-

jamrozik/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Peir:1996:ICP

[1540] Jih-Kwon Peir, Windsor W. Hsu, Honesty Young, and Shauchi Ong. Improving cache performance with balanced tag and data paths. ACM SIGPLAN Notices, 31(9):268–278, September 1996. CODEN SIN-ODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/237090/p268-

peir/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Verghese:1996:OSS

[1541] Ben Verghese, Scott Devine, Anoop Gupta, and Mendel Rosenblum. Operating system support for improving data locality on CC-NUMA compute servers. ACM SIGPLAN Notices, 31 (9):279–289, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/237090/p279-

verghese/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Anonymous:1996:AI

[1542] Anonymous. Author index. ACM SIG-PLAN Notices, 31(9):290, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Anonymous:1996:ARP

[1543] Anonymous. ASPLOS-VI reviewers. ACM SIGPLAN Notices, 31(9): v, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Anonymous:1996:CCP

[1544] Anonymous. Conference Committee. ACM SIGPLAN Notices, 31(9): iv, September 1996. CODEN SINODQ.

ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Nishimura:1996:EOO

[1545] Susumu Nishimura, Atsushi Ohori, and Keishi Tajima. An equational objectoriented data model and its dataparallel query language. ACM SIG-PLAN Notices, 31(10):1–17, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nishimura:1996:EOD

[1546] Susumu Nishimura, Atsushi Ohori, and Keishi Tajima. An equational objectoriented data model and its dataparallel query language. ACM SIG-PLAN Notices, 31(10):1–??, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cameron:1996:PPR

[1547] Lorette Cameron, Charles Berman, Brian Henderson-Sellers, Laura Hill, Randall Smith, and Sanjiv Gossain. Panel 1: Perspectives on reuse. ACM SIGPLAN Notices, 31(10):10–??, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yelland:1996:CHC

[1548] Phillip M. Yelland. Creating host compliance in a portable framework: A study in the use of existing design patterns. ACM SIGPLAN Notices, 31(10):

18–29, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hassen:1996:FOE

[1549] Saniya Ben Hassen, Irina Athanasiu, and Henri E. Bal. A flexible operation execution model for shared distributed objects. ACM SIGPLAN Notices, 31(10):30–50, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adl-Tabatabai:1996:CRO

[1550] Ali-Reza Adl-Tabatabai, Thomas Gross, and Guei-Yuan Lueh. Code reuse in an optimizing compiler. ACM SIGPLAN Notices, 31(10):51–68, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barrett:1996:MSL

[1551] Kim Barrett, Bob Cassels, Paul Haahr, David A. Moon, Keith Playford, Andrew L. M. Shalit, and P. Tucker Withington. A monotonic superclass linearization for Dylan. ACM SIGPLAN Notices, 31(10):69–82, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dean:1996:VOC

[1552] Jeffrey Dean, Greg DeFouw, David Grove, Vassily Litvinov, and Craig Chambers. Vortex: an optimizing compiler for object-oriented languages. ACM SIGPLAN Notices, 31(10):83–100, October 1996. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

www.cs.washington.edu/research/

projects/cecil/. Discusses performance of Vortex compiler for Cecil, C++, Java, and Modula-3.

Cameron:1996:PRP

[1553] Lorette Cameron, Charles Berman, Brian Henderson-Sellers, Laura Hill, Randall Smith, and Sanjiv Gossain. Perspectives on reuse (panel). ACM SIGPLAN Notices, 31(10):101–103, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dutoit:1996:BOS

[1554] Allen Dutoit, Sean Levy, Douglas Cunningham, and Robert Patrick. The basic object system: Supporting a spectrum from prototypes to hardened code. ACM SIGPLAN Notices, 31(10): 104–121, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bardou:1996:SOD

[1555] Daniel Bardou and Christophe Dony.

Split objects: A disciplined use of delegation within objects. ACM SIGPLAN Notices, 31(10):122–137, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

R:1996:RSO

[1556] Hernán Astudillo R. Reorganizing split objects. ACM SIGPLAN Notices, 31(10):138–149, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Astudillo:1996:RSO

[1557] Hernan R. Astudillo. Reorganizing split objects. ACM SIGPLAN Notices, 31(10):138–??, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kleindienst:1996:LLI

[1558] Jan Kleindienst, František Plášil, and Petr Tůma. Lessons learned from implementing the CORBA persistent object service. ACM SIGPLAN Notices, 31(10):150–167, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wallace:1996:SEO

[1559] Evan Wallace and Kurt C. Wallnau. A situated evaluation of the object management group’s (OMG) object management architecture (OMA). ACM SIGPLAN Notices, 31(10):168– 178, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tip:1996:SCH

[1560] Frank Tip, Jong-Deok Choi, John Field, and G. Ramalingam. Slicing class hierarchies in C++. ACM SIG-PLAN Notices, 31(10):179–197, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moller:1996:FLD

[1561] Ralf Möller. A functional layer for description logics: Knowledge representation meets object-oriented programming. ACM SIGPLAN Notices, 31(10):

198–213, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gil:1996:EAN

[1562] Joseph Gil and David H. Lorenz.

Environmental acquisition: A new inheritance-like abstraction mechanism. ACM SIGPLAN Notices, 31(10): 214–231, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chisholm:1996:SIH

[1563] Paul S. R. Chisholm, Larry Constantine, Ward Cunningham, Luke Hohmann, and Norman Kerth. Soft issues and hard problems in software development (panel). ACM SIGPLAN Notices, 31(10):232–234, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chisholm:1996:PSI

[1564] Paul S. R. Chisholm, Larry Constantine, Ward Cunningham, Luke Hohmann, and Norman Kerth. Panel 2: Soft issues and hard problems in software development. ACM SIGPLAN Notices, 31(10):232–??, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moore:1996:AIH

[1565] Ivan Moore. Automatic inheritance hierarchy restructuring and method refactoring. ACM SIGPLAN Notices, 31(10):235–250, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dicky:1996:ACI

[1566] H. Dicky, C. Dony, M. Huchard, and T. Libourel. On automatic class insertion with overloading. ACM SIG-PLAN Notices, 31(10):251–267, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steyaert:1996:RCM

[1567] Patrick Steyaert, Carine Lucas, Kim Mens, and Theo D’Hondt. Reuse contracts: Managing the evolution of reusable assets. ACM SIGPLAN Notices, 31(10):268–285, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1996:OAP

[1568] Steven D. Fraser, Alistair Cockburn, Jim Coplien, Larry Constantine, Dave West, and Leo Brajkovich. OO anthropology (panel): crossing the chasm. ACM SIGPLAN Notices, 31(10):286– 291, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1996:POA

[1569] Steven D. Fraser, Alistair Cockburn, Leo Brajkovich, Jim Coplien, Larry Constantine, and Dave West. Panel 3: OO anthropology: Crossing the chasm. ACM SIGPLAN Notices, 31(10):286– ??, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Diwan:1996:SEA

[1570] Amer Diwan, J. Eliot B. Moss, and Kathryn S. McKinley. Simple and effective analysis of statically-typed objectoriented programs. ACM SIGPLAN Notices, 31(10):292–305, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Driesen:1996:DCV

[1571] Karel Driesen and Urs Hölzle. The direct cost of virtual function calls in C++. ACM SIGPLAN Notices, 31(10): 306–323, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:1996:FSA

[1572] David F. Bacon and Peter F. Sweeney.

Fast static analysis of C++ virtual function calls. ACM SIGPLAN Notices, 31(10):324–341, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ecklund:1996:CCU

[1573] Earl F. Ecklund, Jr., Lois M. L.

Delcambre, and Michael J. Freiling. Change cases: Use cases that identify future requirements. ACM SIGPLAN Notices, 31(10):342–358, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanHilst:1996:URC

[1574] Michael VanHilst and David Notkin.

Using role components to implement collaboration-based designs. ACM SIGPLAN Notices, 31(10):359–369, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klarlund:1996:FDC

[1575] Nils Klarlund, Jari Koistinen, and Michael I. Schwartzbach. Formal design constraints. ACM SIGPLAN Notices, 31(10):370–383, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shan:1996:OSW

[1576] Yen-Ping Shan, Tom Morgan, Phil Proudfoot, Jim Thompson, John Tibetts, and Allen Woolfrey. Object on the server: are we ready? (panel). ACM SIGPLAN Notices, 31(10):384– 388, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shan:1996:POS

[1577] Yen-Ping Shan, Tom Morgan, Phil Proudfoot, Jim Thompson, John Tibbets, and Allen Woolfrey. Panel 4: Objects on the server: Are we ready? ACM SIGPLAN Notices, 31(10):384– ??, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sefika:1996:AOV

[1578] Mohlalefi Sefika, Aamod Sane, and Roy H. Campbell. Architectureoriented visualization. ACM SIGPLAN Notices, 31(10):389–405, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sefika:1996:AV

[1579] Mohlalefi Sefika, Aamod Sane, and Roy H. Campbell. Architectureoriented visualization. ACM SIGPLAN Notices, 31(10):389–??, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Durham:1996:FRT

[1580] Alan M. Durham and Ralph E. Johnson. A framework for run-time systems and its visual programming language. ACM SIGPLAN Notices, 31(10):406– 420, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Durham:1996:FRS

[1581] Alan M. Durham and Ralph E. Johnson. A framework for run-time systems and its visual programming language. ACM SIGPLAN Notices, 31(10):406– ??, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Attali:1996:SBV

[1582] Isabelle Attali, Denis Caromel, Sidi O.

Ehmety, and Sylvain Lippi. Semanticbased visualization for parallel objectoriented programming. ACM SIG-PLAN Notices, 31(10):421–440, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Attali:1996:SVP

[1583] Isabelle Attali, Denis Caromel, Sidi O.

Ehmety, and Sylvain Lippi. Semanticbased visualization for parallel objectoriented programming. ACM SIG-PLAN Notices, 31(10):421–??, October 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Booch:1996:TMR

[1584] Grady Booch, Steven Fraser, Robert C. Martin, Steven J. Mellor, Michael Lee, Steven Garone, Martin Fowler, Douglas C. Schmidt, and Marie Lenzi. Translation: myth or reality? (panel). ACM SIGPLAN Notices, 31(10):441– 443, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1996:PDT

[1585] Steven Fraser, Grady Booch, Steven Garone, Robert Martin, Martin Fowler, Steven J. Mellor, Douglas C. Schmidt, Michael Lee, and Marie Lenzi. Panel 5 (debate): Translation: Myth or reality? ACM SIGPLAN Notices, 31(10): 441–??, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonar:1996:CIP

[1586] Jeffrey Bonar, Nancy Lehrer, Kelly Looney, Randy Schnier, James Russel, Ted Selker, and Stewart Nickolas. Components on the Internet (panel). ACM SIGPLAN Notices, 31(10):444– 448, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonar:1996:PCI

[1587] Jeffrey Bonar, Nancy Lehrer, Kelly Looney, Stewart Nickolas, Randy Schnier, James Russell, and Ted Selker. Panel 6: Components on the Internet. ACM SIGPLAN Notices, 31(10): 444–??, October 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:1996:JBC

[1588] Joseph Bergin. Java as a better C++.

ACM SIGPLAN Notices, 31(11):21– 27, November 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:1996:OTC

[1589] Joseph Bergin. Object technology in the classroom: Java as a better C++. ACM SIGPLAN Notices, 31(11):21– 27, November 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marin:1996:EUA

[1590] Mike A. Marin. Effective use of assertions in C++. ACM SIGPLAN Notices, 31(11):28–32, November 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:LPSb

[1591] Preston Briggs. Literate programming: Sparse matrix multiplication. ACM SIGPLAN Notices, 31(11):33– 37, November 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1996:SMMb

[1592] Preston Briggs. Sparse matrix multiplication. ACM SIGPLAN Notices, 31 (11):33–37, November 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kniesel:1996:WDS

[1593] Günter Kniesel, Thilo Kielmann, Athanasios Demiris, and Milena Shteto. 5th workshop for doctoral students in object-oriented systems. ACM SIGPLAN Notices, 31(11):39– 48, November 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agarwal:1996:SDR

[1594] Rakesh Agarwal, Giorgio Bruno, and Marco Torchiano. Static, dynamic and run-time modeling of compound classes. ACM SIGPLAN Notices, 31 (11):49–55, November 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vandenBroek:1996:MEP

[1595] P. M. van den Broek and K. G. van den Berg. Musical equational programs: a functional approach. ACM SIGPLAN Notices, 31(11):56–65, November 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dotzel:1996:POS

[1596] Günter Dotzel and Hartmut Goebel. Porting the Oberon system to AlphaAXP. ACM SIGPLAN Notices, 31 (11):66–73, November 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Werther:1996:MPC

[1597] Ben Werther and Damian Conway. A modest proposal: C++ resyntaxed. ACM SIGPLAN Notices, 31(11):74– 82, November 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pugh:1996:RO

[1598] John Pugh. A report from OOP-SLA. ACM SIGPLAN Notices, 31(12): 1–2, December 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noble:1996:EWP

[1599] James Noble and Antero Taivalsaari.

ECOOP ’96 Workshop on Prototype Based Object Oriented Programming. ACM SIGPLAN Notices, 31(12):12– 17, December 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sharp:1996:PPS

[1600] Helen Sharp, Mary Lynn Manns, Phil McLaughlin, Maximo Prieto, and Mahesh Dodani. Pedagogical patterns — successes in teaching object technology: a workshop from OOPSLA ’96. ACM SIGPLAN Notices, 31(12):18–21, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dodgson:1996:ECS

[1601] David Dodgson. Evolution of the C++ standard library. ACM SIGPLAN Notices, 31(12):22–26, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:1996:BNT

[1602] Ron K. Cytron. Book nook: Out of their minds by Dennis Shasha. ACM SIGPLAN Notices, 31(12):27–28, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:1996:RTM

[1603] Ron K. Cytron. Review: Out of their minds by Shasha and Lezere. ACM SIGPLAN Notices, 31(12):27–28, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1996:WTF

[1604] Paul Frenger. A whirlwind tour of Forth resources. ACM SIGPLAN Notices, 31(12):29–32, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deBry:1996:RTW

[1605] Roger K. deBry. Report of TC-2 working groups: minutes of the IFIP TC-2 meeting July 13–14, 1996, Oxford, UK. ACM SIGPLAN Notices, 31(12): 33–38, December 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Houser:1996:MCT

[1606] Chris Houser. Manual and compiler for the terse and modular language DEM. ACM SIGPLAN Notices, 31(12):41– 51, December 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knudsen:1996:UOO

[1607] Jørgen Lindskov Knudsen and Ole Lehrmann Madsen. Using object-orientation as a common basis for system development education. ACM SIGPLAN Notices, 31(12):52–62, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Callegarin:1996:RIL

[1608] Giuseppe Callegarin. Requirements for an imperative language to host logic programming in a seamless way. ACM SIGPLAN Notices, 31(12):63–68, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sorkin:1996:SCM

[1609] Arthur Sorkin. Some comments on Morel and Renvoise’s “Global Optimization by Suppression of Partial Redundancies”. ACM SIGPLAN Notices, 31(12):69–72, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferrante:1996:ICS

[1610] Richard D. Ferrante and James R. Allard. Introducing a CPS style optimizer into an existing compiler. ACM SIG-PLAN Notices, 31(12):73–79, December 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Syropoulos:1996:NTCb

[1611] Apostolos Syropoulos. A Note On Type Checking Linear Functional Languages. ACM SIGPLAN Notices, 31(12):80– 83, December 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1996:APO

[1612] Anonymous. Advance program October 6–10, 1996 — San Jose, California. ACM SIGPLAN Notices, ?? (??):??, ???? 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1997:LCa

[1613] Barbara Ryder. Letter from the chair. ACM SIGPLAN Notices, 32(1):1, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trott:1997:PLP

[1614] Peter Trott. Programming languages: Past, present, and future: Sixteen prominent computer scientists assess our field. ACM SIGPLAN Notices, 32 (1):14–57, January 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Extended interview with Al V. Aho, Fran Allen, Jack Dennis, Jeanne Ferrante, Adele Goldberg, Ken Kennedy, Barbara Liskov, David B. MacQueen, Jean Sammet, Ravi Sethi, Burton J. Smith, Guy Steele, Jr., Bjarne Stroustrup, Andrew Tanenbaum, Mark Wegman, and Will Wulf.

Hankin:1997:PSS

[1615] Chris Hankin, Hanne Riis Nielson, and Jens Palsberg. Position statements on strategic directions for research on programming languages. ACM SIG-PLAN Notices, 32(1):59–65, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cardelli:1997:GC

[1616] Luca Cardelli. Global computation. ACM SIGPLAN Notices, 32(1):66–68, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Consel:1997:PAB

[1617] Charles Consel. Program adaptation based on program transformation. ACM SIGPLAN Notices, 32(1):69–72, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cousot:1997:PAA

[1618] Patrick Cousot. Program analysis:

The abstract implementation perspective. ACM SIGPLAN Notices, 32(1): 73–76, January 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanus:1997:IDP

[1619] Michael Hanus. Integration of declarative paradigms: Benefits and challenges. ACM SIGPLAN Notices, 32(1): 77–79, January 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harper:1997:MB

[1620] Robert Harper and John C. Mitchell.

ML and beyond. ACM SIGPLAN Notices, 32(1):80–85, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeMetayer:1997:PAS

[1621] Daniel Le Metayer. Program analysis for software engineering: New applications, new requirements, new tools. ACM SIGPLAN Notices, 32(1):86–88, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Metayer:1997:PAS

[1622] Daniel Le Métayer. Program analysis for software engineering: new applications, new requirements, new tools. ACM SIGPLAN Notices, 32(1):86–88, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nielson:1997:PPA

[1623] Flemming Nielson. Perspectives on program analysis. ACM SIGPLAN Notices, 32(1):89–91, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odersky:1997:CTS

[1624] Martin Odersky. Challenges in type systems research. ACM SIGPLAN Notices, 32(1):92–93, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paige:1997:FDP

[1625] R. Paige. Future directions in program transformations. ACM SIG-PLAN Notices, 32(1):94–98, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pettorossi:1997:FDP

[1626] Alberto Pettorossi and Maurizio Proietti. Future directions in program transformation. ACM SIGPLAN Notices, 32(1):99–102, January 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reynolds:1997:BM

[1627] John C. Reynolds. Beyond ML. ACM SIGPLAN Notices, 32(1):103–105, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Riecke:1997:PLS

[1628] Jon G. Riecke. Programming language semantics: The description of computational structures. ACM SIGPLAN Notices, 32(1):106–109, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1997:PPC

[1629] Barbara G. Ryder. A position paper on compile-time program analysis. ACM SIGPLAN Notices, 32(1):110–114, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:1997:NPF

[1630] David A. Schmidt. On the need for a popular formal semantics. ACM SIG-PLAN Notices, 32(1):115–116, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Volpano:1997:PPL

[1631] Dennis Volpano. Provably-secure programming languages for remote evaluation. ACM SIGPLAN Notices, 32(1): 117–119, January 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Volpano:1997:PSP

[1632] Dennis Volpano. Provably-secure programming languages for remote evaluation. ACM SIGPLAN Notices, 32(1): 117–119, January 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilhelm:1997:PAT

[1633] Reinhard Wilhelm. Program analysis: A toolmaker’s perspective. ACM SIG-PLAN Notices, 32(1):120–121, January 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1997:LCb

[1634] Barbara Ryder. Letter from the chair. ACM SIGPLAN Notices, 32 (2):1, February 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1997:SA

[1635] Barbara Ryder. SIGPLAN awards. ACM SIGPLAN Notices, 32(2):2, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alfredsen:1997:OOF

[1636] Knut Alfredsen and Bjørn Sæther. An object oriented framework for creating models in hydrology. ACM SIG-PLAN Notices, 32(2):16–19, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:1997:BNC

[1637] Ron K. Cytron. Book nook: Compiler Construction by Niklaus Wirth. ACM SIGPLAN Notices, 32(2):20–21, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:1997:RCC

[1638] Ron K. Cytron. Review: Compiler Construction by Wirth. ACM SIG-PLAN Notices, 32(2):20–21, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1997:IBF

[1639] Paul Frenger. An introduction to the Beetle Forth virtual processor. ACM SIGPLAN Notices, 32(2):22–25, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomas:1997:IBF

[1640] Reuben R. Thomas. An introduction to the Beetle Forth virtual processor. ACM SIGPLAN Notices, 32(2): 22–25, February 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yongquiang:1997:DIP

[1641] Sun Yongquiang, Lin Kai, and Shen Li. The design and implementation of a program development system based on rewriting method. ACM SIGPLAN Notices, 32(2):27–34, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yonqiang:1997:DIP

[1642] Sun Yonqiang, Lin Kai, and Shen Li.

The design and implementation of a program development system based on rewriting method. ACM SIGPLAN Notices, 32(2):27–34, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ejiogu:1997:DVC

[1643] Lem O. Ejiogu. On diminishing the vibrant confusion in software metrics. ACM SIGPLAN Notices, 32(2): 35–38, February 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ruckert:1997:CPP

[1644] Martin Ruckert. Conservative pretty printing. ACM SIGPLAN Notices, 32 (2):39–44, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deshpande:1997:SPA

[1645] G. Deshpande, T. Pearse, and P. Oman. Software portability annotated bibliography. ACM SIGPLAN Notices, 32(2):45–53, February 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vandenBrand:1997:REN

[1646] Mark van den Brand, Paul Klint, and Chris Verhoef. Re-engineering needs generic programming language terminology. ACM SIGPLAN Notices, 32(2): 54–61, February 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:1997:LVC

[1647] Mary Lou Soffa. Letter from the Vice-Chair for Conferences. ACM SIG-PLAN Notices, 32(3):1, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:1997:SSW

[1648] Mary Lou Soffa. SIGPLAN sponsorship of workshops. ACM SIG-PLAN Notices, 32(3):2, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:1997:SST

[1649] Mary Lou Soffa. SIGPLAN sponsorship of technical meetings. ACM SIG-PLAN Notices, 32(3):3–4, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:1997:MHI

[1650] Mary Lou Soffa. Meetings held incooperation with SIGPLAN. ACM SIGPLAN Notices, 32(3):5, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmnn:1997:CC

[1651] Seth Bergmnn. Conference corner.

ACM SIGPLAN Notices, 32(3):6–19, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1997:CPM

[1652] Anonymous. Call for participation:

MASPLAS ’97. ACM SIGPLAN Notices, 32(3):20, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Houlder:1997:CSA

[1653] Tom Houlder. C++ for scientific applications of iterative nature. ACM SIGPLAN Notices, 32(3):21–26, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1997:GIGa

[1654] Henry G. Baker. Garbage in/garbage out: When bad programs happen to good people. ACM SIGPLAN Notices, 32(3):27–31, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1997:WBP

[1655] Henry G. Baker. When bad programs happen to good people. ACM SIGPLAN Notices, 32(3):27–31, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1997:TCI

[1656] A. Michael Berman. Technical correspondence: Information for authors. ACM SIGPLAN Notices, 32(3):32, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1997:WCA

[1657] Bruce J. MacLennan. “who cares about elegance?” The role of aesthetics in programming language design. ACM SIGPLAN Notices, 32(3):33–37, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winkler:1997:PAU

[1658] Jürgen F. H. Winkler and Stefan Kauer. Proving assertions is also useful. ACM SIGPLAN Notices, 32(3):38–41, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cicekli:1997:IBS

[1659] Ilyas Cicekli. An intelligent backtracking schema in a logic programming environment. ACM SIGPLAN Notices, 32(3):42–49, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:1997:NAC

[1660] Jiajun Chen and Guoliang Zheng.

NDC++: an approach to concurrent extension of C++. ACM SIG-PLAN Notices, 32(3):50–56, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiajun:1997:NAC

[1661] Chen Jiajun and Zheng Guoliang.

NDC++: An approach to concurrent extension of C++. ACM SIG-PLAN Notices, 32(3):50–56, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Verschaeren:1997:NPF

[1662] Dennis Verschaeren, Annie Cuyt, and Brigitte Verdonk. On the need for predictable floating-point arithmetic in the programming languages Fortran 90 and C/C++. ACM SIGPLAN Notices, 32(3):57–64, March 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1997:LCc

[1663] Barbara Ryder. Letter from the Chair. ACM SIGPLAN Notices, 32(4):1, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1997:SAA

[1664] Barbara Ryder. SIGPLAN awards announced. ACM SIGPLAN Notices, 32 (4):2–15, April 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rada:1997:IET

[1665] Roy Rada and James Schoening. IEEE efforts towards standards in educational technology. ACM SIGPLAN Notices, 32(4):16–18, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Briggs:1997:PPP

[1666] Preston Briggs. Predicting the performance of parallel machines. ACM SIGPLAN Notices, 32(4):19–21, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1997:FWE

[1667] Henry G. Baker. I have a feeling we’re not in emerald city anymore. ACM SIGPLAN Notices, 32(4):22–26, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1997:GIGb

[1668] Henry G. Baker. Garbage in/garbage out: I have a feeling we’re not in Emerald City anymore. ACM SIGPLAN Notices, 32(4):22–26, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1997:RCT

[1669] Paul Frenger. Robot control techniques: Part one: a review of robotics languages. ACM SIGPLAN Notices, 32 (4):27–31, April 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1997:RRL

[1670] Paul Frenger. A review of robotics languages. ACM SIGPLAN Notices, 32 (4):27–31, April 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Discusses PostScript, Ghostscript, Tps (transportable PostScript), Java, and others.

Kelleghan:1997:MRS

[1671] James P. Kelleghan. Monitoring remote sensing devices in C++. ACM SIGPLAN Notices, 32(4):32–36, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wise:1997:CMR

[1672] G. Bowden Wise. C++: monitoring remote sensing devices in C++. ACM SIGPLAN Notices, 32(4):32–36, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zimmermann:1997:NBA

[1673] Chris Zimmermann. A note on before and after metamethods. ACM SIGPLAN Notices, 32(4):37–46, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rapin:1997:BSO

[1674] Charles Rapin. Block structured object programming. ACM SIGPLAN Notices, 32(4):47–54, April 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Proebsting:1997:STG

[1675] Todd A. Proebsting. Simple translation of goal-directed evaluation. ACM SIGPLAN Notices, 32(5):1– 6, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p1-proebsting/

.

Dolby:1997:AIA

[1676] Julian Dolby. Automatic inline allocation of objects. ACM SIGPLAN Notices, 32(5):7–17, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p7-dolby/

.

Ramalingam:1997:MLA

[1677] G. Ramalingam and Harini Srinivasan. A member lookup algorithm for C++. ACM SIGPLAN Notices, 32 (5):18–30, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p18-ramalingam/

.

Wagner:1997:IAR

[1678] Tim A. Wagner and Susan L. Graham.

Incremental analysis of real programming languages. ACM SIGPLAN Notices, 32(5):31–43, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p31-wagner/

.

Eide:1997:FFO

[1679] Eric Eide, Kevin Frei, Bryan Ford, Jay Lepreau, and Gary Lindstrom. Flick: A flexible, optimizing IDL compiler. ACM SIGPLAN Notices, 32 (5):44–56, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p44-eide/

.

Collberg:1997:RIM

[1680] Christian S. Collberg. Reverse interpretation + mutation analysis = automatic retargeting. ACM SIG-PLAN Notices, 32(5):57–70, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

                                        proceedings/pldi/258915/p57-collberg/

.

Diniz:1997:DFE

[1681] Pedro C. Diniz and Martin C. Rinard. Dynamic feedback: An effective technique for adaptive computing. ACM SIGPLAN Notices, 32 (5):71–84, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/258915/p71-diniz/

.

Ammons:1997:EHP

[1682] Glenn Ammons, Thomas Ball, and James R. Larus. Exploiting hardware performance counters with flow and context sensitive profiling. ACM SIGPLAN Notices, 32 (5):85–96, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/258915/p85-ammons/

.

Clinger:1997:GGC

[1683] William D. Clinger and Lars T. Hansen. Generational garbage collection and the radioactive decay model. ACM SIGPLAN Notices, 32 (5):97–108, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/258915/p97-clinger/

.

Poletto:1997:TSF

[1684] Massimiliano Poletto, Dawson R.

Engler, and M. Frans Kaashoek. tcc: A system for fast, flexible, and high-level dynamic code generation. ACM SIGPLAN Notices, 32(5): 109–121, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p109-poletto/

.

Goodwin:1997:IDA

[1685] David W. Goodwin. Interprocedural dataflow analysis in an executable optimizer. ACM SIGPLAN Notices, 32 (5):122–133, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p122-goodwin/

.

Ayers:1997:AI

[1686] Andrew Ayers, Richard Schooler, and Robert Gottlieb. Aggressive inlining. ACM SIGPLAN Notices, 32(5): 134–145, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p134-ayers/

.

Bodik:1997:ICB

[1687] Rastislav Bodík, Rajiv Gupta, and Mary Lou Soffa. Interprocedural conditional branch elimination. ACM SIGPLAN Notices, 32(5):146– 158, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p146-bodik/

.

Bodik:1997:PDC

[1688] Rastislav Bodík and Rajiv Gupta. Partial dead code elimination using slicing transformations. ACM SIGPLAN Notices, 32(5):159–170, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p159-bodik/

.

Hashemi:1997:EPM

[1689] Amir H. Hashemi, David R. Kaeli, and Brad Calder. Efficient procedure mapping using cache line coloring. ACM SIGPLAN Notices, 32(5): 171–182, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p171-hashemi/

.

Young:1997:NIB

[1690] Cliff Young, David S. Johnson, David R. Karger, and Michael D. Smith. Near-optimal intraprocedural branch alignment. ACM SIG-PLAN Notices, 32(5):183–193, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p183-young/

.

Young:1997:NOI

[1691] Cliff Young, David S. Johnson, Michael D. Smith, and David R. Karger. Near-optimal intraprocedural branch alignment. ACM SIG-PLAN Notices, 32(5):183–193, May 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eichenberger:1997:EFO

[1692] Alexandre E. Eichenberger and Edward S. Davidson. Efficient formulation for optimal modulo schedulers. ACM SIGPLAN Notices, 32(5): 194–205, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p194-eichenberger/

.

Dussart:1997:MPS

[1693] Dirk Dussart, Rogardt Heldal, and John Hughes. Module-sensitive program specialisation. ACM SIG-PLAN Notices, 32(5):206–214, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p206-dussart/

.

Dussart:1997:MSP

[1694] Dirk Dussart, Rogardt Heldal, and John Hughes. Module-sensitive program specialisation. ACM SIGPLAN Notices, 32(5):206–214, May 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sperber:1997:TPO

[1695] Michael Sperber and Peter Thiemann. Two for the price of one: Composing partial evaluation and compilation. ACM SIGPLAN Notices, 32(5): 215–225, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p215-sperber/

.

Jensen:1997:AVP

[1696] Jakob L. Jensen, Michael E. Jørgensen, Michael I. Schwartzbach, and Nils Klarlund. Automatic verification of pointer programs using monadic second-order logic. ACM SIG-PLAN Notices, 32(5):226–234, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p226-jensen/

.

Flanagan:1997:CSA

[1697] Cormac Flanagan and Matthias Felleisen. Componential set-based analysis. ACM SIGPLAN Notices, 32(5): 235–248, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p235-flanagan/

.

Flanagan:1997:CSB

[1698] Cormac Flanagan and Matthias Felleisen. Componential set-based analysis. ACM SIGPLAN Notices, 32(5): 235–248, May 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Webber:1997:PAU

[1699] Adam Brooks Webber. Program analysis using binary relations. ACM SIGPLAN Notices, 32(5):249–260, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p249-webber/

.

Heintze:1997:LTS

[1700] Nevin Heintze and David McAllester.

Linear-time subtransitive control flow analysis. ACM SIGPLAN Notices, 32 (5):261–272, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p261-heintze/

.

Chow:1997:NAP

[1701] Fred Chow, Sun Chan, Robert Kennedy, Shin-Ming Liu, Raymond Lo, and Peng Tu. A new algorithm for partial redundancy elimination based on SSA form. ACM SIGPLAN Notices, 32(5):273–286, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/pldi/258915/p273-chow/

.

Bergner:1997:SCM

[1702] Peter Bergner, Peter Dahl, David Engebretsen, and Matthew O’Keefe. Spill code minimization via interference region spilling. ACM SIGPLAN Notices, 32(5):287–295, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/258915/p287-bergner/

.

Lueh:1997:CCD

[1703] Guei-Yuan Lueh and Thomas Gross. Call-cost directed register allocation. ACM SIGPLAN Notices, 32(5):296– 307, May 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lueh:1997:CDR

[1704] Guei-Yuan Lueh and Thomas Gross. Call-cost directed register allocation. ACM SIGPLAN Notices, 32(5):296– 307, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/258915/p296-lueh/

.

Lu:1997:RPC

[1705] John Lu and Keith D. Cooper. Register promotion in C programs. ACM SIGPLAN Notices, 32(5):308– 319, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p308-lu/

.

Taura:1997:FGM

[1706] Kenjiro Taura and Akinori Yonezawa.

Fine-grain multithreading with minimal compiler support — a cost effective approach to implementing efficient multithreading languages. ACM SIG-PLAN Notices, 32(5):320–333, May 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taura:1997:FMM

[1707] Kenjiro Taura and Akinori Yonezawa.

Fine-grain multithreading with minimal compiler support — A cost effective approach to implementing efficient multithreading languages. ACM SIGPLAN Notices, 32(5):320– 333, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/258915/p320-taura/

.

Chandra:1997:DDS

[1708] Rohit Chandra, Ding-Kai Chen, Robert Cox, Dror E. Maydan, Nenad Nedeljkovic, and Jennifer M. Anderson. Data distribution support on distributed shared memory multiprocessors. ACM SIGPLAN Notices, 32(5): 334–345, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

                                        proceedings/pldi/258915/p334-chandra/

.

Kodukula:1997:DCM

[1709] Induprakas Kodukula, Nawaaz Ahmed, and Keshav Pingali. Data-centric multi-level blocking. ACM SIGPLAN Notices, 32(5):346–357, May 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kodukula:1997:DMB

[1710] Induprakas Kodukula, Nawaaz Ahmed, and Keshav Pingali. Data-centric multi-level blocking. ACM SIG-PLAN Notices, 32(5):346–357, May 1997. CODEN SINODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/258915/p346-kodukula/

.

Ernst:1997:CC

[1711] Jens Ernst, William Evans, Christopher W. Fraser, Steven Lucco, and Todd A. Proebsting. Code compression. ACM SIGPLAN Notices, 32(5): 358–365, May 1997. CODEN SIN-ODQ. ISBN 0-89791-907-6. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/258915/p358-ernst/

.

Pollock:1997:SDF

[1712] Lori Pollock. The SIG discretionary fund: Role and process. ACM SIG-PLAN Notices, 32(6):1–??, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1997:CCC

[1713] Anonymous. Conference corner: Calendar. ACM SIGPLAN Notices, 32 (6):2–??, June 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Selvaraj:1997:IDF

[1714] Asokan R. Selvaraj and Debasish Ghosh. Implementation of a database factory. ACM SIGPLAN Notices, 32 (6):14–18, June 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1997:FRL

[1715] Paul Frenger. Forth as a robotics language: Part two. ACM SIG-PLAN Notices, 32(6):19–22, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1997:CCS

[1716] Henry G. Baker. COMFY — comfortable set of control primitives for machine language programming. ACM SIGPLAN Notices, 32(6):23–27, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1997:GIGc

[1717] Henry G. Baker. Garbage in/garbage out: COMFY — A comfortable set of control primitives for machine language programming. ACM SIGPLAN Notices, 32(6):23–??, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culwin:1997:JCC

[1718] Fintain Culwin. Java in the computing curriculum conference. ACM SIG-PLAN Notices, 32(6):28–32, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culwin:1997:RJC

[1719] Fintan Culwin. A report on the ‘Java in the Computing Curriculum Conference’ held at South Bank University, London, on 4th February 1997. ACM SIGPLAN Notices, 32(6):28–32, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carter:1997:CDF

[1720] Debbie K. Carter and Albert D. Baker. Considerations in developing a formally-based visual programming language reference manual: A case study on the SLAM II language. ACM SIGPLAN Notices, 32(6):34–39, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vetter:1997:CSA

[1721] J. S. Vetter. Computational steering annotated bibliography. ACM SIG-PLAN Notices, 32(6):40–44, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:1997:FAT

[1722] Myung Ho Kim. A functional approach to type constraints of generic definitions. ACM SIGPLAN Notices, 32 (6):45–52, June 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zheng:1997:GMA

[1723] Naixiao Zheng, Hongjun Zheng, and Zongyan Qiu. Garment: a mechanism for abstraction and encapsulation of languages. ACM SIGPLAN Notices, 32(6):53–60, June 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naixiao:1997:GMA

[1724] Zhang Naixiao, Zheng Honjun, and Qiu Zongyan. Garment: A mechanism for abstraction and encapsulation of languages. ACM SIGPLAN Notices, 32 (6):53–??, June 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:1997:CVS

[1725] Wen-Ke Chen, Jia-Su Sun, and Zhi-Min Tang. Consummating virtuality to support more polymorphism in C++. ACM SIGPLAN Notices, 32(6):61–68, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frisbie:1997:FPL

[1726] Ralph Frisbie, Richard Hendrickson, and Michael Metcalf. The F programming language. ACM SIGPLAN Notices, 32(6):69–74, June 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zambonelli:1997:HAM

[1727] Franco Zambonelli. How to achieve modularity in distributed object allocation. ACM SIGPLAN Notices, 32 (6):75–82, June 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Subhlok:1997:NMI

[1728] Jaspal Subhlok and Bwolen Yang. A new model for integrated nested task and data parallel programming. ACM SIGPLAN Notices, 32(7):1–12, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hu:1997:HPF

[1729] Y. Charlie Hu, S. Lennart Johnsson, and Shang-Hua Teng. High Performance Fortran for highly irregular problems. ACM SIGPLAN Notices, 32 (7):13–24, July 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Narlikar:1997:SEI

[1730] Girija J. Narlikar and Guy E. Blelloch. Space-efficient implementation of nested parallelism. ACM SIG-PLAN Notices, 32(7):25–36, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Narlikar:1997:SIN

[1731] Girija J. Narlikar and Guy E. Blelloch. Space-efficient implementation of nested parallelism. ACM SIG-PLAN Notices, 32(7):25–??, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:1997:DPA

[1732] Xingbin Zhang and Andrew A. Chien.

Dynamic pointer alignment: Tiling and communication optimizations for parallel pointer-based computations. ACM SIGPLAN Notices, 32(7):37–47, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:1997:CSD

[1733] Honghui Lu, Alan L. Cox, Sandhya Dwarkadas, Ramakrishnan Rajamony, and Willy Zwaenepoel. Compiler and software distributed shared memory support for irregular applications. ACM SIGPLAN Notices, 32(7):48–56, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fu:1997:STE

[1734] Cong Fu and Tao Yang. Space and time efficient execution of parallel irregular computations. ACM SIG-PLAN Notices, 32(7):57–68, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bianchini:1997:IPP

[1735] Ricardo Bianchini, Enrique V. Carrera, and Leonidas Kontothanassis. The interaction of parallel programming constructs and coherence protocols. ACM SIGPLAN Notices, 32(7):69–79, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raghavachari:1997:ALM

[1736] Mukund Raghavachari and Anne Rogers. Ace: Linguistic mechanisms for customizable protocols. ACM SIG-PLAN Notices, 32(7):80–89, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amza:1997:TBF

[1737] Cristiana Amza, Alan Cox, Karthick Rajamani, and Willy Zwaenepoel. Tradeoffs between false sharing and aggregation in software distributed shared memory. ACM SIGPLAN Notices, 32(7):90–99, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chandra:1997:OCH

[1738] Satish Chandra and James R. Larus. Optimizing communication in HPF programs on fine-grain distributed shared memory. ACM SIGPLAN Notices, 32(7):100–111, July 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rinard:1997:EFG

[1739] Martin Rinard. Effective fine-grain synchronization for automatically parallelized programs using optimistic synchronization primitives. ACM SIG-PLAN Notices, 32(7):112–123, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rinard:1997:EFS

[1740] Martin Rinard. Effective fine-grain synchronization for automatically parallelized programs using optimistic synchronization primitives. ACM SIG-PLAN Notices, 32(7):112–??, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sodan:1997:ENN

[1741] Angela Sodan, Guang R. Gao, Olivier Maquelin, Jens-Uwe Schultz, and Xin-Min Tian. Experiences with nonnumeric applications on multithreaded architectures. ACM SIGPLAN Notices, 32(7):124–135, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sodan:1997:ENA

[1742] Angela Sodan, Guang R. Gao, Oliver Maquelin, Jens-Uwe Schultz, and Xin-Min Tian. Experiences with nonnumeric applications on multithreaded architectures. ACM SIGPLAN Notices, 32(7):124–??, July 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hascoet:1997:APC

[1743] Laurent Hascoët. Automatic placement of communications in meshpartitioning parallelization. ACM SIG-PLAN Notices, 32(7):136–144, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:1997:PBF

[1744] Bwolen Yang and David R. O’Hallaron.

Parallel breadth-first BDD construction. ACM SIGPLAN Notices, 32(7): 145–156, July 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:1997:PBB

[1745] Bwolen Yang and David R. O’Hallaron. Parallel breadth-first BDD construction. ACM SIGPLAN Notices, 32 (7):145–??, July 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gu:1997:EEA

[1746] Junjie Gu, Zhiyuan Li, and Gyungho Lee. Experience with efficient array data flow analysis for array privatization. ACM SIGPLAN Notices, 32(7): 157–167, July 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coelho:1997:CDM

[1747] Fabien Coelho. Compiling dynamic mappings with array copies. ACM SIGPLAN Notices, 32(7):168–179, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prasanna:1997:CPM

[1748] G. N. Srinivasa Prasanna. Compilation of parallel multimedia computations — extending retiming theory and Amdahl’s Law. ACM SIG-PLAN Notices, 32(7):180–192, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhou:1997:RCC

[1749] Yuanyuan Zhou, Liviu Iftode, Jaswinder Pal Singh, Kai Li, Brian R. Toonen, Ioannis Schoinas, Mark D. Hill, and David A. Wood. Relaxed consistency and coherence granularity in DSM systems: A performance evaluation. ACM SIGPLAN Notices, 32(7):193–205, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frens:1997:ABM

[1750] Jeremy D. Frens and David S. Wise.

Auto-blocking matrix-multiplication or tracking BLAS3 performance from source code. ACM SIGPLAN Notices, 32(7):206–216, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:1997:ARP

[1751] Dongming Jiang, Hongzhang Shan, and Jaswinder Pal Singh. Application restructuring and performance portability on shared virtual memory and hardware-coherent multiprocessors. ACM SIGPLAN Notices, 32 (7):217–229, July 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lim:1997:PIC

[1752] Beng-Hong Lim, Chi-Chao Chang, Grzegorz Czajkowski, and Thorsten Von Eicken. Performance implications of communication mechanisms in allsoftware global address space systems. ACM SIGPLAN Notices, 32(7):230– 239, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:1997:SMP

[1753] Zhichen Xu, James R. Larus, and Barton P. Miller. Shared-memory performance profiling. ACM SIG-PLAN Notices, 32(7):240–251, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:1997:SPP

[1754] Zhichen Xu, James R. Larus, and Barton P. Miller. Shared-memory performance profiling. ACM SIGPLAN Notices, 32(7):240–??, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiang:1997:IPS

[1755] Dongming Jiang and Jaswinder Pal Singh. Improving parallel shear-warp volume rendering on shared address space multiprocessors. ACM SIG-PLAN Notices, 32(7):252–263, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taura:1997:EGC

[1756] Kenjiro Taura and Akinori Yonezawa. An effective garbage collection strategy for parallel programming languages on large scale distributed-memory machines. ACM SIGPLAN Notices, 32 (7):264–275, July 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frank:1997:LMC

[1757] Matthew I. Frank, Anant Agarwal, and Mary K. Vernon. LoPC: Modeling contention in parallel algorithms. ACM SIGPLAN Notices, 32(7):276–287, July 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Banerjee:1997:MPT

[1758] Anindya Banerjee. A modular, polyvariant and type-based closure analysis. ACM SIGPLAN Notices, 32(8): 1–10, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dimock:1997:STF

[1759] Allyn Dimock, Robert Muller, Franklyn Turbak, and J. B. Wells. Strongly typed flow-directed representation transformations (extended abstract). ACM SIGPLAN Notices, 32(8):11–24, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bell:1997:TDD

[1760] Jeffrey M. Bell, Françoise Bellegarde, and James Hook. Type-driven defunctionalization. ACM SIGPLAN Notices, 32(8):25–37, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bell:1997:TD

[1761] Jeffrey M. Bell, Françoise Bellegarde, and James Hook. Type-driven defunctionalization. ACM SIGPLAN Notices, 32(8):25–??, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gasser:1997:SRC

[1762] Kirsten L. Solberg Gasser, Flemming Nielson, and Hanne Riis Nielson. Systematic realisation of control flow analyses for CML. ACM SIGPLAN Notices, 32(8):38–51, August 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Erwig:1997:FPG

[1763] Martin Erwig. Functional programming with graphs. ACM SIGPLAN Notices, 32(8):52–65, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Okasaki:1997:CDE

[1764] Chris Okasaki. Catenable double-ended queues. ACM SIGPLAN Notices, 32 (8):66–74, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Okasaki:1997:CDQ

[1765] Chris Okasaki. Catenable double-ended queues. ACM SIGPLAN Notices, 32 (8):66–??, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fahndrich:1997:SCP

[1766] Manuel Fähndrich and John Boyland. Statically checkable pattern abstractions. ACM SIGPLAN Notices, 32(8): 75–84, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shao:1997:FRA

[1767] Zhong Shao. Flexible representation analysis. ACM SIGPLAN Notices, 32 (8):85–98, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ashley:1997:EFA

[1768] J. Michael Ashley. The effectiveness of flow analysis for inlining. ACM SIG-PLAN Notices, 32(8):99–111, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blume:1997:LSH

[1769] Matthias Blume and Andrew W. Appel. Lambda-splitting: a higher-order approach to cross-module optimizations. ACM SIGPLAN Notices, 32(8): 112–124, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blume:1997:LHA

[1770] Matthias Blume and Andrew W. Appel. Lambda-splitting: A higherorder approach to cross-module optimizations. ACM SIGPLAN Notices, 32 (8):112–??, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crary:1997:FIH

[1771] Karl Crary. Foundations for the implementation of higher-order subtyping. ACM SIGPLAN Notices, 32(8): 125–135, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marlow:1997:PSS

[1772] Simon Marlow and Philip Wadler. A practical subtyping system for Erlang. ACM SIGPLAN Notices, 32(8):136– 149, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heintze:1997:CSB

[1773] Nevin Heintze and David McAllester.

On the complexity of set-based analysis. ACM SIGPLAN Notices, 32(8):

150–163, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heintze:1997:CSA

[1774] Nevin Heintze and David McAllester. On the complexity of set-based analysis. ACM SIGPLAN Notices, 32(8): 150–??, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hu:1997:TCE

[1775] Zhenjiang Hu, Hideya Iwasaki, Masato Takeichi, and Akihiko Takano. Tupling calculation eliminates multiple data traversals. ACM SIGPLAN Notices, 32(8):164–175, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chin:1997:BIM

[1776] Wei-Ngan Chin and Masami Hagiya. A bounds inference method for vectorbased memoization. ACM SIGPLAN Notices, 32(8):176–187, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lawall:1997:GDO

[1777] Julia L. Lawall and Harry G. Mairson. On global dynamics of optimal graph reduction. ACM SIGPLAN Notices, 32(8):188–195, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Armstrong:1997:DE

[1778] Joe Armstrong. The development of Erlang. ACM SIGPLAN Notices, 32 (8):196–203, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dussart:1997:TSI

[1779] Dirk Dussart, John Hughes, and Peter Thiemann. Type specialisation for imperative languages. ACM SIGPLAN Notices, 32(8):204–216, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kagawa:1997:CRS

[1780] Koji Kagawa. Compositional references for stateful functional programming. ACM SIGPLAN Notices, 32(8): 217–226, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Launchbury:1997:MSA

[1781] John Launchbury and Amr Sabry.

Monadic state: Axiomatization and type safety. ACM SIGPLAN Notices, 32(8):227–238, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Draves:1997:IBA

[1782] Scott Draves. Implementing bitaddressing with specialization. ACM SIGPLAN Notices, 32(8):239–250, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Draves:1997:IBS

[1783] Scott Draves. Implementing bitaddressing with specialization. ACM SIGPLAN Notices, 32(8):239–??, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Claessen:1997:SGP

[1784] Koen Claessen, Ton Vullinghs, and Erik Meijer. Structuring graphical paradigms in TkGofer. ACM SIG-PLAN Notices, 32(8):251–262, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:1997:FRA

[1785] Conal Elliott and Paul Hudak. Functional reactive animation. ACM SIG-PLAN Notices, 32(8):263–273, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shivers:1997:AMO

[1786] Olin Shivers. Automatic management of operating-system resources. ACM SIGPLAN Notices, 32(8):274–279, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ungureanu:1997:FMD

[1787] Cristian Ungureanu and Benjamin Goldberg. Formal models of distributed memory management. ACM SIGPLAN Notices, 32(8):280–291, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hicks:1997:MCC

[1788] Michael W. Hicks, Jonathan T. Moore, and Scott M. Nettles. The measured cost of copying garbage collection mechanisms. ACM SIGPLAN Notices, 32(8):292–305, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aerts:1997:GTF

[1789] Kris Aerts and Karel De Vlaminck. A GUI on top of a functional language (poster). ACM SIGPLAN Notices, 32 (8):308, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Claessen:1997:GCP

[1790] Koen Claessen. Graphs in compilation (poster). ACM SIGPLAN Notices, 32 (8):309, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Claessen:1997:GC

[1791] Koen Claessen. Graphs in compilation. ACM SIGPLAN Notices, 32(8): 309–??, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1997:DMF

[1792] Byron Cook and John Launchbury.

Disposable memo functions (extended abstract). ACM SIGPLAN Notices, 32 (8):310, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dornan:1997:ADT

[1793] Chris Dornan. Affordable dynamic types (poster). ACM SIGPLAN Notices, 32(8):311, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duris:1997:FPA

[1794] Etienne Duris. Functional programming and attribute grammars deforestation (poster). ACM SIGPLAN Notices, 32(8):312, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krasnogor:1997:PFM

[1795] Natalio Krasnogor, Pablo E. Martínez López, Pablo Mocciola, and David Pelta. Protein folding meets functional programming (poster). ACM SIGPLAN Notices, 32(8):313, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krasnogor:1997:MSF

[1796] Natalio Krasnogor, Pablo E. Martínez López, Pablo Mocciola, and David Pelta. Modelling string folding with G2L grammars (poster). ACM SIG-PLAN Notices, 32(8):314, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lassila:1997:FME

[1797] Eero Lassila. A functional macro expansion system for optimizing code generation: gaining context-sensitivity without losing confluence (poster). ACM SIGPLAN Notices, 32(8):315, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McAdam:1997:BMP

[1798] Bruce J. McAdam. BigTypes in ML (poster). ACM SIGPLAN Notices, 32 (8):316, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McAdam:1997:BM

[1799] Bruce J. McAdam and Stephen Gilmore. BigTypes in ML. ACM SIG-PLAN Notices, 32(8):316–??, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pareja-Flores:1997:SFT

[1800] Cristóbal Pareja-Flores and J. Ángel Velázquez-Iturbide. Synthesis of functions by transformations and constraints (poster). ACM SIGPLAN Notices, 32(8):317, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Richard:1997:ISU

[1801] G. Richard, F. Saubion, and A. Tellez-Arenas. Intensional sets using explicit substitutions (poster). ACM SIGPLAN Notices, 32(8):318, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sabelfeld:1997:SSA

[1802] Viktor Sabelfeld and Andrei Sabelfeld.

Simple semantic analysis problems for functional programs (poster). ACM SIGPLAN Notices, 32(8):319, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spiliopoulou:1997:DPP

[1803] Eleni Spiliopoulou, Ian Holyer, and Neil Davies. Distributed programming, a purely functional approach (poster). ACM SIGPLAN Notices, 32(8):320, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taha:1997:MSPa

[1804] Walid Taha and Tim Sheard. Multistage programming (poster). ACM SIGPLAN Notices, 32(8):321, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taha:1997:MP

[1805] Walid Taha and Tim Sheard. Multistage programming. ACM SIGPLAN Notices, 32(8):321–??, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deWailly:1997:SFP

[1806] Guilhem de Wailly. λ-flow (poster): a parallel functional synchronous dataflow language. ACM SIGPLAN Notices, 32(8):322, August 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ostvold:1997:ISR

[1807] Bjarte M. Østvold. Inductive synthesis of recursive functional programs (poster abstract). ACM SIGPLAN Notices, 32(8):323, August 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wailly:1997:LPF

[1808] Guilhem De Wailly. l-flow: A parallel functional synchronous dataflow language. ACM SIGPLAN Notices, 32 (8):377–??, August 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1997:LFC

[1809] Barbara G. Ryder. Letter from the Former Chair. ACM SIGPLAN Notices, 32 (9):1, September 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1997:ARA

[1810] Barbara G. Ryder. 1996–1997 annual report on ACM SIGPLAN. ACM SIG-PLAN Notices, 32(9):2–4, September 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1997:CCC

[1811] Seth Bergmann. Conference corner:

Calendar. ACM SIGPLAN Notices, 32(9):5–14, September 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:1997:CCH

[1812] Tim Bergin. Conference corner: History of Programming Languages Conference II: HOPL II — closing panel: The history of programming: Does our present past have a future? ACM SIG-PLAN Notices, 32(9):15–37, September 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergin:1997:HIC

[1813] Tim Bergin and Michael S. Mahoney.

HOPL II: closing panel: The history of programming (panel): does our present past have a future? ACM SIGPLAN Notices, 32(9):15–37, September 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1997:MCC

[1814] Wei Li and Gansheng Li. Making C++ concurrent and real-time. ACM SIG-PLAN Notices, 32(9):39–46, September 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See correction [1966].

Mansouri:1997:PAB

[1815] F. Z. Mansouri and C. A. Higgins. Prolog: An annotated bibliography. ACM SIGPLAN Notices, 32(9):47– 53, September 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casimir:1997:BF

[1816] Rommert J. Casimir. Basic forever. ACM SIGPLAN Notices, 32(9):54– 56, September 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lu:1997:MMD

[1817] Pei Lu, Dachuan Yu, and Jian Lu. A method for modifying dynamically classes in the object-oriented dynamic programming environment. ACM SIG-PLAN Notices, 32(9):57–60, September 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pei:1997:MMD

[1818] Lu Pei, Yu Dachuan, and Lu Jian. A method for modifying dynamically classes in the object-oriented dynamic programming environment. ACM SIG-PLAN Notices, 32(9):57–60, September 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jianjua:1997:MCN

[1819] Jianjua Zhao, Jiajun Chen, and Guoliang Zheng. Message conversion and a new type system for OO model. ACM SIGPLAN Notices, 32(9):61– 67, September 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:1997:MCN

[1820] Jianhua Zhao, Jiajun Chen, and Guoliang Zheng. Message conversion and a new type system for OO model. ACM SIGPLAN Notices, 32(9):61– 67, September 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qiu:1997:SHN

[1821] Zongyan Qiu. Static head nodes in the operation on dynamic structures: a useful programming technique. ACM SIGPLAN Notices, 32(9): 68–71, September 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mezini:1997:MCB

[1822] Mira Mezini. Maintaining the consistency and behavior of class libraries during their evolution. ACM SIG-PLAN Notices, 32(10):1–21, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mezini:1997:MCC

[1823] Mira Mezini. Maintaining the consistency of class libraries during their evolution. ACM SIGPLAN Notices, 32 (10):1–21, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Price:1997:AMR

[1824] Margaretha W. Price and Steven A. Demurjian, Sr. Analyzing and measuring reusability in object-oriented designs. ACM SIGPLAN Notices, 32(10): 22–33, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weihe:1997:RAS

[1825] Karsten Weihe. Reuse of algorithms: Still a challenge to object-oriented programming. ACM SIGPLAN Notices, 32(10):34–48, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agesen:1997:ATP

[1826] Ole Agesen, Stephen N. Freund, and John C. Mitchell. Adding type parameterization to the Java language. ACM SIGPLAN Notices, 32(10):49–65, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyland:1997:PMI

[1827] John Boyland and Giuseppe Castagna. Parasitic methods: An implementation of multi-methods for Java. ACM SIG-PLAN Notices, 32(10):66–76, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Litman:1997:MDC

[1828] Diane Litman, Peter F. Patel-Schneider, and Anil Mishra. Modeling dynamic collections of interdependent objects using path-based rules. ACM SIGPLAN Notices, 32(10):77–92, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1997:LOO

[1829] Lougie Anderson, Mike Carey, Ken Jacobs, Erin Kinikin, and David Maier. Looking for the objects in objectrelational DBMSs (panel). ACM SIG-PLAN Notices, 32(10):93, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1997:PLO

[1830] Lougie Anderson, Mike Carry, Ken Jacobs, Erin Kinikin, and David Maier. Panel 1: Looking for the objects in object-relational DBMs. ACM SIG-PLAN Notices, 32(10):93, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Franklin:1997:FSD

[1831] Michael Franklin and Stanley Zdonik.

A framework for scalable disseminationbased systems. ACM SIGPLAN Notices, 32(10):94–105, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zdonik:1997:FSD

[1832] Stan Zdonik. A framework for scalable dissemination-based systems. ACM SIGPLAN Notices, 32(10):94–105, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rao:1997:ELC

[1833] Bindu Rama Rao, Chad Edwards, Ted Linden, Regan Moore, and Mark Seager. Exploring largeness, complexity and scalability from the OOT perspective (panel). ACM SIGPLAN Notices, 32(10):106–107, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rao:1997:PEL

[1834] Bindu Rama Rao, Chad Edwards, Ted Linden, Reagan Moore, and Mark Seager. Panel 2: Exploring largeness, complexity and scalability from the OOT perspective. ACM SIGPLAN Notices, 32(10):106–107, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harel:1997:MAO

[1835] David Harel. On modeling and analyzing object behavior. ACM SIG-PLAN Notices, 32(10):108, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). No paper in volume, but table of contents includes an entry for this invited talk.

Grove:1997:CGC

[1836] David Grove, Greg DeFouw, Jeffrey Dean, and Craig Chambers. Call graph construction in object-oriented languages. ACM SIGPLAN Notices, 32 (10):108–124, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zendra:1997:EDD

[1837] Olivier Zendra, Dominique Colnet, and Suzanne Collin. Efficient dynamic dispatch without virtual function tables: The small Eiffel compiler. ACM SIG-PLAN Notices, 32(10):125–141, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vitek:1997:ETI

[1838] Jan Vitek, R. Nigel Horspool, and Andreas Krall. Efficient type inclusion tests. ACM SIGPLAN Notices, 32(10): 142–157, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coplien:1997:CEC

[1839] Jim Coplien, Luke Hohmann, Norm Kerth, John Rae-Grant, and Eileen Strider. Changing the engine of the car? (panel): while driving 60 miles an hour! ACM SIGPLAN Notices, 32(10): 158–161, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coplien:1997:PCE

[1840] Jim Coplien, Luke Hohmann, Norm Kerth, Jon Rae-Grant, and Eileen Strider. Panel 3: Changing the engine of the car? while driving 60 miles an hour! ACM SIGPLAN Notices, 32(10): 158–161, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiczales:1997:AOP

[1841] Gregor Kiczales. Aspect oriented programming. ACM SIGPLAN Notices, 32(10):162, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). No paper in volume, but table of contents includes an entry for this invited talk.

Tokoro:1997:SAN

[1842] Mario Tokoro. Software agents and network-based software — new research directions. ACM SIGPLAN Notices, 32(10):162, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). No paper in volume, but table of contents includes an entry for this invited talk.

Hudson:1997:GCW

[1843] Richard L. Hudson, Ron Morrison, J. Eliot B. Moss, and David S. Munro. Garbage collecting the world: One car at a time. ACM SIGPLAN Notices, 32(10):162–175, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hayes:1997:ENF

[1844] Barry Hayes. Ephemerons: A new finalization mechanism. ACM SIG-PLAN Notices, 32(10):176–183, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1997:DPR

[1845] Timothy H. Harrison, David L. Levine, and Douglas C. Schmidt. The design and performance of a real-time CORBA event service. ACM SIG-PLAN Notices, 32(10):184–200, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coleman:1997:PUL

[1846] Derek Coleman, John Artim, Viktor Ohnjec, Erick Rivas, Jim Rumbaugh, and Rebecca Wirfs-Brock. Panel 4: UML: The language of software blueprints? ACM SIGPLAN Notices, 32(10):201–205, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coleman:1997:UPL

[1847] Derek Coleman, Viktor Ohnjec, John Artim, Erick Rivas, Jim Rumbaugh, and Rebecca Wirfs-Bracle. UML (panel): the language of blueprints for software? ACM SIGPLAN Notices, 32(10):201–205, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eliot:1997:COL

[1848] J. Eliot and B. Moss. Compiling objectoriented languages: Achievements and promise. ACM SIGPLAN Notices, 32 (10):206, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). No paper in volume, but table of contents includes an entry for this invited talk.

Lorenz:1997:TDP

[1849] David H. Lorenz. Tiling design patterns — A case study using the interpreter pattern. ACM SIGPLAN Notices, 32(10):206, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). No paper in volume, but table of contents includes an entry for this invited talk.

Mendelsohn:1997:CHF

[1850] Noah Mendelsohn. Components: How far can we go? ACM SIGPLAN Notices, 32(10):206–217, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Riehle:1997:CDP

[1851] Dirk Riehle. Composite design patterns. ACM SIGPLAN Notices, 32(10): 218–228, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haines:1997:OIA

[1852] Matthew Haines. An open implementation analysis and design for lightweight threads. ACM SIGPLAN Notices, 32 (10):229–242, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oler:1997:MRD

[1853] Jon Oler, Gary Lindstrom, and Terence Critchlow. Migrating relational data to an ODBMS: Strategies and lessons from a molecular biology experience. ACM SIGPLAN Notices, 32(10):243– 252, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alagic:1997:OOM

[1854] Suad Alagic. The ODMG object model:

Does it make sense? ACM SIGPLAN Notices, 32(10):253–270, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tip:1997:CHS

[1855] Frank Tip and Peter F. Sweeney. Class hierarchy specialization. ACM SIG-PLAN Notices, 32(10):271–285, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Volanschi:1997:DSO

[1856] Eugen N. Volanschi, Charles Counsel, Gilles Muller, and Crispin Cowan. Declarative specialization of objectoriented programs. ACM SIGPLAN Notices, 32(10):286–300, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1997:OIL

[1857] Laura Hill, Bruce Anderson, Adele Goldberg, Gregor Kiczales, Colin Scott, and Kevin Tyson. The OT idea life-cycle (panel): from eureka! to shrink wrap. ACM SIGPLAN Notices, 32(10):301–303, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1997:POL

[1858] Laura Hill, Bruce Anderson, Adele Goldberg, Gregor Kiczales, Colin Scott, and Kevin Tyson. Panel 5: The OT life-cycle: From Eureka! to shrink wrap. ACM SIGPLAN Notices, 32(10): 301–303, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lencevicius:1997:QBD

[1859] Raimondas Lencevicius, Urs Hölzle, and Ambuj K. Singh. Query-based debugging of object-oriented programs. ACM SIGPLAN Notices, 32(10):304– 317, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lencevicius:1997:QDO

[1860] Raimondas Lencevicius, Urs Holzle, and Ambuj K. Singh. Query-based debugging of object-oriented programs. ACM SIGPLAN Notices, 32(10):304– 317, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ingalls:1997:BFS

[1861] Dan Ingalls, Ted Kaehler, John Maloney, Scott Wallace, and Alan Kay. Back to the future: the story of Squeak, a practical Smalltalk written in itself. ACM SIGPLAN Notices, 32(10):318– 326, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kent:1997:CDV

[1862] Stuart Kent. Constraint diagrams: visualizing invariants in object-oriented models. ACM SIGPLAN Notices, 32 (10):327–341, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1997:BHP

[1863] Steven Fraser, Kent Beck, Grady Booch, Jim Coplien, Ralph Johnson, and Bill Opdyke. Beyond the hype (panel): do patterns and frameworks reduce discovery costs? ACM SIG-PLAN Notices, 32(10):342–344, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1997:PBH

[1864] Steven D. Fraser, Kent Beck, Grady Booch, Jim Coplien, Ralph Johnson, and Bill Opdyke. Panel 6: Beyond the hype: Do patterns and frameworks reduce discovery costs? ACM SIGPLAN Notices, 32(10):342–344, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1997:AI

[1865] Anonymous. Author index. ACM SIGPLAN Notices, 32(10):345, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dodani:1997:PES

[1866] Mahesh Dodani. Panel: Educator’s symposium report. ACM SIGPLAN Notices, 32(10):345, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper did not appear in the printed proceedings, but does appear in the table of contents, where is says that the paper will appear in the Addendum.

Felten:1997:SNF

[1867] Edward W. Felten. Securing the networked future: Can objects help? ACM SIGPLAN Notices, 32(10):345, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper did not appear in the printed proceedings, but does appear in the table of contents.

Kay:1997:KAC

[1868] Alan Kay. Keynote address: The computer revolution hasn’t happened yet! ACM SIGPLAN Notices, 32 (10):345, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper did not appear in the printed proceedings, but does appear in the table of contents.

Kriens:1997:PDR

[1869] Peter Kriens. Panel: DesignFest report. ACM SIGPLAN Notices, 32 (10):345, October 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper did not appear in the printed proceedings, but does appear in the table of contents, where is says that the paper will appear in the Addendum.

Marcus:1997:PMW

[1870] Bob Marcus. Panel: Mid-year workshops report. ACM SIGPLAN Notices, 32(10):345, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper did not appear in the printed proceedings, but does appear in the table of contents.

Anonymous:1997:TI

[1871] Anonymous. Title index. ACM SIG-PLAN Notices, 32(10):346, October 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:1997:LNE

[1872] Mary Lou Soffa. Letter from the newly elected Chair. ACM SIGPLAN Notices, 32(11):1–??, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1997:CP

[1873] Anonymous. Calls for papers.

ACM SIGPLAN Notices, 32(11):2– ??, November 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mueller:1997:ASW

[1874] Frank Mueller. The 3rd ACM SIG-PLAN Workshop on Languages, Compilers, and Tools for Real-Time Systems. ACM SIGPLAN Notices, 32(11): 15–16, November 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mueller:1997:WRL

[1875] Frank Mueller. Workshop report:

LCTES ’97. ACM SIGPLAN Notices, 32(11):15–??, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weisert:1997:PEP

[1876] Conrad Weisert. Point-extent pattern for dimensioned numeric classes.

ACM SIGPLAN Notices, 32(11):17– 20, November 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weisert:1997:CT

[1877] Conrad Weisert. C++ toolbox. ACM SIGPLAN Notices, 32(11):17– ??, November 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1997:VFF

[1878] Paul Frenger. Venturing Forth with a flashlite. ACM SIGPLAN Notices, 32(11):21–24, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1997:FR

[1879] Paul Frenger. Forth report. ACM SIG-PLAN Notices, 32(11):21–??, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1997:CC

[1880] Henry G. Baker. The COMFY 6502 compiler. ACM SIGPLAN Notices, 32 (11):25–30, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kruijer:1997:BNS

[1881] H. S. M. Kruijer. Book nook: The Search for Simplicity Essays in Parallel Programming by Per Brinch Hansen. ACM SIGPLAN Notices, 32(11):31– 32, November 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1997:RSS

[1882] Per Brinch Hansen. Review of the search for simplicity essays in parallel programming. ACM SIGPLAN Notices, 32(11):31–??, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pronk:1997:SEM

[1883] C. Pronk, R. J. Sutcliffe, M. Schönhacker, and A. Wiedemann. Standardized extensions to Modula-2. ACM SIGPLAN Notices, 32(11):34–48, November 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kanada:1997:WPR

[1884] Yasusi Kanada. Web pages that reproduce themselves by JavaScript. ACM SIGPLAN Notices, 32(11):49– 56, November 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jiazhong:1997:DSS

[1885] Zhang Jiazhong, Wang Yanbing, and Zheng Mingchun. Design and static semantics of algorithm language L. ACM SIGPLAN Notices, 32(11):57– 64, November 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heldal:1997:PES

[1886] Rogardt Heldal and John Hughes. Partial evaluation and separate compilation. ACM SIGPLAN Notices, 32(12): 1–11, December 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Asai:1997:PEC

[1887] Kenichi Asai, Hidehiko Masuhara, and Akinori Yonezawa. Partial evaluation of call-by-value λ-calculus with side-effects. ACM SIGPLAN Notices, 32(12):12–21, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sheard:1997:TDL

[1888] Tim Sheard. A type-directed, online, partial evaluator for a polymorphic language. ACM SIGPLAN Notices, 32(12):22–35, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sheard:1997:TOP

[1889] T. Sheard. A type-directed, online, partial evaluator for a polymorphic language. ACM SIGPLAN Notices, 32(12):22–??, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gengler:1997:SAP

[1890] Marc Gengler and Matthieu Martel. Self-applicable partial evaluation for the pi-calculus. ACM SIGPLAN Notices, 32(12):36–46, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jagannathan:1997:SPE

[1891] S. Jagannathan, M. Gengler, and M. Martel. Self-applicable partial evaluation for pi-calculus. ACM SIGPLAN Notices, 32(12):36–??, December 1997.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marinescu:1997:PET

[1892] Mihnea Marinescu and Benjamin Goldberg. Partial-evaluation techniques for concurrent programs. ACM SIGPLAN Notices, 32(12):47–62, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marinescu:1997:PTC

[1893] M. Marinescu and B. Goldberg.

Partial-evaluation techniques for concurrent programs. ACM SIGPLAN Notices, 32(12):47–??, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hornof:1997:ABT

[1894] Luke Hornof and Jacques Noyé. Accurate binding-time analysis for imperative languages: flow, context, and return sensitivity. ACM SIGPLAN Notices, 32(12):63–73, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sahlin:1997:ABA

[1895] D. Sahlin, L. Hornof, and Jacques Noye. Accurate binding-time analysis for imperative languages: Flow, context and return sensitivity. ACM SIG-PLAN Notices, 32(12):63–??, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Melski:1997:ISC

[1896] David Melski and Thomas Reps. Interconvertibility of set constraints and context-free language reachability. ACM SIGPLAN Notices, 32(12): 74–89, December 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danvy:1997:LDT

[1897] Olivier Danvy and Ulrik P. Schultz. Lambda-dropping: transforming recursive equations into programs with block structure. ACM SIGPLAN Notices, 32(12):90–106, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Danvy:1997:LTR

[1898] O. Danvy and U. P. Schultz. Lambdadropping: Transforming recursive equations into programs with block structure. ACM SIGPLAN Notices, 32 (12):90–??, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beshers:1997:GEV

[1899] Clifford Beshers and Steven Feiner. Generating efficient virtual worlds for visualization using partial evaluation and dynamic compilation. ACM SIG-PLAN Notices, 32(12):107–115, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muller:1997:SPE

[1900] Gilles Muller, Eugen-Nicolae Volanschi, and Renaud Marlet. Scaling up partial evaluation for optimizing the Sun commercial RPC protocol. ACM SIGPLAN Notices, 32(12):116– 126, December 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Augustsson:1997:PEA

[1901] Lennart Augustsson. Partial evaluation in aircraft crew planning. ACM SIG-PLAN Notices, 32(12):127–136, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Etalle:1997:TSC

[1902] Sandro Etalle, Maurizio Gabbrielli, and Elena Marchiori. A transformation system for CLP with dynamic scheduling and CCP. ACM SIGPLAN Notices, 32(12):137–150, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palsberg:1997:TSC

[1903] Jens Palsberg, S. Etalle, M. Gabbrielli, and E. Marchiori. A transformation system for CLP with dynamic scheduling and CCP. ACM SIGPLAN Notices, 32(12):137–??, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alpuente:1997:SLF

[1904] M. Alpuente, M. Falaschi, P. Julián, and G. Vidal. Specialization of lazy functional logic programs. ACM SIG-PLAN Notices, 32(12):151–162, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grant:1997:ADR

[1905] Brian Grant, Markus Mock, Matthai Philipose, Craig Chambers, and Susan J. Eggers. Annotation-directed run-time specialization in C. ACM SIGPLAN Notices, 32(12):163–178, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:1997:ARS

[1906] Yanhong Annie Liu, B. Grant, M. Mock, M. Philipose, C. Chambers, and S. Eggers. Annotation-directed run-time specialization in C. ACM SIGPLAN Notices, 32(12):163–??, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Debray:1997:RBP

[1907] Saumya Debray. Resource-bounded partial evaluation. ACM SIGPLAN Notices, 32(12):179–192, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Debray:1997:RPE

[1908] S. Debray. Resource-bounded partial evaluation. ACM SIGPLAN Notices, 32(12):179–??, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nielson:1997:PFM

[1909] Flemming Nielson and Hanne Riis Nielson. Prescriptive frameworks for multi-level lambda-calculi. ACM SIG-PLAN Notices, 32(12):193–202, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mogensen:1997:PFD

[1910] Torben Mogensen, F. Nielson, and H. Riis Nielson. A prescriptive framework for designing multi-level lambdacalculi. ACM SIGPLAN Notices, 32 (12):193–??, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taha:1997:MSPb

[1911] Walid Taha and Tim Sheard. Multistage programming with explicit annotations. ACM SIGPLAN Notices, 32 (12):203–217, December 1997. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taha:1997:MPE

[1912] W. Taha and T. Sheard. Multi-stage programming with explicit annotations. ACM SIGPLAN Notices, 32(12): 203–??, December 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henglein:1998:ECN

[1913] Fritz Henglein. Executive committee news. ACM SIGPLAN Notices, 33 (1):1–2, January 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1998:CCCa

[1914] Seth Bergmann. Conference corner:

Calendar. ACM SIGPLAN Notices, 33 (1):3–13, January 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CPL

[1915] Anonymous. Call for papers: LCR98 — Fourth Workshop on Languages, Compilers, and Run-Time Systems for Scalable Computers. ACM SIG-PLAN Notices, 33(1):14, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeBry:1998:ITC

[1916] Roger K. DeBry. IFIP Technical Committee on Software, TC-2. ACM SIG-PLAN Notices, 33(1):15–21, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gluche:1998:CT

[1917] Dieter Gluche, Dietmar Kühl, and Karsten Weihe. C++ toolbox. ACM SIGPLAN Notices, 33(1):22–29, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1998:YCLa

[1918] Henry G. Baker. You could learn a lot from a quadratic: overloading considered harmful. ACM SIGPLAN Notices, 33(1):30–39, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nedjah:1998:MDL

[1919] Nadia Nedjah. Minimal deterministic left-to-right pattern-matching automata. ACM SIGPLAN Notices, 33 (1):40–47, January 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anquetil:1998:EOO

[1920] Nicolas Anquetil and Jean Vaucher.

Expressing opposition in the object model, first step towards negation. ACM SIGPLAN Notices, 33(1):48–55, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McDowell:1998:UJC

[1921] C. E. McDowell and E. A. Baldwin. Unloading Java classes that contain static fields. ACM SIGPLAN Notices, 33(1):56–60, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xiaocong:1998:SCL

[1922] Fan Xiaocong, Xu Dianxiang, Hou Jianmin, and Zheng Guoliang. SPLAW: A computable language for agentoriented programming. ACM SIG-PLAN Notices, 33(1):61–69, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:GG

[1923] Anonymous. Garbage In/Garbage Out. ACM SIGPLAN Notices, 33 (1):??, January 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:LT

[1924] Anonymous. Language tips. ACM SIGPLAN Notices, 33(1):??, January 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:TC

[1925] Anonymous. Technical Correspondence. ACM SIGPLAN Notices, 33 (1):??, January 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kessler:1998:ECN

[1926] Robert Kessler. Executive committee news. ACM SIGPLAN Notices, 33 (2):1, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1998:CCCb

[1927] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 33 (2):2–13, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CAS

[1928] Anonymous. Conference announcement: Systems Implementation 2000. ACM SIGPLAN Notices, 33(2):14, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:WRP

[1929] Anonymous. Workshop report: Parallelism and Implementation Technology for (Constraint) Logic Programming. ACM SIGPLAN Notices, 33(2): 15–17, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pontelli:1998:PIT

[1930] Enrico Pontelli. Parallelism and implementation technology for (constraint) logic programming. ACM SIGPLAN Notices, 33(2):15–17, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Everitt:1998:CCO

[1931] Cass W. Everitt, John van der Zwaag, and Robert J. Moorhead. COST: common object support toolkit. ACM SIG-PLAN Notices, 33(2):18–24, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Everitt:1998:CTC

[1932] Cass W. Everitt, John van der Zwaag, and Robert J. Moorhead. C++ toolbox: COST: Common Object Support Toolkit. ACM SIGPLAN Notices, 33 (2):18–24, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:1998:AHD

[1933] Philip Wadler. An angry half-dozen.

ACM SIGPLAN Notices, 33(2):25– 30, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:1998:FPA

[1934] Philip Wadler. Functional programming: An angry half-dozen. ACM SIG-PLAN Notices, 33(2):25–30, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:FR

[1935] Paul Frenger. Forth report: Observations on the EuroForth ’97 Conference. ACM SIGPLAN Notices, 33(2):

31–33, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:OEC

[1936] Paul Frenger. Observations on the EuroForth ’97 conference. ACM SIG-PLAN Notices, 33(2):31–33, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1998:YCLb

[1937] Henry G. Baker. You could learn a lot from a quadratic: II: Digital dentistry. ACM SIGPLAN Notices, 33(2): 34–39, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:IAa

[1938] Anonymous. Information for authors. ACM SIGPLAN Notices, 33 (2):40, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bielikova:1998:LPP

[1939] Mária Bieliková and Pavol Návrat. Learning programming in Prolog using schemata. ACM SIGPLAN Notices, 33(2):41–47, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kitchen:1998:FAT

[1940] Andrew Kitchen. A formal approach to the teaching of programming language concepts. ACM SIGPLAN Notices, 33(2):48–55, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dotzel:1998:BO

[1941] Günter Dotzel and Hartmut Goebel.

64 bit Oberon. ACM SIGPLAN Notices, 33(2):56–58, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rangarajan:1998:WTC

[1942] K. Rangarajan and A. Balasubramaniam. When are two classes equivalent? ACM SIGPLAN Notices, 33(2): 59–64, February 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:1998:CSS

[1943] Baowen Xu. Comments on several syntax rules in Ada95. ACM SIGPLAN Notices, 33(2):65–67, February 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1998:ECN

[1944] Peter Lee. Executive committee news:

SIGPLAN supports your professional activities. ACM SIGPLAN Notices, 33(3):1, March 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1998:CC

[1945] Seth Bergmann. Conference corner.

ACM SIGPLAN Notices, 33(3):2–13, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CPS

[1946] Anonymous. Call for papers: SIGADA ’98. ACM SIGPLAN Notices, 33(3):14, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CPI

[1947] Anonymous. Call for papers: ICFP ’98. ACM SIGPLAN Notices, 33(3):15, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salus:1998:HN

[1948] Peter H. Salus. Historical notes: Languages of the Net. ACM SIG-PLAN Notices, 33(3):16–20, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salus:1998:LN

[1949] Peter H. Salus. Languages of the Net. ACM SIGPLAN Notices, 33(3):16–20, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:FRG

[1950] Paul Frenger. Forth report: The growing machine, a pre-Forth language implementation. ACM SIG-PLAN Notices, 33(3):21–23, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:GMP

[1951] Paul Frenger. The growing machine, a pre-Forth language implementation. ACM SIGPLAN Notices, 33(3):21–23, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1998:MMM

[1952] Henry G. Baker. March Möbius madness with a polynomial PostScript, march 32, 1998. ACM SIGPLAN Notices, 33(3):24–35, March 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1998:ARP

[1953] A. Michael Berman. Amendment to review of Per Brinch Hansen, The Search for Simplicity — Essays in Parallel Programming. ACM SIG-PLAN Notices, 33(3):36, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1998:TCIa

[1954] A. Michael Berman. Technical correspondence: Information for authors. ACM SIGPLAN Notices, 33(3):37, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prather:1998:MMP

[1955] Ronald E. Prather. A modular mathematical programming language. ACM SIGPLAN Notices, 33(3):38–56, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1998:EHP

[1956] Per Brinch Hansen. An evaluation of High Performance Fortran. ACM SIGPLAN Notices, 33(3):57–64, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The author strongly criticizes HPF, and remarks “HPF has severe conceptual and practical limitations.”.

Hansen:1998:EMP

[1957] Per Brinch Hansen. An evaluation of the Message-Passing Interface. ACM SIGPLAN Notices, 33(3):65–72, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The author criticizes MPI, and remarks “MPI … lack[s] the elegance and security that can only by checked by a parallel programming language.”.

AzanonEsteire:1998:JST

[1958] Oscar Azañón Esteire and Juan Manual Cueva Lovelle. J — set of tools for native code generation for the Java Virtual Machine. ACM SIG-PLAN Notices, 33(3):73–79, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Esteire:1998:STN

[1959] Oscar Azañón Esteire and Juan Manuel Cueva Lovelle. Set of tools for native code generation for the Java virtual machines. ACM SIG-PLAN Notices, 33(3):73–79, March 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cytron:1998:ECN

[1960] Ron Cytron. Executive Committee news. ACM SIGPLAN Notices, 33(4):1–??, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Anonymous:1998:EA

[1961] Anonymous. ECOOP ’98 announcement. ACM SIGPLAN Notices, 33 (4):14–??, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Marcus:1998:OMW

[1962] Robert Marcus. OOPSLA midyear workshop. ACM SIGPLAN Notices, 33(4):15–??, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Jaeschke:1998:CPL

[1963] Rex Jaeschke. The C programming language. ACM SIGPLAN Notices, 33(4):16, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Appel:1998:SFP

[1964] Andrew W. Appel. SSA is functional programming. ACM SIGPLAN Notices, 33(4):17–20, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Frenger:1998:TT

[1965] Paul Frenger. The talking toaster. ACM SIGPLAN Notices, 33(4):21–25, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.rowan.edu/sigplan/apr199~1.
                                        htm.

Li:1998:CMC

[1966] Gansheng Li. Correction to “Making C++ Concurrent and Real-time”. ACM SIGPLAN Notices, 33(4):27, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.rowan.edu/sigplan/apr199~1.

htm. See [1814].

Creak:1998:GTN

[1967] G. Alan Creak. Garbage: Two new structures. ACM SIGPLAN Notices, 33(4):28–29, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.rowan.edu/
                                        sigplan/apr199~1.htm.

Berman:1998:SPS

[1968] A. Michael Berman. Selected papers from the Second United Kingdom: Java in the Computing Curriculum (JICC) Conference. ACM SIG-PLAN Notices, 33(4):30, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1998:SSS

[1969] A. Michael Berman. Special section: Selected papers from the Second UK Java in the Computing Curriculum Conference. ACM SIGPLAN Notices, 33(4):30–??, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Culwin:1998:SSS

[1970] Fintan Culwin. Special section: selected papers from the Java in the computing curriculum conference. ACM SIGPLAN Notices, 33(4):31–33, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culwin:1998:EJJ

[1971] Fintan Culwin. Editorial — justifying Java? ACM SIGPLAN Notices, 33(4):31–??, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Martin:1998:JGB

[1972] Peter Martin. Java, the good, the bad and the ugly. ACM SIGPLAN Notices, 33(4):34–39, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Gibbons:1998:SPJ

[1973] Jeremy Gibbons. Structured programming in Java. ACM SIGPLAN Notices, 33(4):40–43, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Cornelius:1998:UCJ

[1974] Barry Cornelius. Using CORBA and JDBC to produce three-tier systems. ACM SIGPLAN Notices, 33 (4):44–52, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Burton:1998:KLK

[1975] Peter Burton. Kinds of language, kinds of learning. ACM SIGPLAN Notices, 33(4):53–61, April 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.rowan.edu/
sigplan/apr199~1.htm.

Bodik:1998:CRR

[1976] Rastislav Bodík, Rajiv Gupta, and Mary Lou Soffa. Complete removal of redundant expressions. ACM SIGPLAN Notices, 33(5):1–14, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p1-bodik/

.

Sastry:1998:NAS

[1977] A. V. S. Sastry and Roy D. C. Ju. A new algorithm for scalar register promotion based on SSA form. ACM SIGPLAN Notices, 33 (5):15–25, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

proceedings/pldi/277650/p15-sastry/

.

Chow:1998:RPS

[1978] Fred Chow, Robert Kennedy, Shin-Ming Liu, Raymond Lo, and Peng Tu. Register promotion by sparse partial redundancy elimination of loads and stores. ACM SIGPLAN Notices, 33(5):26–37, May 1998. CO-DEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.cs.virginia.edu/
pldi98/program.html.

Lo:1998:RPS

[1979] Raymond Lo, Fred Chow, Robert Kennedy, Shin-Ming Liu, and Peng Tu. Register promotion by sparse partial redundancy elimination of loads and stores. ACM SIGPLAN Notices, 33 (5):26–37, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p26-lo/.

Rivera:1998:DTE

[1980] Gabriel Rivera and Chau-Wen Tseng.

Data transformations for eliminating conflict misses. ACM SIGPLAN Notices, 33(5):38–49, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p38-rivera/

.

Lewis:1998:IEF

[1981] E. Christopher Lewis, Calvin Lin, and Lawrence Snyder. The implementation and evaluation of fusion and contraction in array languages. ACM SIGPLAN Notices, 33 (5):50–59, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p50-lewis/

.

Paek:1998:SAA

[1982] Yunheung Paek, Jay Hoeflinger, and David Padua. Simplification of array access patterns for compiler optimizations. ACM SIGPLAN Notices, 33 (5):60–71, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p60-paek/

.

Ammons:1998:IDF

[1983] Glenn Ammons and James R. Larus. Improving data-flow analysis with path profiles. ACM SIGPLAN Notices, 33 (5):72–84, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p72-ammons/

.

Aiken:1998:POC

[1984] Alexander Aiken, Manuel Fahndrich, Jeffery S. Foster, and Zhendong Su. Partial online cycle elimination in inclusion constraint graphs. ACM SIG-PLAN Notices, 33(5):85–96, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.cs.virginia.edu/
pldi98/program.html.

Fahndrich:1998:POC

[1985] Manuel Fähndrich, Jeffrey S. Foster, Zhendong Su, and Alexander Aiken. Partial online cycle elimination in inclusion constraint graphs. ACM SIGPLAN Notices, 33(5):85– 96, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p85-fahndrich/

.

Hasti:1998:USS

[1986] Rebecca Hasti and Susan Horwitz. Using static single assignment form to improve flow-insensitive pointer analysis. ACM SIGPLAN Notices, 33(5): 97–105, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p97-hasti/

.

Diwan:1998:TBA

[1987] Amer Diwan, Kathryn S. McKinley, and J. Eliot B. Moss. Type-based alias analysis. ACM SIGPLAN Notices, 33 (5):106–117, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p106-diwan/

.

Sastry:1998:EIF

[1988] S. Subramanya Sastry, Subbarao Palacharla, and James E. Smith. Exploiting idle floating-point resources for integer execution. ACM SIG-PLAN Notices, 33(5):118–129, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p118-sastry/

.

Yang:1998:IPB

[1989] Minghui Yang, Gang-Ryung Uh, and David B. Whalley. Improving performance by branch reordering. ACM SIGPLAN Notices, 33(5): 130–141, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p130-yang/

.

Traub:1998:QSL

[1990] Omri Traub, Glenn Holloway, and Michael D. Smith. Quality and speed in linear-scan register allocation. ACM SIGPLAN Notices, 33(5): 142–151, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p142-traub/

.

Fessant:1998:ICA

[1991] Fabrice Le Fessant, Ian Piumarta, and Marc Shapiro. An implementation of complete, asynchronous, distributed garbage collection. ACM SIG-PLAN Notices, 33(5):152–161, May 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeFessant:1998:ICA

[1992] Fabrice Le Fessant, Ian Piumarta, and Marc Shapiro. An implementation of complete, asynchronous, distributed garbage collection. ACM SIGPLAN Notices, 33(5):152–161, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p152-le_
fessant/.

Cheng:1998:GSC

[1993] Perry Cheng, Robert Harper, and Peter Lee. Generational stack collection and profile-driven pretenuring. ACM SIGPLAN Notices, 33(5): 162–173, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p162-cheng/

.

Clinger:1998:PTR

[1994] William D. Clinger. Proper tail recursion and space efficiency. ACM SIGPLAN Notices, 33(5):174–185, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

                                        proceedings/pldi/277650/p174-clinger/

.

Adve:1998:UIS

[1995] Vikram Adve and John Mellor-Crummey. Using integer sets for dataparallel program analysis and optimization. ACM SIGPLAN Notices, 33(5): 186–198, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p186-adve/

.

Zhu:1998:COP

[1996] Yingchun Zhu and Laurie J. Hendren. Communication optimizations for parallel C programs. ACM SIGPLAN Notices, 33(5):199–211, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p199-zhu/

.

Frigo:1998:ICM

[1997] Matteo Frigo, Charles E. Leiserson, and Keith H. Randall. The implementation of the Cilk-5 multithreaded language. ACM SIGPLAN Notices, 33 (5):212–223, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p212-frigo/

.

Wickline:1998:RTC

[1998] Philip Wickline, Peter Lee, and Frank Pfenning. Run-time code generation and Modal-ML. ACM SIG-PLAN Notices, 33(5):224–235, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p224-wickline/

.

Flatt:1998:UCM

[1999] Matthew Flatt and Matthias Felleisen.

Units: cool modules for HOT languages. ACM SIGPLAN Notices, 33 (5):236–248, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p236-flatt/

.

Xi:1998:EAB

[2000] Hongwei Xi and Frank Pfenning.

Eliminating array bound checking through dependent types. ACM SIGPLAN Notices, 33(5):249–257, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p249-xi/

.

Bacon:1998:TLF

[2001] David F. Bacon, Ravi Konuru, Chet Murthy, and Mauricio Serrano. Thin locks: Featherweight synchronization for Java. ACM SIGPLAN Notices, 33 (5):258–268, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p258-bacon/

.

Agesen:1998:GCL

[2002] Ole Agesen, David Detlefs, and J. Eliot B. Moss. Garbage collection and local variable type-precision and liveness in Java Virtual Machines. ACM SIGPLAN Notices, 33 (5):269–279, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p269-agesen/

.

Adl-Tabatabai:1998:FEC

[2003] Ali-Reza Adl-Tabatabai, Michał Cierniak, Guei-Yuan Lueh, Vishesh M. Parikh, and James M. Stichnoth. Fast, effective code generation in a just-in-time Java compiler. ACM SIGPLAN Notices, 33(5):280–290, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p280-adl-
tabatabai/.

Piumarta:1998:ODT

[2004] Ian Piumarta and Fabio Riccardi. Optimizing direct-threaded code by selective inlining. ACM SIGPLAN Notices, 33(5):291–300, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p291-piumarta/

.

Ayers:1998:SCM

[2005] Andrew Ayers, Stuart de Jong, John Peyton, and Richard Schooler. Scalable cross-module optimization. ACM SIGPLAN Notices, 33(5):301– 312, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p301-ayers/

.

Gay:1998:MME

[2006] David Gay and Alex Aiken. Memory management with explicit regions. ACM SIGPLAN Notices, 33 (5):313–323, May 1998. CODEN SIN-ODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p313-gay/

.

Sweeney:1998:SDD

[2007] Peter F. Sweeney and Frank Tip. A study of dead data members in C++ applications. ACM SIGPLAN Notices, 33(5):324–332, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/277650/p324-sweeney/

.

Necula:1998:DIC

[2008] George C. Necula and Peter Lee. The design and implementation of a certifying compiler. ACM SIGPLAN Notices, 33(5):333–344, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p333-necula/

.

Colby:1998:ACO

[2009] Christopher Colby, Patrice Godefroid, and Lalita Jategaonkar Jagadeesan. Automatically closing open reactive programs. ACM SIGPLAN Notices, 33(5):345–357, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/277650/p345-colby/

.

Berman:1998:SNW

[2010] A. Michael Berman. SIGPLAN Notices — where we are, where we’re going. ACM SIGPLAN Notices, 33(6):1–??, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:C

[2011] Anonymous. Calendar. ACM SIG-PLAN Notices, 33(6):3–??, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:PCPa

[2012] Anonymous. POPL99 call for papers. ACM SIGPLAN Notices, 33 (6):15, June 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryder:1998:M

[2013] Barbara G. Ryder. MASPLAS98.

ACM SIGPLAN Notices, 33(6):16, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Felleisen:1998:DPO

[2014] M. Felleisen, R. B. Findler, M. Flatt, and S. Krishnamurthi. The DrScheme project: an overview. ACM SIG-PLAN Notices, 33(6):17–23, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Felleisen:1998:FPD

[2015] M. Felleisen, R. B. Findler, M. Flatt, and S. Krishnamurthi. Functional programming: The DrScheme Project: An overview. ACM SIGPLAN Notices, 33 (6):17–??, June 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:FSN

[2016] Paul Frenger. Forth in space, or, so NEAR yet so far out. ACM SIG-PLAN Notices, 33(6):24–26, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:FFS

[2017] Paul Frenger. Forth: Forth in space, or, so NEAR yet so far. ACM SIG-PLAN Notices, 33(6):24–??, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1998:YCLc

[2018] Henry G. Baker. You could learn a lot from a quadratic: Newton squares the circle. ACM SIGPLAN Notices, 33 (6):27–31, June 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1998:GIG

[2019] Henry G. Baker. Garbage in/garbage out: You could learn a lot from a quadratic: Newton squares the circle. ACM SIGPLAN Notices, 33(6):27–??, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kokol:1998:FSR

[2020] Peter Kokol and Janez Brest. Fractal structure of random programs. ACM SIGPLAN Notices, 33(6):33–38, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kobol:1998:FSR

[2021] Peter Kobol and Janez Brest. Fractal structure of random programs. ACM SIGPLAN Notices, 33(6):33–??, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aridor:1998:DAL

[2022] Yariv Aridor, Shimon Cohen, and Amiram Yehudai. Design of an Actor language for implicit parallel programming. ACM SIGPLAN Notices, 33 (6):39–47, June 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rangarajan:1998:UCE

[2023] K. Rangarajan and P. Eswar. Understanding class evolution through atomic changes. ACM SIGPLAN Notices, 33(6):48–60, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanBatenburg:1998:AEC

[2024] F. H. D. van Batenburg, Ellen de Ridder, and Joseph de Kerf. APL Extended Compared With Other Languages According to Halstead’s Theory. ACM SIGPLAN Notices, 33(6):54, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ancona:1998:SHS

[2025] Massimo Ancona and Walter Cazzola.

Static header as sentinel. ACM SIG-PLAN Notices, 33(6):61–64, June 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snelting:1998:CAN

[2026] Gregor Snelting. Concept analysis — A new framework for program understanding (invited paper). ACM SIG-PLAN Notices, 33(7):1–10, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:1998:ECA

[2027] Sean Zhang, Barbara G. Ryder, and William A. Landi. Experiments with combined analysis for pointer aliasing.

ACM SIGPLAN Notices, 33(7):11–18, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tice:1998:ONA

[2028] Caroline Tice and Susan L. Graham. OPTVIEW: A new approach for examining optimized code. ACM SIG-PLAN Notices, 33(7):19–26, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dor:1998:DME

[2029] Nurit Dor, Michael Rodeh, and Mooly Sagiv. Detecting memory errors via static pointer analysis (preliminary experience). ACM SIGPLAN Notices, 33 (7):27–34, July 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krinke:1998:SST

[2030] Jens Krinke. Static slicing of threaded programs. ACM SIGPLAN Notices, 33 (7):35–42, July 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chase:1998:RSA

[2031] Melissa P. Chase, Steven M. Christey, David R. Harris, and Alexander S. Yeh. Recovering software architecture from multiple source code analyses. ACM SIGPLAN Notices, 33(7):43–50, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naumovich:1998:ECD

[2032] Gleb Naumovich, Lori A. Clarke, and Leon J. Osterweil. Efficient composite data flow analysis applied to concurrent programs. ACM SIGPLAN Notices, 33 (7):51–58, July 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reiss:1998:SVD

[2033] Steven P. Reiss. Software visualization in the desert environment. ACM SIGPLAN Notices, 33(7):59–66, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jeffery:1998:LAP

[2034] Clinton Jeffery, Wenyi Zhou, Kevin Templer, and Michael Brazell. A lightweight architecture for program execution monitoring. ACM SIG-PLAN Notices, 33(7):67–74, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Waddell:1998:VPH

[2035] Oscar Waddell and J. Michael Ashley. Visualizing the performance of higher-order programs. ACM SIG-PLAN Notices, 33(7):75–82, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrold:1998:EIP

[2036] Mary Jean Harrold, Gregg Rothermel, Rui Wu, and Liu Yi. An empirical investigation of program spectra. ACM SIGPLAN Notices, 33(7):83–90, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zorn:1998:IO

[2037] Benjamin Zorn. An invitation to OOP-SLA ’98. ACM SIGPLAN Notices, 33 (8):1–2, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1998:CCCc

[2038] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 33 (8):3–12, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:PPC

[2039] Anonymous. PLDI ’99 preliminary call for papers. ACM SIGPLAN Notices, 33(8):13, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1998:FCP

[2040] Seth Bergmann. FOOL 6 call for papers. ACM SIGPLAN Notices, 33(8): 14–15, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pontelli:1998:PSI

[2041] Enrico Pontelli. Programming with {SETS} — an international forum of discussion. ACM SIGPLAN Notices, 33(8):16, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McNamee:1998:DTM

[2042] Paul McNamee and Marty Hall. Developing a tool for memoizing functions in C++. ACM SIGPLAN Notices, 33 (8):17–22, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:1998:WNO

[2043] Philip Wadler. Why no one uses functional languages. ACM SIGPLAN Notices, 33(8):23–27, August 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:TFF

[2044] Paul Frenger. A tribute to FIG-Forth.

ACM SIGPLAN Notices, 33(8):28–31, August 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1998:LE

[2045] Per Brinch Hansen. Letter to the editor. ACM SIGPLAN Notices, 33(8): 32–33, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Offner:1998:PBH

[2046] Carl D. Offner. Per Brinch Hansen’s concerns about High Performance Fortran. ACM SIGPLAN Notices, 33(8): 34–39, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schilling:1998:OCE

[2047] Jonathan L. Schilling. Optimizing away C++ exception handling. ACM SIG-PLAN Notices, 33(8):40–47, August 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledru:1998:JIL

[2048] Pascal Ledru. JSpace: implementation of a Linda system in Java. ACM SIGPLAN Notices, 33(8):48–50, August 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledru:1998:TSJ

[2049] Pascal Ledru. A traveling salesman in Java. ACM SIGPLAN Notices, 33 (8):51–56, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:1998:SA

[2050] Mary Lou Soffa. 1998 SIGPLAN awards. ACM SIGPLAN Notices, 33 (9):1–3, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1998:CCCd

[2051] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 33(9):4–15, September 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:PCPb

[2052] Anonymous. PLDI ’99 call for papers. ACM SIGPLAN Notices, 33(9): 16, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:IAb

[2053] Anonymous. ISMM ’98 announcement. ACM SIGPLAN Notices, 33(9): 17, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scherlis:1998:EST

[2054] William L. Scherlis and Neil D. Jones.

The European scene: Two clusters of programming languages conferences. ACM SIGPLAN Notices, 33(9):18– 19, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scherlis:1998:TCP

[2055] William L. Scherlis. Two clusters of programming language conferences. ACM SIGPLAN Notices, 33(9):18– 19, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:FMA

[2056] Paul Frenger. Forth: MUCKing around with Forth. ACM SIGPLAN Notices, 33(9):20–24, September 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:MAF

[2057] Paul Frenger. MUCKing around with Forth. ACM SIGPLAN Notices, 33(9): 20–24, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1998:TCIb

[2058] A. Michael Berman. Technical correspondence: Information for authors. ACM SIGPLAN Notices, 33(9): 25, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:1998:RRA

[2059] N. I. Adams IV, D. H. Bartley, G. Brooks, R. K. Dybvig, D. P. Friedman, R. Halstead, C. Hanson, C. T. Haynes, E. Kohlbecker, D. Oxley, K. M. Pitman, G. J. Rozas, G. L. Steele, Jr., G. J. Sussman, M. Wand, and H. Abelson. Revised5 report on the algorithmic language Scheme. ACM SIGPLAN Notices, 33(9):26– 76, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kelsey:1998:RRA

[2060] Richard Kelsey, William Clinger, and Jonathan Rees. Revised^{5} report on the algorithmic language Scheme. ACM SIGPLAN Notices, 33(9):26– 76, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). With H. Abelson, N. I. Adams, IV, D. H. Bartley, G. Brooks, R. K. Dybvig, D. P. Friedman, R. Halstead, C. Hanson, C. T. Haynes, E. Kohlbecker, D. Oxley, K. M. Pitman, G. J. Rozas, G. L. Steele, Jr., G. J. Sussman, and M. Wand.

Bantchev:1998:PMM

[2061] Boyko B. Bantchev. Putting more meaning in expressions. ACM SIG-PLAN Notices, 33(9):77–83, September 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McDowell:1998:RGJ

[2062] C. E. McDowell. Reducing garbage in Java. ACM SIGPLAN Notices, 33(9): 84–86, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnstone:1998:RIB

[2063] Adrian Johnstone and Elizabeth Scott.

rdp — an iterator-based recursive descent parser generator with tree promotion operators. ACM SIGPLAN Notices, 33(9):87–94, September 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:1998:EBC

[2064] Fuqing Yang, Hong Mei, Wanghong Yuan, Qiong Wu, and Yao Guo. Experiences in building C++ front end. ACM SIGPLAN Notices, 33(9):95– 102, September 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vogt:1998:FPP

[2065] Christopher J. Vogt. Floating point performance of Common Lisp. ACM SIGPLAN Notices, 33(9):103–107, September 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Atkinson-Abutridy:1998:DIG

[2066] John A. Atkinson-Abutridy and Anita A. Ferreira-Cabrera. The design and implementation of the GILENA natural language interfaces specification system. ACM SIGPLAN Notices, 33(9):108–117, September 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dolby:1998:EAO

[2067] Julian Dolby and Andrew A. Chien. An evaluation of automatic object inline allocation techniques. ACM SIG-PLAN Notices, 33(10):1–20, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Czajkowski:1998:JRA

[2068] Grzegorz Czajkowski and Thorsten von Eicken. JRes: A resource accounting interface for Java. ACM SIGPLAN Notices, 33(10):21–35, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liang:1998:DCL

[2069] Sheng Liang and Gilad Bracha. Dynamic class loading in the Java Virtual Machine. ACM SIGPLAN Notices, 33 (10):36–44, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Artim:1998:OOP

[2070] John Artim, Charlie Bridgeford, Lillian Christman, James Coplien, Mary Beth Rosson, Stanley Taylor, and Rebecca Wirfs-Brock. Object-oriented practice in 1998 (panel): does it help or hinder collaboration? ACM SIGPLAN Notices, 33(10):45–47, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Artim:1998:POP

[2071] John Artim, Charlie Bridgeford, Lilian Christman, James Coplien, Mary Beth Rosson, Stanley Taylor, and Rebecca Wirfs-Brock. Panel 1: Object-oriented practice in 1998: Does it help or hinder collaboration? ACM SIGPLAN Notices, 33(10):45–47, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:1998:OTF

[2072] David G. Clarke, John M. Potter, and James Noble. Ownership types for flexible alias protection. ACM SIGPLAN Notices, 33(10):48–64, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simons:1998:BCS

[2073] Anthony J. H. Simons. Borrow, copy or steal?: loans and larceny in the orthodox canonical form. ACM SIG-PLAN Notices, 33(10):65–83, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bouraqadi-Saadani:1998:SMP

[2074] Noury M. N. Bouraqadi-Saâdani, Thomas Ledoux, and Fred Rivard. Safe metaclass programming. ACM SIG-PLAN Notices, 33(10):84–96, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mezini:1998:APC

[2075] Mira Mezini and Karl Lieberherr.

Adaptive plug-and-play components for evolutionary software development. ACM SIGPLAN Notices, 33(10):97– 116, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mezini:1998:APP

[2076] Mira Mezini and Karl Lieberherr. Adaptive plug-and-play components for evolutionary software development. ACM SIGPLAN Notices, 33(10):97– 116, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Riehle:1998:RMB

[2077] Dirk Riehle and Thomas Gross. Role model based framework design and integration. ACM SIGPLAN Notices, 33(10):117–133, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Agerbo:1998:HPB

[2078] Ellen Agerbo and Aino Cornils. How to preserve the benefits of design patterns. ACM SIGPLAN Notices, 33(10): 134–143, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leino:1998:DGS

[2079] K. Rustan M. Leino. Data groups: Specifying the modification of extended state. ACM SIGPLAN Notices, 33(10): 144–153, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rossie:1998:LOE

[2080] Jonathan G. Rossie, Jr. Logical observable entities. ACM SIGPLAN Notices, 33(10):154–165, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gianantonio:1998:LCO

[2081] Pietro Di Gianantonio, Furio Honsell, and Luigi Liquori. A lambda calculus of objects with self-inflicted extension. ACM SIGPLAN Notices, 33(10): 166–178, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bak:1998:NCJ

[2082] Lars Bak, John Duimovich, Jesse Fang, Scott Meyer, and David Ungar. The new crop of Java virtual machines (panel). ACM SIGPLAN Notices, 33(10):179–182, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ungar:1998:PNC

[2083] David Ungar, Lars Bak, Jesse Fang, John Duimovich, and Scott Meyer. Panel 2: The new crop of Java Virtual Machines. ACM SIGPLAN Notices, 33(10):179–182, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bracha:1998:MFS

[2084] Gilad Bracha, Martin Odersky, David Stoutamire, and Philip Wadler. Making the future safe for the past: Adding genericity to the Java programming language. ACM SIGPLAN Notices, 33(10):183–200, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cartwright:1998:CGR

[2085] Robert Cartwright and Guy L. Steele, Jr. Compatible genericity with runtime types for the Java programming language. ACM SIGPLAN Notices, 33(10):201–215, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Solorzano:1998:PPJ

[2086] Jose H. Solorzano and Suad Alagić. Parametric polymorphism for Java: A reflective solution. ACM SIGPLAN Notices, 33(10):216–225, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cockburn:1998:QTA

[2087] Alistair Cockburn and Martin Fowler. Question time! about use cases. ACM SIGPLAN Notices, 33(10):226– 229, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fowler:1998:PQT

[2088] Martin Fowler, Ivar Jacobson, Bruce Anderson, Alistair Cockburn, and Ian Graham. Panel 3: Question time! about use cases. ACM SIGPLAN Notices, 33(10):226–229, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McAuliffe:1998:VFF

[2089] Mark L. McAuliffe, Michael J. Carey, and Marvin H. Solomon. Vclusters: A flexible, fine-grained object clustering mechanism. ACM SIGPLAN Notices, 33(10):230–243, October 1998.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guerraoui:1998:SSO

[2090] Rachid Guerraoui, Pascal Felber, Benoît Garbinato, and Karim Mazouni. System support for object groups. ACM SIGPLAN Notices, 33(10):244– 258, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bertino:1998:EOO

[2091] Elisa Bertino and Giovanna Guerrini.

Extending the ODMG object model with composite objects. ACM SIG-PLAN Notices, 33(10):259–270, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walker:1998:VDS

[2092] Robert J. Walker, Gail C. Murphy, Bjorn Freeman-Benson, Darin Wright, Darin Swanson, and Jeremy Isaak. Visualizing dynamic software system information through high-level models. ACM SIGPLAN Notices, 33(10):271– 283, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Olken:1998:OLL

[2093] Frank Olken, Hans-Arno Jacobsen, Chuck McParland, Mary Ann Piette, and Mary F. Anderson. Object lessons learned from a distributed system for remote building monitoring and operation. ACM SIGPLAN Notices, 33(10): 284–295, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harris:1998:LOO

[2094] Jacob Harris and Vivek Sarkar. Lightweight object-oriented shared variables for distributed applications on the Internet. ACM SIGPLAN Notices, 33(10):296–309, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harris:1998:LOS

[2095] Jacob Harris and Vivek Sarkar. Lightweight object-oriented shared variables for distributed applications on the Internet. ACM SIGPLAN Notices, 33(10):296–309, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freund:1998:TSO

[2096] Stephen N. Freund and John C. Mitchell. A type system for object initialization in the Java bytecode language. ACM SIGPLAN Notices, 33(10):310–328, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jacobs:1998:RAJ

[2097] Bart Jacobs, Joachim van den Berg, Marieke Huisman, Martijn van Berkum, Ulrich Hensel, and Hendrik Tews. Reasoning about Java classes: preliminary report. ACM SIGPLAN Notices, 33(10):329–340, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Drossopoulou:1998:WJB

[2098] Sophia Drossopoulou, David Wragg, and Susan Eisenbach. What is Java binary compatibility? ACM SIGPLAN Notices, 33(10):341–358, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cleal:1998:PPM

[2099] David Cleal, Luke Hohmann, Chet Hendrickson, Allen B. Davis, and John Nolan. Panel 4: The project management game. ACM SIGPLAN Notices, 33(10):359–361, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buchi:1998:CTJ

[2100] Martin Büchi and Wolfgang Weck.

Compound types for Java. ACM SIG-PLAN Notices, 33(10):362–373, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavens:1998:MDD

[2101] Gary T. Leavens and Todd D. Millstein. Multiple dispatch as dispatch on Tuples. ACM SIGPLAN Notices, 33(10):374–387, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Litvinov:1998:CBP

[2102] Vassily Litvinov. Contraint-based polymorphism in Cecil: towards a practical and static type system. ACM SIG-PLAN Notices, 33(10):388–411, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Litvinov:1998:CPC

[2103] Vassily Litvinov. Constraint-based polymorphism in Cecil: Towards a practical and static type system. ACM SIGPLAN Notices, 33(10):388–411, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lea:1998:BDS

[2104] Doug Lea, David Forslund, Tom Barry, Don Vines, Rajendra Raj, and Ashutosh Tiwary. Building distributed systems (panel). ACM SIGPLAN Notices, 33(10):412–416, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lea:1998:PBD

[2105] Doug Lea, Tom Barry, Don Vines, David W. Forslund, Rajendra Raj, and Ashutosh Tiwary. Panel 5: Building distributed systems. ACM SIGPLAN Notices, 33(10):412–416, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1998:PSR

[2106] Steven Fraser, Maggie Davis, Martin Griss, Luke Hohmann, Ian Hopper, Rebecca Joos, and Bill Opdyke. Panel 6: Software reuse: Nemesis or Nirvana? ACM SIGPLAN Notices, 33(10):417– 420, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1998:SRP

[2107] Steven Fraser, Maggie Davis, Martin Griss, Luke Hohmann, Ian Hopper, Rebecca Joos, and Bill Opdyke. Software reuse (panel): nemesis or nirvana? ACM SIGPLAN Notices, 33(10):417– 420, October 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:AI

[2108] Anonymous. Author index. ACM SIGPLAN Notices, 33(10):421, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:TI

[2109] Anonymous. Title index. ACM SIG-PLAN Notices, 33(10):422, October 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:1998:CCM

[2110] Keith D. Cooper and Timothy J.

Harvey. Compiler-controlled memory. ACM SIGPLAN Notices, 33 (11):2–11, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p2-cooper/

.

Seidl:1998:SHO

[2111] Matthew L. Seidl and Benjamin G.

Zorn. Segregating heap objects by reference behavior and lifetime. ACM SIGPLAN Notices, 33 (11):12–23, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p12-seidl/

.

Strout:1998:SIS

[2112] Michelle Mills Strout, Larry Carter, Jeanne Ferrante, and Beth Simon. Schedule-independent storage mapping for loops. ACM SIGPLAN Notices, 33(11):24–33, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p24-strout/

.

Sodani:1998:EAI

[2113] Avinash Sodani and Gurindar S. Sohi. An empirical analysis of instruction repetition. ACM SIG-PLAN Notices, 33(11):35–45, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p35-sodani/

.

Lee:1998:STS

[2114] Walter Lee, Rajeev Barua, Matthew Frank, Devabhaktuni Srikrishna, Jonathan Babb, Vivek Sarkar, and Saman Amarasinghe. Space-time scheduling of instruction-level parallelism on a raw machine. ACM SIG-PLAN Notices, 33(11):46–57, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/

proceedings/asplos/291069/p46-lee/

.

Hammond:1998:DSS

[2115] Lance Hammond, Mark Willey, and Kunle Olukotun. Data speculation support for a chip multiprocessor. ACM SIGPLAN Notices, 33 (11):58–69, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p58-hammond/

.

VanMeter:1998:VNV

[2116] Rodney Van Meter, Gregory G.

Finn, and Steve Hotz. VISA: Netstation’s Virtual Internet SCSI Adapter. ACM SIGPLAN Notices, 33(11):71–80, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p71-van_
meter/.

Acharya:1998:ADP

[2117] Anurag Acharya, Mustafa Uysal, and Joel Saltz. Active disks: Programming model, algorithms and evaluation. ACM SIGPLAN Notices, 33 (11):81–91, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p81-acharya/

.

Gibson:1998:CEH

[2118] Garth A. Gibson, David F. Nagle, Khalil Amiri, Jeff Butler, Fay W. Chang, Howard Gobioff, Charles Hardin, Erik Riedel, David Rochberg, and Jim Zelenka. A cost-effective, high-bandwidth storage architecture. ACM SIGPLAN Notices, 33(11):92–103, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p92-gibson/

.

Machanick:1998:HST

[2119] Philip Machanick, Pierre Salverda, and Lance Pompe. Hardware-software trade-offs in a direct Rambus implementation of the RAMpage memory hierarchy. ACM SIGPLAN Notices, 33(11):105–114, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p105-
machanick/.

Roth:1998:DBP

[2120] Amir Roth, Andreas Moshovos, and Gurindar S. Sohi. Dependence based prefetching for linked data structures. ACM SIGPLAN Notices, 33 (11):115–126, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p115-
roth/.

Weissman:1998:PCS

[2121] Boris Weissman. Performance counters and state sharing annotations: a unified approach to thread locality. ACM SIGPLAN Notices, 33 (11):127–138, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p127-
weissman/.

Calder:1998:CCD

[2122] Brad Calder, Chandra Krintz, Simmi John, and Todd Austin. Cacheconscious data placement. ACM SIGPLAN Notices, 33(11):139–149, November 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:80/pubs/
citations/proceedings/asplos/291069/
p139-calder/.

Le:1998:OET

[2123] Bich C. Le. An out-of-order execution technique for runtime binary translators. ACM SIGPLAN Notices, 33(11):151–158, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p151-
le/.

Krintz:1998:OET

[2124] Chandra Krintz, Brad Calder, Han Bok Lee, and Benjamin G. Zorn. Overlapping execution with transfer using non-strict execution for mobile programs. ACM SIGPLAN Notices, 33(11):159–169, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p159-
krintz/.

Stark:1998:VLP

[2125] Jared Stark, Marius Evers, and Yale N. Patt. Variable length path branch prediction. ACM SIGPLAN Notices, 33(11):170–179, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p170-
stark/.

Verghese:1998:PIS

[2126] Ben Verghese, Anoop Gupta, and Mendel Rosenblum. Performance isolation: Sharing and isolation in shared-memory multiprocessors. ACM SIGPLAN Notices, 33(11): 181–192, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p181-
verghese/.

Chen:1998:UMA

[2127] Yuqun Chen, Angelos Bilas, Stefanos N. Damianakis, Cezary Dubnicki, and Kai Li. UTLB: A mechanism for address translation on network interfaces. ACM SIGPLAN Notices, 33(11):193–204, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p193-
chen/.

Pai:1998:LAR

[2128] Vivek S. Pai, Mohit Aron, Gaurav Banga, Michael Svendsen, Peter Druschel, Willy Zwaenepoel, and Erich Nahum. Locality-aware request distribution in cluster-based network servers. ACM SIGPLAN Notices, 33(11):205–216, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p205-
pai/.

Temam:1998:IOL

[2129] Olivier Temam. Investigating optimal local memory performance. ACM SIGPLAN Notices, 33(11): 218–227, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p218-
temam/.

Ghosh:1998:PMA

[2130] Somnath Ghosh, Margaret Martonosi, and Sharad Malik. Precise miss analysis for program transformations with caches of arbitrary associativity. ACM SIGPLAN Notices, 33 (11):228–239, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/

                                        proceedings/asplos/291069/p228-
                                        ghosh/.

Peir:1998:CDM

[2131] Jih-Kwon Peir, Yongjoon Lee, and Windsor W. Hsu. Capturing dynamic memory reference behavior with adaptive cache topology. ACM SIGPLAN Notices, 33(11):240–250, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/asplos/291069/p240-
                                        peir/.

Citron:1998:AMM

[2132] Daniel Citron, Dror Feitelson, and Larry Rudolph. Accelerating multimedia processing by implementing memoing in multiplication and division units. ACM SIGPLAN Notices, 33(11):252–261, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/291069/p252-
                                        citron/.

Fu:1998:VSS

[2133] Chao-Ying Fu, Matthew D. Jennings, Sergei Y. Larin, and Thomas M. Conte. Value speculation scheduling for high performance processors. ACM SIGPLAN Notices, 33 (11):262–271, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/291069/p262-
                                        fu/.

Ranganathan:1998:ESD

[2134] Narayan Ranganathan and Manoj Franklin. An empirical study of decentralized ILP execution models. ACM SIGPLAN Notices, 33 (11):272–281, November 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p272-
ranganathan/.

Schnarr:1998:FOP

[2135] Eric Schnarr and James R. Larus. Fast out-of-order processor simulation using memoization. ACM SIGPLAN Notices, 33(11):283–294, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p283-
schnarr/.

Jacob:1998:LSM

[2136] Bruce L. Jacob and Trevor N. Mudge.

A look at several memory management units, TLB-refill mechanisms, and page table organizations. ACM SIGPLAN Notices, 33(11):295–306, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/asplos/291069/p295-
jacob/.

Ranganathan:1998:PDW

[2137] Parthasarathy Ranganathan, Kourosh Gharachorloo, Sarita V. Adve, and Luiz André Barroso. Performance of database workloads on sharedmemory systems with out-of-order processors. ACM SIGPLAN Notices, 33(11):307–318, November 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/asplos/291069/p307-
                                        ranganathan/.

Anonymous:1998:AVR

[2138] Anonymous. ASPLOS-VIII reviewers. ACM SIGPLAN Notices, 33(11): v–vi, November 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CC

[2139] Anonymous. Conference committee. ACM SIGPLAN Notices, 33(11): iv, November 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soffa:1998:SSJ

[2140] Mary Lou Soffa and David Notkin. SIGPLAN and SIGSOFT joint efforts. ACM SIGPLAN Notices, 33(12):1, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1998:CCCe

[2141] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 33 (12):2–13, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CPAa

[2142] Anonymous. Call for papers: ACM SIGPLAN-SIGSOFT Workshop on Program Analysis for Software Tools and Engineering (PASTE ’99). ACM SIGPLAN Notices, 33(12):14–15, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CPAb

[2143] Anonymous. Call for papers: 1999 ACM International Conference on Supercomputing. ACM SIGPLAN Notices, 33(12):15, December 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1998:EYO

[2144] Brent W. Benson, Jr. Eat your own dog food. ACM SIGPLAN Notices, 33(12): 16–18, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1998:JRR

[2145] Brent W. Benson, Jr. Java reflections: Reflection: Eat your own dog food. ACM SIGPLAN Notices, 33(12): 16–19, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1998:OTD

[2146] Chris Clark. Overlapping token definitions. ACM SIGPLAN Notices, 33(12): 20–24, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1998:PPP

[2147] Chris Clark. Practical parsing patterns: Overlapping token definitions. ACM SIGPLAN Notices, 33(12):20– 24, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:FRM

[2148] Paul Frenger. Forth report: Mind.Forth: Thoughts on artificial intelligence and Forth. ACM SIGPLAN Notices, 33 (12):25–31, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1998:MFT

[2149] Paul Frenger. Mind.Forth: thoughts on artificial intelligence and Forth. ACM SIGPLAN Notices, 33(12):25–31, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1998:TCIc

[2150] A. Michale Berman. Technical correspondence: Information for authors. ACM SIGPLAN Notices, 33(12):32, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakharov:1998:TCP

[2151] Alex Sakharov. Technical correspondence: Processing recursive data types in Java. ACM SIGPLAN Notices, 33 (12):33, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Svensson:1998:IBG

[2152] Kurt Svensson. Index in BNF grammars. ACM SIGPLAN Notices, 33 (12):34, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paleri:1998:SAP

[2153] Vineeth Kumar Paleri, Y. N. Srikant, and Priti Shankar. A simple algorithm for partial redundancy elimination. ACM SIGPLAN Notices, 33(12): 35–43, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Curcio:1998:ASA

[2154] Igor D. D. Curcio. ASAP — A simple assertion pre-processor. ACM SIG-PLAN Notices, 33(12):44–51, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See correction [2448].

Hatzigeorgiu:1998:LPS

[2155] Nick Hatzigeorgiu and Apostolos Syropoulos. Literate programming and the “Spaniel” method. ACM SIG-PLAN Notices, 33(12):52–56, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comments [2223, 2224].

Hericko:1998:JDO

[2156] Marjan Hericko, Matjaz B. Juric, Ales Zivkovic, Ivan Rozman, Tomaz Domajnko, and Marjan Krisper. Java and distributed object models: An analysis. ACM SIGPLAN Notices, 33(12):

57–65, December 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Budd:1998:FPF

[2157] Timothy A. Budd. Functional programming and the fragile base class problem. ACM SIGPLAN Notices, 33 (12):66–71, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malik:1998:EHL

[2158] Masud Ahmad Malik. Evolution of the high level programming languages: A critical perspective. ACM SIGPLAN Notices, 33(12):72–80, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). This paper was plagiarized from a published book; see editorial apology in [2207].

Havener:1998:EQL

[2159] Charles D. Havener. EQL — the query language you never heard of. ACM SIGPLAN Notices, 33(12):81–88, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Minamide:1999:RCT

[2160] Yasuhiko Minamide and Jacques Garrigue. On the runtime complexity of type-directed unboxing. ACM SIG-PLAN Notices, 34(1):1–12, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Visser:1999:BPO

[2161] Eelco Visser, Zine el Abidine Benaissa, and Andrew Tolmach. Building program optimizers with rewriting strategies. ACM SIGPLAN Notices, 34(1): 13–26, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hannan:1999:HOA

[2162] John Hannan and Patrick Hicks.

Higher-order arity raising. ACM SIG-PLAN Notices, 34(1):27–38, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gustavsson:1999:TBS

[2163] Jörgen Gustavsson. A type based sharing analysis for update avoidance and optimisation. ACM SIGPLAN Notices, 34(1):39–50, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kieburtz:1999:TEM

[2164] Richard B. Kieburtz. Taming effects with monadic typing. ACM SIGPLAN Notices, 34(1):51–62, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:1999:MEM

[2165] Philip Wadler. The marriage of effects and monads. ACM SIGPLAN Notices, 34(1):63–74, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1999:TCF

[2166] Colin J. Taylor. A theory of core fudgets. ACM SIGPLAN Notices, 34(1): 75–85, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hicks:1999:PPL

[2167] Michael Hicks, Pankaj Kakkar, Jonathan T. Moore, Carl A. Gunter, and Scott Nettles. PLAN: a packet language for active networks. ACM SIGPLAN Notices, 34(1):86–93, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Findler:1999:MOO

[2168] Robert Bruce Findler and Matthew Flatt. Modular object-oriented programming with units and mixins. ACM SIGPLAN Notices, 34(1):94–104, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odersky:1999:PVF

[2169] Martin Odersky. Programming with variable functions. ACM SIGPLAN Notices, 34(1):105–116, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mackie:1999:YYA

[2170] Ian Mackie. YALE: Yet Another Lambda Evaluator based on interaction nets. ACM SIGPLAN Notices, 34(1): 117–128, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benton:1999:CSM

[2171] Nick Benton, Andrew Kennedy, and George Russell. Compiling Standard ML to Java bytecodes. ACM SIG-PLAN Notices, 34(1):129–140, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shao:1999:TCM

[2172] Zhong Shao. Typed cross-module compilation. ACM SIGPLAN Notices, 34(1):141–152, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Finne:1999:HDB

[2173] Sigbjorn Finne, Daan Leijen, Erik Meijer, and Simon Peyton Jones. H/Direct: A binary foreign language interface for Haskell. ACM SIGPLAN Notices, 34(1):153–162, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mountjoy:1999:STG

[2174] Jon Mountjoy. The spineless tagless Gmachine, naturally. ACM SIGPLAN Notices, 34(1):163–173, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bjesse:1999:LHD

[2175] Per Bjesse, Koen Claessen, Mary Sheeran, and Satnam Singh. Lava: Hardware design in Haskell. ACM SIG-PLAN Notices, 34(1):174–184, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edalat:1999:LCE

[2176] Abbas Edalat, Peter John Potts, and Philipp Sünderhauf. Lazy computation with exact real numbers. ACM SIG-PLAN Notices, 34(1):185–194, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karczmarczuk:1999:FDC

[2177] Jerzy Karczmarczuk. Functional differentiation of computer programs. ACM SIGPLAN Notices, 34(1):195–203, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moreau:1999:DGC

[2178] Luc Moreau. A distributed garbage collector with diffusion tree reorganisation and mobile objects. ACM SIG-PLAN Notices, 34(1):204–215, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nordlander:1999:PSP

[2179] Johan Nordlander. Pragmatic subtyping in polymorphic languages. ACM SIGPLAN Notices, 34(1):216–227, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pottier:1999:FTI

[2180] François Pottier. A framework for type inference with subtyping. ACM SIG-PLAN Notices, 34(1):228–238, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Augustsson:1999:CLD

[2181] Lennart Augustsson. Cayenne — a language with dependent types. ACM SIGPLAN Notices, 34(1):239–250, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sobel:1999:RC

[2182] Jonathan Sobel and Daniel P. Friedman. Recycling continuations. ACM SIGPLAN Notices, 34(1):251–260, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scholz:1999:ISM

[2183] Enno Scholz. Imperative streams — A monadic combinator library for synchronous programming. ACM SIG-PLAN Notices, 34(1):261–272, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gibbons:1999:UAU

[2184] Jeremy Gibbons and Geraint Jones.

The under-appreciated unfold. ACM SIGPLAN Notices, 34(1):273–279, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hutton:1999:FUP

[2185] Graham Hutton. Fold and unfold for program semantics. ACM SIGPLAN Notices, 34(1):280–288, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:1999:ETM

[2186] Zhe Yang. Encoding types in MLlike languages. ACM SIGPLAN Notices, 34(1):289–300, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crary:1999:IPT

[2187] Karl Crary, Stephanie Weirich, and Greg Morrisett. Intensional polymorphism in type-erasure semantics. ACM SIGPLAN Notices, 34(1):301–312, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shao:1999:ITI

[2188] Zhong Shao, Christopher League, and Stefan Monnier. Implementing typed intermediate languages. ACM SIG-PLAN Notices, 34(1):313–323, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kutzner:1999:NDC

[2189] Arne Kutzner and Manfred Schmidt-Schauß. A non-deterministic call-by-need lambda calculus. ACM SIGPLAN Notices, 34(1):324–335, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:1999:LLB

[2190] Oleg Kiselyov. LAND*: an AND with local bindings, a guarded LET* special form. ACM SIGPLAN Notices, 34 (1):336, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:1999:FSC

[2191] Oleg Kiselyov. Functional style in C++: closures, late binding, and Lambda abstractions. ACM SIG-PLAN Notices, 34(1):337, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McAdam:1999:TEC

[2192] Bruce J. McAdam. Type errors confuse the programmer (poster abstract). ACM SIGPLAN Notices, 34(1):338, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ennals:1999:CTN

[2193] Robert Ennals. Controlled temporal non-determinism for reasoning with a machine of finite speed. ACM SIGPLAN Notices, 34(1):339, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ennals:1999:VT

[2194] Robert Ennals. Verbose typing. ACM SIGPLAN Notices, 34(1):340, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Putnik:1999:ASS

[2195] Zoran Putnik. On application of selfsimilar pictures in education. ACM SIGPLAN Notices, 34(1):341, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Russell:1999:MFA

[2196] Dan Russell. MAP: A functional analysis and design method. ACM SIGPLAN Notices, 34(1):342, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:1999:DLR

[2197] Oleg Kiselyov. A delegation language to request weather products and a scheme of its interpretation. ACM SIGPLAN Notices, 34(1):343, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Am:1999:SAR

[2198] Nancy Am and Young Park. A structured approach to retrieving functions by types. ACM SIGPLAN Notices, 34 (1):344, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

An:1999:SAR

[2199] Nancy An and Young Park. A structured approach to retrieving functions by types. ACM SIGPLAN Notices, 34 (1):344, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ling:1999:FFR

[2200] Gary Shu Ling. Frob — functional robotics. ACM SIGPLAN Notices, 34 (1):345, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marcos:1999:FPA

[2201] Daniel H. Marcos, Pablo E. Martínez López, and Walter A. Risi. A functional programming approach to hypermedia authoring. ACM SIGPLAN Notices, 34 (1):346, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Faggian:1999:TCU

[2202] Claudia Faggian. A term calculus for a unitary approach to normalization. ACM SIGPLAN Notices, 34 (1):347, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Erlich:1999:IDP

[2203] Yan-David Erlich. Implementing design patterns as language constructs. ACM SIGPLAN Notices, 34(1):348, January 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kiselyov:1999:LCN

[2204] Oleg Kiselyov. A lazy CGI namespace in Scheme. ACM SIGPLAN Notices, 34 (1):349, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:1999:ADI

[2205] Yuchen Zhang and Yanhong A. Liu.

Automating derivation of incremental programs. ACM SIGPLAN Notices, 34 (1):350, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haiming:1999:CCR

[2206] Chen Haiming. Combining CFG and recursive functions to get a new language. ACM SIGPLAN Notices, 34 (1):351, January 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1999:LEM

[2207] A. Michael Berman. Letter from the editor: Malik’s “Evolution of the High Level Programming Languages: A Critical Perspective”: A correction and apology. ACM SIG-PLAN Notices, 34(2):1, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Berman:

A=_Michael.html. The editor apologizes for publication of a paper [2158] that was largely plagiarized from a published book.

Burke:1999:AEC

[2208] Michael G. Burke. Activities: Executive Committee news: Panel sessions at SIGPLAN conference. ACM SIGPLAN Notices, 34(2):2–3, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Burke:
Michael_G=.html.

Bergmann:1999:CCCa

[2209] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 34 (2):4–15, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1999:EEA

[2210] Neil D. Jones. EAPLS: the European Association for Programming Languages and Systems. ACM SIG-PLAN Notices, 34(2):19–20, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1999:ESE

[2211] Neil D. Jones. The European scene: EAPLS: the European Association for Programming Languages and Systems. ACM SIGPLAN Notices, 34(2):19–20, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/j/Jones:Neil_
D=.html.

Ball:1999:FPP

[2212] Thomas Ball. Functional programming: Paths between imperative and functional programming languages. ACM SIGPLAN Notices, 34(2):21–25, February 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Ball:Thomas.
html.

Ball:1999:PBI

[2213] Thomas Ball. Paths between imperative and functional programming. ACM SIGPLAN Notices, 34(2):21– 25, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1999:PPPa

[2214] Chris Clark. Practical parsing patterns: What to do with a dangling else. ACM SIGPLAN Notices, 34(2):26–31, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/c/Clark:Chris.
                                        html.

Clark:1999:WDD

[2215] Chris Clark. What to do with a dangling else. ACM SIGPLAN Notices, 34(2):26–31, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1999:CM

[2216] Brent W. Benson, Jr. Closures for the masses. ACM SIGPLAN Notices, 34(2):32–35, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1999:JRI

[2217] Brent W. Benson, Jr. Java reflections: Inner classes: Closures for the masses. ACM SIGPLAN Notices, 34(2):32–35, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/b/Benson_Jr=:

Brent_W=.html. This paper describes recent extensions to the Java language languages to support inner classes and closures.

Frenger:1999:FOT

[2218] Paul Frenger. Forth and the open terminal architecture. ACM SIGPLAN Notices, 34(2):36–38, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:1999:FRF

[2219] Paul Frenger. Forth report: Forth and the Open Terminal Architecture. ACM SIGPLAN Notices, 34 (2):36–39, February 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/f/Frenger:Paul.
html.

Berman:1999:TCIa

[2220] A. Michael Berman. Technical correspondence: Information for authors. ACM SIGPLAN Notices, 34 (2):40, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carlisle:1999:LES

[2221] Martin C. Carlisle. Letters to the Editor: SIGPLAN Notices on language evolution. ACM SIGPLAN Notices, 34 (2):41, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1999:LED

[2222] Michael B. Feldman. Letters to the Editor: DEC 98 SIGPLAN Notices. ACM SIGPLAN Notices, 34(2):41–42, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bigot:1999:LEC

[2223] Peter A. Bigot. Letters to the Editor: Comment on Spaniel paper. ACM SIG-PLAN Notices, 34(2):42–45, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [2155, 2224].

Syropoulos:1999:LER

[2224] Apostolos Syropoulos. Letters to the Editor: Response to Peter A. Bigot. ACM SIGPLAN Notices, 34 (2):46, February 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [2155, 2223].

Spinellis:1999:DPO

[2225] Diomidis Spinellis. Declarative peephole optimization using string pattern matching. ACM SIGPLAN Notices, 34(2):47–50, February 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spinellis:1999:TCD

[2226] Diomidis Spinellis. Technical correspondence: Declarative peephole optimization using string pattern matching. ACM SIGPLAN Notices, 34 (2):47–51, February 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/s/Spinellis:Diomidis.
                                        html.

Xu:1999:TSL

[2227] Dianxiang Xu. A type system of logic objects. ACM SIGPLAN Notices, 34(2):52–56, February 1999.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/x/Xu:Dianxiang.
html.

Larose:1999:CIC

[2228] Martin Larose and Marc Feeley.

A compacting incremental collector and its performance in a production quality compiler. ACM SIG-PLAN Notices, 34(3):1–9, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p1-larose/

.

Azagury:1999:CCM

[2229] Alain Azagury, Elliot K. Kolodner, Erez Petrank, and Zvi Yehudai. Combining card marking with remembered sets: how to save scanning time. ACM SIGPLAN Notices, 34(3):10–19, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p10-azagury/

.

Pirinen:1999:BTI

[2230] Pekka P. Pirinen. Barrier techniques for incremental tracing. ACM SIG-PLAN Notices, 34(3):20–25, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p20-pirinen/

.

Johnstone:1999:MFP

[2231] Mark S. Johnstone and Paul R. Wilson. The memory fragmentation problem: solved? ACM SIG-PLAN Notices, 34(3):26–36, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/286860/p26-johnstone/

.

Chilimbi:1999:UGG

[2232] Trishul M. Chilimbi and James R. Larus. Using generational garbage collection to implement cache-conscious data placement. ACM SIGPLAN Notices, 34(3):37–48, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/286860/p37-chilimbi/

.

Roth:1999:OBC

[2233] David J. Roth and David S. Wise. Onebit counts between unique and sticky. ACM SIGPLAN Notices, 34(3):49–56, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/286860/p49-roth/

.

Moreau:1999:HDR

[2234] Luc Moreau. Hierarchical distributed reference counting. ACM SIG-PLAN Notices, 34(3):57–67, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p57-moreau/

.

Smith:1999:CMC

[2235] Frederick Smith and Greg Morrisett.

Comparing mostly-copying and marksweep conservative collection. ACM SIGPLAN Notices, 34(3):68–78, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p68-smith/

.

Rodriguez-Rivera:1999:NFN

[2236] Gustavo Rodriguez-Rivera, Michael Spertus, and Charles Fiterman. A non-fragmenting non-moving, garbage collector. ACM SIGPLAN Notices, 34(3):79–85, March 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/286860/p79-rodriguez-
rivera/.

Nishanov:1999:GCG

[2237] Gor V. Nishanov and Sibylle Schupp.

Garbage collection in generic libraries. ACM SIGPLAN Notices, 34(3):86–96, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p86-nishanov/

.

Demoen:1999:MMP

[2238] Bart Demoen and Konstantinos Sagonas. Memory management for Prolog with tabling. ACM SIG-PLAN Notices, 34(3):97–106, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p97-demoen/

.

Wallace:1999:BBL

[2239] Malcolm Wallace and Colin Runciman. The bits between the lambdas: binary data in a lazy functional language. ACM SIGPLAN Notices, 34(3):107–117, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/286860/p107-wallace/

.

Lim:1999:MER

[2240] Tian F. Lim, Przemysław Pardyak, and Brian N. Bershad. A memoryefficient real-time non-copying garbage collector. ACM SIGPLAN Notices, 34(3):118–129, March 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/286860/p118-lim/

.

Siebert:1999:GND

[2241] Fridtjof Siebert. Guaranteeing nondisruptiveness and real-time deadlines in an incremental garbage collector. ACM SIGPLAN Notices, 34(3):130–137, March 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/286860/p130-siebert/

.

Hicks:1999:SLO

[2242] Michael Hicks, Luke Hornof, Jonathan T.

Moore, and Scott M. Nettles. A study of large object spaces. ACM SIG-PLAN Notices, 34(3):138–145, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p138-hicks/

.

Kakkad:1999:PRT

[2243] Sheetal V. Kakkad, Mark S. Johnstone, and Paul R. Wilson. Portable run-time type description for conventional compilers. ACM SIG-PLAN Notices, 34(3):146–153, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p146-kakkad/

.

Colnet:1999:CSC

[2244] Dominique Colnet, Philippe Coucaud, and Olivier Zendra. Compiler support to customize the mark and sweep algorithm. ACM SIG-PLAN Notices, 34(3):154–165, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/286860/p154-colnet/

.

Huelsbergen:1999:VCM

[2245] Lorenz Huelsbergen and Phil Winterbottom. Very concurrent mark-&- sweep garbage collection without finegrain synchronization. ACM SIG-PLAN Notices, 34(3):166–175, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/286860/p166-huelsbergen/

.

Larson:1999:MAL

[2246] Per-Åke Larson and Murali Krishnan. Memory allocation for longrunning server applications. ACM SIG-PLAN Notices, 34(3):176–185, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/286860/p176-larson/

.

Tofte:1999:BIR

[2247] Mads Tofte. A brief introduction to regions. ACM SIGPLAN Notices, 34(3):186–195, March 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/286860/p186-tofte/

.

Zorn:1999:ACS

[2248] Ben Zorn. Activities: Corporate support and SIGPLAN. ACM SIG-PLAN Notices, 34(4):1–2, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/z/Zorn:Benjamin_
G=.html.

Zorn:1999:CSS

[2249] Ben Zorn. Corporate support and SIG-PLAN. ACM SIGPLAN Notices, 34 (4):1–2, April 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1999:CCCb

[2250] Seth Bergmann. Conference corner:

Calendar. ACM SIGPLAN Notices, 34 (4):3–15, April 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1999:FCR

[2251] Anonymous. Federated Computing Research Conference: April 30–May 6, 1999, Atlanta, GA. ACM SIG-PLAN Notices, 34(4):16, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1999:APA

[2252] Anonymous. Advance program for ACM SIGPLAN ’99 Conference on Programming Language Design and Implementation: Atlanta Hilton & Towers, Atlanta, Georgia, May 1–4,

1999; http://www.cs.rutgers.edu/

pldi99. ACM SIGPLAN Notices, 34 (4):17–18, April 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1999:BTS

[2253] Chris Clark. Build a tree — save a parse. ACM SIGPLAN Notices, 34 (4):19–24, April 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1999:PPPb

[2254] Chris Clark. Practical parsing patterns: Build a tree — save a parse. ACM SIGPLAN Notices, 34(4):19–24, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/c/Clark:
Chris.html.

Benson:1999:J

[2255] Brent W. Benson, Jr. JavaScript. ACM SIGPLAN Notices, 34(4):25–27, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:1999:JRJ

[2256] Brent W. Benson, Jr. Java reflections: JavaScript — not Java (but just as hot). ACM SIG-PLAN Notices, 34(4):25–27, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Benson_
Jr=:Brent_W=.html.

Frenger:1999:FRP

[2257] Paul Frenger. Forth report: Parallel Forth. ACM SIGPLAN Notices, 34(4):28–32, April 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/f/Frenger:Paul.
html.

Frenger:1999:PF

[2258] Paul Frenger. Parallel Forth. ACM SIGPLAN Notices, 34(4):28–32, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanRoy:1999:AMP

[2259] Peter Van Roy. Announcing the Mozart programming system. ACM SIGPLAN Notices, 34(4):33–34, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

VanRoy:1999:RAM

[2260] Peter Van Roy. Resources: Announcing the Mozart programming system. ACM SIGPLAN Notices, 34(4):33–34, April 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/r/Roy:Peter_Van.
html.

Peterson:1999:FPM

[2261] John Peterson. Functional programming: Microsoft and Yale conclude agreement to license technology for Haskell. ACM SIGPLAN Notices, 34(4):35–36, April 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/p/Peterson:John.

html. Warning: This is an April Fool’s joke.

Peterson:1999:MYC

[2262] John Peterson. Microsoft and Yale conclude agreement to license technology for Haskell. ACM SIGPLAN Notices, 34(4):35–36, April 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berman:1999:TCIb

[2263] A. Michael Berman. Technical correspondence: Information for authors. ACM SIGPLAN Notices, 34(4):37, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1999:JIP

[2264] Per Brinch Hansen. Java’s insecure parallelism. ACM SIGPLAN Notices, 34 (4):38–45, April 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1999:TCJ

[2265] Per Brinch Hansen. Technical correspondence: Java’s insecure parallelism. ACM SIGPLAN Notices, 34(4):38–45, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/h/Hansen:
                                        Per_Brinch.html.

Harris:1999:ESR

[2266] Timothy Harris. Early storage reclamation in a tracing garbage collector. ACM SIGPLAN Notices, 34(4):46–53, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //ftp.informatik.rwth-aachen.de/

dblp/db/indices/a-tree/h/Harris:
Timothy.html.

Pareja-Flores:1999:CED

[2267] Cristóbal Pareja-Flores and J. Ángel Velázquez-Iturbide. Calculating encoding and decoding functions for prefix codes. ACM SIGPLAN Notices, 34(4):54–60, April 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/p/Pareja=
Flores:Crist=oacute=bal.html;
http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/v/Vel=
aacute=zquez=Iturbide:J=_=Aacute=
ngel.html.

Biswas:1999:REP

[2268] Bitan Biswas and Rajib Mall. Reverse execution of programs. ACM SIGPLAN Notices, 34(4):61–69, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Biswas:
Bitan.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/m/Mall:Rajib.html.

Agrawal:1999:NPR

[2269] Gagan Agrawal, Shyamala Murthy, and Chandrashekhar Garud. A novel program representation for interprocedural analysis. ACM SIG-PLAN Notices, 34(4):70–76, April 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/

                                        db/indices/a-tree/a/Agrawal:Gagan.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        g/Garud:Chandrashekar.html;   http:
                                        //ftp.informatik.rwth-aachen.de/
                                        dblp/db/indices/a-tree/m/Murthy:
                                        Shyamala.html.

Chilimbi:1999:CCSa

[2270] Trishul M. Chilimbi, Mark D. Hill, and James R. Larus. Cacheconscious structure layout. ACM SIGPLAN Notices, 34(5):1–12, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/301122/p1-chilimbi/

. See PLDI’99 proceedings [2895].

Chilimbi:1999:CCSb

[2271] Trishul M. Chilimbi, Bob Davidson, and James R. Larus. Cacheconscious structure definition. ACM SIGPLAN Notices, 34(5):13–24, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/301122/p13-chilimbi/

. See PLDI’99 proceedings [2895].

Jones:1999:SIE

[2272] Simon Peyton Jones, Alastair Reid, Fergus Henderson, Tony Hoare, and Simon Marlow. A semantics for imprecise exceptions. ACM SIGPLAN Notices, 34(5):25–36, May 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

PeytonJones:1999:SIE

[2273] Simon Peyton Jones, Alastair Reid, Tony Hoare, Simon Marlow, and Fergus Henderson. A semantics for imprecise exceptions. ACM SIG-PLAN Notices, 34(5):25–36, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p25-peyton_

jones/. See PLDI’99 proceedings [2895].

Fisher:1999:DCM

[2274] Kathleen Fisher and John Reppy. The design of a class mechanism for Moby. ACM SIGPLAN Notices, 34(5):37–49, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p37-fisher/

. See PLDI’99 proceedings [2895].

Crary:1999:WRM

[2275] Karl Crary, Robert Harper, and Sidd Puri. What is a recursive module? ACM SIGPLAN Notices, 34(5):50–63, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p50-crary/

. See PLDI’99 proceedings [2895].

Bodik:1999:LRA

[2276] Rastislav Bodík, Rajiv Gupta, and Mary Lou Soffa. Load-reuse analysis: Design and evaluation. ACM SIGPLAN Notices, 34(5):64–76, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p64-bodik/

. See PLDI’99 proceedings [2895].

Rugina:1999:PAM

[2277] Radu Rugina and Martin Rinard. Pointer analysis for multithreaded programs. ACM SIGPLAN Notices, 34(5):77–90, May 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p77-rugina/

. See PLDI’99 proceedings [2895].

Yong:1999:PAP

[2278] Suan Hsi Yong, Susan Horwitz, and Thomas Reps. Pointer analysis for programs with structures and casting. ACM SIGPLAN Notices, 34(5):91–103, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p91-yong/

. See PLDI’99 proceedings [2895].

Blelloch:1999:BTS

[2279] Guy E. Blelloch and Perry Cheng. On bounding time and space for multiprocessor garbage collection. ACM SIG-PLAN Notices, 34(5):104–117, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p104-blelloch/

. See PLDI’99 proceedings [2895].

Stichnoth:1999:SGC

[2280] James M. Stichnoth, Guei-Yuan Lueh, and Michał Cierniak. Support for garbage collection at every instruction in a Java compiler. ACM SIG-PLAN Notices, 34(5):118–127, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p118-stichnoth/

. See PLDI’99 proceedings [2895].

Rao:1999:SAO

[2281] Amit Rao and Santosh Pande. Storage assignment optimizations to generate compact and efficient code on embedded DSPs. ACM SIG-PLAN Notices, 34(5):128–138, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p128-rao/

. See PLDI’99 proceedings [2895].

Cooper:1999:ECC

[2282] Keith D. Cooper and Nathaniel McIntosh. Enhanced code compression for embedded RISC processors. ACM SIG-PLAN Notices, 34(5):139–149, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p139-cooper/

. See PLDI’99 proceedings [2895].

Vegdahl:1999:UNM

[2283] Steven R. Vegdahl. Using node merging to enhance graph coloring. ACM SIG-PLAN Notices, 34(5):150–154, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p150-vegdahl/

. See PLDI’99 proceedings [2895].

Schlansker:1999:CCB

[2284] Michael Schlansker, Scott Mahlke, and Richard Johnson. Control CPR: A branch height reduction optimization for EPIC architectures. ACM SIG-PLAN Notices, 34(5):155–168, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p155-schlansker/

. See PLDI’99 proceedings [2895].

Frigo:1999:FFT

[2285] Matteo Frigo. A fast Fourier transform compiler. ACM SIGPLAN Notices, 34(5):169–180, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p169-frigo/

. See PLDI’99 proceedings [2895].

Wu:1999:NFD

[2286] Le-Chun Wu, Rajiv Mirani, Harish Patil, Bruce Olsen, and Wen mei W. Hwu. A new framework for debugging globally optimized code. ACM SIGPLAN Notices, 34(5):181–191, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p181-wu/

. See PLDI’99 proceedings [2895].

Foster:1999:TTQ

[2287] Jeffrey S. Foster, Manuel Fähndrich, and Alexander Aiken. A theory of type qualifiers. ACM SIG-PLAN Notices, 34(5):192–203, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p192-foster/

. See PLDI’99 proceedings [2895].

Leung:1999:SSA

[2288] Allen Leung and Lal George. Static single assignment form for machine code. ACM SIGPLAN Notices, 34(5):204–214, May 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p204-leung/

. See PLDI’99 proceedings [2895].

Song:1999:NTT

[2289] Yonghong Song and Zhiyuan Li.

New tiling techniques to improve cache temporal locality. ACM SIG-PLAN Notices, 34(5):215–228, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p215-song/

. See PLDI’99 proceedings [2895].

Ding:1999:ICP

[2290] Chen Ding and Ken Kennedy. Improving cache performance in dynamic applications through data and computation reorganization at run time. ACM SIGPLAN Notices, 34(5):229–241, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/301122/p229-ding/

. See PLDI’99 proceedings [2895].

Fraser:1999:AIM

[2291] Christopher W. Fraser. Automatic inference of models for statistical code compression. ACM SIG-PLAN Notices, 34(5):242–246, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/pldi/301122/p242-fraser/

. See PLDI’99 proceedings [2895].

Pugh:1999:CJC

[2292] William Pugh. Compressing Java class files. ACM SIGPLAN Notices, 34(5):247–258, May 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/301122/p247-pugh/

. See PLDI’99 proceedings [2895].

Larus:1999:WPP

[2293] James R. Larus. Whole program paths. ACM SIGPLAN Notices, 34(5):259–269, May 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/pldi/301122/p259-larus/

. See PLDI’99 proceedings [2895].

Fraser:1999:FSC

[2294] Christopher W. Fraser and Todd A.

Proebsting. Finite-state code generation. ACM SIGPLAN Notices, 34(5):270–280, May 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p270-fraser/

. See PLDI’99 proceedings [2895].

Marlet:1999:EIR

[2295] Renaud Marlet, Charles Consel, and Philippe Boinot. Efficient incremental run-time specialization for free. ACM SIGPLAN Notices, 34(5):281–292, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p281-martlet/

. See PLDI’99 proceedings [2895].

Grant:1999:ERT

[2296] Brian Grant, Matthai Philipose, Markus Mock, Susan J. Eggers, and Craig Chambers. An evaluation of run-time optimizations. ACM SIG-PLAN Notices, 34(5):293–304, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/
proceedings/pldi/301122/p293-grant/

. See PLDI’99 proceedings [2895].

Grant:1999:ESR

[2297] Brian Grant, Matthai Philipose, Markus Mock, Craig Chambers, and Susan J. Eggers. An evaluation of staged run-time optimizations in DyC.

ACM SIGPLAN Notices, 34(5):293– 304, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davidson:1999:NCI

[2298] Jack Davidson, Monica Lam, David Moore, and Mike Smith. National Compiler Infrastructure progress report. ACM SIGPLAN Notices, 34(5): ??, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Was scheduled to appear in PLDI’99 proceedings [2895], but nothing but a table-of-contents entry was given there.

Rhyne:1999:IPQ

[2299] Jim Rhyne, Steven Lucco, Edward Felten, and Guy L. Steele, Jr. Industrial panel: “Trends in the Future of Industrial Computing”. ACM SIG-PLAN Notices, 34(5):??, May 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Was scheduled to appear in PLDI’99 proceedings [2895], but nothing but a table-of-contents entry was given there.

Kessler:1999:ECN

[2300] Robert Kessler. Executive committee news: SIGPLAN conferences and workshops. ACM SIGPLAN Notices, 34(6): 1, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergmann:1999:CCCc

[2301] Seth Bergmann. Conference corner: Calendar. ACM SIGPLAN Notices, 34(6):2–12, June 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1999:CPP

[2302] Anonymous. Call for papers: POPL ’00: The 27th annual ACM SIGPLAN-SIGACT symposium on principles of programming languages. ACM SIG-PLAN Notices, 34(6):13, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1999:CPAa

[2303] Anonymous. Call for papers: APL99:

On track to the 21st Century: International conference on array processing languages. ACM SIGPLAN Notices, 34(6):14, June 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1999:CPAb

[2304] Anonymous. Call for papers: ACM SIGPLAN workshop on dynamic and adaptive compilation and optimization (Dynamo’00). ACM SIGPLAN Notices, 34(6):15, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:1999:CCC

[2305] Annie Liu and Scott D. Stoller.

Conference corner: Conference report: ETAPS ’99. ACM SIG-PLAN Notices, 34(6):16–17, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/l/Liu:Yanhong_
A=.html;   http://ftp.informatik.

                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/s/Stoller:Scott_D=.html.

Clark:1999:KSS

[2306] Chris Clark. Keywords: Scanners and Screeners. ACM SIGPLAN Notices, 34 (6):18–22, June 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1999:PPPc

[2307] Chris Clark. Practical parsing patterns: Keywords: Scanners and screeners. ACM SIGPLAN Notices, 34(6):18–22, June 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/c/Clark:Chris.
                                        html.

Plasmeijer:1999:FPK

[2308] Rinus Plasmeijer and Marko van Eekelen. Functional programming: Keep it clean: A unique approach to functional programming. ACM SIG-PLAN Notices, 34(6):23–31, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/e/Eekelen:Marko_
                                        C=_J=_D=_van.html;   http://ftp.
                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/p/Plasmeijer:
                                        Marinus_J=.html.

Plasmeijer:1999:KIC

[2309] Rinus Plasmeijer and Marko van Eekelen. Keep it clean: a unique approach to functional programming. ACM SIG-PLAN Notices, 34(6):23–31, June 1999.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sadler:1999:FFO

[2310] John Sadler. Ficl, FORML, & Object Forth. ACM SIGPLAN Notices, 34(6):32–35, June 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sadler:1999:FFR

[2311] John Sadler. Forth: Forth report:

Ficl, FORML, & Object Forth. ACM SIGPLAN Notices, 34(6):32–35, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/s/Sadler:John.
html.

Anonymous:1999:TCI

[2312] Anonymous. Technical correspondence: Information for authors. ACM SIGPLAN Notices, 34(6):36, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Panizzi:1999:RTV

[2313] Emanuele Panizzi and Bernardo Pastorelli. On the return types of virtual functions. ACM SIGPLAN Notices, 34 (6):37–42, June 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Panizzi:1999:TCR

[2314] Emanuele Panizzi and Bernardo Pastorelli. Technical correspondence: On the return types of virtual functions. ACM SIGPLAN Notices, 34(6):37–42, June 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/p/Panizzi:Emanuele.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
p/Pastorelli:Bernardo.html.

Wolfe:1999:PRE

[2315] Michael Wolfe. Partial redundancy elimination is not bidirectional. ACM SIGPLAN Notices, 34(6):43–46, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/w/Wolfe:
Michael.html.

Khedker:1999:BDF

[2316] Uday P. Khedker and Dhananjay M. Dhamdhere. Bidirectional data flow analysis: Myths and reality. ACM SIGPLAN Notices, 34(6):47–57, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/d/Dhamdhere:Dhananjay_
M=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/k/Khedker:Uday_P=.html.

Naik:1999:OOS

[2317] Mayur Naik and Rajeev Kumar. Object-oriented symbol management in syntax-directed compiler systems. ACM SIGPLAN Notices, 34(6):58–67, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/k/Kumar:
Rajeev.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/n/Naik:Mayur.html.

Mernik:1999:IMA

[2318] Marjan Mernik, Viljem Žumer, Mitja Lenič, and Enis Avdičaušević. Implementation of multiple attribute grammar inheritance in the tool LISA. ACM SIGPLAN Notices, 34(6):68–75, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/a/Avdicausevic:
Enis.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/l/Lenic:Mitja.html;   http:
//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/m/Mernik:
Marjan.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/z/Zumer:Viljem.html.

Wolfgang:1999:IST

[2319] Paul Wolfgang and Yang Song. Integration of the Standard Template Library and the Microsoft Foundation Class. ACM SIGPLAN Notices, 34(6):76–81, June 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/s/Song:Yang.html;
 http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/w/Wolfgang:
Paul.html.

Botting:1999:APN

[2320] Richard J. Botting. On the application of a popular notation to semantics. ACM SIGPLAN Notices, 34(6):82–83, June 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/b/Botting:Richard_

J=.html. See comment [?].

Kokol:1999:CFS

[2321] Peter Kokol, Vili Podgorelec, Henri Habrias, and Nassim Hadj Rabia. The complexity of formal specifications — assessments by α-metric. ACM SIGPLAN Notices, 34(6):84–88, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/h/Habrias:Henri.
                                        html;   http://ftp.informatik.rwth-
                                        aachen.de/dblp/db/indices/a-tree/
                                        k/Kokol:Peter.html;   http://ftp.
                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/p/Podgorelec:
                                        Vili.html;   http://ftp.informatik.
                                        rwth-aachen.de/dblp/db/indices/
                                        a-tree/r/Rabia:Nassim_Hadj.html.

Anonymous:1999:F

[2322] Anonymous. Forth. ACM SIG-PLAN Notices, 34(6):??, June 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1999:FP

[2323] Anonymous. Functional programming. ACM SIGPLAN Notices, 34 (6):??, June 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1999:PPP

[2324] Anonymous. Practical parsing patterns. ACM SIGPLAN Notices, 34 (6):??, June 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1999:TC

[2325] Anonymous. Technical Correspondence. ACM SIGPLAN Notices, 34 (6):??, June 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cooper:1999:ORC

[2326] Keith D. Cooper, Philip J. Schielke, and Devika Subramanian. Optimizing for reduced code space using genetic algorithms. ACM SIGPLAN Notices, 34(7):1–9, July 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Uh:1999:EEZ

[2327] Gang-Ryung Uh, Yuhong Wang, David Whalley, Sanjay Jinturkar, Chris Burns, and Vincent Cao. Effective exploitation of a zero overhead loop buffer. ACM SIGPLAN Notices, 34 (7):10–19, July 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eckstein:1999:MCL

[2328] Erik Eckstein and Andreas Krall. Minimizing cost of local variables access for DSP-processors. ACM SIG-PLAN Notices, 34(7):20–27, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stotzer:1999:MST

[2329] Eric Stotzer and Ernst Leiss. Modulo scheduling for the TMS320C6x VLIW DSP architecture. ACM SIG-PLAN Notices, 34(7):28–34, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:1999:PBP

[2330] Jörn Schneider and Christian Ferdinand. Pipeline behavior prediction for superscalar processors by abstract interpretation. ACM SIGPLAN Notices, 34(7):35–44, July 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Persson:1999:LMA

[2331] Patrik Persson. Live memory analysis for garbage collection in embedded systems. ACM SIGPLAN Notices, 34 (7):45–54, July 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kim:1999:SGC

[2332] Taehyoun Kim, Naehyuck Chang, Namyun Kim, and Heonshik Shin. Scheduling garbage collector for embedded real-time systems. ACM SIG-PLAN Notices, 34(7):55–64, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sugawara:1999:TBQ

[2333] Tomoyoshi Sugawara and Kosuke Tatsukawa. Table-based QoS control for embedded real-time systems. ACM SIGPLAN Notices, 34(7):65–72, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tomoyoshi:1999:TBQ

[2334] Sugawara Tomoyoshi and Tatsukawa Kosuke. Table-based QoS control for embedded real-time systems. ACM SIGPLAN Notices, 34(7):65–72, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1999:SDC

[2335] Sungyoung Lee, Byeong-Soo Jeong, and Hyon-Woo Seung. A secure dynamic copy protocol in real-time secure database systems. ACM SIG-PLAN Notices, 34(7):73–79, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Park:1999:SDC

[2336] Soo-Yeon Park, Sungyoung Lee, Byung-Soo Jeong, and Hyon Woo Seung. A secure dynamic copy protocol in real-time database systems. ACM SIGPLAN Notices, 34(7):73–79, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryu:1999:PRE

[2337] Minsoo Ryu, Jungkeun Park, Kimoon Kim, Yangmin Seo, and Seongsoo Hong. Performance re-engineering of embedded real-time systems. ACM SIGPLAN Notices, 34(7):80–86, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kang:1999:SST

[2338] D.-I. Kang, R. Gerber, L. Golubchik, J. K. Hollingsworth, and M. Saksena. A software synthesis tool for distributed embedded system design. ACM SIG-PLAN Notices, 34(7):87–95, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Engblom:1999:WSS

[2339] Jakob Engblom. Why SpecInt95 should not be used to benchmark embedded systems tools. ACM SIGPLAN Notices, 34(7):96–103, July 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ryu:1999:ICD

[2340] In Ryu. Issues and challenges in developing embedded software for information appliances and telecommunication terminals. ACM SIGPLAN Notices, 34 (7):104–120, July 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1999:BCA

[2341] Jaejin Lee, David A. Padua, and Samuel P. Midkiff. Basic compiler algorithms for parallel programs. ACM SIGPLAN Notices, 34(8):1–12, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p1-lee/.

Knoop:1999:CME

[2342] Jens Knoop and Bernhard Steffen. Code motion for explicitly parallel programs. ACM SIGPLAN Notices, 34(8):13–24, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
ppopp/301104/p13-knoop/.

McCurdy:1999:ECP

[2343] Collin McCurdy and John Mellor-Crummey. An evaluation of computing paradigms for N-body simulations on distributed memory architectures. ACM SIGPLAN Notices, 34(8):25–36, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
ppopp/301104/p25-mccurdy/.

Liao:1999:SEI

[2344] Shih-Wei Liao, Amer Diwan, Robert P.

Bosch, Jr., Anwar Ghuloum, and Monica S. Lam. SUIF Explorer: an interactive and interprocedural parallelizer. ACM SIGPLAN Notices, 34(8):37–48, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
ppopp/301104/p37-liao/.

Xu:1999:DIT

[2345] Zhichen Xu, Barton P. Miller, and Oscar Naim. Dynamic instrumentation of threaded applications. ACM SIGPLAN Notices, 34(8):49–59, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p49-xu/.

Taura:1999:SMI

[2346] Kenjiro Taura, Kunio Tabata, and Akinori Yonezawa. StackThreads/ MP: integrating futures into calling standards. ACM SIGPLAN Notices, 34(8):60–71, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p60-taura/.

Rugina:1999:APD

[2347] Radu Rugina and Martin Rinard. Automatic parallelization of divide and conquer algorithms. ACM SIGPLAN Notices, 34(8):72–83, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p72-rugina/.

Moon:1999:EPA

[2348] Sungdo Moon and Mary W. Hall. Evaluation of predicated array dataflow analysis for automatic parallelization. ACM SIGPLAN Notices, 34(8):84–95, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p84-moon/.

Scherer:1999:TAP

[2349] Alex Scherer, Honghui Lu, Thomas Gross, and Willy Zwaenepoel. Transparent adaptive parallelism on NOWs using OpenMP. ACM SIGPLAN Notices, 34(8):96–106, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
ppopp/301104/p96-scherer/.

Tang:1999:CRT

[2350] Hong Tang, Kai Shen, and Tao Yang. Compile/run-time support for threaded MPI execution on multiprogrammed shared memory machines. ACM SIGPLAN Notices, 34(8):107–118, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
ppopp/301104/p107-tang/.

Mainwaring:1999:DCV

[2351] Alan M. Mainwaring and David E.

Culler. Design challenges of virtual networks: fast, general-purpose communication. ACM SIGPLAN Notices, 34(8):119–130, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
ppopp/301104/p119-mainwaring/.

Kielmann:1999:MMC

[2352] Thilo Kielmann, Rutger F. H. Hofman, Henri E. Bal, Aske Plaat, and Raoul A. F. Bhoedjang. MagPIe: MPI’s collective communication operations for clustered wide area systems. ACM SIGPLAN Notices, 34 (8):131–140, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p131-kielmann/.

Sundaram-Stukel:1999:PAW

[2353] David Sundaram-Stukel and Mary K. Vernon. Predictive analysis of a wavefront application using LogGP. ACM SIGPLAN Notices, 34(8):141– 150, August 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p141-sundaram-stukel/

.

Bagrodia:1999:PPL

[2354] Rajive Bagrodia, Ewa Deeljman, Steven Docy, and Thomas Phan. Performance prediction of large parallel applications using parallel simulations. ACM SIGPLAN Notices, 34(8):151–162, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p151-bagrodia/.

Subhlok:1999:ANS

[2355] Jaspal Subhlok, Peter Lieu, and Bruce Lowekamp. Automatic node selection for high performance applications on networks. ACM SIGPLAN Notices, 34(8):163–172, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        ppopp/301104/p163-subhlok/.

Maassen:1999:EIJ

[2356] Jason Maassen, Rob van Nieuwpoort, Ronald Veldema, Henri E. Bal, and Aske Plaat. An efficient implementation of Java’s remote method invocation. ACM SIGPLAN Notices, 34(8):173–182, August 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
ppopp/301104/p173-maassen/.

Ramachandran:1999:STM

[2357] Umakishore Ramachandran, Rishiyur S.

Nikhil, Nissim Harel, James M. Rehg, and Kathleen Knobe. Space-time memory: a parallel programming abstraction for interactive multimedia applications. ACM SIGPLAN Notices, 34(8):183–192, August 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
ppopp/301104/p183-ramachandran/.

Longley:1999:WFP

[2358] John Longley. When is a functional program not a functional program? ACM SIGPLAN Notices, 34(9):1–7, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p1-longley. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Aiken:1999:MLM

[2359] Alex Aiken. Modern languages for modern parallel computing (invited lecture, abstract only). ACM SIGPLAN Notices, 34(9):8, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p-aiken. Proceedings of

the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99). Invited talk: abstract only.

Arvind:1999:YCD

[2360] Arvind. You can design microprocessors too: Putting TRS’s to work (invited lecture, abstract only). ACM SIGPLAN Notices, 34(9):8, September 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        fp/317636/p-arvind. Proceedings of

the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99). Invited talk: abstract only.

Leroy:1999:OCM

[2361] Xavier Leroy. Objects, classes and modules in Objective Caml (invited lecture, abstract only). ACM SIGPLAN Notices, 34(9):8, September 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        fp/317636/p-leroy. Proceedings of

the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99). Invited talk: abstract only.

Semmelroth:1999:MEM

[2362] Miley Semmelroth and Amr Sabry. Monadic encapsulation in ML. ACM SIGPLAN Notices, 34(9):8–17, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

www.acm.org/pubs/citations/proceedings/
fp/317636/p8-semmelroth. Proceed-

ings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Ganz:1999:TS

[2363] Steven E. Ganz, Daniel P. Friedman, and Mitchell Wand. Trampolined style. ACM SIGPLAN Notices, 34(9):18–27, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p18-ganz. Proceedings of

the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Okasaki:1999:FES

[2364] Chris Okasaki. From fast exponentiation to square matrices: An adventure in types. ACM SIGPLAN Notices, 34(9):28–35, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p28-okasaki. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Nilsson:1999:TPP

[2365] Henrik Nilsson. Tracing piece by piece:

Affordable debugging for lazy functional languages. ACM SIGPLAN Notices, 34(9):36–47, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p36-nilsson. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Shivers:1999:AHT

[2366] Olin Shivers, James W. Clark, and Roland McGrath. Atomic heap transactions and fine-grain interrupts. ACM SIGPLAN Notices, 34 (9):48–59, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p48-shivers. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Launchbury:1999:EMD

[2367] John Launchbury, Jeffrey R. Lewis, and Byron Cook. On embedding a microarchitectural design language within Haskell. ACM SIGPLAN Notices, 34(9):60–69, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p60-launchbury.  Pro-

ceedings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Hughes:1999:RDD

[2368] John Hughes and Lars Pareto. Recursion and dynamic data-structures in bounded space: Towards embedded ML programming. ACM SIGPLAN Notices, 34(9):70–81, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p70-hughes. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Crary:1999:SPT

[2369] Karl Crary. A simple proof technique for certain parametricity results. ACM SIGPLAN Notices, 34 (9):82–89, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p82-crary. Proceedings of

the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Kfoury:1999:RTE

[2370] Assaf J. Kfoury, Harry G. Mairson, Franklyn A. Turbak, and J. B. Wells. Relating typability and expressiveness in finite-rank intersection type systems (extended abstract). ACM SIGPLAN Notices, 34(9):90–101, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p90-kfoury. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Splawski:1999:TFI

[2371] Zdzisław Spławski and Paweł Urzyczyn. Type fixpoints: Iteration vs. recursion. ACM SIGPLAN Notices, 34(9):102–113, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p102-splawski. Proceed-

ings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Finne:1999:CHH

[2372] Sigbjorn Finne, Daan Leijen, Erik Meijer, and Simon Peyton Jones. Calling Hell from Heaven and Heaven from Hell. ACM SIGPLAN Notices, 34(9):114–125, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p114-finne. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Benton:1999:IWT

[2373] Nick Benton and Andrew Kennedy. Interlanguage working without tears: Blending SML with Java. ACM SIGPLAN Notices, 34 (9):126–137, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p126-benton. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Flatt:1999:PLO

[2374] Matthew Flatt, Robert Bruce Findler, Shriram Krishnamurthi, and Matthias Felleisen. Programming languages as operating systems (or revenge of the Son of the Lisp Machine). ACM SIGPLAN Notices, 34 (9):138–147, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p138-flatt. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Wallace:1999:HXG

[2375] Malcolm Wallace and Colin Runciman. Haskell and XML: Generic combinators or type-based translation? ACM SIGPLAN Notices, 34 (9):148–159, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p148-wallace. Proceed-

ings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Ohori:1999:TIR

[2376] Atsushi Ohori and Nobuaki Yoshida.

Type inference with rank 1 polymorphism for type-directed compilation of ML. ACM SIGPLAN Notices, 34(9):160–171, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p160-ohori. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Glew:1999:TDN

[2377] Neal Glew. Type dispatch for named hierarchical types. ACM SIGPLAN Notices, 34(9):172–182, September 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic).  URL http://www.acm.
                                        org/pubs/citations/proceedings/
                                        fp/317636/p172-glew. Proceedings of

the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

League:1999:RJC

[2378] Christopher League, Zhong Shao, and Valery Trifonov. Representing Java classes in a typed intermediate language. ACM SIGPLAN Notices, 34 (9):183–196, September 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p183-league. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Zdancewic:1999:PPL

[2379] Steve Zdancewic, Dan Grossman, and Greg Morrisett. Principals in programming languages: A syntactic proof technique. ACM SIGPLAN Notices, 34(9):197–207, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        fp/317636/p197-zdancewic.  Pro-

ceedings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Elsman:1999:SIM

[2380] Martin Elsman. Static interpretation of modules. ACM SIGPLAN Notices, 34(9):208–219, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
fp/317636/p208-elsman. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Shao:1999:TMF

[2381] Zhong Shao. Transparent modules with fully syntactic signatures. ACM SIGPLAN Notices, 34(9):220– 232, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic).  URL http://www.acm.
org/pubs/citations/proceedings/
fp/317636/p220-shao. Proceedings of

the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Crary:1999:FTA

[2382] Karl Crary and Stephanie Weirich.

Flexible type analysis. ACM SIGPLAN Notices, 34(9):233–248, September 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic).  URL http://www.acm.
org/pubs/citations/proceedings/
fp/317636/p233-crary. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Chitil:1999:TIB

[2383] Olaf Chitil. Type inference builds a short cut to deforestation. ACM SIGPLAN Notices, 34(9):249–260, September 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic).  URL http://www.acm.
org/pubs/citations/proceedings/
fp/317636/p249-chitil. Proceedings

of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Huch:1999:VEP

[2384] Frank Huch. Verification of Erlang programs using abstract interpretation and model checking. ACM SIGPLAN Notices, 34(9):261– 272, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic).  URL http://www.acm.
org/pubs/citations/proceedings/
fp/317636/p261-huch. Proceedings of

the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Alpuente:1999:SIS

[2385] María Alpuente, Michael Hanus, Salvador Lucas, and Germán Vidal. Specialization of inductively sequential functional logic programs. ACM SIGPLAN Notices, 34(9):273– 283, September 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic).  URL http://www.acm.
org/pubs/citations/proceedings/

fp/317636/p273-alpuente. Proceed-

ings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99).

Choi:1999:EAJ

[2386] Jong-Deok Choi, Manish Gupta, Mauricio Serrano, Vugranam C. Sreedhar, and Sam Midkiff. Escape analysis for Java. ACM SIGPLAN Notices, 34(10):1–19, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p1-choi/.

Blanchet:1999:EAO

[2387] Bruno Blanchet. Escape analysis for object-oriented languages: application to Java. ACM SIGPLAN Notices, 34 (10):20–34, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p20-blanchet/.

Bogda:1999:RUS

[2388] Jeff Bogda and Urs Hölzle. Removing unnecessary synchronization in Java. ACM SIGPLAN Notices, 34 (10):35–46, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p35-bogda/.

Travassos:1999:DDO

[2389] Guilherme Travassos, Forrest Shull, Michael Fredericks, and Victor R.

Basili. Detecting defects in objectoriented designs: using reading techniques to increase software quality. ACM SIGPLAN Notices, 34 (10):47–56, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p47-travassos/.

France:1999:POA

[2390] Robert France. A problem-oriented analysis of basic UML static requirements modeling concepts. ACM SIGPLAN Notices, 34(10):57–69, October 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p57-france/.

Ovlinger:1999:LSR

[2391] Johan Ovlinger and Mitchell Wand. A language for specifying recursive traversals of object structures. ACM SIGPLAN Notices, 34(10):70– 81, October 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p70-ovlinger/.

Vitek:1999:CT

[2392] Jan Vitek and Boris Bokowski. Confined types. ACM SIGPLAN Notices, 34(10):82–96, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/

pubs/citations/proceedings/oops/
320384/p82-vitek/.

Duggan:1999:MTB

[2393] Dominic Duggan. Modular type-based reverse engineering of parameterized types in Java code. ACM SIGPLAN Notices, 34(10):97–113, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
oops/320384/p97-duggan/.

Madsen:1999:SAV

[2394] Ole Lehrmann Madsen. Semantic analysis of virtual classes and tested classes. ACM SIGPLAN Notices, 34 (10):114–131, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p114-madsen/.

Igarashi:1999:FJM

[2395] Atshushi Igarashi, Benjamin Pierce, and Philip Wadler. Featherweight Java: a minimal core calculus for Java and GJ. ACM SIGPLAN Notices, 34 (10):132–146, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p132-igarashi/.

Freund:1999:FFJ

[2396] Stephen N. Freund and John C.

Mitchell. A formal framework for the Java bytecode language and verifier. ACM SIGPLAN Notices, 34 (10):147–166, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p147-freund/.

Rinat:1999:CPO

[2397] Ran Rinat, Menachem Magidor, and Scott F. Smith. Correspondence polymorphism for object-oriented languages. ACM SIGPLAN Notices, 34 (10):167–186, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p167-rinat/.

Whaley:1999:CPE

[2398] John Whaley and Martin Rinard. Compositional pointer and escape analysis for Java programs. ACM SIGPLAN Notices, 34(10):187–206, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
oops/320384/p187-whaley/.

Agesen:1999:EML

[2399] Ole Agesen, David Detlefs, Alex Garthwaite, Ross Knippel, Y. S. Ramakrishna, and Derek White. An efficient meta-lock for implementing ubiquitous synchronization. ACM SIGPLAN Notices, 34(10):207–222, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
oops/320384/p207-agesen/.

Onodera:1999:SLO

[2400] Tamiya Onodera and Kiyokuni Kawachiya.

A study of locking objects with bimodal fields. ACM SIGPLAN Notices, 34(10):223–237, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
oops/320384/p223-onodera/.

Chambers:1999:EMP

[2401] Craig Chambers and Weimin Chen.

Efficient multiple and predicated dispatching. ACM SIGPLAN Notices, 34 (10):238–255, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p238-chambers/.

Sweeney:1999:STE

[2402] Peter F. Sweeney and Joseph (Yossi)

Gil. Space and time-efficient memory layout for multiple inheritance. ACM SIGPLAN Notices, 34(10):256– 275, October 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p256-sweeney/.

Calder:1999:RTD

[2403] Brad Calder, Chandra Krintz, and Urs Hölzle. Reducing transfer delay using Java class file splitting and prefetching. ACM SIGPLAN Notices, 34 (10):276–291, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        320384/p276-calder/.

Tip:1999:PEA

[2404] Frank Tip, Chris Laffra, Peter F. Sweeney, and David Streeter. Practical experience with an application extractor for Java. ACM SIGPLAN Notices, 34(10):292–305, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/320384/p292-tip/.

Hagimont:1999:PEM

[2405] Daniel Hagimont and L. Ismail. A performance evaluation of the mobile agent paradigm. ACM SIGPLAN Notices, 34(10):306–313, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/320384/p306-hagimont/.

Alpern:1999:IJJ

[2406] Bowen Alpern, C. R. Attanasio, Anthony Cocchi, Derek Lieber, Stephen Smith, Ton Ngo, John J. Barton, Susan Flynn Hummel, Janice C. Sheperd, and Mark Mergen. Implementing Jalapeño in Java. ACM SIG-PLAN Notices, 34(10):314–324, Octo-

                                        ber 1999. URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        320384/p314-alpern/.

Alpern:1999:IJN

[2407] Bowen Alpern, C. R. Attanasio, Anthony Cocchi, Derek Lieber, Stephen Smith, Ton Ngo, John J. Barton, Susan Flynn Hummel, Janice C. Sheperd, and Mark Mergen. Implementing jalapeño in Java. ACM SIGPLAN Notices, 34(10):314–324, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clarke:1999:SOD

[2408] Siobhán Clarke, William Harrison, Harold Ossher, and Peri Tarr. Subjectoriented design: towards improved alignment of requirements, design, and code. ACM SIGPLAN Notices, 34 (10):325–339, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p325-clarke/.

Kersten:1999:ACS

[2409] Mik Kersten and Gail C. Murphy. Atlas: a case study in building a Web-based learning environment using aspect-oriented programming. ACM SIGPLAN Notices, 34 (10):340–352, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p340-kersten/.

Kendall:1999:RMD

[2410] Elizabeth A. Kendall. Role model designs and implementations with aspect-oriented programming. ACM SIGPLAN Notices, 34(10):353–369, October 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        320384/p353-kendall/.

Stefanovic:1999:ABG

[2411] Darko Stefanović, Kathryn S. McKinley, and J. Eliot B. Moss. Agebased garbage collection. ACM SIGPLAN Notices, 34(10):370–381, October 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/oops/
                                        320384/p370-stefanovicacute/.

Hosking:1999:MCR

[2412] Antony L. Hosking and Jiawan Chen. Mostly-copying reachability-based orthogonal persistence. ACM SIGPLAN Notices, 34(10):382–398, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/320384/p382-hosking/.

Siek:1999:GGC

[2413] Jeremy G. Siek, Lie-Quan Lee, and Andrew Lumsdaine. The generic graph component library. ACM SIGPLAN Notices, 34(10):399–414, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        oops/320384/p399-siek/.

Evans:1999:ZCA

[2414] Huw Evans and Peter Dickman. Zones, contracts and absorbing changes: an approach to software evolution.

ACM SIGPLAN Notices, 34(10):415– 434, October 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p415-evans/.

Zorzo:1999:DOO

[2415] A. F. Zorzo and R. J. Stroud. A distributed object-oriented framework for dependable multiparty interactions. ACM SIGPLAN Notices, 34 (10):435–446, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p435-zorzo/.

Lipkind:1999:OVL

[2416] Ilya Lipkind, Igor Pechtchanski, and Vijay Karamcheti. Object views: language support for intelligent object caching in parallel and distributed computations. ACM SIGPLAN Notices, 34 (10):447–460, October 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/oops/
320384/p447-lipkind/.

Das:1999:SAL

[2417] Manuvir Das. Static analysis of large programs (invited talk) (abstract only): some experiences. ACM SIGPLAN Notices, 34(11):1, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/

                                        pepm/328690/p1-das/. Invited talk;

abstract only.

Braux:1999:TPE

[2418] Mathias Braux and Jacques Noyé. Towards partially evaluating reflection in Java. ACM SIGPLAN Notices, 34 (11):2–11, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/pepm/
                                        328690/p2-braux/.

Sumii:1999:OOP

[2419] Eijiro Sumii and Naoki Kobayashi. Online-and-offline partial evaluation (extended abstract): a mixed approach. ACM SIGPLAN Notices, 34 (11):12–21, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/pepm/
                                        328690/p12-sumii/. Extended ab-

stract.

Grobauer:1999:SFP

[2420] Bernd Grobauer and Zhe Yang. The second Futamura projection for type-directed partial evaluation. ACM SIGPLAN Notices, 34(11):22– 32, November 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/pepm/
                                        328690/p22-grobauer/.

Pfenning:1999:LFS

[2421] Frank Pfenning. On the logical foundations of staged computation (invited talk). ACM SIGPLAN Notices, 34 (11):33, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/pepm/
328690/p33-pfenning/. Invited talk.

Taha:1999:SRS

[2422] Walid Taha. A sound reduction semantics for untyped CBN multistage computation. Or, the theory of MetaML is non-trivial (extended abstract). ACM SIGPLAN Notices, 34 (11):34–43, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/pepm/
328690/p34-taha/. Extended ab-

stract.

Gallagher:1999:URA

[2423] John P. Gallagher and Julio C. Peralta. Using regular approximations for generalisation during partial evalution. ACM SIGPLAN Notices, 34 (11):44–51, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/pepm/
328690/p44-gallagher/.

Levi:1999:NPS

[2424] Giorgio Levi and Fausto Spoto. Non pair-sharing and freeness analysis through linear refinement. ACM SIGPLAN Notices, 34(11):52–61, November 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/pepm/
328690/p52-levi/.

Chin:1999:CST

[2425] Wei-Ngan Chin and Siau-Cheng Khoo. Calculating sized types. ACM SIGPLAN Notices, 34(11):62– 72, November 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/pepm/
                                        328690/p62-chin/.

Liu:1999:RIW

[2426] Yanhong A. Liu and Scott D. Stoller. From recursion to iteration: what are the optimizations? ACM SIGPLAN Notices, 34(11):73–82, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

                                        org/pubs/citations/proceedings/
                                        pepm/328690/p73-liu/.

Walpole:1999:OSS

[2427] Jonathan Walpole. Operating systems specialization (invited talk) (abstract only): experiences, opportunities and challenges. ACM SIGPLAN Notices, 34(11):83, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.acm.org/
                                        pubs/citations/proceedings/pepm/
                                        328690/p83-walpole/. Invited talk;

abstract only.

Kobayashi:1999:TBU

[2428] Naoki Kobayashi. Type-based useless variable elimination. ACM SIGPLAN Notices, 34(11):84–93, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.

org/pubs/citations/proceedings/
pepm/328690/p84-kobayashi/.

Murphy:1999:PAP

[2429] Brian R. Murphy and Monica S. Lam.

Program analysis with partial transfer functions. ACM SIGPLAN Notices, 34 (11):94–103, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/pepm/
328690/p94-murphy/.

Scholz:1999:SPA

[2430] Berhard Scholz, Johann Blieberger, and Thomas Fahringer. Symbolic pointer analysis for detecting memory leaks. ACM SIGPLAN Notices, 34 (11):104–113, November 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org/
pubs/citations/proceedings/pepm/
328690/p104-scholz/.

Cytron:1999:ALN

[2431] Ron K. Cytron. Activities: Letter from the Newly Elected Chair. ACM SIGPLAN Notices, 34(12):1, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/c/Cytron:
Ron.html.

Chalk:1999:CCJ

[2432] Peter Chalk. Conference corner: Java in the computing curricula. ACM SIG-PLAN Notices, 34(12):9–11, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/c/Chalk:Peter.
                                        html.

Chalk:1999:JCC

[2433] Peter Chalk. Java in the computing curricula. ACM SIGPLAN Notices, 34(12):9–11, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hind:1999:SAI

[2434] Michael Hind. SIGAda ’98: ACM/ SIGAda international conference. ACM SIGPLAN Notices, 34(12):12, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hind:1999:SAS

[2435] Michael Hind. SIGAda ’98: ACM/ SIGAda international conference (report). ACM SIGPLAN Notices, 34(12):12, December 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

                                        informatik.rwth-aachen.de/dblp/
                                        db/indices/a-tree/h/Hind:Michael.
                                        html.

McClain:1999:BAF

[2436] D. McClain. BlurFit: an application of functional programming to scientific analysis. ACM SIGPLAN Notices, 34(12):13–17, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McClain:1999:FPB

[2437] D. McClain. Functional programming: BlurFit — an application of functional programming to scientific analysis. ACM SIGPLAN Notices, 34(12):13–17, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/m/McClain:D=.
html.

Clark:1999:KSI

[2438] Chris Clark. Keywords: special identifier idioms. ACM SIGPLAN Notices, 34(12):18–23, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1999:PPPd

[2439] Chris Clark. Practical parsing patterns: Keywords: Special identifier idiom. ACM SIGPLAN Notices, 34(12):18–23, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/c/Clark:Chris.
html.

Metcalf:1999:FH

[2440] Michael Metcalf. Fortran 90/95/HPF.

ACM SIGPLAN Notices, 34(12):24– 29, December 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Metcalf:1999:RFI

[2441] Michael Metcalf. Resources: Fortran information file. ACM SIGPLAN Notices, 34(12):24–29, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/m/Metcalf:Michael.
html.

Eckert:1999:FFF

[2442] Brad Eckert. Forth: Firmware factory & Forth. ACM SIGPLAN Notices, 34(12):30–33, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/e/Eckert:Brad.
html.

Kiselyov:1999:TCC

[2443] Oleg Kiselyov and Richard J. Botting. Technical correspondence: Comment on “On the Application of a Popular Notation to Semantics” and reply from the author. ACM SIGPLAN Notices, 34(12):35–36, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/
db/indices/a-tree/b/Botting:Richard_
J=.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/k/Kiselyov:Oleg.html. See

[2320].

Normark:1999:PWW

[2444] Kurt Nørmark. Programming World Wide Web pages in Scheme. ACM SIG-PLAN Notices, 34(12):37–46, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/

dblp/db/indices/a-tree/n/N=oslash=
rmark:Kurt.html;   http://www.cs.
auc.dk/~normark/laml/. Describes

a literate programming environment, LAML (Lisp Abstracted Markup Language), for HTML dynamically generated from Scheme programs.

Dai:1999:ABO

[2445] Guilan Dai and Baowen Xu. An Ada-based object-oriented modeling language. ACM SIGPLAN Notices, 34(12):47–56, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Baowen:
Xu.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/d/Dai:Guilan.html.

Yingjun:1999:SPL

[2446] Li Yingjun and Lu Jian. SEIS++: A pattern language for seismic tools construction and integration. ACM SIG-PLAN Notices, 34(12):57–66, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/j/Jian:Lu.
html;   http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/a-tree/
y/Yingjun:Li.html.

Dai:1999:CRT

[2447] Guilan Dai and Baowen Xu. A comparison of real-time object-oriented modeling methods ROOM and OC-TOPUS. ACM SIGPLAN Notices, 34(12):67–71, December 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:/

/ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/b/Baowen:
Xu.html;   http://ftp.informatik.
rwth-aachen.de/dblp/db/indices/
a-tree/d/Dai:Guilan.html.

Curcio:1999:CCI

[2448] Igor D. D. Curcio. Correction to: Curcio, Igor D. D., “ASAP — A Simple Assertion Pre-Processor”, ACM Sigplan Notices 33(12), December 1998, pp. 44–

51. ACM SIGPLAN Notices, 34(12): 72, December 1999. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [2154].

Abbott:1998:DDO

[2449] Mark Abbott and Lalit Kumar Jain. DOVE: Distributed Objects based scientific Visualization Environment. In ACM [2891],

page ??  ISBN ????  LCCN

???? URL http://www.cs.ucsb.edu/
conferences/java98/papers/dove.
pdf;   http://www.cs.ucsb.edu/conferences/
java98/papers/dove.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Akarsu:1998:DJB

[2450] Erol Akarsu, Geoffrey Fox, and Tomasz Haupt. DARP: Java-based data analysis and rapid prototyping environment for distributed high performance computations. In ACM [2891],

page ??  ISBN ????  LCCN

???? URL http://www.cs.ucsb.edu/
conferences/java98/papers/darp.
pdf;   http://www.cs.ucsb.edu/conferences/

java98/papers/darp.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Baratloo:1998:INC

[2451] A. Baratloo, M. Karaul, H. Karl, and Z. M. Kedem. An infrastructure for network computing with Java applets. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/knittingfactory.pdf;   http:
//www.cs.ucsb.edu/conferences/java98/
papers/knittingfactory.ps. Also

published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Bik:1998:PBP

[2452] Aart J. C. Bik and Dennis B. Gannon. A prototype bytecode parallelization tool (javab manual). In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/javab.pdf;   http://www.
cs.ucsb.edu/conferences/java98/

papers/javab.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Bizarro:1998:JJL

[2453] Pedro Bizarro, Luís M. Silva, and João Gabriel Silva. JWarp: A Java library for parallel discrete-event simulations. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/jwarp.pdf;   http://www.
cs.ucsb.edu/conferences/java98/

papers/jwarp.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Boisvert:1998:DNL

[2454] Ronald F. Boisvert, Jack J. Dongarra, Roldan Pozo, Karin A. Remington, and G. W. Stewart. Developing numerical libraries in Java. In ACM [2891], page ?? ISBN ????

                                        LCCN ????  URL http://www.
                                        cs.ucsb.edu/conferences/java98/
                                        papers/jnt.pdf;   http://www.cs.
                                        ucsb.edu/conferences/java98/papers/

jnt.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Boyles:1998:CCE

[2455] Michael Boyles, Rajeev Raje, and Shiaofen Fang. CEV: Collaborative environment for visualization using Java-RMI. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

                                        cs.ucsb.edu/conferences/java98/
                                        papers/cev.pdf;   http://www.cs.
                                        ucsb.edu/conferences/java98/papers/

cev.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Breg:1998:JRP

[2456] Fabian Breg, Shridhar Diwan, Juan Villacis, Jayashree Balasubramanian, Esra Akman, and Dennis Gannon. Java RMI performance and object model interoperability: Experiments with Java/HPC++. In ACM [2891], page ?? ISBN ????

                                        LCCN ????  URL http://www.

cs.ucsb.edu/conferences/java98/
papers/rmi.pdf;   http://www.cs.
ucsb.edu/conferences/java98/papers/

rmi.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Caromel:1998:JFS

[2457] Denis Caromel and Julien Vayssiere.

A Java framework for seamless sequential, multi-threaded, and distributed programming. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/javapp.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/javapp.ps. Also published as

Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Carpenter:1998:HDP

[2458] Bryan Carpenter, Guansong Zhang, Geoffrey Fox, Xinying Li, and Yuhong Wen. HPJava: Data parallel extensions to Java. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/hpjava.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/hpjava.ps. Also published as

Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Chandy:1998:UAL

[2459] K. Mani Chandy, Adam Rifkin, and Eve Schooler. Using announce-listen with global events to develop distributed control systems. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/dcs.pdf;   http://www.cs.
ucsb.edu/conferences/java98/papers/

dcs.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Cowie:1998:JPD

[2460] James Cowie. JTED: Parallel discreteevent simulation in Java. In ACM [2891], page ?? ISBN ???? LCCN

???? URL http://www.cs.ucsb.edu/
conferences/java98/papers/jted.
pdf;   http://www.cs.ucsb.edu/conferences/
java98/papers/jted.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Demaine:1998:CCP

[2461] Erik D. Demaine. Converting C pointers to Java references. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/pointers.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/pointers.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Ferrari:1998:JNP

[2462] Adam J. Ferrari. JPVM: Network parallel computing in Java. In ACM [2891], page ?? ISBN ???? LCCN

???? URL http://www.cs.ucsb.edu/
conferences/java98/papers/jpvm.
pdf;   http://www.cs.ucsb.edu/conferences/
java98/papers/jpvm.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Getov:1998:HPP

[2463] Vladimir Getov, Susan Flynn-Hummel, and Sava Mintchev. High-performance parallel programming in Java: Exploiting native libraries. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/hpjavampi.pdf;   http://
www.cs.ucsb.edu/conferences/java98/
papers/hpjavampi.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Gladychev:1998:CRJ

[2464] Pavel Gladychev, Ahmed Patel, and Donal O’Mahony. Cracking RC5 with Java applets. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/rc5.pdf;   http://www.cs.
ucsb.edu/conferences/java98/papers/

rc5.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Glossner:1998:JSP

[2465] John Glossner, Jesse Thilo, and Stamatis Vassiliadis. Java signal processing: FFTs with bytecodes. In ACM [2891], page ?? ISBN ???? LCCN

???? URL http://www.cs.ucsb.edu/
conferences/java98/papers/jfft.
pdf;   http://www.cs.ucsb.edu/conferences/
java98/papers/jfft.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Gosling:1998:EJN

[2466] James Gosling. Extensions to Java for numerical computation. In ACM [2891], page ?? ISBN ???? LCCN

                                        ???? URL http://java.sun.com/
                                        people/jag/FP.html. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Gray:1998:NLB

[2467] Paul A. Gray and Vaidy S. Sunderam. Native-language-based distributed computing across network and filesystem boundaries. In ACM [2891], page ?? ISBN ???? LCCN

                                        ???? URL http://www.cs.ucsb.edu/
                                        conferences/java98/papers/icet.
                                        pdf;   http://www.cs.ucsb.edu/conferences/
                                        java98/papers/icet.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Hirano:1998:PEP

[2468] Satoshi Hirano, Yoshiji Yasu, and Hirotaka Igarashi. Performance evaluation of popular distributed object technologies for Java. In ACM [2891],

page ??  ISBN ????  LCCN

                                        ???? URL http://www.cs.ucsb.edu/
                                        conferences/java98/papers/dots.
                                        pdf;   http://www.cs.ucsb.edu/conferences/
                                        java98/papers/dots.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Jacob:1998:LSP

[2469] Matthias Jacob, Michael Philippsen, and Martin Karrenbach. Largescale parallel geophysical algorithms in Java: A feasibility study. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/veltran.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/veltran.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Judd:1998:DDO

[2470] Glenn Judd, Mark Clement, and Quinn Snell. DOGMA: Distributed Object Group Management Architecture. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/dogma.pdf;   http://www.
cs.ucsb.edu/conferences/java98/

papers/dogma.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Kahan:1998:HJF

[2471] William Kahan and University of California at Berkeley. How Java’s floatingpoint hurts everyone everywhere. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/javahurt.pdf. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Karl:1998:BGB

[2472] Holger Karl. Bridging the gap between distributed shared memory and message passing. In ACM [2891], page ?? ISBN ????

                                        LCCN ????  URL http://www.
                                        cs.ucsb.edu/conferences/java98/
                                        papers/bridging.pdf;   http://www.
                                        cs.ucsb.edu/conferences/java98/
                                        papers/bridging.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Keren:1998:APP

[2473] Arie Keren and Amnon Barak. Adaptive placement of parallel Java agents in a scalable computing cluster. In ACM [2891], page ?? ISBN ????

                                        LCCN ????  URL http://www.
                                        cs.ucsb.edu/conferences/java98/
                                        papers/agents.pdf;   http://www.
                                        cs.ucsb.edu/conferences/java98/
                                        papers/agents.ps. Also published as

Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Ki:1998:CSD

[2474] Byeongseob Ki and Scott Klasky. Collaborative scientific data visualization. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

                                        cs.ucsb.edu/conferences/java98/
                                        papers/scivis.pdf;   http://www.
                                        cs.ucsb.edu/conferences/java98/
                                        papers/scivis.ps. Also published as

Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Krall:1998:MEH

[2475] Andreas Krall and Mark Probst. Monitors and exceptions: How to implement Java efficiently. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/cacao.pdf;   http://www.
cs.ucsb.edu/conferences/java98/

papers/cacao.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Miller:1998:MVJ

[2476] Kevin Miller. Microsoft VM for Java.

In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/

program.html. Invited talk.

Sarmenta:1998:TBB

[2477] Luis F. G. Sarmenta, Satoshi Hirano, and Stephen A. Ward. Towards Bayanihan: Building an extensible framework for volunteer computing using Java. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/bayanihan.pdf;   http://
www.cs.ucsb.edu/conferences/java98/
papers/bayanihan.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Schwab:1998:AJC

[2478] Matthias Schwab and Joel Schroeder.

Algebraic Java classes for numerical optimization. In ACM [2891],

page ??  ISBN ????  LCCN

???? URL http://www.cs.ucsb.edu/

conferences/java98/papers/jest.
pdf;   http://www.cs.ucsb.edu/conferences/
java98/papers/jest.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Souder:1998:JCB

[2479] Dan Souder, Morgan Herrington, Rajat P. Garg, and Dennis DeRyke. JSPICE: A component-based distributed Java front-end for SPICE. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/jspice.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/jspice.ps. Also published as

Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Takagi:1998:NMP

[2480] Hiromitsu Takagi, Satoshi Matsuoka, Hidemoto Nakada, Satoshi Sekiguchi, Mitsuhisa Satoh, and Umpei Nagashima. Ninflet: a migratable parallel objects framework using Java. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/ninflet.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/ninflet.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Thiruvathukal:1998:RRM

[2481] George K. Thiruvathukal, Lovely S. Thomas, and Andy T. Korczynski. Reflective Remote Method Invocation. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/reflective.pdf;   http://
www.cs.ucsb.edu/conferences/java98/
papers/reflective.ps. Also pub-

lished as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

vanHalderen:1998:FWB

[2482] Berry A. W. van Halderen and Benno J. Overeinder. Fornax: Web-based distributed discrete event simulation in Java. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/fornax.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/fornax.ps. Also published as

Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Weber:1998:PCA

[2483] Andreas Weber, Wolfgang Küchlin, Bernhard Eggers, and Volker Simonis. Parallel computer algebra software as a Web component. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/algebra.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/algebra.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Yalamanchilli:1998:CPJ

[2484] Narendar Yalamanchilli and William Cohen. Communication performance of Java based Parallel Virtual Machines. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/passing.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/passing.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Yelick:1998:THP

[2485] Kathy Yelick, Luigi Semenzato, Geoff Pike, Carleton Miyamoto, Ben Liblit, Arvind Krishnamurthy, Paul Hilfinger, Susan Graham, David Gay, Phil Colella, and Alex Aiken. Titanium: A high-performance Java dialect. In ACM [2891], page ?? ISBN ???? LCCN ???? URL http://www.

cs.ucsb.edu/conferences/java98/
papers/titanium.pdf;   http://www.
cs.ucsb.edu/conferences/java98/
papers/titanium.ps. Also published

as Concurrency: Practice and Experience, 10(11–13), September 1998, CO-DEN CPEXEI, ISSN 1040-3108.

Zhu:1998:SUP

[2486] Jinsong Zhu, Maria Toeroe, Victor Leung, and Son Vuong. Supporting universal personal computing on Internet with Java and CORBA. In ACM [2891], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/upc.pdf;   http://www.cs.
ucsb.edu/conferences/java98/papers/

upc.ps. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

Drissel:1991:CSR

[2487] William E. Drissel. Common sense and real time executives. In Koopman, Jr. [2858], pages 1–6. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p1-drissel/. ACM order

number 817911.

Frenger:1991:OKV

[2488] Paul Frenger. A one key video graphics terminal. In Koopman, Jr. [2858], pages 7–12. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p7-frenger/. ACM order

number 817911.

Hand:1991:CFR

[2489] Tom Hand. Challenges for future realtime operating systems. In Koopman, Jr. [2858], pages 13–16. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p13-hand/. ACM order num-

ber 817911.

Hand:1991:DES

[2490] Tom Hand. Debugging embedded systems implemented in C. In Koopman, Jr. [2858], pages 17–22. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p17-hand/. ACM order num-

ber 817911.

Philip:1991:ECP

[2491] Philip Koopman, Jr. Embedded control as a path to Forth acceptance. In Koopman, Jr. [2858], pages 23–26. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p23-koopman/. ACM order

number 817911.

Lyons:1991:FMF

[2492] Ken Lyons. A fast method for finding an integer square root. In Koopman, Jr. [2858], pages 27–30. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p27-lyons/. ACM order num-

ber 817911.

Mikiten:1991:FRI

[2493] Brian C. Mikiten, Shawn Mikiten, and John L. Orr. A Forth-based real-time in-flight monitoring system. In Koopman, Jr. [2858], pages 31–34. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p31-mikiten/. ACM order

number 817911.

Morgenstern:1991:PF

[2494] Leonard Morgenstern. Pascal to Forth. In Koopman, Jr. [2858], pages 35–36. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p35-morgenstern/. ACM or-

der number 817911.

Orr:1991:AOI

[2495] John L. Orr and Brian C. Mikiten. The acoustic orientation instrument: real-time digital audio control with Forth. In Koopman, Jr. [2858], pages 37–44. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53 1991. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/259965/p37-orr/

. ACM order number 817911.

Pierce:1991:AFE

[2496] Albert Pierce and Elizabeth Pierce.

The application of Forth engines as coprocessors for the Macintosh computer. In Koopman, Jr. [2858], pages 45–48. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p45-pierce/. ACM order

number 817911.

Snow:1991:FDE

[2497] Paul A. Snow and K. Michael Parker. The fifth development environment. In Koopman, Jr. [2858], pages 49–52. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p49-snow/. ACM order num-

ber 817911.

VanWyck:1991:EAC

[2498] Don Van Wyck. The evolution of an access control system. In Koopman, Jr. [2858], pages 53–56. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p53-van_wyck/. ACM order

number 817911.

Wehrer:1991:LFP

[2499] Wayne J. Wehrer. Low-cost, fast, portable embedded system development. In Koopman, Jr. [2858], pages 57–62. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

                                        1991. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        259965/p57-wehrer/. ACM order

number 817911.

Bean:1991:MFW

[2500] Warren Bean. The marketing of Forth workshop report. In Koopman, Jr. [2858], pages 63–70. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

                                        1991. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        259965/p63-bean/. ACM order num-

ber 817911.

Wavrik:1991:UFC

[2501] John J. Wavrik. Use of Forth in a course in computer algebra. In Koopman, Jr. [2858], pages 71–80. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

                                        1991. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        259965/p71-wavrik/. ACM order

number 817911.

Johnsen:1991:SPP

[2502] Charles Johnsen and David L. Fox. The silicon palimpsest: a programming model for electrically reconfigurable processors. In Koopman, Jr. [2858], pages 81–94. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

                                        1991. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        259965/p81-johnsen/. ACM order

number 817911.

Brakefield:1991:CF

[2503] James C. Brakefield. Challenges for Forth. In Koopman, Jr. [2858], pages 95–98. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

                                        1991. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/

259965/p95-brakefield/. ACM or-

der number 817911.

Frenger:1991:FHN

[2504] Paul Frenger. A Forth-based hybrid neuron for neural nets. In Koopman, Jr. [2858], pages 99–102. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p99-frenger/. ACM order

number 817911.

Furman:1991:PF

[2505] Alan T. Furman. Pipes in Forth.

In Koopman, Jr. [2858], pages 103– 114. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53 1991. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/259965/p103-furman/

. ACM order number 817911.

Howerton:1991:FAW

[2506] Charles P. Howerton. Forth is alive and well and living in a hypercube somewhere in Wyoming. In Koopman, Jr. [2858], pages 115–120. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p115-howerton/. ACM order

number 817911.

Sergeant:1991:FPF

[2507] Frank Sergeant. Files in Pygmy Forth. In Koopman, Jr. [2858], pages 121–124. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

1991. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
259965/p121-sergeant/. ACM order

number 817911.

William:1991:SSR

[2508] George William Shaw, II. Sh-BOOM: the sound of the RISC market changing. In Koopman, Jr. [2858], page 125. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53

                                        1991. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        259965/p125-shaw/. ACM order num-

ber 817911.

Caffrey:1992:FSI

[2509] Robert T. Caffrey, J. Thomas Riley, E. Hilsenrath, S. Hynes, J. Hui, and S. Sjolander. Forth in space: interfacing SSBUV, a scientific instrument, to the Space Shuttle. In ACM [2861], pages 1–8. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199200/p1-caffrey/.

DiMeo:1992:FCE

[2510] Frank N. DiMeo. A Forth course for engineers. In ACM [2861], pages 9–11. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199200/p9-dimeo/.

Lee:1992:FSF

[2511] Yong M. Lee and Edward Conjura. First step to Forth engine construction. In ACM [2861], pages 12–16. ISBN ???? LCCN ???? URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/199200/p12-lee/

.

Sakharov:1992:CC

[2512] Alexander Sakharov. A C-to-Forth compiler. In ACM [2861], pages

17–18.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199200/p17-sakharov/.

Gassanenko:1992:AAS

[2513] Michael L. Gassanenko. Alternative approaches to standardization and portability for Forth. In ACM [2861], pages 27–31. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199200/p27-gassanenko/.

Field:1990:LOL

[2514] John Field. On laziness and optimality in lambda interpreters: tools for specification and analysis. In ACM [2853], pages 1–15. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p1-field/.

Lamping:1990:AOL

[2515] John Lamping. An algorithm for optimal lambda calculus reduction. In ACM [2853], pages

16–30.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p16-lamping/.

Abadi:1990:ES

[2516] M. Abadi, P. L. Curien, and J. J.

Levy. Explicit substitutions. In ACM [2853], pages 31–46. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p31-abadi/.

Griffin:1990:FNC

[2517] Timothy G. Griffin. A formulae-as-type notion of control. In ACM [2853], pages 47–58. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p47-griffin/.

Asperti:1990:IFP

[2518] Andrea Asperti, Gian Luigi Ferrari, and Roberto Gorrieri. Implicative formulae in the proofs of computations’ analogy. In ACM [2853], pages 59–71. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p59-asperti/.

Moschovakis:1990:CP

[2519] Yiannis N. Moschovakis. Computable process. In ACM [2853], pages 72–80. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p72-moschovakis/.

Berry:1990:CAM

[2520] Gerard Berry and Gerard Boudol. The chemical abstract machine. In ACM [2853], pages 81–94. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p81-berry/.

Lafont:1990:IN

[2521] Yves Lafont. Interaction nets. In ACM [2853], pages 95–108. ISBN

????  LCCN ????  URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/96709/p95-lafont/

.

Mitchell:1990:TTF

[2522] John C. Mitchell. Toward a typed foundation for method specialization and inheritance. In ACM [2853], pages 109–124. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p109-mitchell/.

Cook:1990:IS

[2523] William R. Cook, Walter Hill, and Peter S. Canning. Inheritance is not subtyping. In ACM [2853], pages

125–135.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p125-cook/.

Graver:1990:TSS

[2524] Justin O. Graver and Ralph E. Johnson. A type system for Smalltalk. In ACM [2853], pages 136–

150.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p136-graver/.

Burn:1990:RBA

[2525] Geoffrey Burn. A relationship between abstract interpretation and projection analysis. In ACM [2853], pages 151–156. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p151-burn/.

Deutsch:1990:DLA

[2526] Alan Deutsch. On determining lifetime and aliasing of dynamically allocated data in higher-order functional specifications. In ACM [2853], pages 157–168. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p157-deutsch/.

Sekar:1990:SDS

[2527] R. C. Sekar, Shaunak Pawagi, and I. V. Ramarkrishnan. Small domains spell fast strictness analysis. In ACM [2853], pages 169–183. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p169-sekar/.

Marlowe:1990:EHA

[2528] Thomas J. Marlowe and Barbara G. Ryder. An efficient hybrid algorithm for incremental data flow analysis. In ACM [2853], pages

184–196.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p184-marlowe/.

Heintze:1990:FPT

[2529] Nevin Heintze and Joxan Jaffar. A finite presentation theorem for approximating logic programs. In ACM [2853], pages 197–209. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p197-heintze/.

William:1990:PTP

[2530] John H. William, Alexander Aiken, and Edward L. Wimmers. Program transformation in the presence of errors. In ACM [2853], pages

210–217.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p210-william/.

Steele:1990:MAP

[2531] Guy L. Steele, Jr. Making asynchronous parallelism safe for the world. In ACM [2853], pages

218–231.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p218-steele/.

Saraswat:1990:CCP

[2532] Vijay A. Saraswat and Martin Rinard. Concurrent constraint programming. In ACM [2853], pages

232–245.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p232-saraswat/.

Ramakrishnan:1990:PLP

[2533] Raghu Ramakrishnan. Parallelism in logic programs. In ACM [2853], pages 246–260. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p246-ramakrishnan/.

Demmers:1990:CGC

[2534] Alan Demmers, Mark Weiser, Barry Hayes, Hans Boehm, Daniel Bobrow, and Scott Shenker. Combining generational and conservative garbage collection: framework and implementations. In ACM [2853], pages 261–269. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p261-demmers/.

Palem:1990:STI

[2535] Krishna Palem and Barbara Simons.

Scheduling time-critical instructions on RISC machines. In ACM [2853], pages 270–280. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p270-palem/.

Ramesh:1990:AIP

[2536] R. Ramesh, I. V. Ramakrishnan, and D. S. Warren. Automata-driven indexing of Prolog clauses. In ACM [2853], pages 281–291. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p281-ramesh/.

Attie:1990:FHM

[2537] Paul C. Attie, Nissim Francez, and Orna Grumberg. Fairness and hyperfairness in multi-party interactions. In ACM [2853], pages

292–305.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p292-attie/.

Gunter:1990:RTP

[2538] Carl A. Gunter. Relating total and partial correctness interpretations of non-deterministic programs. In ACM [2853], pages

306–319.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p306-gunter/.

Russell:1990:ONK

[2539] James R. Russell. On oraclizable networks and Kahn’s principle. In ACM [2853], pages 320–

328.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        96709/p320-russell/.

Stark:1990:RCC

[2540] Eugene W. Stark. On the relations computable by a class of concurrent automata. In ACM [2853], pages 329–340. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p329-stark/.

Harper:1990:HMP

[2541] Robert Harper, John C. Mitchell, and Eugenio Moggi. Higher-order modules and the phase distinction. In ACM [2853], pages 341–

354.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p341-harper/.

Rouaix:1990:SRO

[2542] François Rouaix. Safe run-time overloading. In ACM [2853], pages

355–366.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p355-rouaix/.

Thatte:1990:QT

[2543] Satish Thatte. Quasi-static typing. In ACM [2853], pages 367–

381.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p367-thatte/.

Mairson:1990:DMT

[2544] Harry G. Mairson. Deciding ML typability is complete for deterministic exponential time. In ACM [2853], pages 382–401. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
96709/p382-mairson/.

Sundaresh:1991:TIC

[2545] R. S. Sundaresh and Paul Hudak.

A theory of incremental computation and its application. In ACM [2856], pages 1–13. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p1-sundaresh/.

Consel:1991:SDS

[2546] Charles Consel and Olivier Danvy. Static and dynamic semantics processing. In ACM [2856], pages

14–24.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p14-consel/.

Odersky:1991:HMD

[2547] Martin Odersky. How to make destructive updates less destructive. In ACM [2856], pages 25–36. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p25-odersky/.

Sekar:1991:PLS

[2548] R. C. Sekar, Prateek Mishra, and I. V. Ramakrishnan. On the power and limitation of strictness analysis based on abstract interpretation. In ACM [2856], pages 37–48. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p37-sekar/.

Abramsky:1991:RAS

[2549] Samson Abramsky and Thomas P. Jensen. A relational approach to strictness analysis for higher-order polymorphic functions. In ACM [2856], pages 49–54. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p49-abramsky/.

Ferrante:1991:ACS

[2550] J. Ferrante. Automatic construction of sparse data flow evaluation graphs. In ACM [2856], pages

55–66.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p55-ferrante/.

Pingali:1991:DFG

[2551] Keshav Pingali, Micah Beck, Richard Johnson, Mayan Moudgill, and Paul Stodghill. Dependence flow graphs: an algebraic approach to program dependencies. In ACM [2856], pages 67–78. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p67-pingali/.

Pinter:1991:POP

[2552] Shlomit S. Pinter and Ron Y. Pinter.

Program optimization and parallelization using idioms. In ACM [2856], pages 79–92. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p79-pinter/.

Landi:1991:PAP

[2553] William Landi and Barbara G. Ryder. Pointer-induced aliasing: a problem taxonomy. In ACM [2856], pages 93–103. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p93-landi/.

Amadio:1991:SRT

[2554] Roberto M. Amadio and Luca Cardelli.

Subtyping recursive types. In ACM [2856], pages 104–118. ISBN

????  LCCN ????  URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/99583/p104-amadio/

.

Henglein:1991:CTI

[2555] Fritz Henglein and Harry G. Mairson. The complexity of type inference for higher-order lambda calculi. In ACM [2856], pages 119–

130.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p119-henglein/.

Harper:1991:RCB

[2556] Robert Harper and Benjamin Pierce. A record calculus based on symmetric concatenation. In ACM [2856], pages 131–142. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p131-harper/.

Cai:1991:LMN

[2557] Jiazhen Cai and Robert A. Paige. Look ma, no hashing, and no arrays neither. In ACM [2856], pages 143–154. ISBN ???? LCCN ???? URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/99583/p143-cai/

.

Clinger:1991:MW

[2558] William Clinger. Macros that work. In ACM [2856], pages

155–162.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p155-clinger/.

Duba:1991:TFC

[2559] Bruce Duba, Robert Harper, and David MacQueen. Typing first-class continuations in ML. In ACM [2856], pages 163–173. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p163-duba/.

Queinnec:1991:DEC

[2560] Christian Queinnec and Bernard Serpette. A dynamic extent control operator for partial continuations. In ACM [2856], pages

174–184.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p174-queinnec/.

Sitaram:1991:MCC

[2561] Dorai Sitaram and Matthias Felleisen. Models of continuations without continuations. In ACM [2856], pages 185–196. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p185-sitaram/.

Lucco:1991:PPC

[2562] Steven Lucco and Oliver Sharp. Parallel programming with coordination structures. In ACM [2856], pages

197–208.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p197-lucco/.

Joung:1991:CFM

[2563] Yuh-Jzer Joung and Scott A. Smolka.

Coordinating first-order multiparty interactions. In ACM [2856], pages

209–220.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p209-joung/.

Kleinman:1991:CDL

[2564] Alon Kleinman, Yael Moscowitz, Amir Pnueli, and Ehud Sharpio. Communication with directed logic variables. In ACM [2856], pages

221–232.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        99583/p221-kleinman/.

Crank:1991:PLC

[2565] Erik Crank and Matthias Felleisen. Parameter-passing and the lambda calculus. In ACM [2856], pages

233–244.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        99583/p233-crank/.

Riecke:1991:FAT

[2566] Jon G. Riecke. Fully abstract translations between functional languages. In ACM [2856], pages

245–254.  ISBN ????  LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        99583/p245-riecke/.

Maranget:1991:ODW

[2567] Luc Maranget. Optimal derivations in weak lambda-calculi and in orthogonal term rewriting systems. In ACM [2856], pages 255–269. ISBN ???? LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        99583/p255-maranget/.

Mitchell:1991:ESM

[2568] John Mitchell, Sigurd Meldal, and Neel Madhav. An extension of standard ML modules with subtyping and inheritance. In ACM [2856], pages 270–278. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p270-mitchell/.

Aiken:1991:STI

[2569] Alex Aiken and Brian Murphy. Static type inference in a dynamically typed language. In ACM [2856], pages

279–290.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p279-aiken/.

Leroy:1991:PTI

[2570] Xavier Leroy and Pierre Weis. Polymorphic type inference and assignment. In ACM [2856], pages

291–302.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p291-leroy/.

Jouvelot:1991:ART

[2571] Pierre Jouvelot and David Gifford. Algebraic reconstruction of types and effects. In ACM [2856], pages

303–310.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p303-jouvelot/.

Pratt:1991:MCG

[2572] Vaughn Pratt. Modeling concurrency with geometry. In ACM [2856], pages 311–322. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p311-pratt/.

Abadi:1991:LVC

[2573] Martin Abadi and Gordon Plotkin.

A logical view of composition and refinement. In ACM [2856], pages

323–332.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p323-abadi/.

Saraswat:1991:SFC

[2574] Vijay A. Saraswat, Martin Rinard, and Prakash Panangaden. The semantic foundations of concurrent constraint programming. In ACM [2856], pages 333–352. ISBN ???? LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p333-saraswat/.

Henzinger:1991:TPM

[2575] Tom Henzinger, Zohar Manna, and Amir Pnueli. Temporal proof methodologies for real-time systems. In ACM [2856], pages

353–366.  ISBN ????  LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
99583/p353-henzinger/.

Wadler:1992:EFP

[2576] Philip Wadler. The essence of functional programming. In ACM [2860], pages 1–14. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p1-wadler/. ACM order num-

ber 54990.

Gonthier:1992:GOL

[2577] Georges Gonthier, Martín Abadi, and Jean-Jacques Lévy. The geometry of optimal lambda reduction. In ACM [2860], pages 15–26. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/

pubs/citations/proceedings/plan/
143165/p15-gonthier/. ACM order

number 54990.

Filinski:1992:LC

[2578] Andrzej Filinski. Linear continuations.

In ACM [2860], pages 27–38. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p27-filinski/. ACM order

number 54990.

Lang:1992:GCW

[2579] Bernard Lang, Christian Queinnec, and José Piquer. Garbage collecting the world. In ACM [2860], pages 39–50. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p39-lang/. ACM order num-

ber 54990.

Edelson:1992:MCC

[2580] Daniel R. Edelson. A mark-and-sweep collector C++. In ACM [2860], pages 51–58. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p51-edelson/. ACM order

number 54990.

Ball:1992:OPT

[2581] Thomas Ball and James R. Larus. Optimally profiling and tracing programs. In ACM [2860], pages 59–70. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p59-ball/. ACM order num-

ber 54990.

Nielson:1992:BFP

[2582] Hanne Riis Nielson and Flemming Nielson. Bounded fixed point iteration. In ACM [2860], pages 71–82. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p71-nielson/. ACM order

number 54990.

Cousot:1992:IDS

[2583] Patrick Cousot and Radhia Cousot. Inductive definitions, semantics and abstract interpretations. In ACM [2860], pages 83–94. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p83-cousot/. ACM order

number 54990.

Barbuti:1992:MPC

[2584] Roberto Barbuti, Michael Codish, Roberto Giacobazzi, and Giorgio Levi. Modelling Prolog control. In ACM [2860], pages 95–104. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p95-barbuti/. ACM order

number 54990.

Rinard:1992:SFJ

[2585] Martin C. Rinard and Monica S. Lam. Semantic foundations of Jade. In ACM [2860], pages 105–118. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p105-rinard/. ACM order

number 54990.

Berry:1992:SMC

[2586] Dave Berry, Robin Milner, and David N. Turner. A semantics for ML concurrency primitives. In ACM [2860], pages 119–129. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p119-berry/. ACM order

number 54990.

Chow:1992:CAP

[2587] Jyh-Herng Chow and William Ludwell Harrison, III. Compile-time analysis of parallel programs that share memory. In ACM [2860], pages 130–141. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p130-chow/. ACM order num-

ber 54990.

Joung:1992:CSC

[2588] Yuh-Jzer Joung and Scott A. Smolka.

A comprehensive study of the complexity of multiparty interaction. In ACM [2860], pages 142–153. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p142-joung/. ACM order

number 54990.

Ohori:1992:CMM

[2589] Atsushi Ohori. A compilation method for ML-style polymorphic record calculi. In ACM [2860], pages 154– 165. ISBN 0-89791-453-8. LCCN QA76.7 .A15 1992. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/143165/p154-ohori/

. ACM order number 54990.

Remy:1992:TRC

[2590] Didier Rémy. Typing record concatenation for free. In ACM [2860], pages 166–176. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p166-remy/. ACM order num-

ber 54990.

Leroy:1992:UOP

[2591] Xavier Leroy. Unboxed objects and polymorphic typing. In ACM [2860], pages 177–188. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p177-leroy/. ACM order

number 54990.

Tofte:1992:PSH

[2592] Mads Tofte. Principal signatures for higher-order program modules. In ACM [2860], pages 189–199. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p189-tofte/. ACM order

number 54990.

Cosmo:1992:TIT

[2593] Robero Di Cosmo. Type isomorphisms in a type-assignment framework. In ACM [2860], pages 200–210. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p200-cosmo/. ACM order

number 54990.

Farnum:1992:PTA

[2594] Charles Farnum. Pattern-based tree attribution. In ACM [2860], pages 211–222. ISBN 0-89791-453-8. LCCN QA76.7 .A15 1992. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/143165/p211-farnum/

. ACM order number 54990.

Farrow:1992:CAG

[2595] R. Farrow, T. J. Marlowe, and D. M.

Yellin. Composable attribute grammars: support for modularity in translator design and implementation. In ACM [2860], pages 223–234. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p223-farrow/. ACM order

number 54990.

Bates:1992:RSL

[2596] Joseph Bates and Alon Lavie. Recognizing substrings of LR(k) languages in linear time. In ACM [2860], pages 235–245. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p235-bates/. ACM order

number 54990.

Gupta:1992:GDP

[2597] Rajiv Gupta. Generalized dominators and post-dominators. In ACM [2860], pages 246–257. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p246-gupta/. ACM order

number 54990.

Jorgensen:1992:GCL

[2598] Jesper Jørgensen. Generating a compiler for a lazy language by partial evaluation. In ACM [2860], pages 258– 268. ISBN 0-89791-453-8. LCCN QA76.7 .A15 1992. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/143165/p258-jorgensen/

. ACM order number 54990.

Nirkhe:1992:PEH

[2599] Vivek Nirkhe and William Pugh. Partial evaluation of high-level imperative programming languages with applications in hard real-time systems. In ACM [2860], pages 269–280. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p269-nirkhe/. ACM order

number 54990.

Ma:1992:PS

[2600] QingMing Ma. Parametricity as subtyping. In ACM [2860], pages 281–

292. ISBN 0-89791-453-8. LCCN QA76.7 .A15 1992. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/143165/p281-ma/

. ACM order number 54990.

Lincoln:1992:AAT

[2601] Patrick Lincoln and John C. Mitchell. Algorithmic aspects of type inference with subtypes. In ACM [2860], pages 293–304. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        143165/p293-lincoln/. ACM order

number 54990.

Pierce:1992:BQU

[2602] Benjamin C. Pierce. Bounded quantification is undecidable. In ACM [2860], pages 305–315. ISBN 0-89791-453-8. LCCN QA76.7 .A15

                                        1992. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/

143165/p305-pierce/. ACM order

number 54990.

Bruce:1992:PMS

[2603] Kim Bruce and John C. Mitchell. PER models of subtyping, recursive types and higher-order polymorphism. In ACM [2860], pages 316–327. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p316-bruce/. ACM order

number 54990.

Cartwright:1992:OSF

[2604] Robert Cartwright and Matthias Felleisen. Observable sequentiality and full abstraction. In ACM [2860], pages 328–342. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p328-cartwright/. ACM or-

der number 54990.

Clarke:1992:MCA

[2605] Edmund M. Clarke, Orna Grumberg, and David E. Long. Model checking and abstraction. In ACM [2860], pages 343–354. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
143165/p343-clarke/. ACM order

number 54990.

Jagadeesan:1992:ASH

[2606] Radha Jagadeesan and Keshav Pingali. Abstract semantics for a higherorder functional language with logic variables. In ACM [2860], pages 355– 366. ISBN 0-89791-453-8. LCCN QA76.7 .A15 1992. URL http:/

/www.acm.org:80/pubs/citations/

                                        proceedings/plan/143165/p355-jagadeesan/

. ACM order number 54990.

Snir:1993:CAP

[2607] Marc Snir. Computer architectures and programming models for scalable parallel computing (abstract). In ACM [2863], page 1. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p1-snir/. ACM order number

549930.

Maydan:1993:AFA

[2608] Dror E. Maydan, Saman P. Amarasinghe, and Monica S. Lam. Arraydata flow analysis and its use in array privatization. In ACM [2863], pages 2–15. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p2-maydan/. ACM order num-

ber 549930.

Chatterjee:1993:AAA

[2609] Siddhartha Chatterjee, John R. Gilbert, Robert Schreiber, and Shang-Hua Teng. Automatic array alignment in data-parallel programs. In ACM [2863], pages 16–28. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p16-chatterjee/. ACM or-

der number 549930.

Ning:1993:NFR

[2610] Qi Ning and Guang R. Gao. A novel framework of register allocation for software pipelining. In ACM [2863], pages 29–42. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p29-ning/. ACM order num-

ber 549930.

Odersky:1993:CNA

[2611] Martin Odersky, Dan Rabin, and Paul Hudak. Call by name, assignment, and the lambda calculus. In ACM [2863], pages 43–56. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p43-odersky/. ACM order

number 549930.

Weeks:1993:OAP

[2612] Stephen Weeks and Matthias Felleisen.

On the orthogonality of assignments and procedures in Algol. In ACM [2863], pages 57–70. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p57-weeks/. ACM order num-

ber 549930.

Jones:1993:IFP

[2613] Simon L. Peyton Jones and Philip Wadler. Imperative functional programming. In ACM [2863], pages 71–84. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p71-peyton_jones/. ACM

order number 549930.

Berry:1993:CRP

[2614] G. Berry, S. Ramesh, and R. K. Shyamasundar. Communicating reactive processes. In ACM [2863], pages 85–98. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p85-berry/. ACM order num-

ber 549930.

Ohori:1993:SCP

[2615] Atsushi Ohori and Kazuhiko Kato. Semantics for communication primitives in a polymorphic language. In ACM [2863], pages 99–112. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p99-ohori/. ACM order num-

ber 549930.

Doligez:1993:CGG

[2616] Damien Doligez and Xavier Leroy. A concurrent, generational garbage collector for a multithreaded implementation of ML. In ACM [2863], pages 113–123. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p113-doligez/. ACM order

number 549930.

Lawall:1993:SSC

[2617] Julia L. Lawall and Olivier Danvy. Separating stages in the continuationpassing style transformation. In ACM [2863], pages 124–136. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p124-lawall/. ACM order

number 549930.

Wand:1993:SCB

[2618] Mitchell Wand. Specifying the correctness of binding-time analysis. In ACM [2863], pages 137–143. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p137-wand/. ACM order num-

ber 549930.

Launchbury:1993:NSL

[2619] John Launchbury. A natural semantics for lazy evaluation. In ACM [2863], pages 144–154. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p144-launchbury/. ACM or-

der number 549930.

Reynolds:1993:ILR

[2620] John C. Reynolds. An introduction to logical relations and parametric polymorphism (abstract). In ACM [2863], pages 155–156. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p155-reynolds/. ACM order

number 549930.

Abadi:1993:FPP

[2621] Martín Abadi, Luca Cardelli, and Pierre-Louis Curien. Formal parametric polymorphism. In ACM [2863], pages 157–170. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p157-abadi/. ACM order

number 549930.

OHearn:1993:RPL

[2622] P. W. O’Hearn and R. D. Tennent. Relational parametricity and local variables. In ACM [2863], pages 171–184. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/158511/p171-o_

hearn/. ACM order number 549930.

Riecke:1993:ARC

[2623] Jon G. Riecke and Ramesh Subrahmanyam. Algebraic reasoning and completeness in typed languages. In ACM [2863], pages 185–195. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p185-riecke/. ACM order

number 549930.

Klarlund:1993:GT

[2624] Nils Klarlund and Michael I. Schwartzbach. Graph types. In ACM [2863], pages 196–205. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p196-klarlund/. ACM order

number 549930.

Harper:1993:EPC

[2625] Robert Harper and Mark Lillibridge. Explicit polymorphism and CPS conversion. In ACM [2863], pages 206–219. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p206-harper/. ACM order

number 549930.

Leroy:1993:PNR

[2626] Xavier Leroy. Polymorphism by name for references and continuations. In ACM [2863], pages 220–231. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/158511/p220-leroy/

. ACM order number 549930.

Choi:1993:EFI

[2627] Jong-Deok Choi, Michael Burke, and Paul Carini. Efficient flowsensitive interprocedural computation of pointer-induced aliases and side effects. In ACM [2863], pages 232–245. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p232-choi/. ACM order num-

ber 549930.

Yi:1993:AGM

[2628] Kwangkeun Yi and Williams Ludwell Harrison, III. Automatic generation and management of interprocedural program analyses. In ACM [2863], pages 246–259. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p246-yi/. ACM order number

549930.

Srinivasan:1993:SSA

[2629] Harini Srinivasan, James Hook, and Michael Wolfe. Static single assignment for explicitly parallel programs. In ACM [2863], pages 260–272. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p260-srinivasan/. ACM or-

der number 549930.

Lakhotia:1993:CCM

[2630] Arun Lakhotia. Constructing call multigraphs using dependence graphs. In ACM [2863], pages 273–284. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/158511/p273-lakhotia/

. ACM order number 549930.

Bruce:1993:STC

[2631] Kim B. Bruce. Safe type checking in a statically-typed object-oriented programming language. In ACM [2863], pages 285–298. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p285-bruce/. ACM order

number 549930.

Pierce:1993:OPR

[2632] Benjamin C. Pierce and David N.

Turner. Object-oriented programming without recursive types. In ACM [2863], pages 299–312. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p299-pierce/. ACM order

number 549930.

Mairson:1993:CLM

[2633] Harry G. Mairson. A constructive logic of multiple subtyping. In ACM [2863], pages 313–324. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p313-mairson/. ACM order

number 549930.

Leivant:1993:SFP

[2634] Daniel Leivant. Stratified functional programs and computational complexity. In ACM [2863], pages 325–333. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/158511/p325-leivant/

. ACM order number 549930.

Marriott:1993:ROC

[2635] Kimball G. Marriott and Peter J.

Stuckey. The 3 R’s of optimizing constraint logic programs: refinement, removal and reordering. In ACM [2863], pages 334–344. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/

                                        pubs/citations/proceedings/plan/
                                        158511/p334-marriott/. ACM order

number 549930.

VillemontedelaClergerie:1993:LSI

[2636] E. Villemonte de la Clergerie. Layer sharing: an improved structure-sharing framework. In ACM [2863], pages 345– 358. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/158511/p345-de_

la_clergerie/. ACM order number 549930.

Bossi:1993:DLP

[2637] A. Bossi, M. Bugliesi, M. Gabbrielli, G. Levi, and M. C. Meo. Differential logic programming. In ACM [2863], pages 359–370. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p359-bossi/. ACM order

number 549930.

Adl–Tabatabai:1993:EVI

[2638] Ali-Reza Adl-Tabatabai and Thomas Gross. Evicted variables and the interaction of global register allocation and symbolic debugging. In ACM [2863], pages 371–383. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p371-adl-tabatabai/. ACM

order number 549930.

Bates:1993:IPT

[2639] Samuel Bates and Susan Horwitz. Incremental program testing using program dependence graphs. In ACM [2863], pages 384–396. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p384-bates/. ACM order

number 549930.

Dhamdhere:1993:CBD

[2640] Dhananjay M. Dhamdhere and Uday P.

Khedker. Complexity of bi-directional data flow analysis. In ACM [2863], pages 397–408. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p397-dhamdhere/. ACM or-

der number 549930.

Nipkow:1993:TCT

[2641] Tobias Nipkow and Christian Prehofer.

Type checking type classes. In ACM [2863], pages 409–418. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p409-nipkow/. ACM order

number 549930.

Kozen:1993:ERS

[2642] Dexter Kozen, Jens Palsberg, and Michael I. Schwartzbach. Efficient recursive subtyping. In ACM [2863], pages 419–428. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p419-kozen/. ACM order

number 549930.

Gay:1993:SIA

[2643] Simon J. Gay. A sort inference algorithm for the polyadic π-calculus. In ACM [2863], pages 429–438. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/158511/p429-gay/

. ACM order number 549930.

Shao:1993:SR

[2644] Zhong Shao and Andrew W. Ael. Smartest recompilation. In ACM [2863], pages 439–450. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        158511/p439-shao/. ACM order num-

ber 549930.

Codish:1993:CAM

[2645] Michael Codish, Saumya K. Debray, and Roberto Giacobazzi. Compositional analysis of modular logic programs. In ACM [2863], pages 451–464. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/158511/p451-codish/

. ACM order number 549930.

Aponte:1993:ERT

[2646] Maria Virginia Aponte. Extending record typing to type parametric modules with sharing. In ACM [2863], pages 465–478. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

                                        1993. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/

158511/p465-aponte/. ACM order

number 549930.

Lee:1993:QSS

[2647] Shinn-Der Lee and Daniel P. Friedman. Quasi-static scoping: sharing variable bindings across multiple lexical scopes. In ACM [2863], pages 479–492. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p479-lee/. ACM order num-

ber 549930.

Consel:1993:TNP

[2648] Charles Consel and Olivier Danvy. Tutorial notes on partial evaluation. In ACM [2863], pages 493–501. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/158511/p493-consel/

. ACM order number 549930.

Ramalingam:1993:CBI

[2649] G. Ramalingam and Thomas Reps.

A categorized bibliography on incremental computation. In ACM [2863], pages 502–510. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
158511/p502-ramalingam/. ACM or-

der number 549930.

Diwan:1994:MSP

[2650] Amer Diwan, David Tarditi, and Eliot Moss. Memory subsystem performance of programs using copying garbage collection. In ACM [2870], pages 1–14. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p1-diwan/.

Ramsey:1994:CTB

[2651] Norman Ramsey. Correctness of trapbased breakpoint implementations. In ACM [2870], pages 15–24. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p15-ramsey/.

Agrawal:1994:DSB

[2652] Hiralal Agrawal. Dominators, super blocks, and program coverage. In ACM [2870], pages 25–34. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p25-agrawal/.

Garrigue:1994:TPL

[2653] Jacques Garrigue and Hassan Aït-Kaci. The typed polymorphic labelselective [lambda]-calculus. In ACM [2870], pages 35–47. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p35-garrigue/.

Odersky:1994:FTL

[2654] Martin Odersky. A functional theory of local names. In ACM [2870], pages 48–59. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p48-odersky/.

Lescanne:1994:LLJ

[2655] Pierre Lescanne. From [lambda][sigma] to [lambda]v: a journey through calculi of explicit substitutions. In ACM [2870], pages 60–69. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p60-lescanne/.

Doligez:1994:PUG

[2656] Damien Doligez and Georges Gonthier.

Portable, unobtrusive garbage collection for multiprocessor systems. In ACM [2870], pages 70–83. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p70-doligez/.

Nielson:1994:HCP

[2657] Hanne Riis Nielson and Flemming Nielson. Higher-order concurrent programs with finite communication topology (extended abstract). In ACM [2870], pages 84–97. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p84-nielson/.

deBoer:1994:PCC

[2658] Frank S. de Boer, Maurizio Gabbrielli, Elena Marchiori, and Catuscia Palamidessi. Proving concurrent constraint programs correct. In ACM [2870], pages 98–108. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p98-de_boer/.

Leroy:1994:MTM

[2659] Xavier Leroy. Manifest types, modules, and separate compilation. In ACM [2870], pages 109–122. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p109-leroy/.

Harper:1994:TAH

[2660] Robert Harper and Mark Lillibridge. A type-theoretic approach to higherorder modules with sharing. In ACM [2870], pages 123–137. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p123-harper/.

Katiyar:1994:TSP

[2661] Dinesh Katiyar, David Luckham, and John Mitchell. A type system for prototyping languages. In ACM [2870], pages 138–150. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p138-katiyar/.

Castagna:1994:DBQ

[2662] Giusee Castagna and Benjamin C. Pierce. Decidable bounded quantification. In ACM [2870], pages 151–

162. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/174675/p151-castagna/

.

Aiken:1994:STC

[2663] Alexander Aiken, Edward L. Wimmers, and T. K. Lakshman. Soft typing with conditional types. In ACM [2870], pages 163–173. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p163-aiken/.

Thatte:1994:ASI

[2664] Satish R. Thatté. Automated synthesis of interface adapters for reusable classes. In ACM [2870], pages 174– 187. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/174675/p174-thatte/

.

Tofte:1994:ITC

[2665] Mads Tofte and Jean-Pierre Talpin.

Implementation of the typed call-by-value [lambda]-calculus using a stack of regions. In ACM [2870], pages 188–201. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/174675/p188-tofte/

.

Hankin:1994:DAT

[2666] Chris Hankin and Daniel Le Métayer.

Deriving algorithms from type inference systems: application to strictness analysis. In ACM [2870], pages 202–212. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/174675/p202-hankin/

.

Henglein:1994:FOB

[2667] Fritz Henglein and Jesper Jørgensen.

Formally optimal boxing. In ACM [2870], pages 213–226. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p213-henglein/.

Cortesi:1994:CAD

[2668] Agostino Cortesi, Baudouin Le Charlier, and Pascal Van Hentenryck. Combinations of abstract domains for logic programming. In ACM [2870], pages 227–239. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p227-cortesi/.

Marriott:1994:ALP

[2669] Kim Marriott, María José García de la Banda, and Manuel Hermenegildo. Analyzing logic programs with dynamic scheduling. In ACM [2870], pages 240–253. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/174675/p240-marriott/

.

Qian:1994:HEL

[2670] Zhenyu Qian. Higher-order equational logic programming. In ACM [2870], pages 254–267. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p254-qian/.

Antoy:1994:NNS

[2671] Sergio Antoy, Rachid Echahed, and Michael Hanus. A needed narrowing strategy. In ACM [2870], pages 268–279. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/174675/p268-antoy/

.

Proebsting:1994:DPS

[2672] Todd A. Proebsting and Christopher W. Fraser. Detecting pipeline structural hazards quickly. In ACM [2870], pages 280–286. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p280-proebsting/.

Ramalingam:1994:IAM

[2673] G. Ramalingam and Thomas Reps.

An incremental algorithm for maintaining the dominator tree of a reducible flowgraph. In ACM [2870], pages 287–296. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p287-ramalingam/.

Weise:1994:VDG

[2674] Daniel Weise, Roger F. Crew, Michael Ernst, and Bjarne Steensgaard. Value dependence graphs: representation without taxation. In ACM [2870], pages 297–310. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p297-weise/.

Maslov:1994:LAD

[2675] Vadim Maslov. Lazy array dataflow dependence analysis. In ACM [2870], pages 311–325. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p311-maslov/.

Cleaveland:1994:OFV

[2676] Rance Cleaveland and Daniel Yankelevich. An operational framework for value-passing processes. In ACM [2870], pages 326–338. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p326-cleaveland/.

Bloom:1994:CCH

[2677] Bard Bloom. CHOCOLATE: Calculi of Higher Order COmmunication and LAmbda TErms (preliminary report). In ACM [2870], pages 339–347. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p339-bloom/.

Honda:1994:CRM

[2678] Kohei Honda and Nobuko Yoshida. Combinatory representation of mobile processes. In ACM [2870], pages 348– 360. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/174675/p348-honda/

.

Paakki:1994:MEF

[2679] Jukka Paakki. Multi-pass execution of functional logic programs. In ACM [2870], pages 361–374. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        174675/p361-paakki/.

Boyland:1994:CTA

[2680] John Boyland and Susan L. Graham. Composing tree attributions. In ACM [2870], pages 375–388. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p375-boyland/.

Muller:1994:SCA

[2681] Robert Muller. A staging calculus and its application to the verification of translators (preliminary report). In ACM [2870], pages 389–396. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p389-muller/.

Calder:1994:RIF

[2682] Brad Calder and Dirk Grunwald. Reducing indirect function call overhead in C++ programs. In ACM [2870], pages 397–408. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p397-calder/.

DeBosschere:1994:CFS

[2683] Koen De Bosschere, Saumya Debray, David Gudeman, and Sampath Kannan. Call forwarding: a simple interprocedural optimization technique for dynamically typed languages. In ACM [2870], pages 409–420. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p409-de_bosschere/.

Feigen:1994:RT

[2684] Lawrence Feigen, David Klaholz, Robert Casazza, and Xing Xue. The revival transformation. In ACM [2870], pages 421–434. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p421-feigen/.

Wand:1994:SLC

[2685] Mitchell Wand and Paul Steckler. Selective and lightweight closure conversion. In ACM [2870], pages 435–

445. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/174675/p435-wand/

.

Filinski:1994:RM

[2686] Andrzej Filinski. Representing monads. In ACM [2870], pages 446–

457. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/174675/p446-filinski/

.

Hatcliff:1994:GAC

[2687] John Hatcliff and Olivier Danvy. A generic account of continuation-passing styles. In ACM [2870], pages 458–

471. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/174675/p458-hatcliff/

.

L:1994:BIC

[2688] Guy L. Steele, Jr. Building interpreters by composing monads. In ACM [2870], pages 472–492. ISBN 0-89791-636-0. LCCN QA76.7 .A15

1994. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
174675/p472-steele/.

Riecke:1995:ISE

[2689] Jon G. Riecke and Ramesh Viswanathan.

Isolating side effects in sequential languages. In ACM [2877], pages 1– 12. ISBN 0-89791-692-1. LCCN QA 76.7 A11 1995. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/199448/p1-riecke/

. ACM order number: 549950.

Brookes:1995:SAD

[2690] Stephen Brookes and Denis Dancanet. Sequential algorithms, deterministic parallelism, and intensional expressiveness. In ACM [2877], pages 13–24. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p13-brookes/. ACM order

number: 549950.

Reynolds:1995:UFC

[2691] John C. Reynolds. Using functor categories to generate intermediate code. In ACM [2877], pages 25–36. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p25-reynolds/. ACM order

number: 549950.

Duesterwald:1995:DCI

[2692] Evelyn Duesterwald, Rajiv Gupta, and Mary Lou Soffa. Demand-driven computation of interprocedural data flow. In ACM [2877], pages 37–48. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p37-duesterwald/. ACM or-

der number: 549950.

Reps:1995:PID

[2693] Thomas Reps, Susan Horwitz, and Mooly Sagiv. Precise interprocedural dataflow analysis via graph reachability. In ACM [2877], pages 49–61. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p49-reps/. ACM order num-

ber: 549950.

Sreedhar:1995:LTA

[2694] Vugranam C. Sreedhar and Guang R. Gao. A linear time algorithm for placing [phi]-nodes. In ACM [2877], pages 62–73. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p62-sreedhar/. ACM order

number: 549950.

Altucher:1995:EFM

[2695] Rita Z. Altucher and William Landi. An extended form of must alias analysis for dynamic allocation. In ACM [2877], pages 74–84. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p74-altucher/. ACM order

number: 549950.

Emerson:1995:RAR

[2696] E. Allen Emerson and Kedar S. Namjoshi. Reasoning about rings. In ACM [2877], pages 85–94. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p85-emerson/. ACM order

number: 549950.

Bouajjani:1995:VIS

[2697] Ahmed Bouajjani, Rachid Echahed, and Peter Habermehl. Verifying infinite state processes with sequential and parallel composition. In ACM [2877], pages 95–106. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p95-bouajjani/. ACM order

number: 549950.

Bloom:1995:SOS

[2698] Bard Bloom. Structured operational semantics as a specification language. In ACM [2877], pages 107–117. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p107-bloom/. ACM order

number: 549950.

Dubois:1995:EP

[2699] Catherine Dubois, François Rouaix, and Pierre Weis. Extensional polymorphism. In ACM [2877], pages 118–129. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p118-dubois/. ACM order

number: 549950.

Harper:1995:CPU

[2700] Robert Harper and Greg Morrisett.

Compiling polymorphism using intensional type analysis. In ACM [2877], pages 130–141. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p130-harper/. ACM order

number: 549950.

Leroy:1995:AFF

[2701] Xavier Leroy. Applicative functors and fully transparent higher-order modules. In ACM [2877], pages 142–153. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p142-leroy/. ACM order

number: 549950.

Biswas:1995:HFT

[2702] Sandip K. Biswas. Higher-order functors with transparent signatures. In ACM [2877], pages 154–163. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p154-biswas/. ACM order

number: 549950.

Vorobyov:1995:SDE

[2703] Sergei G. Vorobyov. Structural decidable extensions of bounded quantification. In ACM [2877], pages 164–175. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p164-vorobyov/. ACM order

number: 549950.

Hoang:1995:LBT

[2704] My Hoang and John C. Mitchell. Lower bounds on type inference with subtypes. In ACM [2877], pages 176–185. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p176-hoang/. ACM order

number: 549950.

Hofmann:1995:PS

[2705] Martin Hofmann and Benjamin Pierce. Positive subtyping. In ACM [2877], pages 186–197. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p186-hofmann/. ACM order

number: 549950.

Mackie:1995:GIM

[2706] Ian Mackie. The geometry of interaction machine. In ACM [2877], pages 198–208. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p198-mackie/. ACM order

number: 549950.

Flanagan:1995:SFU

[2707] Cormac Flanagan and Matthias Felleisen. The semantics of future and its use in program optimization. In ACM [2877], pages 209–220. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p209-flanagan/. ACM order

number: 549950.

Sands:1995:TCL

[2708] David Sands. Total correctness by local improvement in program transformation. In ACM [2877], pages 221–232. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p221-sands/. ACM order

number: 549950.

Ariola:1995:CLC

[2709] Zena M. Ariola, John Maraist, Martin Odersky, Matthias Felleisen, and Philip Wadler. A call-by-need lambda calculus. In ACM [2877], pages 233–246. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p233-ariola/. ACM order

number: 549950.

Dawson:1995:UFE

[2710] S. Dawson, C. R. Ramakrishnan, I. V. Ramakrishnan, K. Sagonas, S. Skiena, T. Swift, and D. S. Warren. Unification factoring for efficient execution of logic programs. In ACM [2877], pages 247–258. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p247-dawson/. ACM order

number: 549950.

Schauser:1995:SCP

[2711] Klaus E. Schauser, David E. Culler, and Seth C. Goldstein. Separation constraint partitioning: a new algorithm for partitioning non-strict programs into sequential threads. In ACM [2877], pages 259–271. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p259-schauser/. ACM order

number: 549950.

Saraswat:1995:DTC

[2712] Vijay A. Saraswat, Radha Jagadeesan, and Vineet Gupta. Default timed concurrent constraint programming. In ACM [2877], pages 272–285. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p272-saraswat/. ACM order

number: 549950.

Cardelli:1995:LDS

[2713] Luca Cardelli. A language with distributed scope. In ACM [2877], pages 286–297. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p286-cardelli/. ACM order

number: 549950.

Bailey:1995:FMS

[2714] Mark W. Bailey and Jack W. Davidson. A formal model and specification language for procedure calling conventions. In ACM [2877], pages 298–310. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p298-bailey/. ACM order

number: 549950.

Plevyak:1995:OSE

[2715] John Plevyak, Xingbin Zhang, and Andrew A. Chien. Obtaining sequential efficiency for concurrent object-oriented languages. In ACM [2877], pages 311–321. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p311-plevyak/. ACM order

number: 549950.

Proebsting:1995:OAC

[2716] Todd A. Proebsting. Optimizing an ANSI C interpreter with superoperators. In ACM [2877], pages 322–332. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p322-proebsting/. ACM or-

der number: 549950.

Liang:1995:MTM

[2717] Sheng Liang, Paul Hudak, and Mark Jones. Monad transformers and modular interpreters. In ACM [2877], pages 333–343. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p333-liang/. ACM order

number: 549950.

King:1995:SDS

[2718] David J. King and John Launchbury. Structuring depth-first search algorithms in Haskell. In ACM [2877], pages 344–354. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p344-king/. ACM order num-

ber: 549950.

Sansom:1995:TSP

[2719] Patrick M. Sansom and Simon L. Peyton Jones. Time and space profiling for non-strict, higher-order functional languages. In ACM [2877], pages 355–366. ISBN 0-89791-692-1. LCCN QA 76.7 A11

                                        1995. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        199448/p355-sansom/. ACM order

number: 549950.

Palsberg:1995:TSE

[2720] Jens Palsberg and Patrick O’Keefe. A type system equivalent to flow analysis.

In ACM [2877], pages 367–378. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p367-palsberg/. ACM order

number: 549950.

Field:1995:PPS

[2721] John Field, G. Ramalingam, and Frank Tip. Parametric program slicing. In ACM [2877], pages 379–392. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p379-field/. ACM order

number: 549950.

Jagannathan:1995:UTF

[2722] Suresh Jagannathan and Stephen Weeks. A unified treatment of flow analysis in higher-order languages. In ACM [2877], pages 393–407. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p393-jagannathan/. ACM

order number: 549950.

Castagna:1995:CDB

[2723] Giusee Castagna and Benjamin C.

Pierce. Corrigendum: decidable bounded quantification. In ACM [2877], page 408. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
199448/p408-castagna/. ACM order

number: 549950.

Ghiya:1996:ITD

[2724] Rakesh Ghiya and Laurie J. Hendren. Is it a tree, a DAG, or a cyclic graph? A shape analysis for heap-directed pointers in C. In ACM [2885], pages 1–15. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p1-ghiya/. ACM order num-

ber: 549960.

Sagiv:1996:SSP

[2725] Mooly Sagiv, Thomas Reps, and Reinhard Wilhelm. Solving shapeanalysis problems in languages with destructive updating. In ACM [2885], pages 16–31. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p16-sagiv/. ACM order num-

ber: 549960.

Steensgaard:1996:PAA

[2726] Bjarne Steensgaard. Points-to analysis in almost linear time. In ACM [2885], pages 32–41. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p32-steensgaard/. ACM or-

der number: 549960.

Jim:1996:WPT

[2727] Trevor Jim. What are principal typings and what are they good for? In ACM [2885], pages 42–53. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p42-jim/. ACM order num-

ber: 549960.

Odersky:1996:PTA

[2728] Martin Odersky and Konstantin Läufer. Putting type annotations to work. In ACM [2885], pages 54– 67. ISBN 0-89791-769-3. LCCN QA76.7 A15 1996. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/237721/p54-odersky/

. ACM order number: 549960.

Jones:1996:UPS

[2729] Mark P. Jones. Using parameterized signatures to express modular structure. In ACM [2885], pages 68– 78. ISBN 0-89791-769-3. LCCN QA76.7 A15 1996. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/237721/p68-jones/

. ACM order number: 549960.

Jackson:1996:FCS

[2730] Daniel Jackson, Somesh Jha, and Craig A. Damon. Faster checking of software specifications by eliminating isomorphs. In ACM [2885], pages 79–90. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p79-jackson/. ACM order

number: 549960.

Govindarajan:1996:ORC

[2731] Kannan Govindarajan, Bharat Jayaraman, and Surya Mantha. Optimization and relaxation in constraint logic languages. In ACM [2885], pages 91–103. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p91-govindarajan/. ACM

order number: 549960.

Pienger:1996:PVI

[2732] Nicholas Pienger. Pure versus impure Lisp. In ACM [2885], pages 104– 109. ISBN 0-89791-769-3. LCCN QA76.7 A15 1996. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/237721/p104-pippenger/

. ACM order number: 549960.

Asperti:1996:CB

[2733] Andrea Asperti. On the complexity of beta-reduction. In ACM [2885], pages 110–118. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p110-asperti/. ACM order

number: 549960.

Proebsting:1996:FF

[2734] Todd A. Proebsting and Scott A. Watterson. Filter fusion. In ACM [2885], pages 119–130. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p119-proebsting/. ACM or-

der number: 549960.

Engler:1996:CLH

[2735] Dawson R. Engler, Wilson C. Hsieh, and M. Frans Kaashoek. ’C: a language for high-level, efficient, and machineindependent dynamic code generation. In ACM [2885], pages 131–144. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p131-engler/. ACM order

number: 549960.

Consel:1996:GAR

[2736] Charles Consel and François Noël. A general approach for run-time specialization and its application to C. In ACM [2885], pages 145–156. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p145-consel/. ACM order

number: 549960.

Liu:1996:DAI

[2737] Yanhong A. Liu, Scott D. Stoller, and Tim Teitelbaum. Discovering auxiliary information for incremental computation. In ACM [2885], pages 157–170. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p157-liu/. ACM order num-

ber: 549960.

Birkedal:1996:RIN

[2738] Lars Birkedal, Mads Tofte, and Magnus Vejlstrup. From region inference to von Neumann machines via region representation inference. In ACM [2885], pages 171–183. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p171-birkedal/. ACM order

number: 549960.

Ashley:1996:PFF

[2739] J. Michael Ashley. A practical and flexible flow analysis for higher-order languages. In ACM [2885], pages 184–194. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p184-ashley/. ACM order

number: 549960.

Colby:1996:TPA

[2740] Christopher Colby and Peter Lee.

Trace-based program analysis. In ACM [2885], pages 195–207. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p195-colby/. ACM order

number: 549960.

George:1996:IRC

[2741] Lal George and Andrew W. Ael. Iterated register coalescing. In ACM [2885], pages 208–218. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p208-george/. ACM order

number: 549960.

Hoover:1996:GMS

[2742] Roger Hoover and Kenneth Zadeck. Generating machine specific optimizing compilers. In ACM [2885], pages 219–229. ISBN 0-89791-769-3. LCCN QA76.7 A15 1996. URL http:/

                                        /www.acm.org:80/pubs/citations/
                                        proceedings/plan/237721/p219-hoover/

. ACM order number: 549960.

Kurlander:1996:MCI

[2743] Steven M. Kurlander and Charles N. Fischer. Minimum cost interprocedural register allocation. In ACM [2885], pages 230–241. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p230-kurlander/. ACM or-

der number: 549960.

Danvy:1996:TPE

[2744] Olivier Danvy. Type-directed partial evaluation. In ACM [2885], pages 242–257. ISBN 0-89791-769-3. LCCN QA76.7 A15 1996. URL http:

                                        //www.acm.org:80/pubs/citations/

proceedings/plan/237721/p242-danvy/

. ACM order number: 549960.

Davies:1996:MAS

[2745] Rowan Davies and Frank Pfenning. A modal analysis of staged computation. In ACM [2885], pages 258–270. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p258-davies/. ACM order

number: 549960.

Minamide:1996:TCC

[2746] Yasuhiko Minamide, Greg Morrisett, and Robert Harper. Typed closure conversion. In ACM [2885], pages 271–283. ISBN 0-89791-769-3. LCCN QA76.7 A15 1996. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/237721/p271-minamide/

. ACM order number: 549960.

Fegaras:1996:RCD

[2747] Leonidas Fegaras and Tim Sheard. Revisiting catamorphisms over datatypes with embedded functions (or, programs from outer space). In ACM [2885], pages 284–294. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p284-fegaras/. ACM order

number: 549960.

PeytonJones:1996:CH

[2748] Simon Peyton Jones, Andrew Gordon, and Sigbjorn Finne. Concurrent Haskell. In ACM [2885], pages 295–308. ISBN 0-89791-769-3. LCCN QA76.7 A15 1996. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/237721/p295-peyton_

jones/. ACM order number: 549960.

Greiner:1996:PTP

[2749] John Greiner and Guy E. Blelloch. A provably time-efficient parallel implementation of full speculation. In ACM [2885], pages 309–321. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p309-greiner/. ACM order

number: 549960.

Gupta:1996:SAR

[2750] Manish Gupta and Edith Schonberg. Static analysis to reduce synchronization costs in data-parallel programs. In ACM [2885], pages 322–332. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p322-gupta/. ACM order

number: 549960.

Niehren:1996:FCC

[2751] Joachim Niehren. Functional computation as concurrent computation. In ACM [2885], pages 333–343. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p333-niehren/. ACM order

number: 549960.

Honda:1996:CP

[2752] Kohei Honda. Composing processes. In ACM [2885], pages 344–357. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p344-honda/. ACM order

number: 549960.

Kobayashi:1996:LP

[2753] Naoki Kobayashi, Benjamin C. Pierce, and David N. Turner. Linearity and the pi-calculus. In ACM [2885], pages 358–371. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p358-kobayashi/. ACM or-

der number: 549960.

Fournet:1996:RCJ

[2754] Cédric Fournet and Georges Gonthier.

The reflexive CHAM and the joincalculus. In ACM [2885], pages 372–

385. ISBN 0-89791-769-3. LCCN QA76.7 A15 1996. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/237721/p372-fournet/

. ACM order number: 549960.

Gordon:1996:BFC

[2755] Andrew D. Gordon and Gareth D.

Rees. Bisimilarity for a first-order calculus of objects with subtyping. In ACM [2885], pages 386–395. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p386-gordon/. ACM order

number: 549960.

Abadi:1996:IOO

[2756] Martín Abadi, Luca Cardelli, and Ramesh Viswanathan. An interpretation of objects and object types. In ACM [2885], pages 396–409. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
237721/p396-abadi/. ACM order

number: 549960.

Hughes:1996:PCR

[2757] John Hughes, Lars Pareto, and Amr Sabry. Proving the correctness of reactive systems using sized types. In ACM [2885], pages 410–423. ISBN 0-89791-769-3. LCCN QA76.7 A15

                                        1996. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        237721/p410-hughes/. ACM order

number: 549960.

Shapiro:1997:FAF

[2758] Marc Shapiro and Susan Horwitz. Fast and accurate flow-insensitive pointsto analysis. In ACM [2888], pages 1–14. ISBN 0-89791-853-3. LCCN QA 76.7 A11 1997. URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/263699/p1-shapiro/

.

Ruf:1997:PDA

[2759] Erik Ruf. Partitioning dataflow analyses using types. In ACM [2888], pages 15–26. ISBN 0-89791-853-3. LCCN QA 76.7 A11 1997. URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/263699/p15-ruf/

.

Fradet:1997:ST

[2760] Pascal Fradet and Daniel Le Métayer. Shape types. In ACM [2888], pages 27–39. ISBN 0-89791-853-3. LCCN QA 76.7 A11 1997. URL http:

                                        //www.acm.org:80/pubs/citations/
                                        proceedings/plan/263699/p27-fradet/

.

Remy:1997:OMS

[2761] Didier Rémy and Jérôme Vouillon. Objective ML: a simple objectoriented extension of ML. In ACM [2888], pages 40–53. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p40-remy/.

Chen:1997:RYO

[2762] Chih-Ping Chen and Paul Hudak. Rolling your own mutable ADT — a connection between linear types and monads. In ACM [2888], pages 54–66. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p54-chen/.

Apt:1997:SIP

[2763] Krzysztof R. Apt and Andrea Schaerf.

Search and imperative programming. In ACM [2888], pages 67–79. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p67-apt/.

Hanus:1997:UCM

[2764] Michael Hanus. A unified computation model for functional and logic programming. In ACM [2888], pages 80–93. ISBN 0-89791-853-3. LCCN QA 76.7 A11 1997. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/263699/p80-hanus/

.

Middeldorp:1997:CNC

[2765] Aart Middeldorp. Call by need computations to root-stable form. In ACM [2888], pages 94–105. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p94-middeldorp/.

Necula:1997:PC

[2766] George C. Necula. Proof-carrying code.

In ACM [2888], pages 106–119. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p106-necula/.

Plezbert:1997:DQT

[2767] Michael P. Plezbert and Ron K. Cytron. Does ‘just in time’ = ‘better late than never’? In ACM [2888], pages 120–131. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p120-plezbert/.

Bank:1997:PTJ

[2768] Joseph A. Bank, Andrew C. Myers, and Barbara Liskov. Parameterized types for Java. In ACM [2888], pages 132–145. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p132-bank/.

Odersky:1997:PJT

[2769] Martin Odersky and Philip Wadler. Pizza into Java: translating theory into practice. In ACM [2888], pages 146–159. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p146-odersky/.

Hogstedt:1997:DIT

[2770] Karin Högstedt, Larry Carter, and Jeanne Ferrante. Determining the idle time of a tiling. In ACM [2888], pages 160–173. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/

pubs/citations/proceedings/plan/
263699/p160-hogstedt/.

Godefroid:1997:MCP

[2771] Patrice Godefroid. Model checking for programming languages using VeriSoft. In ACM [2888], pages 174–186. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p174-godefroid/.

Diniz:1997:STP

[2772] Pedro Diniz and Martin Rinard. Synchronization transformations for parallel computing. In ACM [2888], pages 187–200. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p187-diniz/.

Lim:1997:MPM

[2773] Amy W. Lim and Monica S. Lam. Maximizing parallelism and minimizing synchronization with affine transforms. In ACM [2888], pages 201–214. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p201-lim/.

Ong:1997:CFF

[2774] C.-H. L. Ong and C. A. Stewart. A Curry–Howard foundation for functional computation with control. In ACM [2888], pages 215–227. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p215-ong/.

Boudol:1997:CDS

[2775] Gérard Boudol. The π-calculus in direct style. In ACM [2888], pages 228–242. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p228-boudol/.

Pierce:1997:BEP

[2776] Benjamin C. Pierce and Davide Sangiorgi. Behavioral equivalence in the polymorphic pi-calculus. In ACM [2888], pages 242–255. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p242-pierce/.

Palamidessi:1997:CEP

[2777] Catuscia Palamidessi. Comparing the expressive power of the synchronous and the asynchronous π-calculus. In ACM [2888], pages 256–265. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p256-palamidessi/.

Cardelli:1997:PFL

[2778] Luca Cardelli. Program fragments, linking, and modularization. In ACM [2888], pages 266–277. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p266-cardelli/.

Rehof:1997:MTA

[2779] Jakob Rehof. Minimal typings in atomic subtyping. In ACM [2888], pages 278–291. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p278-rehof/.

Saibi:1997:TAT

[2780] Amokrane Saïbi. Typing algorithm in type theory with inheritance. In ACM [2888], pages 292–301. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p292-saibi/.

Bourdoncle:1997:TCH

[2781] François Bourdoncle and Stephan Merz. Type checking higher-order polymorphic multi-methods. In ACM [2888], pages 302–315. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p302-bourdoncle/.

Cousot:1997:TAI

[2782] Patrick Cousot. Types as abstract interpretations. In ACM [2888], pages 316–331. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p316-cousot/.

Nielson:1997:ICF

[2783] Flemming Nielson and Hanne Riis Nielson. Infinitary control flow analysis: a collecting semantics for closure analysis. In ACM [2888], pages 332–345. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p332-nielson/.

Lesens:1997:AVP

[2784] David Lesens, Nicolas Halbwachs, and Pascal Raymond. Automatic verification of parameterized linear networks of processes. In ACM [2888], pages 346–357. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p346-lesens/.

Deutsch:1997:CEA

[2785] Alain Deutsch. On the complexity of escape analysis. In ACM [2888], pages 358–371. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p358-deutsch/.

Biswas:1997:DSA

[2786] Sandip K. Biswas. A demanddriven set-based analysis. In ACM [2888], pages 372–385. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p372-biswas/.

Wand:1997:DSU

[2787] Mitchell Wand and Gregory T. Sullivan. Denotational semantics using an operationally-based term model. In ACM [2888], pages 386–399. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p386-wand/.

Seidl:1997:CSH

[2788] H. Seidl and M. H. Sørensen. Constraints to stop higher-order deforestation. In ACM [2888], pages 400–413. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p400-seidl/.

Pettorossi:1997:RNW

[2789] Alberto Pettorossi, Maurizio Proietti, and Sophie Renault. Reducing nondeterminism while specializing logic programs. In ACM [2888], pages 414–427. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p414-pettorossi/.

Sands:1997:SRP

[2790] David Sands. From SOS rules to proof principles: an operational metatheory for functional languages. In ACM [2888], pages 428–441. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p428-sands/.

Kennedy:1997:RPU

[2791] Andrew J. Kennedy. Relational parametricity and units of measure. In ACM [2888], pages 442–455. ISBN 0-89791-853-3. LCCN QA 76.7 A11

1997. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
263699/p442-kennedy/.

Paige:1997:HLR

[2792] Robert Paige and Zhe Yang. High level reading and data structure compilation. In ACM [2888], pages 456–469. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p456-paige/.

Jansson:1997:PPP

[2793] Patrik Jansson and Johan Jeuring. PolyP — a polytypic programming language extension. In ACM [2888], pages 470–482. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p470-jansson/.

Jones:1997:FPT

[2794] Mark P. Jones. First-class polymorphism with type inference. In ACM [2888], pages 483–496. ISBN 0-89791-853-3. LCCN QA 76.7 A11

                                        1997. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        263699/p483-jones/.

Nipkow:1998:JTS

[2795] Tobias Nipkow and David von Oheimb. Java_{light} is type-safe-definitely. In ACM [2892], pages 161–170. ISBN 0-89791-979-3. LCCN QA76.7 .A15

                                        1998. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        268946/p161-nipkow/. ACM order

number: 549981.

Hannan:1998:HU

[2796] John Hannan and Patrick Hicks. Higher-order unCurrying. In ACM [2892], pages 1–11. ISBN 0-89791-979-3. LCCN QA76.7 .A15

                                        1998. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        268946/p1-hannan/. ACM order num-

ber: 549981.

Debray:1998:AAE

[2797] Saumya Debray, Robert Muth, and Matthew Weiert. Alias analysis of executable code. In ACM [2892], pages 12–24. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p12-debray/. ACM order

number: 549981.

Blanchet:1998:EAC

[2798] Bruno Blanchet. Escape analysis: correctness proof, implementation and experimental results. In ACM [2892], pages 25–37. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p25-blanchet/. ACM order

number: 549981.

Schmidt:1998:DFA

[2799] David A. Schmidt. Data flow analysis is model checking of abstract interpretations. In ACM [2892], pages 38–48. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p38-schmidt/. ACM order

number: 549981.

Jones:1998:BGC

[2800] Simon Peyton Jones, Mark Shields, John Launchbury, and Andrew Tolmach. Bridging the gulf: a common intermediate language for ML and Haskell. In ACM [2892], pages 49–61. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http:/

/www.acm.org:80/pubs/citations/

                                        proceedings/plan/268946/p49-peyton_

jones/. ACM order number: 549981.

Ariola:1998:CMS

[2801] Zena M. Ariola and Amr Sabry. Correctness of monadic state: an imperative call-by-need calculus. In ACM [2892], pages 62–74. ISBN 0-89791-979-3. LCCN QA76.7 .A15

                                        1998. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        268946/p62-ariola/. ACM order

number: 549981.

Minamide:1998:FRD

[2802] Yasuhiko Minamide. A functional representation of data structures with a hole. In ACM [2892], pages 75–84. ISBN 0-89791-979-3. LCCN QA76.7 .A15

                                        1998. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        268946/p75-minamide/. ACM order

number: 549981.

Morrisett:1998:SFT

[2803] Greg Morrisett, David Walker, Karl Crary, and Neal Glew. From system F to typed assembly language. In ACM [2892], pages 85–97. ISBN 0-89791-979-3. LCCN QA76.7 .A15

                                        1998. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        268946/p85-morrisett/. ACM order

number: 549981.

Barthou:1998:MSE

[2804] Denis Barthou, Albert Cohen, and Jean-François Collard. Maximal static expansion. In ACM [2892], pages 98– 106. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http:/

                                        /www.acm.org:80/pubs/citations/

proceedings/plan/268946/p98-barthou/

. ACM order number: 549981.

Knobe:1998:ASF

[2805] Kathleen Knobe and Vivek Sarkar. Array SSA form and its use in parallelization. In ACM [2892], pages 107– 120. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/268946/p107-knobe/

. ACM order number: 549981.

Ghiya:1998:PPA

[2806] Rakesh Ghiya and Laurie J. Hendren.

Putting pointer analysis to work. In ACM [2892], pages 121–133. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p121-ghiya/. ACM order

number: 549981.

Ball:1998:EPV

[2807] Thomas Ball, Peter Mataga, and Mooly Sagiv. Edge profiling versus path profiling: the showdown. In ACM [2892], pages 134–148. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p134-ball/. ACM order num-

ber: 549981.

Stata:1998:TSJ

[2808] Raymie Stata and Martín Abadi. A type system for Java bytecode subroutines. In ACM [2892], pages 149– 160. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/268946/p149-stata/

. ACM order number: 549981.

Flatt:1998:CM

[2809] Matthew Flatt, Shriram Krishnamurthi, and Matthias Felleisen. Classes and mixins. In ACM [2892], pages 171–183. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p171-flatt/. ACM order

number: 549981.

Collberg:1998:MCR

[2810] Christian Collberg, Clark Thomborson, and Douglas Low. Manufacturing cheap, resilient, and stealthy opaque constructs. In ACM [2892], pages 184–196. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p184-collberg/. ACM order

number: 549981.

Palsberg:1998:PFI

[2811] Jens Palsberg and Christina Pavlopoulou. From polyvariant flow information to intersection and union types. In ACM [2892], pages 197–208. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p197-palsberg/. ACM order

number: 549981.

Jensen:1998:IPC

[2812] Thomas P. Jensen. Inference of polymorphic and conditional strictness properties. In ACM [2892], pages 209– 221. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/268946/p209-jensen/

. ACM order number: 549981.

DeFouw:1998:FIC

[2813] Greg DeFouw, David Grove, and Craig Chambers. Fast interprocedural class analysis. In ACM [2892], pages 222–236. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p222-defouw/. ACM order

number: 549981.

Bodik:1998:PVA

[2814] Rastisalv Bodík and Sadun Anik.

Path-sensitive value-flow analysis. In ACM [2892], pages 237–251. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p237-bodik/. ACM order

number: 549981.

Pierce:1998:LTI

[2815] Benjamin C. Pierce and David N.

Turner. Local type inference. In ACM [2892], pages 252–265. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p252-pierce/. ACM order

number: 549981.

Nishimura:1998:STD

[2816] Susumu Nishimura. Static typing for dynamic messages. In ACM [2892], pages 266–278. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p266-nishimura/. ACM or-

der number: 549981.

Schubert:1998:SUT

[2817] Aleksy Schubert. Second-order unification and type inference for Churchstyle polymorphism. In ACM [2892], pages 279–288. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p279-schubert/. ACM order

number: 549981.

Shields:1998:DTS

[2818] Mark Shields, Tim Sheard, and Simon Peyton Jones. Dynamic typing as staged type inference. In ACM [2892], pages 289–302. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p289-shields/. ACM order

number: 549981.

Asperti:1998:PBR

[2819] Andrea Asperti and Harry G. Marison. Parallel beta reduction is not elementary recursive. In ACM [2892], pages 303–315. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p303-asperti/. ACM order

number: 549981.

Hu:1998:PCF

[2820] Zhenjiang Hu, Masato Takeichi, and Wei-Ngan Chin. Parallelization in calculational forms. In ACM [2892], pages 316–328. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/268946/p316-hu/

. ACM order number: 549981.

Jagannathan:1998:SLI

[2821] Suresh Jagannathan, Peter Thiemann, Stephen Weeks, and Andrew Wright.

Single and loving it: must-alias analysis for higher-order languages. In ACM [2892], pages 329–341. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p329-jagannathan/. ACM

order number: 549981.

Aiken:1998:BI

[2822] Alexander Aiken and David Gay. Barrier inference. In ACM [2892], pages 342–354. ISBN 0-89791-979-3. LCCN QA76.7 .A15 1998. URL http:/

/www.acm.org:80/pubs/citations/
proceedings/plan/268946/p342-aiken/

. ACM order number: 549981.

Smith:1998:SIF

[2823] Geoffrey Smith and Dennis Volpano.

Secure information flow in a multithreaded imperative language. In ACM [2892], pages 355–364. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p355-smith/. ACM order

number: 549981.

Heintze:1998:SCP

[2824] Nevin Heintze and Jon G. Riecke.

The SLam calculus: programming with secrecy and integrity. In ACM [2892], pages 365–377. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
268946/p365-heintze/. ACM order

number: 549981.

Riely:1998:TLD

[2825] James Riely and Matthew Hennessy.

A typed language for distributed mobile processes (extended abstract). In ACM [2892], pages 378–390. ISBN 0-89791-979-3. LCCN QA76.7 .A15

                                        1998. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        268946/p378-riely/. ACM order

number: 549981.

Leroy:1998:SPT

[2826] Xavier Leroy and François Rouaix. Security properties of typed applets. In ACM [2892], pages 391–403. ISBN 0-89791-979-3. LCCN QA76.7 .A15

                                        1998. URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        268946/p391-leroy/. ACM order

number: 549981.

Eidorff:1999:ATT

[2827] Peter Harry Eidorff, Fritz Henglein, Christian Mossin, Henning Niss, Morten Heine Sørensen, and Mads Tofte. AnnoDomini: from type theory to Year 2000 conversion tool. In ACM [2901], pages 1–14. ISBN 1-58113-095-3. LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        292540/p1-eidorff/.

Wansbrough:1999:OUP

[2828] Keith Wansbrough and Simon Peyton Jones. Once upon a polymorphic type. In ACM [2901], pages 15–28. ISBN 1-58113-095-3. LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        292540/p15-wansbrough/.

Kobayashi:1999:QT

[2829] Naoki Kobayashi. Quasi-linear types. In ACM [2901], pages 29– 42. ISBN 1-58113-095-3. LCCN

                                        ???? URL http://www.acm.org:80/

pubs/citations/proceedings/plan/
292540/p29-kobayashi/.

Moran:1999:ILC

[2830] Andrew Moran and David Sands.

Improvement in a lazy context: an operational theory for call-by-need. In ACM [2901], pages 43– 56. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p43-moran/.

Yelland:1999:CAJ

[2831] Phillip M. Yelland. A compositional account of the Java virtual machine. In ACM [2901], pages 57– 69. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p57-yelland/.

OCallahn:1999:SCT

[2832] Robert O’Callahn. A simple, comprehensive type system for Java bytecode subroutines. In ACM [2901], pages 70–78. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p70-o_callahn/.

Cardelli:1999:TMA

[2833] Luca Cardelli and Andrew D. Gordon. Types for mobile ambients. In ACM [2901], pages 79– 92. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p79-cardelli/.

Riely:1999:TPT

[2834] James Riely and Matthew Hennessy. Trust and partial typing in open systems of mobile agents. In ACM [2901], pages 93–

104. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p93-riely/.

Sagiv:1999:PSA

[2835] Mooly Sagiv, Thomas Reps, and Reinhard Wilhelm. Parametric shape analysis via 3-valued logic. In ACM [2901], pages 105–118. ISBN 1-58113-095-3. LCCN ???? URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/292540/p105-sagiv/

.

Ramalingam:1999:ASI

[2836] G. Ramalingam, John Field, and Frank Tip. Aggregate structure identification and its application to program analysis. In ACM [2901], pages 119–

132. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p119-ramalingam/.

Chatterjee:1999:RCI

[2837] Ramkrishna Chatterjee, Barbara G. Ryder, and William A. Landi. Relevant context inference. In ACM [2901], pages 133–146. ISBN 1-58113-095-3. LCCN ???? URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/292540/p133-chatterjee/

.

Abadi:1999:CCD

[2838] Martín Abadi, Anindya Banerjee, Nevin Heintze, and Jon G. Riecke. A core calculus of dependency. In ACM [2901], pages 147–160. ISBN 1-58113-095-3. LCCN ???? URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/292540/p147-abadi/

.

Kfoury:1999:PDT

[2839] A. J. Kfoury and J. B. Wells.

Principality and decidable type inference for finite-rank intersection types. In ACM [2901], pages 161– 174. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p161-kfoury/.

Filinski:1999:RLM

[2840] Andrzej Filinski. Representing layered monads. In ACM [2901], pages 175–188. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p175-filinski/.

Gupta:1999:SPC

[2841] Vineet Gupta, Radha Jagadeesan, and Prakash Panangaden. Stochastic processes as concurrent constraint programs. In ACM [2901], pages 189– 202. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p189-gupta/.

Waddell:1999:ESS

[2842] Oscar Waddell and R. Kent Dybvig.

Extending the scope of syntactic abstraction. In ACM [2901], pages 203– 215. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p203-waddell/.

Xi:1999:DTP

[2843] Hongwei Xi and Frank Pfenning. Dependent types in practical programming. In ACM [2901], pages 214– 227. ISBN 1-58113-095-3. LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        292540/p214-xi/.

Myers:1999:JPM

[2844] Andrew C. Myers. JFlow: practical mostly-static information flow control. In ACM [2901], pages 228– 241. ISBN 1-58113-095-3. LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        292540/p228-myers/.

Ertl:1999:OCS

[2845] M. Anton Ertl. Optimal code selection in DAGs. In ACM [2901], pages 242–249. ISBN 1-58113-095-3. LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        292540/p242-ertl/.

Glew:1999:TLM

[2846] Neal Glew and Greg Morrisett. Typesafe linking and modular assembly language. In ACM [2901], pages 250– 261. ISBN 1-58113-095-3. LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        292540/p250-glew/.

Crary:1999:TMM

[2847] Karl Crary, David Walker, and Greg Morrisett. Typed memory management in a calculus of capabilities. In ACM [2901], pages 262– 275. ISBN 1-58113-095-3. LCCN

                                        ???? URL http://www.acm.org:80/
                                        pubs/citations/proceedings/plan/
                                        292540/p262-crary/.

Pessaux:1999:TAU

[2848] François Pessaux and Xavier Leroy.

Type-based analysis of uncaught exceptions. In ACM [2901], pages 276–

290. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p276-pessaux/.

Wand:1999:CSU

[2849] Mitchell Wand and Igor Siveroni. Constraint systems for useless variable elimination. In ACM [2901], pages 291–302. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p291-wand/.

Ruckert:1999:CG

[2850] Martin Ruckert. Continuous grammars. In ACM [2901], pages 303–

310. ISBN 1-58113-095-3. LCCN

???? URL http://www.acm.org:80/
pubs/citations/proceedings/plan/
292540/p303-ruckert/.

Collberg:1999:SWM

[2851] Christian Collberg and Clark Thomborson. Software watermarking: models and dynamic embeddings. In ACM [2901], pages 311–324. ISBN 1-58113-095-3. LCCN ???? URL http:

//www.acm.org:80/pubs/citations/
proceedings/plan/292540/p311-collberg/

.

ACM:1990:PAS

[2852] ACM, editor. Proceedings of the ACM SIGPLAN ’90 Conference on Programming Language Design and Implementation, White Plains, New York, June 20–22, 1990, volume 25(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1990. CO-DEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7.S53 1990.

ACM:1990:PPS

[2853] ACM, editor. POPL ’90. Proceedings of the seventeenth annual ACM symposium on Principles of programming languages, January 17–19, 1990, San Francisco, CA. ACM Press, New York, NY, USA, 1990. ISBN ???? LCCN ????

ACM:1990:PSA

[2854] ACM, editor. Proceedings of the Second ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, PPOPP, Seattle, Washington, March 14–16, 1990, volume 25(3) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, March 1990. ISBN 0-89791-350-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA 76.7 A1095 v.25 no.3 1990.

ACM:1991:PAS

[2855] ACM, editor. Proceedings of the ACM SIGPLAN ’91 Conference on Programming Language Design and Implementation, Toronto, Ontario, Canada, June 26–28, 1991, volume 26(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1991. ISBN 0-89791-428-7. LCCN QA76.7.S53 1991. ACM order number: 548910.

ACM:1991:PPE

[2856] ACM, editor. POPL ’91. Proceedings of the eighteenth annual ACM symposium on Principles of programming languages, January 21–23, 1991, Orlando, FL. ACM Press, New York, NY, USA, 1991. ISBN ???? LCCN ????

ACM:1991:PTA

[2857] ACM, editor. Proceedings of the Third ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, PPOPP: Williamsburg, Virginia, April 21–24, 1991, volume 26(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1991. ISBN 0-89791-390-6. LCCN QA76.7 .S54 v.26:7.

Koopman:1991:PST

[2858] Philip J. Koopman, Jr., editor. The proceedings of the second and third annual workshops for the ACM Special Interest Group on Forth : SIGForth ’90, February 16–18, 1990, Dallas, Texas … SIGForth ’91, March 7–9, 1991, San Antonio, Texas. ACM Press, New York, NY, USA, 1991. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53 1991. ACM order number 817911.

ACM:1992:AVP

[2859] ACM, editor. ASPLOS-V proceedings. ACM Press, New York, NY, USA, 1992. ISBN 0-89791-534-8 (paperback), 0-89791-535-6 (hard cover). LCCN QA76.9.A73I573 1992. SIG-GRAPH Computer architecture news, volume 20, Special issue, October 1992. SIGOPS Operating systems review, volume 26, special issue, October 1992. SIGPLAN ACM SIGPLAN notices, volume 27, number 9, September 1992.

ACM:1992:CRN

[2860] ACM, editor. Conference record of the Nineteenth Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Albuquerque, New Mexico, January 19–22, 1992. ACM Press, New York, NY, USA, 1992. ISBN 0-89791-453-8. LCCN QA76.7 .A15

1992. URL http://www.acm.org/
pubs/contents/proceedings/plan/
143165/index.html. ACM order num-

ber 54990.

ACM:1992:FPF

[2861] ACM, editor. FORTH ’92. Proceedings of the fourth annual workshop on Forth from the classroom to the real world, March 5–7, 1992, Kansas City, MO. ACM Press, New York, NY, USA, 1992. ISBN ???? LCCN ????

ACM:1992:PAS

[2862] ACM, editor. Proceedings of the ACM SIGPLAN ’92 Conference on Programming Language Design and Implementation, San Francisco, California, June 17–19, 1992, volume 27(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1992. CODEN SIN-ODQ. ISBN 0-89791-475-9 (soft cover), 0-89791-476-7 (hard cover). ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). ACM order number: 548920.

ACM:1993:CRT

[2863] ACM, editor. Conference record of the Twentieth Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Charleston, South Carolina, January 10–13, 1993. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15

1993. URL http://www.acm.org/
pubs/contents/proceedings/plan/
158511/index.html. ACM order num-

ber 549930.

ACM:1993:PASa

[2864] ACM, editor. Proceedings of the ACM SIGPLAN Symposium on Partial Evaluation and Semantics-Based Program Manipulation. PEPM’93. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-594-1. LCCN QA76.7.S97 1993.

ACM:1993:ASH

[2865] ACM, editor. ACM SIGPLAN HOPL-II. 2nd ACM SIGPLAN History of Programming Languages Conference (Preprints), volume 28(3) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, March 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1993:FAS

[2866] ACM, editor. Fourth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, volume 28(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1993:PASb

[2867] ACM, editor. Proceedings of the ACM SIGPLAN ’93 Conference on Programming Language Design and Implementation: Albuquerque, New Mexico, June 23–25, 1993, volume 28(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7.A152 v. 28, no. 6.

ACM:1993:PFA

[2868] ACM, editor. Proceedings of the Fourth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, PPOPP: San Diego, California, May 19–22, 1993, volume 28(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-589-5. LCCN QA76.7 .S54 v. 28:7.

ACM:1993:PSA

[2869] ACM, editor. Preprints: the Second ACM SIGPLAN History of Programming Languages Conference (HOPL-II): April 20–23, 1993, Cambridge, Massachusetts, USA, volume 28(3) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-570-4. LCCN QA76.7 .S54 v.28:3.

ACM:1994:CRP

[2870] ACM, editor. Conference record of POPL ’94, 21st ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: Portland, Oregon, January 17–21, 1994. ACM Press, New York, NY, USA, 1994. ISBN 0-89791-636-0. LCCN QA76.7 .A15

                                        1994. URL http://www.acm.org/
                                        pubs/contents/proceedings/plan/
                                        174675/index.html.

ACM:1994:PAS

[2871] ACM, editor. Proceedings of ACM SIG-PLAN Workshop on Partial Evaluation and Semantics-Based Program Manipulation. Univ. Melbourne, Melbourne, Vic., Australia, ???? 1994.

ACM:1994:AWI

[2872] ACM, editor. ACM Workshop on Interface Definition Languages. Portland, OR, USA, 20 January, 1994, volume 29(8) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, August 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1994:SIC

[2873] ACM, editor. Sixth International Conference on Architectural Support for Programming Languages and Operating Systems (ASPLOS-VI). San Jose, CA, USA, 4–7 October, 1994, volume 29(11) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, November 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1994:ASC

[2874] ACM, editor. ACM SIGPLAN ’94 Conference on Programming Language Design and Implementation (PLDI). Orlando, FL, USA, 20–24 June, 1994, volume 29(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1994:NAC

[2875] ACM, editor. Ninth Annual Conference on Object-Oriented Programming Systems, Languages, and Applications. OOPSLA ’94. Portland, OR, USA, 23– 27 October, 1994, volume 29(10) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1995:ASW

[2876] ACM SIGPLAN Workshop on Intermediate Representations (IR’95), volume 30(3) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, March 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1995:CRP

[2877] ACM, editor. Conference record of POPL ’95, 22nd ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: San Francisco, California, January 22–25, 1995. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.org/
pubs/contents/proceedings/plan/
199448/index.html. ACM order num-

ber: 549950.

ACM:1995:OCO

[2878] ACM, editor. OOPSLA ’95: Conference on Object-Oriented Programming Systems, Languages, and Applications: conference proceedings. October 15–19, 1995, Austin, Texas, volume 30(10) of ACM SIGPLAN Notices. Addison-Wesley and ACM Press, Reading, MA, USA and New York, NY, USA, 1995. ISBN 0-89791-703-0, 0-201-87810-0. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.64.C67 1995. ACM order number 548951.

Bilow:1995:OAP

[2879] Steven Craig Bilow and Patricia S.

Bilow, editors. OOPSLA ’95: Addendum to the Proceedings, volume 6(4) of OOPS Messenger. ACM Press, New York, NY, USA, 1995. ISBN 0-89791-721-9. ISSN 1055-6400. LCCN QA 76.64 O56 v.6 no.4 1995.

Ferrante:1995:FAS

[2880] Jeanne Ferrante and David Padua, editors. Fifth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, PPoPP, volume 30(8) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gerber:1995:ASW

[2881] Richard Gerber and Thomas Marlowe, editors. ACM SIGPLAN 1995 Workshop on Languages, Compilers, and Tools for Real-Time Systems, volume 30(11) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, November 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirfs-Brock:1995:OTA

[2882] Rebecca J. Wirfs-Brock and Mary E. S.

Loomis, editors. OOPSLA ’95: Tenth Annual Conference on Object-Oriented Programming Systems, Languages and Applications, volume 30(10) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1995. CODEN SIN-ODQ. ISBN 0-201-87810-0. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wall:1995:ASC

[2883] David W. Wall, editor. ACM SIG-PLAN ’95 Conference on Programming Language Design and Implementation (PLDI), volume 30(6) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, June 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1996:AVP

[2884] ACM, editor. ASPLOS-VII proceedings / Seventh International Conference on Architectural Support for Programming Languages and Operating Systems, Cambridge, Massachusetts, October 1–5, 1996, volume 31(9) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Copublished as SIGPLAN Notices 31(9), September 1996, as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

ACM:1996:CRP

[2885] ACM, editor. Conference record of POPL ’96, 23rd ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: St. Petersburg Beach, Florida, 21–24 January 1996. ACM Press, New York, NY, USA, 1996. ISBN 0-89791-769-3. LCCN QA76.7 A15

1996. URL http://www.acm.org/
pubs/contents/proceedings/plan/
237721/index.html. ACM order num-

ber: 549960.

ACM:1996:PASa

[2886] ACM, editor. Proceedings of the ACM SIGPLAN ’96 Conference on Programming Language Design and Implementation (PLDI), Philadelphia, PA May 21–24, 1996, Sponsored by the Special Interest Group on Programming Languages, Association for Computing Machinery, volume 31(5) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, May 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1996:PASb

[2887] ACM, editor. Proceedings of the ACM SIGPLAN ’96 International Conference on Functional Programming (ICFP ’96), May 24–26, 1996, volume 31(5) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1997:CRP

[2888] ACM, editor. Conference record of POPL ’97, the 24th ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Paris, France, 15–17 January 1997. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-853-3. LCCN QA 76.7

A11 1997. URL http://www.acm.org/
pubs/contents/proceedings/plan/
263699/index.html.

ACM:1997:PAS

[2889] ACM, editor. Proceedings of the ACM SIGPLAN ’97 Conference on Programming Language Design and Implementation (PLDI), Las Vegas, Nevada, 15– 18 June 1997, volume 32(5) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-907-6. LCCN QA 76.7 A1095 v.32 no.5 1997.

ACM:1997:PFT

[2890] ACM, editor. POPL the first ten years: ACM Symposium on Principles of Programming Languages proceedings compendium, 1973–1983. ACM Press, New York, NY, USA, 1997. ISBN 0-89791-947-5. LCCN ???? CD ROM only.

ACM:1998:AWJ

[2891] ACM, editor. ACM 1998 Workshop on Java for High-Performance Network Computing. ACM Press, New York, NY, USA, 1998. ISBN ???? LCCN

????  URL http://www.cs.ucsb.
edu/conferences/java98/program.

html. Also published as Concurrency: Practice and Experience, 10(11–13), September 1998, CODEN CPEXEI, ISSN 1040-3108.

ACM:1998:CRP

[2892] ACM, editor. Conference record of POPL ’98: the 25th ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium, San Diego, California, 19– 21 January 1998. ACM Press, New York, NY, USA, 1998. ISBN 0-89791-979-3. LCCN QA76.7 .A15

1998. URL http://www.acm.org/
pubs/contents/proceedings/plan/
268946/index.html. ACM order num-

ber: 549981.

ACM:1998:PASa

[2893] ACM, editor. Proceedings of the ACM SIGPLAN ’98 Conference on Programming Language Design and Implementation (PLDI), Montréal, Canada, 17– 19 June 1998, volume 33(5) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.33 no.5.

ACM:1998:PASb

[2894] ACM, editor. Proceedings of the 1998 ACM SIGPLAN/SIGSOFT Workshop on Program Analysis for Software Tools and Engineering, volume 33(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1998.

ACM:1999:PASa

[2895] ACM, editor. Proceedings of the ACM SIGPLAN ’99 Conference on Programming Language Design and Implementation (PLDI ’99), Atlanta, Georgia, 2–4 May 1999, volume 34(5) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1999. ISBN 1-58113-094-5. LCCN ????

ACM:1999:PASb

[2896] ACM, editor. Proceedings of the ACM SIGPLAN 1999 Workshop on Languages, Compilers, and Tools for Embedded Systems (LCTES’99), Atlanta, Georgia, USA, May 5, 1999, volume 34(7) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, July 1999. ISBN ???? LCCN ????

ACM:1999:PASc

[2897] ACM, editor. Proceedings of the 1999 ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (PPOPP’99), Atlanta, Georgia, 4–6 May 1999, volume 34(8) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, August 1999. ISBN ???? LCCN ????

ACM:1999:PASd

[2898] ACM, editor. Proceedings of the ACM SIGPLAN International Conference on Functional Programming (ICFP ’99): September 27–29, 1999, Paris, France, volume 34(9) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, September 1999. ISBN 1-58113-111-9. LCCN ???? ACM order number: 565990.

ACM:1999:PASe

[2899] ACM, editor. Proceedings of the 1999 ACM SIGPLAN Conference on Object-Oriented Programming Systems, Languages & Applications (OOPSLA ’99), Denver, Colorado, USA, November 1– 5, 1999, volume 34(10) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, 1999. ISBN ???? LCCN ????

ACM:1999:PASf

[2900] ACM, editor. Proceedings of the 2000 ACM SIGPLAN Workshop on Partial Evaluation and Semantics-Based Program Manipulation (PEPM ’00), Boston, Massachusetts, USA, January 22–23, 2000, volume 34(10) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1999. ISBN 1-58113-201-8. LCCN ????

ACM:1999:PPA

[2901] ACM, editor. POPL ’99. Proceedings of the 26th ACM SIGPLAN-SIGACT on Principles of programming languages, January 20–22, 1999, San Antonio, TX, ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1999. ISBN 1-58113-095-3. LCCN

????  URL http://www.acm.org/
pubs/contents/proceedings/plan/
292540/index.html.

ACM:2000:PAS

[2902] ACM, editor. Proceedings of the 2000 ACM SIGPLAN workshop on Partial evaluation and semantics-based program manipulation January 22– 23, 2000, Boston, MA USA, volume 34(11) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 2000. ISBN ???? LCCN

????  URL http://www.acm.org/
pubs/contents/proceedings/pepm/
328690/.